A Complete Bibliography of Publications in Marine Ecology Progress Series: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

23 April 2025 Version 1.07

Title word cross-reference

106 [CSH89]. ^{12} [SGSD86]. ^{13} [CS84, SW85, SGSD86, ZCF89]. ^{137} [CSH89]. 13N [FHC88]. ^{14} [BBL84, Cam85, Gei88, KB84, LT87b, LC82, SP84a, SKS^{+}84]. ^{15} [CS84, Gar84, KP89, SG87a]. ^{210} [FBCH83]. ^{237} [PH81]. ^{238} [CSH89, HC80]. 239 [MG85b]. ^{239+240} [CSH89, HC80]. ^{241} [CSH89, FBCH83, HC80, MG85b, MGT89]. ^{3} [Hol88]. ^{3−} [JS85]. ^{51} [BBL84, SFN82]. ^{∘} [JP85]. _{2} [JBS82, JS85, JB86, ORK^{+}86]. _{3} [LT87a, ZCF89]. _{4} [LT87a]. _{4}^{+} [LBH89]. _{4}^{−} [ZCF89]. _{4}^{2−} [JS85]. : [BBL84, LT87a, MC88, OA88]. δ [SW85]. γ [TTH83]. K [Sha88a]. μ [LD85]. n [OA88].

-Glutamyl [TTH83]. -systems [Sha88a].

0-group [KBv89, vdVB86, van86b, vB87].

1 [WAN84]. 10th [Kin88]. 14 [ORK^{+}86]. 1960s [Bre86]. 1980 [LSB85].

1980s [Bre86]. 1983 [HC86a]. 1984 [Ric85a]. 1985 [ESR^{+}88]. 1986 [BNVP87, Gra86b, Gra86c].

22-year [KS84].

450 [Liv88, SWG88].

5 [SKS^{+}84, Tam89]. 5’-hexachlorobiphenyl [SKS^{+}84]. 5’-nucleotid-ase [Tam89]. 50th [Kin88].

6040 [TC81].

A. [SM86]. A2 [Ger84]. Aarhus [JAS88]. abalone [TB85]. abalones [TC88]. abandoned [DS86a]. Aberrant [Gie81]. Aberystwyth [KB87]. Abilities [Las81]. Ability [GB89, WM82, Phi85]. abiotic [HBJ^{+}88]. above [Bre86]. aboveground [PB88]. Abra [AGA89, ID88]. Abramovitch [Cos82]. Abrasion [Mil89]. Abridged [ABB80]. Abrolhos [CHAS83]. abscission [Ger89]. absence [CTG86]. Absorption [BBL84, SFN82, LlJC83, NL86, SPK^{+}87]. Abudefduf [Rus87]. Abundance [BMZZ88, Cam82, CC82, Hod85, SACC88, SKM85, STM88, WR81, BNVP87, Beu87, DG88a, FL88, FL86b, HC86a, KPJ84, KMM88, KP89, LM85, McC89c, MS89, OQ83, Ov87, RO87a, Rus84a, Rus84b, Ryl83, SG85, SD85, Sch87b, Wei89a, Zel85, dB86, BSv81]. abundances [JT85, McG88]. Abundant [Rei81a]. Acanthaster [DRB86, DE88, Gly85, Ols85, SSN87, WWB86, Wil86a]. Acanthasterplanci [ME89]. Acanthopagrus [Pol84]. Acanthurus [FMM87, MP88b]. Acartia [HR87, KMH85, PS88, STW89, SM86, TC81, Tis89, TR87, Uch88, WR89]. accessibility [CL84b]. acclimated [Moe85]. Acclimation [BNF81, OSB85]. accounting [EM89]. accumulate [MM86]. accumulating [SJNL89]. Accumulation [Cha88, FBCH83, FBT84, KDT^{+}88, NLBFD80, RSWJ80, WR82, WR86, HBJ^{+}88, PWM85]. Accumulations [LRH82]. accuracy [DL86a]. Accurate [JBS82, EMHS86, LMO86]. ace [Bro86]. Acetabularia [LB81, Yon80]. acetic [MMS89]. Acetylene [Cap82, SG87a]. acid [ALP84, KZ85, LSB85, Man89, MMS89, RS87, SGB89, YD84]. Acids [JK80a, JK80b, Jør82, SHH81, TW80, ARGL87, CC89, FB86, FF89, FB85, FF86b, Fuh87, GP88b, HK86, LVdVdV85, MJPH^{+}88, MCA89, PMJH84, RHP89, VB85a, VB85b]. Acropora [Cro84]. across [AW86, DCMS89, KEK^{+}84, KBNO85, Rus84a, SF89]. Actinia [HTC84, OTC82, SCT87]. Actiniaria [SBR87]. action [KMH85, KMR87, MB85, NVHHH88, Tam87]. Active [Arm88a, Arm88b, NM89, Pal88, WB86]. actively [DS84]. Activities [SC82, Båm88, GMK85, Hop83, Kin86, LM87a, Sch88a, Thr86, VB85a]. Activity [BAG80b, DM82, HO80, Has81, JP87, KOO80, PGBN81, Pau81, Sch80a, SLO82, AE88, AV87a, AH87, BA83, CA86, CA87, Dob83, EZ89, Gra86a, GMQ^{+}88, Her88, HSB88, JBAH83, Kaa86, LDDH88, LSP84, Liv88, MR86, MR88a, Moe85, MR84, MERW84, MK85, NS87, NK86, OC89, PLM89, PDJ86, Pie83, Ric85a, RA87, SMB85, Sch89a, Tam89, TR89, TTH83, VL88, VBFS85]. Acute [TR87, LL83, War84]. acutifrons [CS89a]. Adaptation [AZ85, CG87, WU82, Bar89, Ber84, Kro88, RO84, Ver89]. Adaptations [FA83, PCR80, CC87b, Fen82c]. Adaptive [SC85, GSC83, TE86a, TH84]. addition [ET84, EMA88, Lev85]. Additions [OFGG82, FER85]. Adenochromes [MG81]. Adenosine [AA81]. adenylate [KS83, TR89]. adherent [HL85]. adjacent [ABT89, DKMR83, Epi88, NVLL88, WB83b, WS88]. Adriatic [HM87, Her88, HPF89]. adspersus [BL88, HO80, LB86]. Adult [OTK80, SS80a, CJS87, MKG88]. Adults [Bre82, Keo89, Qui88]. Advection [McG89b, AAK^{+}88, GP85, MW85, Sch87b]. advective [LH88]. Adventitious [Van89]. aeglefinus [SM85]. aeola [Hic81]. aequispina [Slo85]. aequizonata [CVF89]. Aerobic [KA85, AH84a]. aerobic/anaerobic [AH84a]. Aerobically [HT81]. aestiva [CT85]. Affecting [Lee81, Lan83, PWM85]. affects [Jen85]. affinis [BK82]. affinity [BR85]. Africa [Mac88, BS80c, BS83b, Don87, McC88, Sch89a]. African [GLB87, Rüg89, Sch83b, TR89, Whi89]. After [DB82, BCd84, Bar84, MAWM83, MGT89, NPVW86, RG85, TRL89, Van88a]. Ag [Dah86, FBT84, HL85, JCL86]. against [HDFG88]. Agaricia [Van83]. Age [BAG80b, Lon81a, De 84, DB88, Lam87, PN88, Pet86a, TN86, Wei89a]. age-0 [PN88]. Age-Specific [Lon81a]. agent [Ger84]. Agents [RSWJ80, TE86b]. aggregate [TH84]. aggregates [Cof84, DS86a, Rei83]. Aggregation [SdT86, BP88, Bid88, HH89]. Aggregations [Sch80a, Sch82b, Her88, LPV87, Oka86]. Aggregative [Sch82a]. aggressive [LCS86]. agree [FNH83]. agreement [FF86b]. Aiptasia [Cla85]. air [TL84]. al [AS87]. al-Arab [AS87]. al. [Gra83, JNCS90, Mil90]. ALA [KOO80]. ALA-D [KOO80]. Alaska [DD88, Hei89]. Alaskan [DS86c, FSZ89, SS89b]. alba [AGA89, ID88]. Alcyonacea [BL87b]. alcyonaceans [LCS86]. Alexandrium [Han89]. Alga [BL80, DBFM80, Gun80, OF82, AV87a, BGD87, BH87, CT89, D’A85a, HBJ^{+}88, Jer86, Lün88, NPMW87, RBA86, RBA88, Van89]. Algae [BAG80b, BAG80a, DB82, Gun80, KSR80, SvDB81, WU82, dB81b, DM83, DJTL88, FBCH83, GPCW88, HH84b, Jer85, KM85, ME88, MWM86, NVHHH88, PF86, PST87, Qui88, RG89, Sam83, SACC88, SP89, Thu84a, TRL89, WW88a, WS83, WP88]. algae-free [Thu84a]. Algal [DW87, KSS85, SMO81, SU87, Sch80b, WU82, Bal84, Col86, CH89, FHC88, Hei89, Hig85, Rus83b, SO83a, Sch84a, SR87, SB87, Sto85, TC88, WC88, Zel88, ZM84, Hir86, LTH87, Sam83, TKS^{+}85]. Alkalophilic [MT80]. Alkalotolerant [MT80]. allocation [GMK85, HSB^{+}84]. allochthonous [TP85]. Allometric [JP88]. allometry [Pet86a]. Alloteuthis [RSM88]. Allowable [SA82]. allozyme [FG87]. along [BHJ^{+}89, BK85a, DD88, Kin89, Kro88, Sch88c, Sch88b, SdT86]. Alongside [Rei81a]. aloricate [Tis89]. Alpha [FPRW82]. Alterations [Blo80, Low88, Nel89]. altered [Rye88]. Alternative [Cha82]. alterniflora [DK86, KLN88, NL86, OK89, WBVS86]. alticostata [Fle87]. Alvinella [GH86]. Am [CSH89, Dah86, FBCH83, HC80, MG85b, MGT89]. ambient [RL83, ZK85]. Ambulacra [ZH81]. America [SO83a]. American [GNBO85, HL89, MP86, NJ83, Rii88b, SK82, SHB^{+}83, SC82, WH85]. americanus [BL88, Buc82, BTH^{+}84, CLB88, EC87a, FKW81, FHKR85, Gar89, HPV^{+}87, HL89, LB86, MP86, SHB^{+}83]. Americium [GF82, MG81, MGC82, BTFF85, FT86, MG85b, MGT89, VAF83]. Americium-241 [MG81, VAF83]. Amino [JK80a, JK80b, Jør82, Man89, SHH81, TW80, ALP84, ARGL87, CC89, FB86, FF89, FB85, FF86b, Fuh87, GP88b, HK86, LVdVdV85, LSB85, MJPH^{+}88, PMJH84, RHP89, VB85a, VB85b]. aminoacids [ALAH85]. Ammodytes [BTH^{+}84, MP86, SHB^{+}83]. Ammonia [SA82, AA89, Reg84]. Ammonium [GTG81, PHT85, Pro87, Sel88, WYG85, ARGL87, ES87, FHC88, GD84b, HM86, HGS89, HSD87, JMG84, JP85, JM85a, JM85b, KHHB86, KW88, KP89, Le 86, PM88, VB85a, VB85b, VGJ89, WC88]. ’Amoco [Jac80, Bod88, RG85]. Amonardia [Nil87]. among [BH83a, BTQS88, BP86, CLB88, DSG^{+}89, HC86b, JJN85, JCL86, KSVB89, LC88, SS86a, SS89a, SHPH86, WZ85]. amorphous [Her88]. Ampelisca [CA83]. Ampeliscidae [SB82]. Amphipod [CM87, MR81, NK82, OOS82, OS85, Shu80, AA89, AS87, CA83, Col85, DeW87, Jen85, OOF83, SMB85]. Amphipoda [BC88, HK84, Jac87, SM87a, SB82, TAK88, WR81]. Amphipods [NGV^{+}82, Sto80b, AT86, HDFG88, Kaa86, Rus87]. amputated [GC85]. amylase [Båm88]. Anadara [Bro82]. Anaerobic [KCS89, AH84a, HS86, KA85]. Anaerobically [HT81]. Anaerobiosis [JK80b]. analog [Bar89]. analysed [VK85, WJ87]. Analyses [Shu80, AS88, Bro88, Dec88, KWW^{+}88, MUC84, PGJ83, PCM86, PG88]. Analysing [FCW82, LP88]. Analysis [ACF^{+}80, Ang86, BCF82, Bro82, BNF81, DP81, GAC^{+}88, Ham82a, HWC^{+}88, MR81, OWP80, War88a, Yan82, BGD87, BY83, CG88, EMHS86, GP85, GL88b, ID88, JH87, Jon89, KLE85, Kv88, KBv89, KS87a, KM85, Lam86, MB84b, PLT87, RB84, Sch83c, Sha88b, TP85, Uch88, YY89, BT85, TL87b]. Analysls [KC81]. analyzed [SSA^{+}86]. Analyzer [GGV89]. Anasterias [BB84, Bla84]. anatomy [MP88b]. Anaulus [TB86, WPW88]. Anchoa [LH87b]. anchovy [AI88, JF88, LH87b, SDRN87, TKT86, Wro84]. Anemone [OTC82, BM86b, Cla85, HTC84, HVK85, SCT87]. anemones [PE89]. Anemura [AZ85]. angiosperm [Thu84a, Thu84b]. anglica [HKd88]. Animal [SED81]. Animal-Sediment [SED81]. Animals [Ten82, Cam85, Pie87, TN85, TN86]. annelid [Bar89]. Annelida [GH86, JK80a, JK80b, Kri83b, Kri83c]. Annelida. [GMT89]. Annelids [ML80]. Annual [BL87a, BSF84, FHKR85, SS80a, SS80b, FMM87, JS88, KGS87, NG86, Ov87, SO85, TP89, Wal85a]. annulata [LlJC82, LlJC83]. anomalies [MD87]. Anomura [Zel85]. Anonyx [SM87a]. anoxic [GG88, JB86, SC84b]. Antagonistic [RBA88]. Antarctic [AB89, BB84, BS89b, Dan83, FHS88, KHHB86, KS87b, MS84, Mor85b, Oli84, PSMS87, Pia89, RTK^{+}89, Rei87a, Spi87, SHH^{+}88, Tar81, Wat86, WtD89]. Antarctica [HHE89, HH89, KGS87, NK89, NvB89, PKMS85, SPK^{+}87, CB88, Law86]. antarcticus [Cla82, Dan83]. Antenna [RR82]. antennatus [CS89b]. antennule [GC85]. Anthozoa [SBR87]. anthropogenic [BCA89]. anti [TE86a]. anti-predator [TE86a]. antifouling [Wah89a]. antifreeze [FHKR85]. antillarum [BCd84, HY88, WC88, dB86, dB87b]. Antilles [dB81b]. Antimicrobial [WM82]. Aphotic [PRS^{+}82]. aplacophoran [Sch87a]. Aplysia [Kem81, Paw89]. apochlorotic [CS86, DCJS85]. apparant [BT85]. Apparatus [DB81a, Kv88, Ign87]. Apparent [Fuh81, SWG88, KB84]. appearance [Kin88]. appendages [PG88]. appendicularian [DD85, DT85, DP87b]. appendicularians [DP87a]. Appendix [BAC^{+}88c, BAC^{+}88b, BAC^{+}88a, BAC^{+}88d]. Application [DO86, TL85a, AV87a, BLR86, BMGB88, YY89, dHS88]. applied [DJ84]. Approach [CM82, KC81, AM88, BWH87, Fuh87, HA87, LDER83, Ste89]. April [Ric85a]. Aqaba [SF89]. Aquaculture [SA81]. Aquarium [SFHU82]. aquatic [ABAI87, Col86, Jen87, RT85, TP85, Thu84b]. Arab [AS87]. Arabian [BM86a, BSSW89]. araneus [ALPS83, Ang87, JA85, KA84]. araucana [CA83]. Archaeogastropod [Hic81]. archipelago [JAN85]. Archosargus [DD80]. Arcidae [Bro82]. Arctic [Car87, Hsi88, AA89, CM82, CSW84, CH89, DGHK82, DS86c, DSG^{+}89, GL89b, KC82, LDG87, LSP84, MS88, Par87a, SPL^{+}84, SP85, SACC88, WA87]. Arctica [MW83]. arcuatus [Pau81]. Area [CL82, HO80, Wei82, And86, BSSW89, ESR^{+}88, JT85, Joc89, KMR^{+}88, Ric85a, Sch83b, SM85]. Areas [LRH82, CS84, Möl86, PR84, Pih85, RGH89, Rüg89, SV85a]. Arenicola [MAWM83, MGC82, Rei88]. Argopecten [BEM87, Wol88, WRC83]. argus [HB86]. Ariidae [YALD88]. aristeid [CS89b]. Aristeus [CS89b]. Arm [Gif88]. Armandia [Tam85]. aromatic [HMLF88, STP84]. Arosa [HAOC^{+}86, LNF81, VP85]. Arothron [GR89]. arquata [ZE89]. arsenic [PD86, SC85]. Artemia [VS80, Ama83, GB89, LTS86, LS84a, SJSO82, Sor80, VS80, VCS81]. Artifact [SA81]. Artifactual [Sto86]. Artificial [OE81, BL87b, Lan89, McG89a, RZ85, Van88a, VC86]. Artificially [Bro82]. arval [TL87a]. Ascidiacea [BBBM80]. Ascidian [BSv81, Cow81, SvDB81, vDBS81, Dav89, KD86, Klu84, Rii88a, SN88]. Ascidians [Sto80a, Par84, Sch83a, SW86a]. Ascidiella [Sch83a]. Ascophyllum [JS87a]. ase [Tam89]. Asexual [Las84, OTC82, ME88, Oli84, Wah85]. Ashmole [BBG^{+}87]. Asian [KAK88]. Aspects [Ger89, LHT88, LBH89, Pea82, Shi80, Tan88, VTHVd81, AH84b, Jør83, Wah89a]. aspersa [Sch83a]. assay [BA83, CA86, CA87, FDM86, Joh88, SG87a]. assaying [Mac86c].

Assemblage [CM82, KC82, Shu80, TW82, KB88, Lan89, PCW86, RA89, SRTM83]. Assemblages [AH81, Che82, Din82a, EAM87, Gol85, Hei89, HHKA88, JS86, JS87a, JP88, Jos87, KWG85, KSVB89, Lin88, SLP83]. assess [AAH87]. assessing [Fro87]. Assessment [FWE89, BNF^{+}89, DO86, KAK88, LP88]. Assimilation [Gri80a, Gri80b, Jør82, ARGL87, AGA89, Col84, DL86a, Gle86, GAA89, HL85, KN83, Klu84, Mø83, Rud88, TL85a]. Associated [BCFS82, Col80, Gie81, GL82, Ham80, LNF81, LNB81, Sto80b, DM83, DL83, FHS88, Gly85, GA83, GGA85, GB86, Jac87, Lew84, McC89a, MJ84, MW87, OB88c, Pos86, Sha83, TB89, TP89, TN86, TKS^{+}85]. associates [SG84]. Association [MG81, OS88, SvDB81, SS89a]. Associations [WU80]. assortment [Paw89]. assuming [SWH84]. Astarte [TS88b]. Asterina [Rum89]. Asteroidea [BB84, Sch80a]. ater [SNL82, SFN82]. Atherinidae [Ben84]. Atlantic [BCCH84, Gas87, Le 86, Mee85, SB82, SK85, WR81, BAH87, BTQS88, Che82, Che84, DCJS85, Fev89, FG88, Fui89, GMS85, GD81, GL89a, GOAYH86, HL84, HKS86, LP87, Mit87, NPVW86, PN88, PA84, Sha89, SP88, ZW89c]. Atlantis [Wei82]. Atolls [She80]. atomic [MC88]. ATP [NK89, PGBN81, TR89]. ATP-biomass [TR89]. Attached [BA81, ASSF87, ES84, VBFS85]. attacks [SB85a]. attributes [GAC^{+}88, HWC^{+}88, War88a]. August [ESR^{+}88]. auklets [SR85]. Aulacomya [SNL82, SFN82]. auratus [KC81, KOO80]. Aurelia [Grö88, Grö89, Kee87]. aurita [Grö88, Grö89, Kee87]. austral [BS89b]. Australia [BK81, Bra81, BW81, CH88, CHAS83, HJ84, HSD87, JG89, O’B88b, RDD88, SV85a, SV85b, SP89, WW88a, WY82, YLH86]. Australian [BW83, JH80a, JH80b, Ken87a, Ken87b, LM85, RK82]. australis [O’B88b, Pol84, You81]. Austrocochlea [MUC84]. autocorrelation [MB89]. Automatic [Aik81]. Autonomous [GGV89]. autoradiographic [MG85b, MGT89]. Autotrophic [RHS89, Joc89, MW87, STM88]. autumn [BFP87, GMS85, Gra87]. autumn-migrating [BFP87]. autumnal [Ger89]. auxin [MMS89]. Availability [Gri80a, Sor80, Ten82, BTH^{+}84, CC84, Eva87, Kel88, KH88a, LOvM89, LNS^{+}87, MT85a, MT85b, MR88b, SRL^{+}86, ZK85, ZE89]. available [BS83b, FD89]. averaged [HSB^{+}84]. Aves [WLB84]. Avian [Blo80, HHE89]. Avicennia [AK88, BSC85]. avoidance [AS85, NPVW86, Rub85]. Away [Kem81]. axenic [Man89].

B [DL85]. bachmani [WLB84]. Baciliariophyceae [BP84a]. Background [BCG88]. backreef [WYG85]. Backwaters [NGV^{+}82]. Bacteria [AA81, ABM81, AH81, BA81, Gas85, Gie81, KK89, MT80, MRF80, Rie82a, SHH81, ASSF87, ALH86, Bau85, BMZZ88, BAN89, BT85, BR85, DSS^{+}85, DS84, DNF87, ES84, FA83, GG88, HL85, HHKA88, IEK86, JP87, Kem87, Kem88, Kim85, KNH86, LDER83, LDH83, LN89, LD85, LNL83, MT87, MF88, MF89b, Mon84, NBT88, OS88, PT89, PST87, PHL^{+}86, SDRN84, SN85, SAS87, SS86b, SA89, SGSD86, TR89, TIS85, TKP86, Thi87, VDM89, VGJ89, Wei89b, WC83]. bacteria-sized [MF88]. Bacterial [CFP88, DKMR83, Dem85, FNH83, Fen82d, HSdWM88, KPJ84, KH88b, KSVB89, Li82, MP81, VB85a, VB85b, Wol82, Alo85, ABT89, Bau85, BNF^{+}89, FHC88, FSCP89, FST85b, HL85, HVK85, Hol88, JB86, Kem87, KMAM84, KS87b, LV86, Mø83, NBT88, PLT87, PLM89, RNJ^{+}84, RA87, Rüg89, TL85b, TNR86, KGS87]. bacterial-mat [JB86]. bacterial-sized [NBT88]. bacterio [SSPA89]. bacterio-plankton [SSPA89]. Bacterioplankton [AFHA84, Bjø86, Fuh81, HAHA84, MP88a, TGA89, Alb83, FEHA85, GH83, HM87, HF88, HMLF88, LVdVdV85, LP87, LNS^{+}87, SSA^{+}86, Tam89, VDL88, BF87]. bacterium [GMB87, NNA84]. bacterivores [WH88]. bacterivorous [BB87, PST87]. Bacterivory [Kem88, SRS89, SSPA89]. bactivorous [TIS85]. badionotus [Sv80a]. bag [KBv^{+}83b]. bahamensis [DG88b]. bahia [LS84a]. Bail [Sam82]. Bail-Out [Sam82]. bait [KC89, MAWM83]. Baja [LL81, PE89]. Balaena [HBFA80]. Balance [LH82, NFG82, Pea81a, Bat86, Fis86, Klu84, WAN84]. balanoides [RSWJ80]. Balanus [SW82b]. balthica [Bd89, BMRN84, CL86, HL85, JCL86, Mee85, Ola89, Rei83]. Baltic [GG88, TS88b, BFP87, Dah86, Gas85, GKOW87, JAN85, Joc89, KAE81, KEK^{+}84, KKKW86, KE87, KK89, Lar84, LL83, OWKG85, RGH89, SS87, SPSS89]. Bamfield [McG89b]. banana [SV85a, SV85b]. bands [BL87a, LG88]. Bank [HPV^{+}87, KHW88, Col85, Kan84, LVP^{+}89, ML80, NPSV87, PN88, SM85]. Banks [PH83, SGG87, SHIP86, SPH87, BP85]. Barbados [HY88]. barbicornis [HK84]. bare [Lew84, Mac86b]. Barents [RS87]. Barkley [Rum89]. Barnacle [Nel81b, RSWJ80, KB87, Pag86]. Barnacles [PR88, Jer86]. barren [EC87a, Gar89]. barrens [Him86]. Barrier [AA84, AG82, CLS^{+}82, DCMS89, Din82a, Din82b, Rid88, Rus84a, Rus84b, SR87, She85, WH82, Wil86a, HLS83]. Based [KNL82a, Dou83, DNF87, HL84, NL84, Pea85, SR85, TP85]. basic [Wah89a]. Basin [GD81, HH87, KM88b, SLE^{+}89, CC89, LVdVdV85, LV86, WWD^{+}88, Bri85, Par87b, Roy89, SN89]. basis [SVK83, Sul84]. basrensis [AS87]. bass [Che89]. batch [GKA88, RWK83]. bathyal [GL89a]. Bathybembix [Hic81]. bathymetric [OD89]. Batillus [Shi80]. Batophora [BL80, Mor83]. Bay [CE89, DS86a, EMG89, GD81, Ham82a, Ham82b, HH89, Hir86, Hir87, Hod85, Hsi88, JG89, LHF84, LDG87, NT85, Oka86, PM88, SM86, Tra88, Ued87, WW88a, BEM87, JAS88, KM87a, KM87b, KM89b, LH87b, TGA89, WRC83, AHF87, AC88, And86, Bre86, CL86, CDL87, Clo82, CC84, Epi88, FK84, GG82, GT87, HMF85, HMH89, HHD^{+}84, HP86a, HLH87, Jac87, JT85, KS84, Kel88, KH88b, Law86, Lon81b, MF88, Nic84, OB88c, RHS89, RD89, SPH87, TS88b, TÓ89, Ver86a, Ver86b]. bays [Don87]. Beach [KNL82a, Leb82, HP86b, MH84, NM89, vM87]. beaches [Bod88, MM86]. beadlet [SCT87]. beads [NBT88]. bearing [CVF89]. Beaufort [BC88, BTL86, CSW84, Car87]. Bed [Car82, NFG82, Asm87, FD89, SDRN84, SN85, SN88, STM86, SK84, SP84b, Wil88, ZM84]. Bedford [HH87, KM88b, Par87b, Roy89]. beds [BTL86, BBR88, Hic86, How85, Jac87, MKG88, OB88c, Pos86, RO87b]. before [BCd84, Har85, HH87, Hoe88, MH89]. Behavior [NS80, OSB85, OFGG82, OWP80, RYD84, AP83, CC87b, HH84a, Mor87, PL89, SC85, SH84, Ste86, SHH^{+}88, TC88, Van86a, WP87]. Behavioral [FWB89, OOF83, SVK83, Sul84]. Behaviour [AAH87, HS86, Jen81, Jen82, vDBS81, AAN87, Ans89, AI88, AS83, ER87, Han89, Hic86, JF88, Jon89, KSM88, LH85, MW83, MK85, O’B88a, O’B88b, Pie83, TE86a, vdVB86]. Behavioural [MR81]. Belgian [BF87, LMO86, LM87a]. Belize [HLS83, LC83]. below [FH83, RS87]. belowground [PB88]. belts [MJ84]. beneath [Car87, MS84]. Benguela [FC82, AG81, Abr83, Car82, PAH89, PWMl89, Pro85, RM88, SF85, Stu86, VDL88]. bennetti [ZH81]. benoiti [SCF^{+}82]. Benthic [BCF82, CD81, Gas87, Gra87, GP82, HPF89, HW82, Jac80, KBNO85, LWG^{+}81, MCM84, MGC82, MPR85, MBP^{+}86, NN85, OSM82, PB81, Rüg89, SB82, Spa80, Sto80a, TR89, TD81b, WU80, de 80, dB81b, ARGL87, AH84a, Bar88a, BAC^{+}88c, Bje88, Bro86, Col87, CYC83, DA84, DeW87, ER87, Eme89, FK86, FER85, GHS85, GL89a, GC83, GAC^{+}88, GMF88, GFM89, GM89, Gri88, HP86a, HWC^{+}88, Hug84, JS87b, Kem88, KAK88, LM85, Mac88, Mar84, MAWM83, MR88b, MGM86, PGJ83, PL85, Pie87, PH83, Pos86, Pos87, Ryg85, Sch83c, SHPH86, Sch88b, Sei87, SO85, Sha88a, Sha88b, Sha83, SLP83, SO87, SG87b, Sto89, SP84b, SCSD86, VAF83, VP85, VA85b, WA87, WCGG86, WE88, WZ83a, WF83, GFM89, GM89]. Benthic-pelagic [KBNO85]. Benthodytes [BCS84]. Benthogone [BCS84, TMBG85]. Benthos [Clo82, RTW81, SS80b, ST82, DOB^{+}89, LDMT86, Rei87a, RB84, War88a, ZWO89]. Benthosema [Sam88]. Bering [Bai89, DCNS84, GMF88, GFM89, GM89, GCBM82, LBH89, Sch82b, SR85, SRL^{+}86, Wal85b]. Bermuda [HL82, Sv80a]. Bermudian [SKS^{+}84]. Bertalanffy [DJ84, Pat88]. beryllina [Ben84]. Best [VP88]. Best-estimates [VP88]. Between [ABM81, BS80d, Din81, FC82, Sto80a, WB80, And86, Bai89, BBH^{+}88, DE87, DA84, DAC^{+}83, FF86a, FJ88b, FF86b, Fuh87, Grö88, HS85, Har86, HTC84, HHE89, HL89, HG86, Jer85, Jer86, JMSK86, KM88a, KMH85, KMR87, KW88, LCS86, Law89, LSD85, LCP84, LNL83, MUC84, McC89a, McC84, MERW84, Pag86, PT89, PST87, Pol84, RMO88, Rud88, SW86a, SP86b, SSA^{+}86, SF89, SG88b, TD89, VB85a, WDS^{+}84, Wil87, WS88, WW88b, WW83]. biases [HKB87]. bicarbonate [Gei88]. Bight [Cus89, Hov89, JAS88, Joc89, Lan83, SB82, Bar86, Bar88b, FEHA85, Gas87, KLE85, LSB85, OS80, Sch87b, SCSD86, TB85]. bimaculatus [Amb82]. binder [DL85]. Binding [NL81, DL85]. bio [GMF88, Joh88]. bio-assay [Joh88]. bio-mass [GMF88]. Biochemical [Add88b, Add88a, Buc82, ELK84, OTC82, CG87, MCA89, PCM86, SDMN88]. bioculata [HBJ^{+}88]. Biodegradation [KNL82a]. biodeposition [KE87]. Bioenergetic [CL84a]. Bioenergetics [KMH85, Fen82b]. Bioerosion [MR88b, HLS83]. Biogenic [LvB87, AN88]. Biogeochemistry [HL82, Sei87]. biogeographic [SG84]. bioherms [FH83]. Biokinetics [BTFF85, GF82, FT86]. ’biological [CG88, FK86, FB85, OFGG82, PSB^{+}81, PHL^{+}89, BCG88, Col86, DK85, GMQ^{+}88, HKS86, Mil85, MPR85, Möl86, OWKG85, TC87, dKB88a]. biologically [HA87]. Biology [BAC^{+}88c, BAC^{+}88b, FMM87, WF80, CFH89, FJ88a, GGA85, Him86, Ste89, TMBG85, Wil89]. bioluminescence [ABJS88, BNF^{+}89, LGS^{+}89]. Biomass [BvdEH^{+}82, Car82, CH88, Clo82, HH89, LDDH88, VN82, dB81b, ASMV88, AW86, ALPS83, CSDZ86, CC84, Col84, DG88a, DRCA83, Dua89, FSCP89, GG88, GHS85, Gra86a, HGS89, KSS85, Kel88, KNH86, LVdVdV85, LV86, LT87a, MB84b, NL84, Red83, RNJ^{+}84, RNH^{+}88, Sch83c, Sch88c, Sch88b, Spr84c, TR89, TGA89, VB85a, VB85b, Bar88b]. biomass-size [Sch88b]. Biomechanical [BGD87]. biomonitors [PR88]. biosynthesis [KNH86]. biosynthetic [BA83]. biota [HG86]. Biotic [Rei83, WZ85, BB85, HBJ^{+}88]. biotope [BLR86]. bioturbated [ES87]. bioturbation [Rei88]. biphenyl [STP84]. Biphenyls [ITBW81]. bird [Gra84, RM88]. Birds [Sch81b, AV87b, Eva87, NG86]. birostratus [TB86, WPW88]. Birth [SS82]. Biscayne [LH87b]. bisphosphate [SP85]. bispinis [Dan83]. Bivalve [Gri80a, Bd89, CL86, CSW84, CVF89, CP87, DSS^{+}85, Jen85, JCL86, JFK^{+}86, LN89, Man89, Man88, MB89, Rei83, RR87, RNK87, VAF83, WW83, WWL83]. Bivalves [KM81, Leb82, DO86, HNR^{+}88, HH84c, MD87, Rii88b, ZJ83]. Bivalvia [AGA89, Bro82, GG82, LlJC82, LlJC83, Luc84, Mee85, SC84a, Sha89, TO88, TS88b]. Black [BSF84, Gri80b, BL87a, Kro88, SH84, WLB84]. blade [Ger89]. blades [Rus83a, Rus83b]. bleaching [Har85]. bloom [Bar84, BNF^{+}89, BF87, BHJ^{+}89, Coh85, FSZ89, FHS88, GKA88, Gra87, HH87, KS87a, LVdVdV85, LM87a, NK89, Rie89, Roy89, Sch83a, SPK^{+}87, SS87, TRL89, IBMD81]. blooms [BM86a, HK86, Par87b, RS87, ZW89b]. Blue [HS81, OF82, PGV^{+}89, BCA89, EMG89, HLH87, HK84, LOvM89, Ov87, Tho84, WAH89b]. Blue-Green [OF82]. bluefish [OSB85]. Bly [WWD^{+}88]. Boccardia [BK81]. bodies [EM89, KHE^{+}83]. Body [BS89a, SC82, ELK84, Klu84, LS88, Mac88]. Bombay [BSSW89]. boobies [And89a]. border [SMN86]. bordered [AZ85]. Borderland [CC89]. boreal [Arm88b, Bam83, Joc89, Möl86]. boring [PB89]. borne [OA88]. Botfylloides [BBBM80]. both [BT85]. Bothnian [TW82]. Botryllus [BBBM80]. bottle [GD84a]. Bottom [Eva82, OOS82, Rei81b, SS80a, SS80b, Amb84, Amb86b, CA85, Gri88, GKM^{+}87, KN84, LVP^{+}89, Mac86a, MPR85, dB86]. Bottoms [Thi81, ET85]. boulders [McG88]. bound [HL85]. Boundary [Rei81b]. Bow [LC83]. Box [PLT87]. boxcore [Dem85]. Brachiopod [TW88]. Brachyura [Cho86, McD82, Pau81, BS83a, FW86, HC86b]. brachyuran [Law89, Sul84]. Brackish [BS80d, HO80, TM80, AH86, HS86, Hop83, PAN89, RBA83, TAK88]. Brackish-Water [TM80, HS86, RBA83]. Branchial [GHCH81, MG81]. branchialis [Ans89]. branching [Van89]. Bransfield [HHE89, LvB87, NK89]. Brazil [Jac87]. breadths [Pea85]. Bream [DD80, WF80, Pol84]. Breeding [Din81, FC82, Kle80, VS82, Rum89, TP89]. Bremerhaven [FSBOG84]. Brevoortia [FAG89, GHC83]. brightwellii [HH84a]. Bristol [BK82, CW82, NLBFD80, WU80]. British [Alb83, BA81, JB86, Lar86, McG89b, PSB^{+}81, Rum89, BCFS82, SBR84, TW88]. Brittany [RG85, RGLK87, Vid87]. brittle [ME88]. brood [Mau88]. Brooded [OTC82]. brooding [BB84, Hig85]. Brown [DBFM80, Gun80, BS83b, CT89, DJTL88, Gle86, KD84, Lün88, NPMW87, SP89, TE86a, Tra88, Van89]. Browns [HPV^{+}87]. Browsing [NK82, Las85]. Brunswick [FHKR85]. Bryozoa [Rub85]. bryozoan [Keo89, PB85b]. bryozoans [Lan89]. Buccinum [HGH87]. budget [Abo84, Bar88b, BS83b, DS86c, FB87, NG86, RR87]. budgets [HB85, SLE^{+}89]. Bugula [Keo89]. building [Sch84b]. bullkelp [CH88]. buoyancy [NPVW86]. burrow [DG88a, EZ89]. Burrowing [WP87, LB86, Pos86]. Burrows [LB86, Rei81a, KM83]. butter [GB84]. butterflyfish [Las85].

C [BLK88, SGSD86, Liv88, BBL84, Cam85, CS84, DL86a, Gei88, Hic84, KB84, LT87b, LC82, MT87, ORK^{+}86, SW85, SP84a, SKS^{+}84, Thi87, VA85b, SGSD86, JP85]. C- [BLK88]. C-14 [ORK^{+}86]. C-bicarbonate [Gei88]. C-derived [LT87b]. C-Formaldehyde [LC82]. C-labelled [SKS^{+}84]. C. [Bro80, LN86, Pau81]. C/ [SGSD86]. Cadiz [Ama83, Bod88, Jac80, RG85]. Cadmium [Fis89, Kui81, LGK87, MKS80, MS88, NLBFD80, NL81, RSWJ80, SHG89, WR82, WU82, BV85, Bje88, Cha88, Fis83, Fis86, Fis88, HBJ^{+}88, KGV84, RWK83, RBA86, RBA88, TAK88, TR87, WR86, WRKD84]. Cadmium- [NL81]. cage [PW87]. caging [SW84, Sto86]. Caisson [FSBOG84]. calanoid [Abo84, AM88, DGK88, Fro87]. Calanoida [PS88, PL89, STW89]. Calanus [BR83, HM86, Hir83, HSRM86, Pet86b, Pet88, RHP89, Tan88]. calcification [Cro84]. Calcium [GCK82]. calculated [DW83]. Caledonia [BLR86]. calibration [SG87a, dHS88]. Calidris [BFP87]. Calidrisalpina [BFP87]. California [FEHA85, LWE87, PE89, SCSD86, BK85b, CC89, DS86a, EHF^{+}81, Gun80, JJN85, KFPH88, KB88, LWG^{+}81, LL81, McC84, MH89, NT85, Ohm88, Oka86, OQ83, PH87, SLE^{+}89, SB85a, TB85, TC88, De 84]. californica [BGD87, Paw89]. californicus [OFGL88]. californiensis [Pos86]. caligid [NPSV87]. Callianassa [DG88a, KM83, Pos86]. Callinectes [HLH87, HS81, Ov87, Pau81]. Callionymus [Sog84]. Caloric [Ten82]. calyculus [Yon80]. CAMP [AA81]. camtschatica [SS89b]. Can [SC82]. Canada [BA81, EC87a, Hsi88, JNCS89, JS87a, Roy89, STM86, TW88, ZW89b]. Canadian [LDG87, SACC88, MS88, Par87a, WA87]. Cancer [HL89, Joh81a, Joh81b, Joh82, LH85, Law89, MGT89, Thr86]. Cannibalism [DGK88, KO87, Bai89, KSM88]. canopies [Ken88]. canopy [FF86a, HH84b, SO83a]. canopy-forming [HH84b]. canyon [KCFS89, KCS89, WBAN85]. capacities [AAK^{+}88]. Capacity [DL82, NL81, BP84a, BP85, FDM86, KWG85, LSP84, SP85]. Cape [Don87, JF88]. Capelin [Ste89, MFS83, TL87b, TL87a]. capensis [GLB87, JF88, SDRN87]. capistratus [Las85]. capitata [Alo85, Cam85, CT84, GMT89, HT81, Pet84, PT84, Tsu87]. capitatus [Sto85]. Capitella [Alo85, Cam85, CT84, GMT89, HT81, Pet84, PT84, Tsu87]. Capitellidae [Tsu87]. capricorni [GB86, You81]. capture [Bar88a, BB84, Pau83, ZE89]. captured [KHE^{+}83]. Carbohydrates [IBMD81, EBH^{+}83, EB86, LDER83]. Carbon [Abo84, ABM81, Bra87, CSS82, DSW89, FD89, FT82, GTG81, GCBM82, KNL82a, Mor87, Rei81b, RMZ82, RR87, SP84a, SH89, WS88, de 80, ABT89, BB87, BAN89, BR83, Cam85, CSS83, CD88, DL86a, Dou83, DS86c, EWM87, ER87, FB87, Ger89, Gle86, GM89, HK86, HP89, HB85, HHN^{+}83, JMG84, Kem88, KLN88, KK89, LB87, LT87b, LBH89, Mø83, NL84, NL86, PT89, Pet89, PBP85, RWK83, Red83, Rei87a, Riv89, RD89, SN85, SP86b, SW85, SLE^{+}89, SHIP86, STM86, TL85a]. carbon-bacteria [BAN89]. Carbon-controlled [SH89]. carbon-specific [Dou83]. carbonate [KCS89, OC89]. carboxylase [LSP84, SP85]. carboxylating [MLDT88]. Carcinus [AAH87, AAN87, BT88, BV85, CL88, CH84, RN89, ap 86]. Cardiidae [GG82]. Caribbean [Ger84, Las85, PF89, SO87, Van88a, Wil87, Wil88]. Carnivore [RK82]. Carnivorous [SF80]. carnivory [SE85]. Carolina [BTL86, Het89, DCB^{+}86, Leb82, MW85, OK89, WWD^{+}88, WS88]. carotenoid [KM85, Nel89]. Carpentaria [SV85a, SV85b]. Carrie [LC83]. carrying [AAK^{+}88]. case [Eml86, KMDS85, PHL^{+}89]. Caswell [GL88b, LP88]. Catalonian [BCA89]. Catastrophic [SS80b, Lam86]. catfishes [YALD88]. catharus [WH87]. Caulerpa [WBAN85]. Caulerpales [PF86]. causal [Sch83c]. cause [CT84]. Causes [DB82, HTL85, BB85, Ger86]. causing [Sch83a]. Cavanaugh [Gra86b]. Cavern [Rei81b]. Cay [LC83]. Cays [KHW88]. CCD [ABJS88]. Cd [ELK84, FBT84, HL85, RSWJ80]. Ce [Dah86]. celebrates [Kin88]. Cell [BAG80b, BB87, BH83a, CC86a, CC88a, CG87, FNH83, HH84a, JM85a, TL85a, TB86]. cell-specific [FNH83]. Cells [Fuh81, BP86, CC86a, Moo88c, PHT85, Pet89, RNJ^{+}84]. Cellular [Moo88b, Moo88a, AFHA84, FBCH83, LT87b, Low88, VMM^{+}88]. Cellular- [Moo88b, Moo88a]. Cellulase [Moe85]. Cellulose [WM82]. cellulosic [KLN88]. Celtic [BR83, JP84, JOP86, JP87]. census [JAN85]. Centimeter [MF89b]. Central [Shu80, Wei82, BHP^{+}86, CB88, CC87b, HLH87, MP88a, MH89, OS84a, PH87, Rid88, Rus84a, Rus84b, SO84, SR87, She85, SKM85, WH87, Wil86a, CC82, GG88, MS89, SMO81]. Centropages [CT85, Tis89, Van86a, WR89]. cephalic [PG88]. cephalid [OOF83]. Cephalopholis [SF89]. Cephalopod [Amb82, GHCH81, GF82, MG81]. Cerastoderma [Bro80, Jen85, LG88, MG85b]. Cerithium [LN86]. Certain [ABB80]. Chaetodon [Las85]. chaetognath [Øre86]. Chaetognatha [Kim84, Pea80, Pea81a]. chaetognaths [Øre87]. Chaetopterus [Rii89]. Chagos [She80]. chain [AGCD86, FK84, GCD87, Mon83, NR89a, VB85b]. chains [CSD88]. chalcogramma [Bai89, DCNS84, Pau83]. chamber [Bau85]. Chang [NVLL88]. Chang-jiang [NVLL88]. Change [GP82, Dal85, PGJ83, MM81]. changeover [FMM87]. Changes [ALPS83, BCFS82, FK84, FF89, FL86a, FST85a, Mag88, OTK80, Pos86, SGB89, VN82, You81, ZW89b, dB86, Auf88, BB85, ELK84, GC83, ID88, KH88a, Kri83c, Lar86, MN87, MCA89, MT84, NR89b, Pia89, PCW86, Reg84, RG85, RNJ^{+}84, RZ85, STW89, SCSD86, TAK88, WW83, ZK85, FST85b, PSMS87]. Changing [HL80, NK81, CL86, JT85, KHE^{+}83]. channel [LV86, BK82, CW82, HHN^{+}83, ID88, NLBFD80, Øre86, WU80]. Chanos [KBU85]. Characteristics [Lee81, AFHA84, BA87, Dan83, DTLN85, HGS89, JKMR84, ME88, SN89, Slo85, TB86]. Characterization [NK86, RG82, BHJ^{+}89, CVF89, TE86b]. charge [ABJS88, KS83, TR89]. charge-coupled [ABJS88]. Chatognatha [Pea82]. Cheilodactylidae [McC89b]. Cheilodactylus [McC89b]. Chelating [RSWJ80]. Chemical [CLS^{+}82, FG83, HDFG88, Jen81, OFGG82, Par84, PF86, SG88a, Sto80a, ALPS83, Auf88, DK85, FB85, FF86b, Ger84, GKOW87, HG86, KWW^{+}88, OWKG85, PH86, RGH89, Riv89, SP89, SDMN88, WZ83b, WP88, ZK85]. Chemically [HSRM86]. Chemically-mediated [HSRM86]. chemicals [SWG88]. Chemistry [BAC^{+}88a]. chemo [JMG84]. chemo-lithotrophic [JMG84]. chemoautotrophic [SGSD86]. Chemodetection [HGH87]. chemolithotrophic [VGJ89]. Chemosensory [SAS87]. chemostat [PST87, TP85, Thi87]. Chesapeake [FK84, HMF85, HMH89, HLH87, Lon81b, MF88, OB88c, RHS89]. Chile [CB88, CC87b, MP88a, MJ84, MS89, OS84a, OS84b, SMO81, SO83b, SO84, VCS84, WAN84]. chilensis [WAN84]. China [IEK86, Lat86, NVLL88]. Chione [MB89]. chitwoodi [FT82]. chl [LDKR88]. chl-a [LDKR88]. Chlamys [Fev89, RZ85, SC84a]. Chlorinated [OS80, STP84]. Chlorine [RRB^{+}80]. Chlorine-Produced [RRB^{+}80]. Chlorophyceae [BP84a, RR82]. Chlorophyll [Aik81, BAG80b, CD81, KM85, LDG87, Smi81, SF86, de 80, AH84b, BP85, DE87, Dec88, ESR^{+}88, FL86a, JH87, JD85, MB84a, MF89b, NVHHH88, PD88, Red83, Roy89, RHP89]. Chlorophyll-a [CD81, de 80, BP85, FL86a, JD85, MB84a, MF89b, Red83]. chlorophyll-derived [PD88]. chlorophyll-enriched [DE87]. chlorophyll-fluorescence [AH84b]. chlorophyll-pigment [Dec88]. Chlorophyll-to-Cell [BAG80b]. chlorophyll-type [Roy89, RHP89]. Chlorophyllase [JH87]. Chlorophyllous [RLB89]. Chlorophylls [ABM81]. Chlorophyta [Mor83, WBAN85]. chlorophyte [FF89]. choanoflagellate [GL88a]. Choice [Sto80b]. choreotrichous [SRS89]. Chorismus [Cla82]. Choromytilus [Gri80a, Gri80b, SDRN84, SN85]. Chromadorita [Jen81, Jen82]. chromatography [KM85, VK85]. Chromium [ZJ83, BMRN84]. Chronic [LGP88, OFGG82, SM87b, LL83]. chronically [CSH83]. Chronology [HB81]. Chroococcoid [IM86, IM88]. Chrysaora [FK84]. Chrysochromulina [BHJ^{+}89, EM89]. Chrysophrys [KC81, WF80]. Chthamalus [KB87]. Chukchi [GMF88, GFM89, GM89]. chuss [SLO82]. Ciliary [IR81, JKMR84]. Ciliate [CTG86, FJ88a, GH83, Han89]. Ciliated [SED81, ASSF87]. Ciliates [ABB80, Gas85, Hei88, Jon86, Jon89, Kem88, SRS89, STM88, Tis89, TNR86]. ciliatus [CL84b]. Ciliophora [Din81, FJ88a, Gif85, Jon86]. Ciona [Sch83a]. Circa [Has81]. Circa-Tidal [Has81]. Circadian [CK89, RN89]. circatidal [RN89]. circulation [HS88a, HMM87, KE87, MH84, SSM^{+}84]. Circumpolar [KZ89]. Cirripedia [SW82b]. cladoceran [JKJ88, PY86]. clam [FS84, LGP88]. clarity [HS88b]. class [De 84, GG88, LDMT86]. classes [JH87, PW87, Par87b]. classification [VK85]. Claude [Mon89]. clausi [Tis89, WR89]. Clava [ACF^{+}80]. clear [HS88b]. Clearance [KSF89, MF88, LNS^{+}87]. Clibanarius [VS82]. climax [Van88b, ESR^{+}88]. Clinocardium [GG82]. Clinus [Gib88]. clocks [RN89]. clone [BSP89]. Close [Fuh87, TL85b]. Closed [SA81]. closely [Van83]. Clue [HBFA80, FS84]. Clupea [Che82, Che84, CL82, KMR87, Lam87, McG89b, MK85, SP88]. Clupeidae [FAG89]. Clupeoid [IR81]. clustering [O’B88a]. Cnidaria [Kee87]. cnidarians [Kar86]. Co [BH83a, BB84, McD82, VB81, Kan84, Yos86a, JBS82, JB86, ORK^{+}86, VGJ89, Dah86]. co-existence [Yos86a]. Co-Occurring [McD82, VB81, Kan84]. Co-operative [BB84]. Co-variation [BH83a]. coagulation [KJ88]. coarctatus [Ang84, JA85]. Coarse [KAE81]. Coarse-Scale [KAE81]. Coast [BS80c, BCA89, BS83b, BSF84, BHJ^{+}89, FL88, GB84, HMM87, Kin89, LNS^{+}87, Mac85b, MKHD85, PPCR86, Sha89, TO88, Vid87, WEHF85]. Coastal [BM86a, EHF^{+}81, HW82, LH82, Liv82, Tar81, ALP84, AS86, AH87, BF87, BBH^{+}88, CT85, DR86, DV83, DOB^{+}89, ES87, FVDBD87, FDM86, Gar84, GKOW87, GD84a, Gri84, HSW89, JAS88, KWG85, KN88, KE87, KN84, KSF89, LMO86, LM87a, McC88, MBP^{+}86, NK86, NN85, O’B88b, Ohm88, OQ83, OWKG85, OQM^{+}87, RNJ^{+}84, RR87, RA87, SO83a, SO83b, SP86a, SG87a, Sel88, SRS89, SdT86, SG87b, Tis89, WC83, YALD88, ZW89b]. Coasts [GW81, Sha88a]. cobble [DW87]. Cochin [NGV^{+}82]. Cockle [GG82, MG85b]. Cocoon [SBR84]. cod [Kan84, LVP^{+}89, MS88, NPVW86, PN88, STP84]. Coelenterata [CLS^{+}82, KQ82, LCS86]. coelenterate [vdV85]. coelenteron [HVK85]. Coenobita [AZ85]. coevolution [DRB86]. cofactors [CA86]. Cohort [LH87b]. Cold [JS87b, BAH87, CA83, KSF89, LP87, Mit87, SMB85]. cold-core [BAH87, LP87, Mit87]. cold-temperate [CA83]. cold-water [SMB85]. collected [VL88]. Collections [SJSO82]. collectors [RZ85]. Colleen [Gra86b]. ’Collisella’ [HD89, PAF86]. Cololabis [OTK80]. colonial [Dav89, GSC83, Kar86, Ryl83, SW86a, Sto86, SB87]. Colonies [BBBM80, LMO86, VA85a]. Colonization [KL84, Mac86b, VC86, DeW87, VDM89, Wah85]. Colony [Cow81, DE89]. Color [BM86a]. colour [HTC84]. Columbia [BA81, BCFS82, Alb83, JB86, Lar86, McG89b, PSB^{+}81, Rum89, SBR84, TW88]. Column [AFF^{+}82, DGHK82, DL82, Sch81a, HM87, HSD87, JB86, WZ83b]. columns [BSSW89, BH83b]. Comanthus [ZH81]. Combinations [MR81]. Combined [BBBM80, BNF81, DD80, IBMD81, PCM86, Bak88, Wah85]. commensal [MUC84]. commensalism [BT85]. Comment [Mil90, JNCS90]. comments [Gol85]. common [AS83, GB86, HTC84, KM89b, SCT87]. Communities [Bou81, Bul82, CW82, GP82, GL82, Kui81, LWG^{+}81, LNB81, LR80, Moo81, OOS82, Tar81, Thi81, WH82, Alo87, Amb84, Amb86b, BG84, CA85, DG88a, Dua89, Eme89, Gal87a, Gal87b, GABM84, GPCW88, GC83, GSC83, GKM^{+}87, Hug84, JS87b, KHW88, KGS87, KBv^{+}83a, KBv^{+}83b, LHF84, Lat86, MB84b, MB85, MW87, NL84, NR89a, OS84b, PKMS85, PSMS87, Par89, PB85a, PGJ83, Pia89, Pos87, PCH85, Rid88, Rüg89, SC85, SSN87, SG85, Sch88b, SPH87, Sor85, SP84b, TKS^{+}85, VS84, WW86, Wil86a, Wil86b]. Community [AG81, BCF82, Blo81, BCFS82, Eva82, GA83, Gri83, HP86b, HFK82, HW82, Jac80, LNF81, LH82, Mac81, MB87, NFG82, OE81, OD89, SW82a, SMO81, SS80a, WY82, WG84, WBJA81, AB89, Amb86a, Ang86, Asm87, Bal84, BAC^{+}88c, BV83, BH86, BY83, BB85, CDL87, CSDZ86, DS86a, FK86, FST85b, GL88b, GAC^{+}88, GFM89, HWC^{+}88, HPF89, ID88, JWC83a, KLE85, KN84, Ken87a, Ken87b, KM89a, KAK88, LH88, MKHD85, NT85, NM89, NN85, Oka86, OWZ89, PLM89, PB85b, Pos86, PDIF85, Sam83, SO83a, SW84, SHPH86, SR87, TN85, TN86, TNR86, Van88b, War88a, War88b, War88c, WCC^{+}88, WB83b, WF83, YY89]. compact [DR86]. Comparative [CLS^{+}82, GMS85, GHC83, HKB87, PY81, WB83b, GPCW88, PR88, GP82]. compare [STM86]. Compared [Fuh81, DR86, NBT88]. comparing [Bre86]. Comparison [AH87, DP87a, ES84, Kri83a, Las81, Mar83b, MCGP82, NVHHH88, Rie89, Tam87, Tow84, BBH^{+}88, DD88, FF86a, HF87, HPC89, KW88, McC89a, Mee85, ORK^{+}86, PCH85, Sha88a, LT87b, MT87]. comparisons [LH85]. Compartmental [LH87a]. compartments [Asm87]. Competent [PRS^{+}82]. Competition [DA84, RO84, Wil87, BT85, BM86b, FH89, Gar89, HC86b, KM89a, Lan89, SW86a, Sma83, Smi86]. Competitive [LCS86, SMO81, Yos86a, Rub85]. competitor [Lev85]. Complex [Lev85, DL85]. Complexation [RSWJ80]. complexing [TE86b, ZW89b, ZW89c]. complexity [RO87a, Rus87, Rye88]. component [IM86]. Components [LNV81, Pea81a, AAN87, BH85, DCJS85, KN83, MW87, dB86]. Composition [Buc82, CC82, CD81, KKT81, KB80, RR82, RS80, Abo84, ALPS83, BV85, CL88, CDL87, CG87, DL86a, DG88a, EAM87, FT86, JP85, KMH85, KS87a, Law86, LTH87, MJPH^{+}88, MCA89, Mø83, Pos87, PCM86, Riv89, SR87, SGB89, TKP86, TN86, WB83b, WZ83b, ZK85]. compositions [RGLK87]. Compound [BSv81, SvDB81, vDBS81]. compounds [SP86b, SAS87, VK85]. comprehensive [Gol85, RT85]. Comrnunities [LL81]. concentrating [DP87a]. Concentration [CSH83, DD80, HC80, Ham80, OWKG85, PGBN81, Abo84, Bam83, BHNI89, Bro86, BR85, JF88, Mor87, RBA86, WR84a]. Concentrations [GHCH81, JK80a, LN81, AHF87, BBH^{+}88, BW88, CL86, DRCA83, ELK84, FF86b, Pau83, RWK83, RGLK87, SG87b, SKIR89, TW88, WR87, WRKD84, ZM84]. Concept [WBJA81, BDKK83, Thi87, CC88b]. conceptual [BGR86, Pal88]. conch [SLMB88, Sto89]. Conchocelis [CSB81]. concomitant [CS86]. Concretion [GCK82]. Condition [NPVW86]. Conditions [GLD82, KB80, MM81, OF82, RHvd81, Auf88, CP84, GG88, LTS86, Low88, NM85, SCSD86, TNR86, WRC83, WF83]. Confirmation [OTC82, SGSD86]. conform [SG85]. conica [TP81]. consensus [PHL^{+}89]. Consequences [GB86]. considerations [FB85]. Consistency [PCW86]. consistent [Thi87]. consists [SBR87]. conspecific [Kee87, Qui88]. Constant [HL80, LL81, DL85, HGD86]. constituents [Bie85, TBRBA85]. constraints [Col86]. constricta [MUC84]. constructed [Alo85]. Construction [dHS88, MGM86]. Consumers [Fen82d, STM86]. Consumption [BS80c, BB80, BNF81, NG86, RS87, SK82, AA89, AGA89, Bau85, BK85a, Col87, Dal85, GAA89, JMSK86, KN83, MW87, PR84, Pih85, RDF^{+}88]. contained [DR86]. containing [BA87, RWK83]. containment [Cus89]. contaminant [SKIR89]. contaminants [Low88]. Contaminated [MGC82, WU82, VL88]. contamination [Auf88, CDL87, Lam86, VMM^{+}88]. Content [Cam82, BNVP87, Fis83, JD85, Kiø89, MC87, Mor87, PBP85, SA89, Uch88, ZIO84]. contents [DW83, DT85, GNBO85, KM83, Tur84]. Continental [CL82, FR80, LWG^{+}81, SB82, YLH86, BC88, Epi88, HJ84, Rus84a, Tur84]. Continuous [Kle80, Bjø86, DTLN85, MF89a]. Continuously [VS82]. contorta [RK81, Rie82b]. contrast [CS86]. Contrasting [And89b, Dav89, Leb82, Ola89, LN86, ZE89]. Contribution [LJIO84, Tis89, IM88, WW88a, WC88]. contributions [WS88]. Control [Clo82, OSM82, RK81, RS80, CH89, Fro87, GLT^{+}86, HVK85, LCF89, Le 86, LT84, Lün88, Mil85]. Controlled [ITBW81, LDER83, SH89]. controlling [SM86]. Controls [CPM89, She82, Ten82, Mon83, SO87]. conventional [RGLK87]. Convergence [TCA82]. Conversion [NLL81, SNL82, BB87, HB85]. Convoluta [DG82a]. cope [BBSS89]. cope-pods [BBSS89]. Copepod [BNVP87, BK82, FP82, HH87, Lon81b, SR85, Ste86, TC81, Abo84, ASMV88, Ans89, Daa87, DCNS84, Geo85, Gil85, HR87, KMH85, Kiø89, Mar84, Mor87, MK85, MR83, OFGL88, Ohm84, Ohm88, PV85, Pet86b, Pet88, TR87, UKTA83, Van86a]. Copepod-based [SR85]. Copepoda [BS89a, BSS89, OFGL88, PS88, PL89, STW89]. Copepoda. [CS89a, HHG84, Nil87, VB81]. Copepods [KKT81, KC82, Kle80, Lon81a, PB81, Rie82a, AM88, AN88, BWH87, CT85, DGK88, DW83, Dec86, Fro87, GC85, GP88b, Hic86, HSRM86, KM89b, PD88, PG88, Sch83b, SM87b, SM86, TB88, Tur84, Uch88, Ued87]. Copper [NLBFD80, NL81, OFGL88, WR82, WU82, AE88, Cha88, OWKG85, OA88, Ryg85, WCC^{+}88, WRKD84, ZW89b, ZW89c]. Copper-Binding [NL81]. copper-complexing [ZW89b, ZW89c]. Coral [Bou81, BY81, BR82, BY83, Bul82, Col80, Din82a, FV82, FH83, KB80, Las81, Lat86, LR80, OE81, Pea81b, Rei81b, Rob82, SW82a, She80, She82, Wil82, AWSB83, AA84, BCd84, BG84, Cof84, CHAS83, Cro84, DE88, Gal87a, Gal87b, Gly85, GA83, GGA85, Gri83, GR89, Hod85, Hop88, HSB88, JWC^{+}83b, JWC83a, Kar86, KMD87, KMM88, LDDH88, Mar83a, McC89a, PB85a, PDJ86, PF89, PDIF85, RB84, Ric85b, Rid88, RO87a, RA89, SS86a, SS89a, Sam83, Sch84b, SR87, She85, SB85b, WB83a, WWB86, WS83, WC88, dB86, dB87a, dB87b]. coral-associated [GA83, GGA85]. coral-reef [BG84, PB85a, dB86]. coralline [WW88a]. corallivorous [GR89]. Corals [CLS^{+}82, Din82b, Hig82, KQ82, Sam82, BL87a, BL87b, DE89, Har85, HF87, HLS83, HGS89, Hug87, LCS86, Ryl83, SSN87, Sie88, SKS^{+}84, Van83, Van88a, Wal85a, WWB86]. Corambe [Yos86a, Yos86b]. cordatum [DJ84]. cordgrass [Het89]. Core [Bra81, JH80a, JH80b, BAH87, GB83, LP87, Mit87]. cores [YD84]. cornutus [Shi80]. Corophiidae [SB82]. Corophium [MGC82, BMRN84, Jen85, MR81, NK82]. Corpuscles [NL81]. Correlates [Thr82, PH86]. correlation [BP85, SP86a]. Correlations [Fai88, LGS^{+}89, VB85a]. corridors [RMO88]. Coscinodiscus [RHP89]. Cost [Rii88a, Rii89, Tho84]. costatum [CSD88, EBH^{+}83, LDER83, Pet89, SF86]. coulteri [KH88a]. Counting [BL80]. counts [Lin88]. coupled [ABJS88, WH88]. coupling [AS86, CS89a, Fuh87, GMF88, GFM89, GM89, KBNO85]. Course [GTG81, BA87, HR89, PCH85]. Courses [Li82]. covariation [KZ89]. cove [Tsu87]. Cover [She80, BG84, Dec86, GKM^{+}87, Ken87b, SPH87]. Coverage [dB81b]. Cr [BBL84, SFN82]. Cr-Labelled [SFN82]. Crab [HS81, SJSO82, Sv80b, VS82, ALPS83, Ang87, AHMB89, BV85, Chr89, CE89, CF84, ELR88, Epi88, EMG89, FWB89, GLNR86, HL89, KO87, KSM88, LB86, LH85, Law89, LOvM89, MGT89, NGR86, RN89, SS89b, SBR84, SVK83, WH87, WP87, aH85, ap 86]. Crabs [LN81, McD82, Pau81, BL88, BM86b, Cho86, CH84, DL86b, FW86, HLH87, Ken89, KC89, KA84, NS87, Ov87, Slo85, Sul84, Wil89, WAH89b, aH85, ap 86]. Crangon [HS86, KD84, PR84, RL83, Reg84]. Crassosfrea [NL86]. Crassostrea [Bro88, LN89, Man89, NJ83, OWZ89, PWM85, SK82, WU82, WH85, ZJ83]. crassus [Din81]. crawling [FL86a]. creek [CSS83, Ov87, WB83b, WS88, WWD^{+}88]. creekbanks [RMO88]. creeks [RO87b, RMO88]. Crepidula [NCG^{+}83]. crevice [BB86]. Crinoidea [ZH81]. crispatus [SED81]. cristatus [DW83]. Criteria [Sv80a, LCP84]. critical [VMM^{+}88]. critique [GW85]. croaker [GOAYH86]. Croix [WYG85]. Crop [KAE81, Law86]. cross [CFP88]. cross-system [CFP88]. Crossostrea [RRB^{+}80]. crown [Ols85]. crown-of-thorns [Ols85]. Cruces [CB88]. Crude [Nel81b, DLT^{+}83]. Crustacea [SMB85, Spa80, AZ85, BC88, DAC^{+}83, DG88a, FW86, GLNR86, GB83, HK84, HC86b, Jay89, McD82, NR89c, O’B88a, O’B88b, SM87a, SW82b, TS88a, TAK88, WH87, WPW87, WPW88, WR84a, WR84b, WR86]. Crustacean [MM81, CS89b, Law89, LS84a, SB85a, WR87]. Crustaceans [BB80, OOS82, Lew84, Mau88, NR89b, vB87]. Cryopreservation [BAG80b, BAG80a]. cryptofauna [KMM88]. Cs [CSH89]. Ctenodiscus [SED81]. Ctenophora [Pur88]. Cu [JCL86, RSWJ80]. Cucumber [Sv80a]. cultivation [RR87, RBA83, WAN84]. cultural [SH89]. Culture [BAG80b, MKS80, OF82, RK81, VS80, EBH^{+}83, Gif85, GKA88, HHKA88, LTS86, Mac86a, MHL85, PW87, RWK83, Thu84a]. cultured [HGD86]. Cultures [AP80, ALP84, AV87a, AFHA84, Bjø86, DA84, HAHA84, HH84a, LT87b, PW87, PST87, RO84]. culturing [Ols85]. cumaceans [Kaa86]. cunner [LB86]. cunners [BL88]. Curacao [BCd84, dB81b, dB87a, dB87b]. curlew [BFP87]. curlews [ZE89]. Current [JH80a, JH80b, Bd89, Stu86, AG81, Abr83, Ohm88, PAH89]. currents [EMG89, FCFP84, JD85, LVP^{+}89]. curve [CP84, HMF85]. cusps [MH84]. Cutleria [DBFM80]. cyanobacteria [HAA^{+}88, IM86, IM88, SS87, BSS89, CC86a]. cyanobacterium [Scr83, TBRBA85]. Cyanophyceae [TBRBA85]. Cycle [ABB80, JT85, RK81, Sie82, VTHVd81, ALP84, Ang87, BSF84, Dec88, EM89, GMT89, HK84, KK89, Reg84, RSM88, Scr83, TR87, WZ83b]. Cycles [SB85b, BEM87, CC88b, CC88c, FMM87, FHKR85, JS85, LDKR88]. Cycling [GCBM82, AGCD86, GCD87, GM89, KBNO85, KHHB86, LH87a, LBH89, RDF^{+}88, SH89]. Cycllc [AA81]. Cyclonic [BAH87, LP87, Mit87]. cyclostoma [SBR84]. Cylones [BP84b]. cymosum [TW81]. Cyphoma [Ger86, HS87, LC88]. Cysts [SJSO82, Sor80, TCA82, VCS81, LTS86]. Cytochemical [Moo88c]. Cytochrome [SC82, Liv88, SWG88]. cytometric [BHJ^{+}89]. cytosolic [JCL86].

D [ARGL87, NNA84, AH81, KOO80]. D- [ARGL87]. D-glucose [NNA84, AH81]. Daily [Mac85b, Wil82, DA85b, HPV^{+}87, KS87a, KN83, SK85]. damage [ME89]. damicornis [SB85b]. Damselfish [KMD87, SW82a, Rus87, Sam83, WS83, Zel88]. Dana [Wat86]. Dark [DM82, JB86]. Darwin [AG82]. dasmotrophy [EM89]. data [AH84b, BAC^{+}88c, BAC^{+}88a, BLR86, SG85, Wei89a, PGJ83]. Davies [HSD87]. davisae [Uch88]. day [Eva87, MR84, tBB83]. daytime [BBSS89, Nic84]. DC2 [BSP89]. DCMU [FITF80]. DCMU-Induced [FITF80]. deactivation [LTS86]. Dead [Col80]. debilis [WR87]. Debris [KNL82a, KNL82b, NLL81, SNL82]. Decapod [Nel81a, GB83, Law89, NR89b, WR87]. Decapoda [GLNR86, Jay89, Ang84, AHMB89, Cho86, DAC^{+}83, DCB82, JA85, NR89c, Pau81, RL83, WR84a, WR84b, WR86, Zel85]. Decapods [Col80, GA83, GGA85]. Decay [HP89, PLM89, WBVS86]. decipiens [KCFS89]. decline [Bar84, Fis89, PH87]. declines [Gar89]. Decomposed [HT81]. decomposing [ZM84]. Decomposition [HKd88, KNL82b, WBVS86, FVDBD87, FST85a, FST85b, KCFS89, NL84, SF86]. decrease [Cam85]. Decreased [KKKW86]. Deep [Hic81, Ise81, PRS^{+}82, Sv80b, Wei82, dB87a, dB87b, BGSS88, BCS84, Bri85, CS89b, Dem85, GH86, Gri88, HH84c, Kaa86, LDMT86, MG85a, RLB89, Sch87a, SBR84, SH85, TMBG85, VAF83, Wil88]. deep-living [Kaa86]. Deep-Sea [Hic81, BCS84, Bri85, Dem85, MG85a, RLB89, Sch87a, TMBG85, VAF83]. deep-seamount [LDMT86]. deep-water [BGSS88, CS89b, Wil88]. Defamation [SS80a]. defecation [Dob83, MW84]. defence [DE88]. Defences [CLS^{+}82]. Defense [Sto80a, FG83, Ger84, HDFG88, PF86, SG88a, Van89]. defenses [SP89, WP88]. defensive [PAF86]. deficiency [SGB89]. Degradation [Bau85, LNB81, NLL81, BH85, BP88, Bid88, PD88, PST87]. Delaware [CE89, CS87, ELR88, Epi88, EMG89, KH88b, PS86, RD89]. delay [SS86b]. delicatula [HDFG88]. Demersal [Mac81, Bue87, JAN85, LJIO84, MN87, SH85]. demersus [Wil85]. demissa [BH88, KLN88, LN89]. democratica [HMD88]. Demographic [Dan83, Geo85, ZW89a]. Demography [Nil87, LS88, RSM88]. Dendraster [Eml86]. Dendrophylliidae [Sch84b]. Denitrification [JAS88, JS88, Sei87, SC84b]. Denmark [JAS88, JS88]. dense [BB88, Pos86]. Densities [BCd84, PB81, BLP87, FCFP84]. Density [BvdEH^{+}82, Kne81, ME88, Sch80a, CH88, CJS87, FH89, HKS86, HVK85, Her88, HGS89, Hug87, Kee87, Lev85, Mac87, Mac88, MB89, MD87, MK85, MWM86, SS89b, TL87a, ZWO89, van86b]. density-dependence [TL87a]. density-dependent [FH89, van86b]. Denslty [KC81]. dentatus [JH80c]. Dependence [KR82, TW80, GPR86, NPMW87, TL87a]. dependent [FH89, Kar86, Kee87, SD85, SP86a, ZE89, van86b]. depleted [TB85]. depletion [BBG^{+}87]. deploying [ZE89]. Deposit [Bri85, Ham82a, Ham82b, Hic81, LlJC82, LC82, NK82, RTW81, SED81, Dob83, DG88b, GAA89, Kem87, Lev85, LS88, LlJC83, Ola89, Pet84, PT84, Pos86, Rei83, TE86b, Wat88]. Deposit-Feeder [LlJC82, Kem87]. Deposit-Feeding [Ham82a, Ham82b, LC82, NK82, RTW81, SED81, GAA89, LlJC83, Ola89, Pet84, Rei83, Wat88]. deposited [GL89a]. Deposition [TCA82, KN84, LML83, Rud88, SBR84]. depressus [McD82, SVK83]. deprivation [GP88a, JM85b]. Depth [CSB81, Smi81, BS89a, BBSS89, DA85b, EZ89, KKKW86, MR84, RA87, SVK83, Sul84]. depuration [CSH83, CSH89]. Derived [Ten82, BP88, Bid88, KLN88, LT87b, PD88]. Dermal [AS83]. Dermasterias [AP83]. describe [MC88]. Description [CM82, Mit87]. desert [AZ85]. desert-bordered [AZ85]. design [CG88, DR86, HBJ^{+}88, KC89]. Desmotrichum [RHvd81]. desolationis [BTQS88]. destratification [TGA89]. destroyed [HSB88]. destruction [SSN87]. Detached [LRH82]. detect [TKT86]. detected [PF83]. Detection [BA83, CA86, SLP83, CS86, Gra83]. determinant [BH86]. Determination [AV87a, FDM86, HPC89, JBS82, MS89, DB88, KS83, SF86]. determinded [RGLK87]. determined [DL85, Red83, RNJ^{+}84]. determining [HH84b, KP89, Sha88b, WDS^{+}84]. Detoxification [PCR80]. Detrital [FT82, Lev85, LS88, Pet89]. Detritus [HT81, RTW81, Ric82, RMZ82, Sch81a, SFN82, Ten82, AH84a, And86, BTL86, Bau85, BP88, Bid88, BMZZ88, FVDBD87, HP89, KA85, LN89, SN85]. Detritus-Like [RMZ82]. Development [Ang84, BK81, EB86, Joh81a, Joh82, Joh88, LH81, MM81, Pet86b, PB85b, SW82b, vdVB86, ASMV88, AHMB89, Bod88, CS89a, Chr89, Daa87, GMS85, Gri83, KHE^{+}83, KBv^{+}83b, KA84, PLM89, PSMS87, SW84, SHB^{+}83, Tan88, TR87, LNB81]. developmental [Tan88]. device [ABJS88]. Diadema [BCd84, HY88, SW82a, WC88, dB86, dB87b]. dianthus [SWSL81]. diapause [LTS86, Mar84]. Diatom [de 80, DA84, Dec86, HK86, HGD86, KS87a, Nil87, PDWF85, RS87, Rie89, SP84a, TB86, WPW88]. Diatoms [AP80, CD81, ALP84, ARGL87, Col84, GPR86, KM88b, Mon84, RO84]. dichotoma [NPMW87]. Dictyocha [TL85a]. Dictyopteris [HDFG88]. Dictyota [NPMW87]. Didemnum [Cow81]. didyma [RNK87]. Diel [BK85b, CC86a, CC88a, DL86a, FEHA85, GL82, HO80, HM86, HHA85, Kaa86, KWG85, KMBSJ87, LDKR88, RNJ^{+}84, STW89, TB86, WB86, AS85, Båm88, CC88b, CC88c, DCJS85, FB87, HMT83, HM87, JG89, Ued87]. diesel [AE88, TAK88]. diesel-oil [TAK88]. Diet [HB81, MN87, Pau81, Pea81a, WPW88, AAH87, Cho86, FMM87, Mac88, SDRN87, SWB89, Tis89, WH87]. Dietary [AB89]. Diets [PA84, EC87a, KFPH88, LJIO84, Nil87, VCS84]. differ [Dec88]. Differences [Bro80, DJ84, LS84b, Pag86, AC85, HS85, Hov89, KBv89, Ken87b, McC84, SWB89, TE86a, TB89, Ver89, WDS^{+}84]. Different [LN81, Ovi81, VS80, VCS81, Yon80, BL88, GKOW87, LS84a, LS84b, Mar83b, Øre87, RGH89, Rüg89, SDY87, WF83]. differentation [DAC^{+}83]. Differential [And89a, CSD88, CYC83, DJTL88, Rei87a, VB81, CT89, HDFG88, LNV81]. differentiation [AB89, Fev89, HTC84]. differing [Ber84]. diffusion [KP89, McG89b]. diflubenzuron [WP87]. Digenea [Shi80]. Digest [WM82]. Digestion [NL86, Tse80, BHNI89, BH87, SU87]. Digestive [SDRN87, CL88, Liv88, Moo88c, RLB89]. digging [MAWM83, Thr86]. digitalis [HD89]. digitata [Rus83a, TL84]. dilution [CC86b, Fuh87, HH85, KB84, LD85]. Dimension [BR82, Mar83a]. dimensional [GP85]. dimensions [KBv^{+}83b]. Dimilin(R) [TC81]. dino [CC88c]. dino-flagellate [CC88c]. Dinoflagellate [TCA82, BP86, Coh85, FF89, GFG89, Han89, HMT83, HSRM86]. Dinoflagellates [Kle80, AS85]. Dinophyceae [HMH89]. Dinophyta [RWK83]. diplocostata [GL88a]. Diplolaimella [FT82]. diploproa [BK80]. Direct [O’B88a, TL85b, BBG^{+}87, KP89, MKG88]. Discharge [Joh80, RC86]. discipes [HHG84]. discovered [WB83a]. Discovery [SC86, Ham82a, Ham82b]. discussion [Mar83a]. disease [SMN86]. dislodgement [WS84]. Dispersal [BS80b, ELR88, Epi88, Keo89, Pal88, Eml86, Hic88, Hig85, PAH89, TB85, dB87a]. dispersants [CL84a]. Disperse [Kem81]. dispersion [CF84, Mor85a, Wah85]. disruption [Oka86]. dissimilatory [ES87]. dissociation [DL85]. Dissodactylus [BS83a]. Dissolved [AA81, BAN89, CHAS83, EHF^{+}81, HK86, IBMD81, Jør82, LNV81, NK89, NLVS80, RMZ82, Tam89, TW80, ABT89, BW88, CSS83, CC89, DPJ87, EBH^{+}83, EB86, Fis88, FB86, FB85, FF86b, Fuh87, GP88b, HA87, LVdVdV85, MR86, OWKG85, PW87, Par87b, PMJH84, RWK83, RHP89, TIS85, VA85b]. distance [LB87, VC86]. distichus [Van89]. distorted [SKIR89]. Distribution [ACF^{+}80, CS80, CE89, CD81, DPJ87, DG82b, FL88, FCW82, Fuh81, FW86, GLD82, Gra81, GB83, HC80, Ham80, Hic86, HB81, Joc89, KC81, Kin89, LVdVdV85, Lew84, LM85, Man88, MD86, MAWM83, Mor85a, PD86, PMJH84, Rus84a, Rus84b, RS80, SW82a, ST82, Sv80a, Smi81, TP81, Thr82, VN82, WU80, Wei82, WR81, ZW89c, AWSB83, Bai89, BSS89, BK85b, BBSS89, CPM89, De 84, DA85b, Don87, EM89, EM87, FBCH83, FT86, FAG89, GG88, Gas87, HPV^{+}87, HMH89, HKS86, HHB88, Hod85, HHN^{+}83, HSB88, Jac87, JG89, JL86, JAN85, JH87, Jen88, Jon89, KAA88, KAE87, KD86, KCS89, KMM88, LC88, LGK87, LVP^{+}89, Mac88, MERW84, MS89, NBR^{+}88, NBMM88, OD89, OQ83, Øre86, PPCR86, RTLR88, RGH89, Ric85a, Rid88, RD89, RA87, SS86a, SG85]. distribution [Sch87b, TW88, WR84b, Wil89, WC83, YLH86, Zel85, dB87a, MS88]. Distributional [CSW84, VCS84]. Distributions [Bra81, CMC89, DCJS85, GPCW88, HHE89, Hob88, KEK^{+}84, KM87a, KM87b, KM89b, LC83, MB84a, PN88, RM88, TRMF89, Ued87, WCGG86, WYG85]. Disturbance [Gra81, Kar86, SS80b, Wil88, ZW82a, ZW82b, DW87, GC85, Gly85, KM89a, LP88, PE89, SLP83, SRTM83, WLB84, ZW89a]. disturbance-induced [ZW89a]. disturbance-mediated [KM89a]. Disturbances [Thi81, Ken87a, Ken87b, WBAN85]. disturbed [BB86, KSS85]. Ditylum [HH84a]. Diurnal [RK82]. diurnally [HH84a]. Divergence [BS80d]. Divergent [McD82]. divers [HS88b]. diversicolor [BMRN84, EZ89, ZE89]. Diversity [Col80, She80, DKMR83, DW87, GL88b, KSS85, LDH83, LN86, PL85, RO87a]. divided [DS84]. dividing [CC86a, DS84, RNJ^{+}84]. dividing-divided [DS84]. Division [Cow81, CC86a, CC88a, HH84a]. DNA [BTH^{+}84, CC88b, CC88c, DPJ87, DRCA83, Hov89, MP81, RNJ^{+}84, WH85]. Do [Kem81, MM86, AAH87, Dec88]. DOC [Vel86]. Does [Clo82, DRB86]. dogwhelks [CH84]. dollar [BS83a, Eml86]. Dollard [Cd83]. domains [Wal85b]. Dominance [FSCP89, PDIF85, OS84a]. Dominant [MM81, ML80, SB82, IEK86, KS84, PH83]. dominated [BF87, BW88, DKMR83, KM89a, LM87a]. domination [D’A85a]. don [CM87]. Donacidae [Leb82]. Donax [Don87]. D’Orb. [RZ85]. dorid [FG83, Yos86a, Yos86b]. Doridella [Yos86a, Yos86b]. dosed [AE88, OA88]. down [ZH81]. downwelling [WF83]. Dr [Mon89]. drag [WS84]. dragonet [Sog84]. Drake [HHE89]. Drift [TB85, FL85, KMR^{+}88, ZM84]. Drift-tube [TB85]. drifting [Bd89, SJNL89]. dry [BR83, SV85a]. due [Hol88, Ken87b, Ken88]. Dumontia [RK81, Rie82b]. Dunaliella [FF89, HBJ^{+}88, RBA86, RBA88, SGB89]. Dunalielta [BP84a]. dunlin [BFP87]. Duration [Lam87, Tse80]. During [DB82, Gun80, IBMD81, Joh82, LNB81, LNV81, AHMB89, BHJ^{+}89, BS89b, Cam85, Col84, Dun86, FF89, GKA88, GKOW87, GD84a, GMB87, Gra87, GMT89, HAHA84, HK86, Har85, HH87, Hoe88, JBAH83, KHE^{+}83, KBv^{+}83a, KBv^{+}83b, LVdVdV85, MN87, MP88a, MH89, NK89, OKSB86, PSMS87, Par87b, PD88, RTLR88, RYD84, Roy89, SGB89, SW86b, TS88a, TGA89, Tra88, TRL89, WBVS86, ZW89b]. Durvillaea [CB88, CH88, Law86]. Durvilleales [CB88]. Dutch [CD81]. Dwelling [Din82a, Din82b, HP86b, Slo85]. Dynamic [TCA82, KMH85, KMR87, Sha88a]. Dynamics [Bar84, BF87, HHD^{+}84, HP86a, JBS82, KN88, LM87a, LHF84, NGV^{+}82, PB88, Sch80a, Tal80, AM88, AGCD86, AS86, BS83a, CS89a, CF84, Col86, DM83, Dav89, Fle87, FVDBD87, Har86, HPF89, HLH87, HK84, Hug84, IM86, JS87a, Jon86, Jos87, KL84, KMR^{+}88, LT87a, MCPW88, MD87, PH87, PWMl89, RHS89, RDD88, SO83b, Smi86, SS87, SKM85, SM87b, TB86, Tsu87, Vel86, VP88, Wei89b, WBVS86, Zel85, ZSA87, dB87b].

Early [IR81, LH81, Van88a, CLB88, Luc84, RTLR88, TB89, van86b]. earthen [BLK88]. East [Bra81, WR81, KAK88, JH80a, JH80b, McC88, NVLL88]. eastern [Bai89, BHP^{+}86, BAH87, CH88, Don87, GR89, KSS85, LHT88, LP87, LVP^{+}89, Mee85, Mit87, Par87a, SSM^{+}84, Sor85, SRL^{+}86]. eating [Ger86]. ebumeus [SW82b]. Echinocardium [DJ84]. Echinoderm [TM80]. Echinodermata [BS83a, Ham82a, Ham82b, HY88, Sch80a, ZH81]. echinoid [LS84b]. Echinoidea [HY88]. Echinoids [Ham82a, Ham82b]. Echinometra [LS84b]. Ecklonia [KNL82a, LNB81, LNV81, NLVS80]. Ecological [AG81, AFF^{+}82, BCS84, BR82, DG82a, GW81, HS85, Joh80, MJ84, OS84a, RG82, RR82, Sv80b, SCSD86, Van88b, And89b, Bod88, CG87, FSCP89, HTC84, MT85a, MT85b, Mar83a, MS88, Sch83c, UP88]. ecologists [MGM86]. Ecology [BSv81, Bla84, Fen82a, Fen82b, Fen82c, Fen82d, FR80, GG82, Her88, KQ82, LVP^{+}89, MM81, Mur84, RBA83, SCF^{+}82, SED81, Sto80a, SvDB81, Tar81, YALD88, Zel85, vDBS81, Ans89, BRG83, BGR86, BMGB88, CSW84, CS89b, Coh85, DT85, FG83, FF86a, GB84, GNBO85, GLNR86, Gra86b, Gra86c, Hop83, Jen87, Klu84, LHT88, McC89b, MP86, Ste89, SK85, Tur84, WB83b, WWL83, vM87, Gra86b, Gra86c, Kin86]. Ecophysiology [TAK88, Bar86, Bar88b]. Ecosystem [BY81, ITBW81, KdC81, LH82, OWP80, AN88, Bra87, CL84a, DLT^{+}83, DOB^{+}89, FB87, HAA^{+}88, KN88, KE87, LDER83, NN85, PST87, SP86a, SF85, SH89, WZ83a, WWD^{+}88, YALD88, CSS82, CSS83]. ecosystem-scale [SH89]. Ecosystems [AG82, JBS82, CFP88, DKMR83, DK85, GMS85, JBAH83, KHE^{+}83, LDH83, TP85, WE88, WBVS86]. Ectocarpales [RW86]. ectocarpoid [Rus83a]. ectoparasites [NPSV87]. edaphic [Mon84]. eddies [Han86]. Eddy [Bra81, JH80a, JH80b, BAH87, GB83, LP87, Mit87]. edge [HC86a]. edible [MG85b, MGT89]. edolis [Ham80]. edule [Bro80, Jen85, LG88, MG85b]. edulis [AAH87, Auf88, BW80, BT88, BHNI89, BTFF85, BTQS88, BNF81, CL88, CL84a, CSH83, CSH89, Dah86, Fis86, Fis88, GT87, GW81, HC80, HSB^{+}84, HB85, HL80, Joh88, KE87, Liv88, Low88, LNS^{+}87, MKG88, NH82, Oka86, RYD84, Spr84c, Spr84d, Spr84a, Spr84b, Tho84, Tra88, TC87, TN86, WDS^{+}84, WJ88, Wil83, ZJ83, ap 86]. edwardsii [Mac85a]. Eelgrass [Cap82, HFK82, VN82, BTL86, BL88, DA85b, KMBSJ87, NG86, SMN86, ZSA87]. Effect [ABT89, BCA89, BV83, Bje88, BMRN84, BR85, CL86, CS80, Cam82, DDR89, FT82, FL86b, GD84a, HL85, JA85, JF88, KAE87, KGV84, Kim85, NGR86, RSWJ80, RBA86, RTW81, RM88, Ryg85, SHCL87, TC81, VGJ89, WM82, WZ83b, Abo84, BNF^{+}89, CTG86, CG88, KN84, KSVB89, Mor87, PKMS85, PPCR86, RBA88, SM87b, TP85, TN85, TN86, VA85a, VCS81]. Effects [AA89, AT86, ASMV88, Amb86a, AH84a, BHNI89, BV85, BY81, Bra81, Bre82, BBBM80, Buc82, BTH^{+}84, BNF81, DLT^{+}83, Dah86, DM83, DD80, EAM87, FCFP84, FKW81, GP88a, Hei89, HR87, HF88, ITBW81, Jac80, JD85, Ken88, Ken89, KC89, KD86, KM89a, Kne81, KHW88, Kui81, KBv^{+}83a, LDMT86, LS88, Lin88, LML83, LR80, MKS80, McG89a, NR89b, OWZ89, Ovi81, PB85a, PB89, PDP87, PT84, Qui88, RL83, Rie82b, RRB^{+}80, RA89, Sam83, SO83a, SW84, SR87, SC84b, SE85, TL84, Van86a, WY82, War88b, War88c, WU82, WS83, ZWO89, ZM84, Add88b, Add88a, BCG88, Cap88a, Cap88b, CB88, Che89, CD88, CP84, DL83, DE88, DJTL88, DV83, Dun86, ES87, FDM86, Gra83, Han89, HH85, HDFG88, Jos87, Ken87a, Keo89, KB84, LL83, MD87]. effects [Mil89, Moo88b, Moo88a, NCG^{+}83, OQ83, OQM^{+}87, Pet84, PE89, PGV^{+}89, PHL^{+}86, Rei87a, RHP89, Rye88, SSN87, SR86, Sma83, Sto86, TR87, TKS^{+}85, Ver89, WCC^{+}88, WR84a, WR87, Wil86a, Zel88, ZIO84, ZSA87, Joh81a, Joh81b, Joh82, JK80b, Ken87b, Kri83b, Kri83c]. efficiencies [HB85, Spr84b]. Efficiency [NLL81, Rii88b, KMR87, Klu84, LDG87, LlJC83, Pea85, Rei87a, Wil83]. Efficient [RG82]. efflux [WYG85]. effort [Cla87, LRS87, Tho84]. egested [FL86a]. Egg [Lee81, Mau88, TC81, Daa87, KMH85, MC87, Pet88, STW89, SM86, TR87]. Eggs [LHH81, FL88, HKS86, Mar84, McG86, STP84]. Egyptian [WEHF85]. eid [OA88]. Eilat [BL87b]. Ekofisk [DLT^{+}83, Nel81b]. Elasipoda [TMBG85]. electrode [dHS88]. Electron [Fuh81, PGBN81, SM84, SC82, AH87, JP84, MB87, Sch88a]. elegans [Dal85, MT84, NR89b, NR89c, Øre87, Pea81a, TS88a, WR82, WR84a, WR84b, WR86]. element [Bro86]. Elements [BP81, GHCH81, FS84, PR88]. Elevated [PDJ86]. elevation [HG86]. elimination [STP84]. ELISPOT [TKT86]. elliptica [TS88b]. elongatum [LHT88]. Elutriator [Hoc81]. embayment [AS85, JKJ88]. Embiotoca [SH84]. Embiotocidae [SH84]. embryo [LGP88, SHB^{+}83]. embryo-larval [LGP88]. Embryology [LH81]. Embryos [JH80c, Joh88, WRC83]. emergence [Arm88a, Arm88b, JG89, Pal88, SLMB88]. Emergent [JG89, TL87b, TL87a]. Emersion [DB82, DG82b, TS88a]. Emigration [SV85a, KB88, TP89]. Emphasis [KC82, FMM87, Ov87]. empirical [CC87a]. employing [FL86b]. Ems [Cd83]. Encapsulated [Bre82]. enclosed [BH83b, DLT^{+}83, GMS85, JBAH83, KHE^{+}83, KBv^{+}83a, KBv^{+}83b, LDH83]. enclosure [EBH^{+}83, EB86]. Enclosures [BDKK83, ITBW81, Kui81, FL86b, FG88, HNR^{+}88, RNH^{+}88]. encounter [VDM89]. encrusting [Lan89, Rub85]. Encystment [TCA82]. endemic [WtD89]. endocrine [NS87]. Endogenous [Has81, LCF89]. endosymbiont [CVF89]. endosymbiont-bearing [CVF89]. endosymbiotic [SGSD86]. Endproducts [PCR80]. Energetic [SC84a]. Energetics [Ang87, Bak88, GL88a, MT85a, MT85b, Spr84c, Spr84d, Spr84a, Spr84b, WJ88, BH81b]. Energy [BB88, BY81, DP81, HS81, Joh82, KdC81, NFG82, Pea81a, SF80, TR89, WA87, Bar88b, BRS86, BS83b, GFS89, Gra87, JM85b, KS83, Klu84, MCM84, MB85, MPR85, NM89, Rii88a, Rii89, Spa80, WAN84]. energy-source [JM85b]. enflata [Kim84]. English [ID88, GB84, HHN^{+}83, Øre86]. Engraulis [AI88, JF88, SDRN87]. enhanced [CA87]. enhancement [FK86]. enriched [BV83, BB87, DE87, KA85]. Enrichment [Ten82, AH84a, And86, CHH83, OOF83, Rei83, Tam89, TKS^{+}85, WE88]. Enteropneusta [DG88b]. entire [Rus84a]. entomon [HS88a]. entrapment [WRC83]. enumeration [BHJ^{+}89, Dou83, DNF87, GGV89]. Environment [Aik81, KC81, OFGG82, Sch80b, TCA82, AC88, BB86, FD89, HK86, Har86, Hob88, Jen88, LML83, MERW84, OQM^{+}87, Pro87, SdT86, Sto89, Tho84, TE86b, WH88, Wro84]. Environmental [ACF^{+}80, FG87, KB80, LCF89, Le 86, MM81, RK81, Sam82, Thr82, Bar88b, BCA89, Bro88, CP84, Liv88, Low88, Moo88c, MT84, RO84, STW89, SWG88, TS88a, VB85a, WDS^{+}84]. Environmentally [LL81]. Environments [Fen82c, NK81, ABAI87, Col86, DPJ87, Fle87, LH87a, PS88, RNJ^{+}84, VBFS85]. enzymes [MLDT88, SDRN87, SDMN88]. ephemeral [FVDBD87]. epi [OS88]. epi-bacteria [OS88]. Epibenthic [GL82, SPH87, WY82, You81, BV83, BB85, MD86, Pih85, SGG87, WW86]. epibenthos [BB86]. Epibiontic [CP87]. epibiosis [Wah89a]. epifauna [How85, LJIO84]. Epifaunal [Sto80b, JS87a, SW84]. epiflora [Rus83a]. Epifluorescence [Fuh81, CS86]. epilithic [LCF89, Mac86c, Mac87, SR87]. epipelagic [VS84]. epiphytal [Rus87]. Epiphyte [RV84, KH88a]. epiphytes [Rus83b]. Epiphytic [ACF^{+}80, KMAM84, WW88a]. Epipsammic [NK82]. Epithelium [HL80]. epizoans [WS84]. equation [DJ84]. equatorial [DE87, Le 86]. equina [SCT87]. equine [HTC84, OTC82]. equiproportional [Tan88]. error [SO85]. Escape [Sam82, BH87]. Esk [HC80]. Especially [TW81]. established [Fis89]. Establishing [Li82]. Establishment [SS80a]. Estimate [BS80c, FITF80, GM87, KM87a, LMO86, RG89]. estimated [Cam85, DJ84, JAN85, KH88b]. estimates [Dou83, DNF87, ER87, GP85, LT87b, NL84, Pat88, PGV^{+}89, SWH84, SK85, TL87b, VP88]. Estimating [CC86b, Che89, DCB82, Gei88, Kel88, Li82, CC87a, Hol88, JMG84, PBP85, SO85]. Estimation [SA82, AM88, Bau85, DS84, DTLN85, JP88, RHP89]. estuaries [CC87a, STW89, WC83]. Estuarine [DS84, DL82, FP82, Lon81a, NGV^{+}82, OFGG82, WU82, ZW82a, ZW82b, de 80, ARGL87, AS85, BBH^{+}88, BK85a, BB85, BMRN84, BR85, DL83, DPJ87, DA84, EC87b, FK86, HP86b, Hic88, HNR^{+}88, HMLF88, JS85, JS88, LNS^{+}87, LM85, MB87, OPN^{+}84, ORK^{+}86, PWM85, PCW86, RNH^{+}88, SSA^{+}86, SSPA89, SW85, TKS^{+}85, WG84, WZ85, WAH89b, WW88b, YD84]. Estuary [BA81, FW86, Jør82, SS80b, Alo87, ABAI87, DK86, ELR88, HP89, HLH87, KH88b, LD89, LDKR88, MHW88, MKD^{+}86, MCPW88, NVLL88, PCM86, RHS89, RG85, Sei87, Sha88b, SP86b, SV85a, SV85b, Whi89, WAN84, CS87, Cd83, CW82, HC80, HKd88, PLT87, PT89, PS86]. ETS [Sch88a, TR89]. ETS-activity [TR89]. Eu [Dah86]. Eucalanus [PL89, Tur84]. Eudendrium [Bar88a]. Eudiplogaster [RBA83]. euhaline [DK86, WWD^{+}88, WS88]. Euphausia [EM87, HHE89, MM81, Mor85b, Ohm84, PAH89, SHH^{+}88, Stu86, Wat86]. Euphausiacea [O’B88b, Stu86]. euphausiid [Nic84, TKT86]. euphausiids [SdT86]. euphotic [CS84, FH83, JP87, RS87]. Euplotes [CSD88, Din81, GH83]. Eupolymnia [GAA89]. Europe [Kro88]. European [MGT89, NH82]. Eurydice [Has81]. euryhaline [SC85]. euryoxic [MB87]. Eurypanopeus [McD82, SVK83]. Eurytemora [BK82, Geo85]. Euterpina [CS89a]. eutrophic [AS86, TGA89]. eutrophicated [Sel88]. Eutrophication [OR80, OSM82, DOB^{+}89, KKKW86, KBNO85, OKSB86, SH89]. Evadne [PY86]. evaluate [HBJ^{+}88]. Evaluation [GM87, LC82, PB87, PY81, Sch80b, BBL84, CA86, Geo85, LS84a, Mac86c, PF89, VMM^{+}88]. evasion [KSM88]. event [And89a, Bar84, HAOC^{+}86, Har85]. events [FS84]. Evidence [Bro80, BS80d, Grö89, MKG88, PAF86, RS80, SFN82, SK84, TB89, TD89, WB80, WH88, BBG^{+}87, CD88, HSB^{+}84, JS87b, ME88, NM89, SW85, SCT87, WCGG86, WAH89b, KM89b]. Evolutinonary [Cha82]. evolution [JS87b]. Evolutionary [Sv80b, FG83]. Ewens [GL88b, LP88]. Ewens/Caswell [GL88b, LP88]. example [Sch83c, SO87]. examples [BMGB88]. exaptation [DE88]. excentricus [Eml86]. Excess [WU82]. excessive [Dun86]. exchange [And86, RD89, SP84a, SHIP86, WWD^{+}88]. excitation [NVHHH88]. Exclosure [Blo80]. exclusion [CB88, Zel88]. excreted [WC88]. Excretion [ABB80, AA89, ALP84, HM86, Reg84, VA85a, VA85b]. Existence [IEK86, Kin88, Yos86a]. exocrines [DL85]. exoenzymatic [Hop83]. exogenous [NPVW86]. Exogone [PH83]. Exological [Shi80]. exopolysaccharide [GMB87]. experiment [CJS87, EBH^{+}83, GAC^{+}88, LDER83, McG88, OPN^{+}84, OKSB86, OQM^{+}87, RYD84, Sch88c, Sch88b, SPSS89, Tam89, TKP86, WCC^{+}88]. Experimental [Bro82, GF82, Jer86, KN84, KC81, Kui81, Mac86c, MUC84, Nel81a, RBR86, RS80, Sie88, SSK80, Auf88, Bra87, DO86, DK85, FC83, FL86b, Grö88, Grö89, HA87, HBJ^{+}88, Hir86, Hir87, Jon89, KBNO85, Liv88, Low88, Moo88c, Rei83, SM87b, SJNL89, TNR86, WE88, WAN84, Zel88]. experimentally [AE88, VAF83]. Experiments [EHF^{+}81, EB86, HF87, KB84, LHF84, LSD85, LD85, MT87, Mon83, NC89, Pet84, dKB88b]. explain [Ebe86]. Explaining [McG88, Pal88]. Exploitation [SN88]. Exploratory [OWP80]. Export [BTL86]. exported [ABT89]. Exposed [KOO80, LN81, SMO81, CSH83, MHL85, MKHD85, SO84]. Exposure [BNF81, BR85, ELK84, KM89a, Liv88, Low88, MB84b, Moo88c, SWG88, TAK88]. exposures [BFMS88]. extent [DE87]. extracellular [FF89, HL85, Lan83, Mø83, SP86b]. Extract [Sch80b, Sch84a]. Extracting [Hoc81, ES84]. extraction [AH86, Mar83b, NW84]. exuded [Thu84b]. Eye [Som80].

f [OA88]. Fabricius [Cam85]. facilitated [DE88]. Fact [SA81]. Factor [Shu80, Jos87, VDM89]. factorial [HBJ^{+}88]. Factors [HH84b, HB86, Lan83, Lee81, Lon81b, PWM85, SM86, Arm88b, BCA89, BB87, Bro88, HKS86, HBJ^{+}88, LT84, Möl86, Sha88b, STW89, WDS^{+}84, ap 86]. Fade [Kem81]. Faecal [SNL82, Mor87, NvB89]. failure [McG88, Tra88]. fall [CC89, FVDBD87]. far [NVHHH88]. far-red [NVHHH88]. fast [Car87]. Fate [Dun86, Kui81, OQM^{+}87, TP85, GFS89, Gar84, Kem87, MKD^{+}86, SKS^{+}84]. fates [SW85]. fatty [MCA89, SGB89]. Fauna [TW81, WY82, You81, CM87, DSG^{+}89, JAN85, Kro88, McC89a, MAWM83, MJ84, Pih85, RT85, Ryg85, SH85, SG87b, TN86, War84]. Faunal [CM82, KC82, MH84, Lam86, MCM84, MB84b, MPR85]. Faviid [KQ82]. FDC [RNJ^{+}84]. Features [KSR80, Wat88]. February [LSB85]. Fecal [Ise81, BNVP87, DT85, ER87, GABM84, Tur84]. Feces [Bro86, Rob82, Nel89]. Fecundity [DCB82, Kne81, LRS87, TC81, BGSS88, Kiø89]. fed [CSD88]. Feeder [LlJC82, BBL84, DG88b, FAG89, Kem87, Lev85, LS88, Pos86, SN85, TE86b]. feeders [Dob83, FD89, JMSK86, LN89, SK84]. Feeding [BH81b, CC82, CS89b, Che82, CS81, CC87b, FKW81, GB84, GNBO85, GLNR86, GD81, Ham82a, Ham82b, HSB^{+}84, Hic81, HB81, Jen87, KM81, Las81, LC82, MF89a, MP86, MR84, MERW84, MK85, NK82, OSM82, Øre87, PL89, PDTH86, Pau81, Pea80, Pea81a, Pea82, RTW81, Rie82a, RK82, Sam88, Sch80a, SCF^{+}82, SED81, Sv80a, Stu86, Tra88, TW83, TC88, WPW87, WPS89, WP88, AAH87, AS83, BHNI89, Bri85, Bro86, CTG86, Cho86, CH84, DD85, DT85, DS86b, EZ89, FL86a, Gal89, Gas87, GHC83, GOAYH86, GAA89, GR89, HH87, HNR^{+}88, JF88, Jen88, Jon86, Jør83, JKMR84, Kan84, KMH85, Kiø89, Klu84, LH85, LlJC83, Mac85b, MD86, Mor87, Mor85b, NPVW86, Ola89, PV85, Pau83, PA84, Pet84, PT84, Rei83, Rii89, RR87, RHP89]. feeding [SM87a, Sch83b, Ste86, SP89, SK85, Tur84, Van86a, Wat88, WPW88, WR89, Wro84, aH85, ap 86]. female [BPH87, BS89a, Jay89]. Fenchel [Gra86c]. ferreirae [AHMB89]. ferrihemoprotein [Moo88c]. ferruginea [BFP87]. Fertile [WB80]. Fertility [LB81]. Fertilized [LHH81]. fiber [Hic84]. fiddler [Chr89, CF84, DL86b, ELR88, WP87]. Field [BS80b, CL82, Cow81, LSD85, Auf88, Bak88, CPM89, DO86, FSBOG84, FM88, Low88, RYD84, VMM^{+}88, Wei89a, Dec88]. Fiji [SN89]. filefish [CL84b]. filiforme [Wil87]. films [CMC89]. filograneus [TM80]. Filter [OSM82, DD85, DP87a, DP87b, FD89, FAG89, JKMR84, JMSK86, Klu84, PHT85, SN85, SK84]. filter-feeder [FAG89, SN85]. filter-feeding [Klu84]. filtered [LD85, SS86b]. filters [Hic84]. Filtration [BNF81, Gri80b, DO86, Rii88b]. Final [Fis88]. Fine [ABAI87, JL86, ID88, Jen88]. Fine-scale [ABAI87, JL86, Jen88]. finmarchicus [Hir83, Tan88]. fire [Bar89]. First [CM82, MN87, Pau83]. Fish [BH81a, BH81b, Bra81, FC82, GABM84, HB81, Kne81, Lee81, LHH81, LRH82, Mac81, MCGP82, Nel81a, NL81, Par89, Rob82, RO87b, SP89, Tar81, TW82, WH82, Yan82, BG84, BK85b, Che84, CPM89, FL85, FH89, FL86b, Gal87a, Gal87b, Gib88, GFG89, GMK85, HL84, Hoe88, HNR^{+}88, JL86, JAN85, Kv88, KM87c, KM88a, Kin89, KB88, KHW88, LHT88, Lin88, LJIO84, Mag88, MG85a, McC89b, McC89c, McG86, MP88b, MJ84, Pea85, PCW86, Pur84, RTK^{+}89, RO87a, Rye88, SS89a, Sam88, Sam83, SSN87, SH84, SP86a, SF85, SO87, SBR84, SH85, Sog84, SPH87, SK85, TE86a, TÓ89, WS83, Wil86a, YLH86]. fisheries [Abr83, Ste89, Zel85]. fishery [WEHF85]. Fishes [Cha82, Liv82, OE81, OQ83, Som80, Thr82, Tse80, Wil82, Bue87, DD84, GCSP88, GHC83, HDFG88, JT85, Mac85b, Mac88, Rus84a, Rus84b, RA89, SS86a, Sha83, SSM^{+}84, SB85a, TCH87, TB89, TB88, Zel88]. fishing [KHW88, RA89]. fishponds [BLK88]. fission [DE89]. fissiparous [ME88]. fits [Pat88]. Five [Gun80, DL86a, HC86b, VK85]. Fixation [Cap82, JB86, SG87a, SHCL87, SC84b, WS83]. Fjord [JS88, KMDS85, Tal80, AAK^{+}88, Eml86, JBAH83, LGS^{+}89, LH88, Slo85, Was83, Was85, Gas85]. fjord-dwelling [Slo85]. fjords [Kaa89, LPV87, SBR84, TW88]. FLA [RG89]. flagellate [CC88c]. flagellates [ALAH85, KK89, LDER83]. Flat [CD81, KdC81, MD86, Rei87b, SR86, Tam85, Tam87, VP85, WZ83a]. flatfishes [GP88a]. flats [KD84]. flatworms [SST89]. flavescens [GNBO85, WP88]. flavolineatum [SO87]. flesus [AE88, OA88, SM84, SWG88, VL88, vdV85]. Flexibility [Sv80b]. Floating [BA81, Hig85, vM87]. flooded [Het89]. flooding [PDP87]. floral [MB84b]. Florida [LH87b, SS80b, SPH87, SC86, TCH87, ZM84]. Flounder [Buc82, FKW81, JH80c, LH81, AE88, CLB88, Col87, FHKR85, Jen88, OA88, SWG88, VL88, vdV85]. flounders [SM84]. Flow [BHJ^{+}89, KNL82a, KdC81, Bra87, Gra87, MCPW88, MB85, MPR85, NL84, PF83, VP88, WK88, WS84]. Flows [DP81, SPSS89]. Fluctuating [LL81, LML83, SRL^{+}86]. Fluctuations [PLM89, de 80, GKOW87, Wal85a]. Fluence [NPMW87]. Fluid [Jør83, DP87b, JKMR84]. flume [CPM89]. Flumes [MGM86]. Fluorescence [BP84a, BP85, FITF80, AH84b, DTLN85, NVHHH88, Ohm88]. fluorescent [CS86, NBT88, PB87]. fluorescently [RG89]. Flux [PM88, SN81, AH87, Car87, ER87, ES87, JWC83a, NvB89, RD89, SG88b, VB85b, WR84a, WR84b]. Fluxes [HW82, ABT89, BW88, BBR88, HAA^{+}88, KM83, LSB85, LvB87, Man89, PLT87]. foliifera [RR82]. Following [NLVS80, SS80a, DE89, HAOC^{+}86, WW83]. Food [BCF82, CS81, Col87, Eva82, Gri80a, JKJ88, NR89a, PR84, Pih85, Rob82, SK85, TW82, VB81, VP88, WU82, ZH81, Abo84, AGCD86, And89b, BB88, BS83b, BTH^{+}84, CTG86, Che89, DCNS84, DP87a, DG88b, FK84, FT86, GP88a, Gas85, GNBO85, Gle86, GCD87, GM89, HHD^{+}84, HP86a, HMD88, HR87, IM86, JF88, KN83, Klu84, LN89, MT85a, MT85b, Mon83, Mor87, NBR^{+}88, NBMM88, OSB85, PT84, SN88, SF85, SW85, Sog84, SKIR89, SR85, SE85, SK84, TL87a, TC88, VB85b, ZWO89, GMF88, SJSO82, Spr84d, TÓ89]. food-concentrating [DP87a]. food-limited [SF85]. food-particle [PT84]. food-resource [SK84]. foods [LlJC83]. Foraging [AP83, GMK85, LH85, MG85a, SHH^{+}88, TB88, FH89, Gra84, HS88b, Mag88, Pie83, Rye88, SH84]. Foraminifera [KM85, KSR80]. Foraminiferal [GL89a]. forced [Sha83]. Forcing [BY81, LCF89]. Fore [dB81b]. Fore-Reef [dB81b]. Foreign [FS84]. Forest [TD81a, AA84, BW83, PH87, VCS84]. forests [PB89, SO83a, SO83b]. form [Hug87]. Formaldehyde [LC82]. formalin [Pur88]. formalin-preserved [Pur88]. Formation [FITF80, Oka86, Rus83a, DE89, Lün88, tBB83]. forming [HH84b]. formosa [Cro84]. Forms [Fen82a, MM81, OK89, TH84, dKB88b]. Formulation [Wro84]. fornicate [NCG^{+}83]. Forskål [AZ85]. Fouling [Moo81, GSC83, Oka86, SG85, Wah89a]. Four [Cam82, SJSO82, AM88, FBT84, FF86a, FW86, GM87, HF88, OA88, Sto88, VCS84]. Fractal [BR82, Mar83a]. Fractals [BRG83]. fractionated [HSW89, LDG87, Pro85]. fractionation [JCL86]. fractions [LL83]. Fragmentation [Hig82, LNB81, LNV81, NLVS80, Kar86, Las84]. fragmented [CSD88]. fragments [BH87]. framework [UP88]. France [KS87a, Vid87]. franciscanus [TD81a]. Francisco [AC88, CL86, CDL87, Clo82, CC84, NT85, Oka86]. Fraser [BA81]. Free [BA81, FT82, IBMD81, Jen81, Jen82, Jør82, VBFS85, ASSF87, Alo87, CC89, FB86, FB85, FF86b, Fuh87, GP88b, HK86, JBAH83, JNCS89, Jen87, LVdVdV85, LPV87, MJPH^{+}88, PMJH84, PM88, RHP89, SRTM83, Thu84a, TW83, vM87]. Free-Floating [BA81, vM87]. Free-Living [FT82, Jen81, Jen82, Alo87, JNCS89, Jen87, LPV87, PM88, SRTM83, TW83]. Freezing [BAG80a, HL80, DDR89]. French [LDDH88, Gal87a, Gal87b]. frequency [CC86a, DS84, Hob88, RNJ^{+}84]. fresh [CFP88]. freshwater [RO87b]. friction [FF86a]. friendly [EMHS86]. Frierfjord [GAC^{+}88, HWC^{+}88, War88a]. Frierfjord/Langesundfjord [GAC^{+}88, HWC^{+}88, War88a]. fringe [Kro88]. fringed [CL84b]. Fringing [Bou81, Bul82, KM89a, RO87a, SF89]. fringing-reef [KM89a]. Frobisher [Hsi88]. fronds [JJN85]. front [KEK^{+}84, PMJH84, TL85b, Vid87]. Frontal [TCA82, Han86, HHN^{+}83, HlWPH84, KMR^{+}88, Man88, PCH85, Ric85a, VDL88]. Fronts [PSB^{+}81, Sch82b, CE89]. Fucoxanthin [WJ87]. Fucus [HC80, JS87a, KKKW86, KL84, Van89]. Fueguian [MJ84]. Fuel [OFGG82, FER85, LL83, SM87b]. fulgens [TB85]. Function [BY81, KB80, MM81, DOB^{+}89, KD84, LS83]. Functional [Fen82a, Som80, TH84, TS88a, FJ88a, HR87]. funding [Kin86]. Fundulus [Kne81, Moe85]. Fundy [And86, GT87, GD81, Nic84]. fungi [BMZZ88]. Further [SCT87]. Future [Mar89b].

gadids [NPSV87]. Gadus [NPVW86, STP84]. gahi [Pat88]. Galapagos [And89a, WPS89]. galbana [Wil83]. Galeta [McG89a]. gallo [BTQS88]. gallo-provincialis [BTQS88]. galloprovincialis [CR84, GW81]. Galway [TÓ89]. gametangia [tBB83]. gamete [Pet86a]. gametes [GFS89]. gametogenesis [KGV84, SB85b]. gamma [SG85]. gammarid [AT86]. Gammaridea [TW81]. Gammaridean [BS80d, SMB85]. Gammarus [AA89, CL84a, TAK88, MGT89]. gardneri [Ger89]. Garfish [RK82, KN83]. Garrison [GG82]. Gas [MW85, RC86, VK85]. Gastric [Tse80]. Gastropod [Bre82, Kem81, BM86b, FJ88b, Ger86, LC88, McC89a, Mor85a, NCG^{+}83, Pet84, PD86, RNK87]. Gastropoda [BS80a, HS87, LSD85, LlJC83, Shi80, Tal80, TE86b]. gastropods [And89b, FG83, LN86, LH85, MUC84]. Gastrosaccus [WPW88]. Gastrotrich [PCR80]. gaudichaudi [WR81]. GEEP [AS88, BFMS88, BAC^{+}88d, BAC^{+}88e, FM88, KWW^{+}88]. gelatinous [Lar86, VA85a]. Gelidium [Jer86]. General [Cha82]. Generated [Kah82, Thr86]. Generating [Sch83c]. Generation [HL82, KJ88]. Genetic [BTQS88, BS80d, DAC^{+}83, Fev89, GLB87, HTC84, LN86, Mee85, MFS83, OTC82, WB80, SCT87]. genetical [BCS84]. Genetics [GW81]. genus [DE89, PL89, PDIF85, SF89]. Geochemistry [Ric82]. Geographic [SG84, DCJS85]. Geographical [SJSO82, VS80, VCS81, DD88, LS84a, Mac88]. Georges [HHD^{+}84, HP86a, Col85, Kan84, LVP^{+}89, ML80, NPSV87, PN88, SM85]. Georgia [Alb83, PSB^{+}81, PAN89, MR84, MERW84, Mor85b, PCH85]. German [SHG89]. Geryon [Sv80b]. Geukensia [BH88, KLN88, LN89]. GF [Hic84]. GF/C [Hic84]. giant [BGSS88, HVK85, JJN85, LDMT86, MT85a, MT85b, Mac86a, VCS84, WK88, ZK85]. gibbosum [Ger86, HS87, LC88]. giganteum [HVK85]. giganteus [DRB86]. Gigartinales [HS85]. gigas [Bro88, Man89, Sto89]. Gill [HL80]. Gill-Epithelium [HL80]. gills [BT88, DSS^{+}85]. glacial [HS88a]. glaciale [Sam88]. glacialis [HBFA80]. glaciations [Ste89]. gladius [SK85]. glands [CL88, Liv88, RLB89]. glass [Hic84]. glass-fiber [Hic84]. Glucose [AH81, NNA84, PST87]. glutamate [CA86, FF89]. glutamine [BA83, CA87, FF89]. glutamine/glutamate [FF89]. Glutamyl [TTH83]. Glutathione [Lee88, VL88]. Glycollate [AHF87]. Gnathostomulida [Rei81a]. Goa [HP86b]. gobiid [Mag88, TE86a]. Gobiidae [WB80]. Gobius [Mag88]. Goby [WB80]. GOGAT [CA86]. golden [Slo85]. gonadal [FMM87]. gonads [LGK87]. Gonostoma [LHT88]. Gonostomatidae [LHT88]. Gonyaulax [GKA88, MHW88]. Gorgonian [Las81, Cof84, Ger84, Ger86, Las84, PF89, WB83a, YY89]. gorgonian-eating [Ger86]. gorgonians [HS87, Las85]. Gosse [SCT87]. Gracilaria [RR82]. gracilis [BS89a, BSS89, BBSS89]. graded [GKM^{+}87]. gradient [Add88b, Add88a, BK85a, Cap88a, Cap88b, DD88, KBNO85, MCM84, Moo88b, Moo88a, Sch88c, Sch88b, War88b, War88c]. Gradients [ACF^{+}80, CL88, DM83, DSG^{+}89, MFP87, OSB85, SMB85]. Grain [Ham82a]. Grain-Size [Ham82a]. Grand [BP85, PH83, SGG87, SHIP86]. granosa [Bro82]. Grant [Tod85]. grapsid [KO87]. grapsus [KO87]. grassbed [Sog84]. gravel [Thr86]. gray [Rus87, SS82]. grazer [NBMM88]. grazers [LC88, NBR^{+}88, NBMM88, SB87, TIS85]. Grazing [ASSF87, Fro87, NBT88, Par87a, TRL89, Ver86a, ALH86, Bat86, CC86b, DJTL88, Dec88, Gal89, Gif88, HM86, Kem87, LDKR88, Mon84, PB87, RG89, Rus84a, Rus84b, Sam83, SR87, SHCL87, Ste86, WS83, dB87a]. Grazlng [BH81a]. Great [KPJ84, AA84, AG82, CLS^{+}82, DCMS89, Din82a, Din82b, Rid88, Rus84a, Rus84b, SR87, She85, WH82, Wil86a]. Green [BL80, Gun80, OF82, Cha88, Him86, PF86, Phi85, SCT87, TB85]. green-lipped [Phi85]. Greenland [SBK88]. gregaria [Zel85]. Gregarious [SWSL81, Grö89]. Gregariousness [Ger86]. Grevelingen [VN82]. Grey [KOO80]. Grey-Mullet [KOO80]. grinding [Kv88]. Grooves [ZH81]. ground [EC87a, Sto88]. Grounds [HBFA80, Gar89]. Groundwater [Joh80]. group [Cus89, KBv89, vdVB86, van86b, vB87]. groups [SHPH86, VK85]. growing [DS84, NM89]. grown [JP85, Mac86a, PW87]. Growth [AP80, AS87, AHMB89, BSv81, Ban82, BW80, BK81, BK80, Bro82, Buc82, Bue87, BNF81, De 84, DD80, Fen82b, FKW81, FPRW82, GT87, HLS83, Hop88, Hov89, HH84c, KBv89, Kne81, KBU85, Lar84, LN81, LD85, NH82, RR82, SN89, Spi87, SLO82, TC81, Ver86b, Wal85b, WJ88, WU82, WK88, WBAN85, Yon80, AC88, AFHA84, AC85, Ang84, Bar86, BHNI89, Bjø86, BH88, BEM87, Bro88, BTH^{+}84, CC86a, CSD88, CS89a, CC88a, CC88b, CR84, CC88c, Che84, CF84, Col86, DCMS89, Dem85, DRCA83, Dou83, DNF87, DJ84, DS86c, Ebe86, Eml86, FB86, FL86b, GM87, GMS85, GPR86, Gei88, GL88a, GKA88, GMB87, HAHA84, Han89, HR89, HF88, Hug87, HSdWM88, IM88, JWC^{+}83b, JM85a, Jon86, JB86, KN88, KMBSJ87, KL84, KM87a, KMR87, KMR^{+}88, Kiø89]. growth [KPJ84, KA84, LVdVdV85, Lar86, LH87b, Lev85, MD87, Pag86, Pat88, Pet86b, PT84, PDWF85, RTK^{+}89, Red83, RBR86, RNJ^{+}84, Riv89, Rum89, Rus83b, SJ88a, SO84, Sch87a, SS86b, SGB89, SHB^{+}83, Spr84c, TO88, TS88b, TKS^{+}85, Van88a, WtD89, Wro84, ZWO89, ZIO84, ZK85, ZSA87, dKB88b, AFHA84, FITF80, GKM^{+}87, MT85a, VS80]. grypus [SS82]. GS [CA87]. Guanine [ABB80, Som80]. guano [BH86]. Guild [Wil82, ET85]. Guinea [MD86]. Gulf [GHC83, Lat86, LL81, YALD88, AW86, BGSS88, GOAYH86, Han86, HHD^{+}84, HP86a, HM87, HPF89, KB88, LHT88, LRS87, McC84, OD89, SMB85, Sha88a, SF89, SV85a, SV85b, Tow84, TRMF89, Tur84, Wil89]. Gulland [McG87]. Gullmar [LH88, Tal80]. Gullmarsfjorden [Øre87]. Gut [MP88b, Uch88, ZH81, BBL84, DW83, KN83, Ohm88, PD88]. Gutless [Gie81, GLD82]. Guts [IR81]. Gyre [Shu80, SM85]. gyre-influenced [SM85].

H [Hol88]. H. [JA85, Sie82]. Habitat [AC85, BWH87, CVF89, GLD82, RO87a, Sch80a, SF89, SWB89, Wei82, Ber84, DW87, HD89, Het89, HLH87, KBU85, Mac86b, MHW88, MH89, Rus87, Rye88, TCH87, TN85, TN86]. Habitats [RK82, SMO81, DR86, EC87a, LS84b, MPR85, Par89, RDD88, SO84, WAH89b]. Habits [CS81, BB84, Cho86, Kan84]. haddock [Kan84, LVP^{+}89, MN87, PN88, SM85]. Haematopus [SHG89, WLB84]. haemoglobin [DSS^{+}85]. Haemulon [SO87]. Hake [Bai81, SLO82]. hakes [GLB87]. halibut [HKS86]. Halichoerus [SS82]. Halichondria [Bar86, Bar88b, CM87, HSdWM88]. Haliclona [OS85]. Halicryptus [OS88]. Halifax [Gif88, JNCS89]. Haliotis [TB85, TC88]. halo [BBG^{+}87]. halocline [Bre86]. Halodule [Lew84, PB88]. Halophila [KCFS89]. halophilic [RBA86]. hamatus [CT85, Tis89, Van86a]. Hampshire [SMN86]. Hand [SS82]. Hand-Rearing [SS82]. Handling [aH85]. haptonema [EM89]. Haptophyceae [VA85a]. Harbor [Boa80]. Harbour [Gif88, JNCS89]. hard [D’A85a, GSC83, Gri88, LGP88]. hard-bottom [Gri88]. hare [Paw89]. harengus [McG89b, Che82, Che84, CL82, KMR87, Lam87, MK85, SP88]. Harpacticoid [Rie82a, BWH87, Dec88, SM87b, TB88]. Harpacticoida [BS89a, BSS89, CS89a, HHG84, Nil87, OFGL88, VB81]. Harpagifer [Dan83]. harrisii [LN81]. harvest [Smi86]. hatchability [LTS86]. Hatching [VCS81, SVK83]. Haul [Boa80]. Hawaii [Gri83, HBFA80, Hod85, LHF84]. Hawaiian [Bie85, Gri88, HC86b]. headland [Pag86]. Heart [MT84]. Hearts [GHCH81, MG81]. Heat [HL80, PGBN81, PF83]. Heavy [NL81, WR87, Kim85]. hebes [PH83]. height [Qui88]. helgolandicus [BR83, HM86, Hir83, RHP89]. Helice [KSM88]. Hemi [KO87]. Hemi-grapsus [KO87]. Hemichordata [DG88b]. Hemifusus [PD86]. hemolymph [BV85]. Hepatic [AE88]. hepatopancreas [CHH83]. Herbivore [RK82, Van89]. herbivore-induced [Van89]. herbivores [BH87]. herbivorous [And89b, Kiø89, PD88, RBA83, Rus84a, Rus84b]. Herbivory [ZIO84, CT89, PH86, SE85]. herbstii [McD82]. herdmani [Geo85]. Hermaphroditic [GG82]. hermatypic [SSN87, Sie88]. Hermit [VS82, BM86b]. Herring [Che82, Che84, FG88, Fui89, GMS85, HMM87, HHB88, KMR87, KMR^{+}88, McG89b, MK85, SP88, TRMF89, CL82]. Heterocapsa [CC88c]. heteroclitus [Kne81, Moe85]. heterogeneity [Ken89, LML83, MF89b, Thr86]. Heterogeneous [Fen82c]. Heterophyidae [TP81]. Heterotrophic [ABM81, Fen82a, Fen82b, Fen82c, Fen82d, HVK85, Jør82, Kle80, Li82, LNB81, LNV81, Sch80b, Sch84a, SWH84, SLN81, ALAH85, CC86b, FF89, JP87, KCFS89, KK89, LNL83, MT87, NBT88, PLM89, SSA^{+}86, SAS87, STM88, TIS85, Thi87, VB85b, Wei89b]. Heterotrophy [NK81, KB84]. heterozygosity [GB89]. heuristic [Pie87]. hexachlorobiphenyl [SKS^{+}84]. Hg [FBT84]. Hierarchical [LT84]. High [Bul82, DS86c, GHCH81, NL81, Sch80a, VA85b, BR85, CHAS83, Cro84, GL89b, KMD87, KM85, MCM84, NM89, PHT85, VK85, BSSW89]. High-Density [Sch80a]. high-latitude [CHAS83, Cro84]. high-resolution [PHT85]. higher [War88a]. Hippoglossoides [SC82]. Hippoglossus [HKS86]. Histo [MGT89, MG85b]. Histo-autoradiographic [MGT89, MG85b]. Histopathological [Auf88, Moo88b, Moo88a]. Histories [Cha82, BPH87, Dav89, Ryl83]. History [Lon81a, McD82, RHvd81, BC88, CA83, CLB88, Col85, DS86a, FS84, Fle87, Kar86, LS84b, MR83, OFGL88, Pat88, SU87, Slo85, TB89, Ver89, Yos86b]. holdfasts [OS84b]. Holothurians [Ham82a, BCS84, TMBG85]. Holothurioidea [TMBG85]. Holothuroids [Ham82b, MD86]. holsatus [Cho86]. Homarus [BL88, EC87a, Gar89, HPV^{+}87, HL89, LB86, MGT89, BS80c, Jay89]. hombergii [Beu87, SR86]. homogenization [ES84]. homomalla [Ger84, PF89]. Hong [PY81, Phi85, ST82]. horizontal [GLT^{+}86, KM87b, KM87a]. host [BS83a, DRB86, DE88, KH88a]. hourly [SO85]. house [DD85]. houses [DS86a]. Houtman [CHAS83]. HPLC [AV87a, KS87a, RGLK87, Roy89, WJ87]. HPLC-measured [Roy89]. HPTLC [WJ87]. Hudson [LDG87, MHW88]. Hudson-Raritan [MHW88]. hudsonica [SM86]. Human [CB88]. humidity [AT86]. humpback [BHP^{+}86, BPH87]. Hurricane [Bal84]. Hurricane-induced [Bal84]. Hyale [HK84]. Hyas [ALPS83, Ang84, Ang87, JA85, KA84]. hybrid [Wil89]. Hybrids [WB80]. Hydrobia [FL86a, FJ88b, Lev85, LSD85, LlJC83, Pet84, Sie82, TE86b]. Hydrobiidae [FL86a]. hydrocarbon [GMB87, WCC^{+}88]. Hydrocarbons [MST81, OS80, HMLF88, STP84]. Hydrodynamic [Spa80]. hydrogen [Scr83]. Hydrographic [Cus89, RGH89]. hydrographic-chemical [RGH89]. Hydrography [KAE81, AHF87, Pia89]. Hydroid [ACF^{+}80, Bar88a, Sch83a]. Hydroides [SWSL81]. hydroids [BM86b]. Hydrology [VDL88]. hydrolytic [MR86]. hydrostatic [FWB89]. hydrothermal [CFH89, GH86, SN89]. Hygophum [SCF^{+}82]. Hymeniaci [CM87]. Hymeniaci-don [CM87]. Hyne [CM87]. Hyperiid [Shu80]. Hyperiidea [WR81]. hypertrophic [Was85]. hypoosmotic [RW86]. Hyporhamphus [KN83, RK82]. hypotheses [Sch83c]. Hypothesis [TD81b, DCB^{+}86, Gar89, GW85, Pea85, Tod85]. Hypotrichida [Din81]. Hypoxanthine [ABB80]. hypoxia [HS86, HS88a, Kri83b, TS88a]. hystrix [HGS89, OS85].

Ice [CM82, DGHK82, KC82, BGL^{+}88, BB86, Car87, CH89, GLT^{+}86, GKM^{+}87, HC86a, Hei89, Hsi88, JS86, KSS85, KS87b, KGS87, LDG87, MLDT88, MS84, PKMS85, PSMS87, SACC88, SPK^{+}87, SBK88, SHH^{+}88, TRL89]. Iceland [Fev89]. ichiharai [Shi80]. ichthyodeterrent [PF89]. Ichthyofauna [BvdEH^{+}82]. ichthyoneuston [TÓ89]. Ichthyoplankton [Whi89, OB88c, Tow84]. idae [SSK80]. identification [GGV89]. II [BH81a, BSv81, BA87, Bar88b, BAG80a, Bid88, BF87, CC88c, Fen82b, FST85b, Gal87b, GFM89, HAHA84, Her88, Hir87, JH80b, Joh81b, JK80b, Ken87b, KM87b, KS87a, Kri83c, LP87, LTH87, LNV81, MT85b, MT87, MERW84, NBMM88, NVHHH88, RGLK87, Rus84b, Spa80, Spr84d, TL87a, TÓ89, Wei82, WS83, ZW82b]. III [AGCD86, BH81b, BAH87, CS84, Fen82c, GM89, Joh82, KM89b, LCS86, Spr84a, SvDB81]. illumination [HHB88]. Ilyanassa [Bre82, LSD85]. Image [GGV89, EMHS86]. image-analysis [EMHS86]. imagery [KB88]. Imaging [RG82]. imbricate [AP83]. Immediate [DV83]. immediately [TRL89]. immersed [Hir86, Hir87]. Immigration [KB88, van86b, SV85b]. immunoassay [TKT86]. Immunochemical [SCH89b]. Impact [Boa80, CC82, Daa87, Gib88, Gif88, HM87, OFGG82, SJNL89, TKP86, vdV85, Bat86, Kem87, Kem88, Kim84, Rei88]. implication [Her88]. Implications [BS80b, ER87, LD85, SW82a, Smi86, And89b, CT84, Chr89, CG87, DP87b, FB85, FSCP89, Kiø89, KNH86, KSF89, MB85, MS88, NM85, Pal88, SG84, TB85, TRMF89]. Importance [Amb86b, Fen82d, Li82, OR80, Pea81a, SF80, SB87, Wil86b, WF83, dKB88a, AS85, BB88, Joc89, LD89, LML83, Mar84, SW86a, Sei87, SGSD86, TB88, Wat88, WDS^{+}84, WAH89b, ap 86, TL87a]. Important [Fen82a, Gib88]. improved [CS86, KP89]. Improvements [KS83]. incident [LS83]. incise [ZW89a]. Incorporation [HT81, KNH86, TW80, Wol82, Cof84, KH88b, MR86, RNJ^{+}84, ZCF89]. Increase [Bre86, FITF80]. increments [Hov89]. incubation [LS83]. incubations [GD84a]. Indaea [HS85]. Independence [BB80, KC89]. independent [Fis83, Kee87]. independently [PBP85]. index [ME89]. India [HP86b]. Indian [Bou81, Law86, SKM85]. indicated [Hov89]. indicates [KKKW86]. indication [SWG88]. Indicative [GP82, PGJ83]. Indicator [BP81, Fis88, FF89, Gri84, Phi85, WH85]. Indicators [PY81, SC82, DRCA83]. Indices [Blo81, VMM^{+}88]. indicus [Emm80]. individual [BP86, CS89a, PT84, WR84b]. Indo [Sch88a]. indole [MMS89]. indole-3-acetic [MMS89]. Induced [Emm80, FITF80, GP82, AAN87, Bal84, BGD87, HSdWM88, LT87a, PGJ83, Reg84, SM84, SLP83, SS86b, TC87, Van89, ZW89a]. induction [BP84a, BP85, BR85, SWG88, BAG80a]. industrial [Lam86]. Inefficient [RHP89]. Infauna [ZW82a, ZW82b, Amb84, Amb86b, DL83, DV83, LJIO84, MFP87, Möl86, VK85, Wil86b]. Infaunal [ST82, CDL87, Dob83, Gra81, Ola89, Pet84, WZ85, ZW89a]. Infection [TP81]. inferred [TRMF89]. infestation [SSN87, Wil86a]. Influence [AP80, Alb83, And86, BG84, Beu87, BSP89, CJS87, Fis86, FT86, FAG89, Fuh81, GL89a, HKS86, HGS89, HL80, LH81, Lin89, LG88, Lon81a, MT85a, MT85b, Mac88, MB84b, NS87, Pos87, RWK83, Riv89, SW82a, Sch88c, Sch88b, SG87b, SG88b, WRKD84, Bro86, EWM87, Geo85, GMK85, HBJ^{+}88, KM89b, Klu84, OSB85, Sto89, dB87a, LTS86]. influenced [MR88a, SM85].

Influences [KBv^{+}83b, MCPW88, SHH81, Dec86, DG88a, Ola89, SV85b]. influencing [Arm88b, Sha88b]. infralittoral [PB85b]. Infrared [JBS82]. Ingestion [Cof84, GFS89, LlJC82, Che84, DW83, Dec86, HR87, JF88, LlJC83, NJ83, TR87]. inhabited [BM86b]. Inhabiting [KC82, OD89, TW83]. Inherent [BP81, SWH84]. Inhibition [Nel81b, WP87]. inhibitor [CC86b]. inhibitors [SSA^{+}86]. Initial [ZW82a, Cam85]. Initiation [Rie82b]. Inlet [CSS82, CSS83, DCB^{+}86, JB86, Lar86, McG89b, WWD^{+}88, FVDBD87, Par87b, RR87, DK86]. inner [BV83, BC88]. Inorganic [HHKA88, KHHB86, Li82, ALAH85, BW88, KM83, LVdVdV85, PD86, PCH85, Was83]. input [AA84, LS88]. Insect [TC81]. insecticide [WP87]. Inshore [Bul82, HMM87, WH82, GABM84, Joc89, RLB89, Tow84]. Institute [Gra86b, Gra86c, Kin86]. instrumentation [AH84b]. intact [GC85]. integer [LL83]. Integrated [Mor85b, VMM^{+}88]. integration [Bar88b, TD89]. integrifolia [TW83]. intense [RA89]. intensities [HGD86, Rüg89]. intensity [CD88, Lam87, VCS81]. intensive [RR87]. Inter [Alo87, Pea82, TB89, ZWO89]. Inter- [Pea82]. Inter-estuary [Alo87]. inter-oceanic [TB89]. inter-specific [ZWO89]. interacting [Par89]. Interaction [Bai89, Law89, MPR85, PST87, WCGG86]. Interactions [ARGL87, FJ88b, FC82, Grö88, Jer85, NPSV87, TIS85, BAN89, BY83, BT85, DD88, DL86b, GCD87, Jer86, LCS86, Lev85, Möl86, MKHD85, PF83, PN88, She85, SSA^{+}86, WZ85, WW88b, CS84]. Interannual [SS89b]. interchange [Whi89]. interface [BAN89, BBR88, ES87, HA87, PKMS85, TRL89]. Interference [DBFM80, LSD85]. intermediate [KM89a]. intermedius [JKJ88, KGV84, LGK87]. intermittent [EMA88]. intermittently [JB86]. Internal [Kah82, Sha83]. International [LTS86, LS84a, SJSO82, VS80, VCS81]. Interoceanic [Ver89]. interplay [LDER83]. interpopulational [ME88]. interpretation [AH84b, BRG83]. Interrelations [KM88a]. Interrelationships [ABM81, MRF80]. Interspecific [Båm88, BM86b, MJPH^{+}88, Smi86, Col84]. Interstitial [TW80, MFP87]. Intertidal [AP80, ABM81, Blo80, DP81, DB82, DG82b, Gra81, Gun80, Kne81, McD82, RMO88, SMO81, AT86, Alo87, AH84a, And86, Arm88a, Arm88b, Asm87, BH86, Bro86, CB88, CSDZ86, D’A85a, DD88, Fai88, FSBOG84, Fle87, FW86, GT87, HD89, HH84b, HTC84, Hic88, HG86, Jer85, JS87a, LCF89, LML83, Mac86b, Mac86c, Mac87, MUC84, MB84b, MB85, Mor85a, MT84, MH89, PAF86, Pie83, Pie87, Qui88, Rei88, Rei83, Rei87b, SHPH86, SCT87, Tam85, Tam87, TH84, TS88a, TN85, TN86, Van89, VP85, dHS88, dKB88a]. Intestinal [NL81, VL88]. intestinalis [Sch83a]. Intra [Pea82, TB89]. intra- [TB89]. Intra-Specific [Pea82]. intracellular [FF89]. Intraspecific [BEM87, MJPH^{+}88, Sma83]. Intrinsic [Ban82]. Introduced [Bre82, NT85]. introduction [Add88b, Cap88a, Moo88b, Sha88a, War88b]. inundation [WS88]. invadens [JNCS89]. inverse [VP88]. inversion [Gra86a]. Invertebrate [OS84b, CYC83, Hug84, PH83, Sto86, SB87]. Invertebrates [Ban82, Sch82a, TD81b, Amb86a, Bje88, Bro86, BMRN84, CHH83, Hig85, Hir87, Lee88, MS88, NG86, Pie83, Sha83, SGSD86, VA85b, WA87, Zel88]. investigating [UP88]. Investigation [BS80b, SKS^{+}84, Jer86, Sie88]. Investigations [DGHK82, FUM84, SM84]. investment [BRS86, RSM88]. Involving [ABB80]. Ioszanoi [WB80]. Ireland [CM87, TÓ89]. Iriomote [SSN87]. Irish [GW81, PH81, TL85b]. IRPE [Gra86b]. irradiance [BS89b, CG87, FDM86, HR89, LS83, MHL85, NM85, PSMS87, PDWF85, Riv89]. irradians [BEM87, WRC83]. Irradiated [LB81]. irroratus [HL89, Joh81a, Joh81b, Joh82]. Isla [SO83b]. Island [Bou81, Bul82, CC82, CC88a, LDDH88, LWE87, MD86, MP86, PAN89, SSN87, VCS84, AA84, Sto85, TN85, AB89, BB84, Bla84, FHKR85, SM86, Tra88]. islandica [Fev89, MW83]. Islands [CHAS83, Hei89, KB80, Lat86, MJ84, TN85, TN86, Gri88, WPS89, WYG85]. Isochrysis [Wil83]. isolated [BT88]. isolation [SCT87, TL85a]. isolepis [GB84]. Isomerase [BS80a]. isometrically [GCSP88]. isomorphic [HS85]. Isopod [SF80, PB89, VAF83]. Isopsetta [GB84]. Isostichopus [Sv80a]. Isotope [HH85, CD88, EWM87, Fuh87, SKIR89, STM86]. Isotope-dilution [HH85]. isotopic [DSG^{+}89, KB84]. Isotopically [BCF82]. Israel [TP81]. items [DCNS84]. IV [Fen82d, GCD87, PKMS85, Spr84b]. Izu [Hir86, Hir87].

J [JNCS90]. jack [HTL85]. Jackass [Wil85]. jacksoni [SH84]. Jamaica [Ham82a, Ham82b, KHW88]. Jamaican [Ryl83]. James [Man88]. Japan [FW86, Hir86, Hir87, JS87b, SSN87, Ued87]. Japanese [AI88, TO88]. japonica [AI88, KM83]. japonicum [TO88]. japonicus [DAC^{+}83]. Jasus [BB88, Mac85a]. Jellett [JNCS90, Mil90]. jellyfish [MWM86]. jiang [NVLL88]. Jolla [Gun80, JJN85]. journal [Kin88]. Joyuda [Sto88]. juliana [Kem81]. June [LSB85, BNVP87]. Just [Kem81]. Juvenile [CS80, Hoe88, SLO82, AC85, Ang84, Bai89, CPM89, Gib88, HB86, Kv88, KBU85, KA84, LVP^{+}89, MF89a, NPSV87, NC89, OSB85, Rum89, SV85a, SLMB88, Sto89, TÓ89, Wal85a, WWB86, WAH89b]. Juveniles [OOS82, Bd89, Keo89, Ov87].

K-systems [Sha88b]. Kaneohe [Hod85, LHF84]. Katodinium [Coh85]. Kattegat [Ric85a]. Kelp [Car82, KNL82a, LNB81, LNV81, NLVS80, NFG82, SLN81, SNL82, SFN82, TD81a, BCCH84, DKMR83, FD89, GM87, Ger89, JJN85, Ken87a, Ken87b, Ken88, MS89, OS84b, PH87, SDRN84, SN85, SN88, STM86, SK84, TW83, VCS84, ZK85]. kelp-bed [SDRN84, SK84]. kelp-dominated [DKMR83]. Kenyan [McC89a]. Kept [RK81]. kerathurus [DAC^{+}83]. Kerguelen [Law86, BTQS88]. Kidney [GCK82, BFP87]. Kiel [Bar86, Bar88b, Bre86, Joc89, KMDS85, LSB85, OS80, Sch87b, TS88b]. kinetic [HPC89]. Kinetics [AH81, GL88a, JK80a, Pet89, AFHA84, Gal89, JM85a, SWH84]. king [SS89b, Slo85]. Kong [PY81, Phi85, ST82]. koreni [DS86b]. Kosterfjorden [Båm88]. krill [EM87, HH89, MR84, MERW84, Mor85b, O’B88b, SHH^{+}88, Wat86].

L [ARGL87]. L-amino [ARGL87]. L. [Cho86, KKKW86, SCT87]. labeled [Cam85, Red83]. Labelled [BCF82, Sch80b, SFN82, MGT89, RG89, Sch84a, SKS^{+}84, VAF83]. Labidocera [CT85]. Laboratory [DB81a, Gif85, KM80, OF82, Pet84, ALPS83, Ang84, CLB88, CT84, KA84, LSD85, Pet86b, Pet88, RDF^{+}88, SHB^{+}83, SG88b]. laboratory-reared [CLB88]. laceration [Cla85, Wah85]. lactuca [SJ88a, SJ88b, WAH89b]. Lacunar [RC86]. Laetmogone [TMBG85]. lag [Col86]. lagged [HH84a]. Lagoon [Ham82a, Ham82b, FVDBD87, Gri84, NN85, OQ83, TGA89, Wil87, BBBM80, Thr82]. lags [Rud88]. Laguna [Sto88]. Laing [MD86]. Lake [VN82]. lalandii [BB88]. lamarcki [Bro80]. Laminana [LNB81, LNV81]. Laminaria [BCCH84, DS86c, Lün88, MKS80, NLVS80, Rus83a, Rus83b, Smi86, SLN81, TL84]. Laminariales [Ger89]. lance [BTH^{+}84, MP86, SHB^{+}83]. lancifolia [PDP87]. land [Was83]. land-locked [Was83]. landlocked [Was85]. Langesundfjord [AE88, Lee88, GAC^{+}88, HWC^{+}88, War88a]. Lanka [Jay89]. lantern [Sam88]. Lanternfishes [SCF^{+}82]. lapillus [LH85]. Large [Jos87, LH88, OE81, AAH87, FL86b, Fro87, FG88, GMS85, HH89, Hop88, McC89c, SCSD86, VA85b, Zel88]. Large-scale [Jos87, LH88]. Larger [KSR80]. larix [D’A85a]. larva [Mar89a]. larvacean [DS86a]. Larvae [Bre82, Che82, IR81, JH80c, Joh81a, Joh81b, Joh82, Kem81, OOS82, Paw89, RRB^{+}80, SWSL81, SJSO82, SSK80, Sv80b, VS80, WU82, ALPS83, BK85b, Che84, Che89, CE89, CYC83, DCNS84, Eml86, ELR88, EMG89, FL88, FH89, FWB89, FG88, Fui89, GB84, GP88a, GMS85, GFG89, GOAYH86, Grö88, Grö89, HMM87, HHB88, Hod85, JL86, KAE87, Kan84, Kee87, Mac85a, Man89, MW83, Man88, MKG88, McG86, McG89b, MP86, NPVW86, Ols85, Pau83, PA84, Pur84, Ric85b, Sha83, SS89b, STP84, Spr84c, Spr84d, Spr84a, Spr84b, SP88, Sul84, TB85, WRC83, YLH86, LH87b]. Larval [Bai81, BK81, Buc82, CL82, DD80, Eml86, LN81, LHH81, MM81, SS80a, SW82b, vDBS81, Ang84, AHMB89, BTH^{+}84, CS89a, Chr89, FL85, FL86b, GCSP88, Gle86, GHC83, HPV^{+}87, Hov89, JA85, Jen88, Kv88, KMR87, KMR^{+}88, KA84, LD89, LGP88, MK85, OWZ89, PT84, SM85, SP88, SVK83, TL87b, TKT86, TRMF89, TC87, TÓ89, Wal85b, Wro84, vdV85, van86b]. larynx [Sch83a]. Late [KS87b, MB84a]. Lateral [MKD^{+}86]. lateralis [Luc84]. latex [NBT88, PHL^{+}86]. latistrigata [Jer85]. Latitude [BK80, Cla87, CHAS83, Cro84]. Latitudinal [JWC^{+}83b, TO88]. latus [STPA88]. Lawrence [Hei89, PLT87, PT89, HHD^{+}84, HP86a, PCM86, SMB85]. layer [HM87, LTH87]. layers [KN88, Kim85, NK86]. leachable [BH85]. leachi [BBBM80]. Lead [KOO80, WY82, Cha88, War84]. Lead-Exposed [KOO80]. leaf [ZIO84]. Learning [CH84]. least [Pat88]. least-squares [Pat88]. leather [AP83]. leaves [BH85, NGR86]. lebouriae [HMH89]. ledge [DE89]. leech [SBR84]. leidyi [Pur88]. Leiostomus [GHC83, GOAYH86]. Length [HMD88, SW82b, BS89a, BR83]. Leptasterias [HGH87]. Lernaeocera [Ans89]. lerruginea [LH81]. Lessonia [MS89, OS84a, SO84]. lethal [BAC^{+}88b]. leucine [KNH86, KH88b]. leukodermatus [Gie81, GLD82]. level [FER85, RL83]. Levels [Yan82, CA85, ME88, OA88, PDJ86, RWK83, SO84, SHG89, War88a, Rus84a]. Libinla [AHMB89]. Life [BC88, CA83, Cha82, Col85, Fle87, Lon81a, McD82, Pat88, RK81, RHvd81, RSM88, Ryl83, Sie82, Slo85, VTHVd81, Yos86b, AZ85, CLB88, Dav89, DS86a, EM89, HK84, Kar86, LS84b, MN87, MR83, OFGL88, SU87, TB89, TR87]. Life-Cycle [VTHVd81, TR87]. Life-Histories [Cha82]. life-history [MR83, SU87]. ligands [ZW89b, ZW89c]. Light [AP80, Mar89a, Pau83, RHvd81, TC87, VCS81, AC88, BBSS89, Che89, CC84, CD88, CP84, DA85b, HMF85, HGD86, KAE87, Kel88, LL83, NS87, PKMS85, PGV^{+}89, SJ88a, SJ88b, SPK^{+}87, Ste86, VGJ89, ZSA87]. Light-induced [TC87]. light-limited [AC88, PGV^{+}89]. light-saturation [CP84]. lignocellulosic [BH85]. Like [RMZ82]. lilljeborgi [Nil87]. Limanda [LH81]. limatula [PAF86]. limatulone [PAF86]. Limit [SA82, BCCH84]. limitation [And89b, Eme89, FWE89, Gri88]. Limitations [Hol88]. limited [AC88, ET84, PGV^{+}89, SF85, SP84a, WC88, ZCF89, ZSA87]. limiting [VDM89]. limits [HH84b, JWC^{+}83b, Pie87]. limpet [Fle87, Jer85, MS89, PAF86, Qui88]. Limpets [NLBFD80, HD89, LWE87]. Lindåspollene [AM88, SW86b, Was83]. line [AZ85]. linear [Pat88]. lines [Ebe86]. link [WH88]. Linuche [CK89]. Liocarcinus [Cho86, aH85, ap 86]. lion [GNBO85]. Lipid [CL88, Cla82, GMB87, JP85, KKT81, TBRBA85, PW87, Par87b, RLB89]. Lipids [OTK80, Ohm88]. lipopolysaccharide [TBRBA85]. lipped [Phi85]. liquid [KM85]. Lithodes [Slo85]. lithotrophic [JMG84]. litoralis [BC88]. litter [ABT89, FVDBD87, WBVS86]. Littoral [BS80c, BvdEH^{+}82, Kro88]. littorale [FP82]. littoralis [PM88, RW86]. littorea [Bak88, FG87, FJ88b, LH85, Liv88]. Littorina [Bak88, FG87, FJ88b, LH85, Liv88]. Live [Mon83, BG84]. Lived [Kem81]. Liver [SC82, BFP87, SM84]. Liverpool [AHF87]. Living [Col80, FT82, Jen81, Jen82, Alo87, AH86, BM86b, JNCS89, Jen87, Kaa86, LPV87, NW84, PM88, SRTM83, TW83]. lmplications [BP81]. lnfauna [Jac80]. lnteractions [SED81]. lntertidal [CD81, LL81]. Load [GCK82]. Lobata [Pur88]. Lobophora [CT89, dB87a, dB87b, dKB88a, dKB88b]. Lobster [BS80c, Gar89, EC87a, HL89, Mac85a, MGT89, STPA88]. lobsters [BB88, BL88, HPV^{+}87, HB86, LB86]. localisation [MGT89]. location [CK89, HMLF88, Lin89]. locked [Was83]. locomotion [Wah85]. Locomotor [DB81a]. log [Don87, Gra83]. log-normal [Gra83]. log-spiral [Don87]. Loligo [Pat88]. Long [BL87b, CC82, DL82, ID88, Kah82, Kem81, NCG^{+}83, OE81, PH87, RG85, SSN87, ELK84, FER85, FHS88, GC83, HMF85, LB87, LH88, Rei87b, SV85b, tBB83, CC88a, FHKR85, MP86]. long-day [tBB83]. long-distance [LB87]. Long-Lived [Kem81]. Long-Term [DL82, OE81, BL87b, ID88, NCG^{+}83, PH87, RG85, SSN87, ELK84, FER85, FHS88, GC83, HMF85, LH88, Rei87b, SV85b]. longiceps [NS80]. longicornis [RHP89]. Longshore [Don87]. longterm [Dah86]. loop [HAA^{+}88, NR89a, Wei89b]. Loss [LB81, Cam85, Dah86, Gei88, ZJ83]. Losses [NLVS80]. Lottia [HD89]. Lough [CM87]. Louisiana [FC83, Sha88b]. Low [JM85a, OFGG82, ELK84, FER85, RWK83, SJ88b, TL84, TW88, WRKD84]. low-level [FER85]. Low-temperature [JM85a]. lower [FK84, MHW88, OB88c]. lowers [WLB84]. lsopod [Has81]. lucens [PAH89, Stu86]. Lucidum [Som80]. luciferase [KS83]. luciferin [KS83]. luciferin-luciferase [KS83]. Lucinoma [CVF89]. lucunter [LS84b]. lugworm [Rei87b]. Lumen [NL81]. Lunar [Has81, JG89]. lysianassid [SM87a]. lysosomal [Moo88c].

m [LD85]. M. [BTQS88, GW81, GLB87]. mackerel [HTL85, PA84]. Maclmage [EMHS86]. Macoma [Bd89, BMRN84, CL86, HL85, JCL86, Mee85, Ola89, Rei83]. macro [MFP87, Qui88, War88a]. macro-algae [Qui88]. macro-benthos [War88a]. macro-infauna [MFP87]. macroalga [Mor83]. macroalgae [ET84, JS87a, SJNL89, WtD89, WAH89b]. macroalgal [EC87a, FWE89, LvB87, Paw89]. Macrobenthic [HFK82, Lew84]. Macrobenthos [Nel81a, Gri84]. Macrobrachium [SSK80]. macrocera [NS87]. ’macrocosm [SH89]. ’macrocosm’- [SH89]. macrocrustacean [RO87b]. macrocrustaceans [Sto85]. Macrocystis [JJN85, MJ84, OS84b, SO83a, SO83b, TW83, VCS84, ZK85]. macrodidal [HP89]. Macrofauna [WU80, GB86, GAC^{+}88, HP86b]. macrofaunal [DG88b, Kem87, Rid88, SG84]. macrofaunal-deposit [DG88b]. Macroflocs [KM80, KM88b]. Macroinvertebrate [GL82]. macroinvertebrates [OD89]. Macronutrient [OK89]. macroparticulate [Bid84]. Macrophyte [KNL82b, LL81, FB87]. Macrophytes [LRH82, MH89, RV84, vM87]. Macroplankton [Pia89, KAA88]. macropredators [BV83]. Macrosetella [BS89a, BSS89]. Macrothallus [Rie82b]. macrozoobenthic [Jos87]. macrozoobenthos [Bre86]. macrozooplankton [Nel89]. made [PG88]. maenas [AAH87, AAN87, BT88, BV85, CL88, CH84, RN89, ap 86]. magellanicus [BGSS88, LRS87, MT85a, MT85b, Mac86a, RLB89]. Magnetic [Bul82]. Maine [BGSS88, LRS87, OD89, SMN86, Tow84, TRMF89]. maintain [KM87b]. Maintaining [Spa80, HD89]. ’maintenance [Thi87, HB85]. Maizuru [Ued87]. Majidae [Ang84, AHMB89, JA85]. Major [BY81, FSZ89, She80, She82, CSS83, MW87, OQ83, KM83, WF80]. male [ZE89]. Mallotus [MFS83, Ste89, TL87b, TL87a]. man [Pur84]. management [Smi86]. managing [SH89]. manganese [RBA88]. mangle [BH85]. mangrove [Alo87, ABT89, AK88, BH85, BMZZ88, BW83, BSC85, BW88, BAN89, DL86b, GLNR86, KAK88, NGR86, PB89, RDD88, SV85a, SV85b, TCH87]. Mangroves [BW81, FVDBD87]. mannitol [RW86]. mapping [BLR86]. marginal [JS86, SBK88]. marginalba [FUM84, Mor85a]. mariae [HMH89]. mariae-lebouriae [HMH89]. mariaelebouriae [HMT83]. marina [MGC82, AK88, BTL86, BSC85, Cap82, DA85b, FF89, HFK82, KMBSJ87, KMAM84, MAWM83, NG86, Rei88, VN82, ZSA87]. Marine [Aik81, ACF^{+}80, AA81, ABB80, AH81, AV87b, BAG80b, BAG80a, BB80, Cam82, DS86a, Eva82, FT82, FPRW82, Gie81, GTG81, Har82, HW82, ITBW81, Jen81, Jen82, JBS82, Kle80, KM80, Kui81, MST81, NLL81, NL81, NK81, OR80, OOS82, Ovi81, PB81, PGBN81, Rie82a, RS80, SS80a, Sch82a, SHH81, Spa80, SNL82, TC81, Thi81, TD81b, Wah89a, ASMV88, Amb84, Amb86b, AH87, ALAH85, AH86, ABJS88, BA87, Bar89, BP84a, BLP87, BGD87, BH83a, BLK88, BP86, BGR86, BA83, BMGB88, CC86a, CHH83, CA87, Col84, CS84, CT85, CYC83, DL86a, DK85, DRCA83, DSG^{+}89, EMHS86, EAM87, EMA88, ET85, FNH83, FS84, FJ88a, FBCH83, FBT84, FB86, FDM86, FH89, FB87, FA83, Gar84, GH83, GL88a, Gif85, GGV89, GFG89, GMB87, Gra86a, GMQ^{+}88]. marine [GC83, Gra86b, Gra86c, GSC83, HAA^{+}88, HK86, HYT^{+}88, Hei89, HHE89, Her88, HGD86, Hob88, HHKA88, HF88, Hug84, HSdWM88, JKJ88, JBAH83, JP85, JM85a, JM85b, Jon86, KN88, KN84, KM87a, KM87b, KM89b, KNH86, KS83, KDT^{+}88, LH87a, LDH83, Lam86, LN86, Lee88, LS84a, LSP84, LB87, LT87b, LPV87, Mar89b, MJPH^{+}88, MMS89, MSS86, MCA89, McG89a, McG86, MS88, MR86, MPR85, MWM86, NBR^{+}88, NBMM88, NCG^{+}83, NNA84, NW84, NBT88, OC89, OQM^{+}87, PV85, Pal88, Par87b, PST87, PD86, PL85, PY86, PR84, Pih85, PDWF85, PHL^{+}86, Rei87a, RO84, Scr83, SG87a, SLP83, SRTM83, SS86b, SA89, SPL^{+}84, SH89, SG87b, Spi87, SGSD86, SM87b, TAK88, TTH83, TE86b, TW83, TNR86, Van88b, VA85b, VS84, WA87, Wat88, WE88, WH88, Wil86b, WJ87]. Marion [BB84, AB89, Bla84]. maritima [Thu84a, Thu84b]. markers [WJ87]. marsh [EWM87, FC83, FCFP84, Gle86, HKd88, Het89, Lin89, MW85, OK89, Ov87, RD89, RO87b, SC84b, WBVS86, WZ83b, WWD^{+}88, WS88, KPJ84, Kne81]. marshallae [Pet86b, Pet88]. marshes [RMO88]. Maryland [Gal89]. Masfjorden [KAA88]. Mass [Ban82, LPV87, BCd84, Bal84, GMF88, Har85, KMDS85, KB88, SLMB88, Sto89, dB86, dB87b]. Mass-Scaled [Ban82]. Massachusetts [PM88, WC83]. massive [DE89, HLS83]. Mastigias [MWM86]. mat [JB86, ZM84]. Material [SN81, SNL82, BSSW89, NJ83, STM86, TP85, Was83, Was85]. Materials [Ham80, BW88]. maternal [GB89]. mathematical [MT87]. Mating [NS80]. Matrices [Hoc81]. Matter [EHF^{+}81, Ise81, LlJC82, LC82, MRF80, OS80, SLN81, Bid84, FST85a, GT87, HAHA84, KE87, LTH87, MCA89, SW86b, TIS85]. Matthew [Hei89]. Maturation [Emm80]. mature [Lam87]. maturity [Jay89, Øre86]. Maxima [CSB81, JS88, KNL82a, LNB81, LNV81, NLVS80]. Maximizer [HS81]. maximum [ESR^{+}88, JS85, Van88a]. maximus [KAE87]. may [Sha83, BNVP87]. May/June [BNVP87]. McMurdo [KGS87, PKMS85, SPK^{+}87]. Meadow [HFK82, Ov87, WB83b]. Meadows [You81, WW88a]. Meal [SM87a]. meandrina [Col80]. Means [Sam82, BH87, Hop83, RNJ^{+}84]. Measure [MP81, BP84a, CC86a, RC86]. measured [Roy89]. Measurement [MSS86, MB89, MR86, BBR88, FB85, GD84a, GGV89, LS83, Mon84, SSPA89, TL85b, VA85a]. Measurements [HW82, How85, PSB^{+}81, BT88, CA87, FF86b, GM87, HH85, Hop83, LlJC83, MHL85, Ohm88, ORK^{+}86, PF83, Rie89]. Measures [KR82, CDL87, FNH83]. measuring [ABJS88, KNH86, PB87]. mechanical [Jør83, JKMR84, ap 86]. Mechanics [RNK87, DP87b]. Mechanism [FITF80, NL81, DRB86, Hic88, HSRM86, RO84, SP88]. Mechanisms [SCF^{+}82, Hig85]. mechanistic [GP86]. Media [WU82, Yon80]. mediated [Gar89, HD89, HA87, HSRM86, KM89a]. mediating [Rus87]. Mediterranean [SRS89, BS80a, BTQS88, LN86, Mac81, MR88a, PB85b, PPCR86, RTLR88, STPA88, VB85a, VB85b, Vel86, WEHF85]. megabenthos [Bri85]. megafauna [SGG87]. megalopa [Ang87]. megalopae [Chr89]. Meganyctiphanes [Nic84]. Megaptera [BHP^{+}86, BPH87]. Meiobenthic [FP82, Dec86, EC87b, Mon84, NC89, Rei87b, Rud88, WCC^{+}88]. meiobenthos [CPM89, DL86b, WG84]. Meiofauna [CM82, FC83, Hoc81, MRF80, PCR80, Alo85, AH86, Arm88a, Arm88b, Arm89, Bod88, DG88a, FCFP84, FER85, Gib88, HWC^{+}88, Hic88, MFP87, Mon83, Pal88, Sog84, WB86, WE88, WZ83a]. Meiofaunal [BS80b, FSZ89, Hic86]. melanochir [KN83, RK82]. Melanogrammus [SM85]. meleagris [GR89]. Melinna [ID88]. Mellita [BS83a]. mellitae [BS83a]. membranes [LD85]. Memoriam [Mon89]. Memory [Cos82]. mendionalis [SN85]. Menidia [Ben84, GD81]. Menippe [Sv80b, Wil89]. mensuration [EMHS86]. Menten [SWH84]. MEPS [Add88b, Cap88a, Kin88, Moo88b, War88b]. Mercenaria [GCK82, LGP88, Pet86a, Sv80b]. Mercury [Din81, SHH81, KBv^{+}83a]. merguiensis [CS81, SV85a, SV85b]. meridionalis [Gri80a, Gri80b, SDRN84]. MERL [Kel88]. Merluccius [Bai81, GLB87]. Mesidotea [HS88a]. Mesochra [Nil87]. mesocosm [BFMS88, Bak88, GAC^{+}88, OPN^{+}84, OKSB86, OQM^{+}87, RDF^{+}88, Sch88c, Sch88b, SPSS89, Tam89, WCC^{+}88]. mesocosms [AE88, DO86, Kel88, ORK^{+}86, SM87b]. Mesodinium [Cra89]. mesohaline [WZ83b]. mesopelagic [GB83, LHT88, WR87]. Mesopodopsis [WPW87]. Mesoscale [EM87]. mesozooplankton [Fro87]. Metabolic [Cam85, GCSP88, HGD86, HGS89, RL83, SSA^{+}86]. Metabolism [HT81, PF83, SC82, SHIP86, AH84a, Bid88, BBR88, FNH83, GHS85, Ger89, HS86, HPF89, Hol88, HSW89, KN84, KA85, KCS89, KHHB86, KAK88, LTH87, MC88, MBP^{+}86, NN85, ORK^{+}86, Riv89, Sor85, Stu86, Thi87]. metabolizing [SDMN88]. Metal [NL81, RTW81, TE86b, WU82, AS88, BAC^{+}88a, BBH^{+}88, CL86, OA88, PWM85, VMM^{+}88, War84, WR87]. Metal-complexing [TE86b]. Metal-Contaminated [WU82]. Metallothionein [OA88]. Metallothioneins [NLBFD80]. Metals [BFP87, PY81, THSW80, WY82, Kim85, OOF83, Phi85, RYD84, SC84b]. Metamorphose [Kem81, Paw89]. metamorphosed [Keo89]. metamorphosis [Ric85b]. Metazoa [WCGG86]. Method [BL80, DCB82, Fuh81, JMG84, RG82, CC88b, DTLN85, Gra83, Hol88, How85, KP89, LS83, MW84, Uch88, WB83a]. Methodology [GP82, JH80b, DL86a, KS83]. methods [BRG83, HPC89, LTS86, Mac86c, Mar83b, SP84a, VP88, aH85]. methylumbelliferyl [Hop83]. methylumbelliferyl-substrates [Hop83]. Metridia [Bat86]. Metridium [Wah85]. Mexican [FVDBD87]. Mexico [GHC83, GOAYH86, LHT88, PE89, Tur84, Wil89, YALD88]. micans [RWK83]. Michaelis [SWH84]. micranthus [Sch84b]. Micro [CC82, Jen88, NLL81, SNL82, ALAH85, Kim85, OS85]. Micro- [Jen88]. micro-flagellates [ALAH85]. micro-layers [Kim85]. Micro-Organisms [NLL81, SNL82]. micro-predator [OS85]. Micro-Zooplankton [CC82]. microalgae [BGL^{+}88, BBL84, CG87, DL83, FB86, GP86, GLT^{+}86, Hsi88, LCF89, LDG87, LM85, Mac86c, MLDT88, PKMS85, SPK^{+}87, SHH^{+}88, VP85]. microalgal [DL85, DL83, PSMS87]. Microarthridion [FP82]. Microautoradiography [Dou83, DNF87, PAN89, RNJ^{+}84]. Microautoradiography-based [Dou83, DNF87]. Microbenthos [DL86b]. Microbes [Alo85, AFF^{+}82, Mil89, PF83]. Microbial [BH85, Bid84, BP88, Bid88, Cam82, CS87, DGHK82, DG88b, FST85a, FST85b, GCBM82, HAA^{+}88, HT81, HSD87, KJ88, KNL82b, LNB81, MMS89, NK81, RMZ82, Sch89a, Bra87, CSDZ86, DS86a, DG88a, GKOW87, GKM^{+}87, HM87, HPF89, HSW89, KN88, KSVB89, KGS87, LHF84, NR89a, NK86, PKMS85, PSMS87, PDJ86, Pet89, Sch88c, TP85, Wei89b, BP88, Bid88, Her88]. microbiological [FB85, FF86b, Rei88, RGH89]. Microbiology [NM89]. Microbiomass [MR88a]. Microcladia [KH88a]. Microcosm [KNL82a, OWP80, PLM89]. Microcosms [Ovi81, ASMV88, EAM87, EMA88]. Microdistribution [MFP87]. microfauna [MS84]. microflagellate [AGCD86, ALH86, DS84, GCD87, SAS87]. Microflagellates [Fen82a, Fen82b, Fen82c, Fen82d, LNL83, NBT88, SHCL87]. microflora [Mac86b, Mac87, SO85]. microfloral [Sha88a, Sha88b]. microgradients [AH87, AK88]. microgrowth [LG88]. Microhabitat [DeW87, HB86]. microheterotrophic [CC89]. microheterotrophs [NL84]. microlayer [PDJ86]. Microlayers [DGHK82, Ign87]. Micromesistius [BCA89]. Micromonas [PDTH86]. micronekton [KAA88, Sch88a]. microneuston [MR88a]. micronucleus [BMGB88]. Microorganisms [NFG82, Sch81a, AS86, HSB88, HMLF88, MR86, PB87]. microparticulate [Bie85]. Microphytobenthos [RGLK87, Cd83, RG85, SG88b].

Microplankton [Har82, HH85, Jen88, MR88a, Pro87, SHIP86, SKM85, SE85, VB85b]. Micropogonias [GHC83, GOAYH86]. microscope [JP84]. microscopical [SM84]. Microscopy [Fuh81, CS86]. microsomal [AE88, Liv88]. Microsphere [SFN82, PB87]. microspheres [CS86]. microstructure [Kv88, KBv89]. microtopography [Gra84]. Microzooplankton [Gal89, Gas85, Gif88, LDKR88, Par87a, RG89, Ver86a, WR89]. Mid [WH82, KB87, Rus84b]. Mid-Shelf [WH82, Rus84b]. mid-Wales [KB87]. Middle [Gas87, SB82]. midine [DNF87]. Midtidal [Boa80]. migrating [BFP87, Kaa89]. migration [AS85, Jen85, KM87b, PAH89, Pol84, SLMB88, Sto89, WB86]. Migrations [Leb82, LD89]. Migratory [BHP^{+}86]. Mileikovsky [Cos82]. miliaris [Daa87]. milkfish [KBU85]. Miller [JNCS90]. mimicry [DD84]. Minas [GD81]. Mineral [KB80]. Mineralization [FT82, HMLF88, BLK88, LBH89, Pet89]. minerals [GABM84]. mini [dHS88]. mini-electrode [dHS88]. miniata [Rum89]. Minimizer [HS81]. minimizes [Blo80]. Minimum [SJ88a, KMDS85]. minutus [Mag88, WB80]. misuse [Gra83]. mitchilli [LH87b]. mitigates [DE88]. mixed [AS85, DA84, FF89, HHN^{+}83, HlWPH84, LD89, LTH87, PST87, PN88]. Mixing [Ovi81, DK85, HHB88, KSVB89, LCP84]. mixotrophic [STM88, SST89]. Mixture [THSW80]. Mn [Dah86]. Mnemiopsis [Pur88]. Mobile [TW81, ET85, Pih85]. mobilities [Pos87]. mode [GB89]. Model [Kah82, VN82, AWSB83, AN88, BT85, CS89a, CL84a, CC88c, CC87a, CP84, DOB^{+}89, Fro87, GP86, GL88b, Hug84, KSF89, LP88, MT87, MR84, Mor85b, NM85, PLT87, Pal88, Pie87, Sha88a, SH89]. modelanalysis [PL85]. Modeling [SFHU82, JLMR88]. Modelling [AGA89, GAA89, LMO86, LDER83]. models [Chr89, DO86, GM87, LH87a, Pat88]. modes [ZE89]. Modified [CA87]. molecular [DPJ87, Mø83]. molle [Cow81]. mollusc [DSS^{+}85, Sch87a, VAF83]. Mollusca [AGA89, CSW84, Sha89, Sie82]. Molluscan [Amb82, Moo88c]. molluscs [LlJC83, aH85]. mollusk [Sma83]. mollusks [Fis83, RR87]. molt [HLH87, LOvM89]. moluccensis [Dav89]. Monacanthidae [CL84b, KM87c]. Monacanthus [CL84b]. Monas [BB87]. Monhystera [VTHVd81]. Monhysteridae [VTHVd81]. Monica [SLE^{+}89]. Monitoring [Dob83, RG82, Bod88, SDMN88]. mono [AE88]. mono-oxygenase [AE88]. monocultures [EB86]. monodisperse [NBT88]. monomictic [TGA89]. Monophosphate [AA81]. monoxide [JMG84]. montagui [KB87]. Monterey [DS86a]. Monthly [Wil82, SO85]. Montrose [FPRW82]. Moorea [LDDH88]. moored [FHS88]. Morbihan [Law86]. mordax [LD89]. Moreton [JG89]. morhua [NPVW86, STP84]. Morocco [Sha89]. Morone [Che89]. Morphological [CT89, JJN85, CC87b, PH86, dKB88b]. Morphology [Fen82a, Som80, Las84, LS84b, MH84, SH84, TH84]. morphometric [BWH87]. morphometries [NPVW86]. morphs [HTC84]. mortalities [BCd84, Bal84]. Mortality [FKW81, Har85, Kne81, Lon81a, Rub85, TD81a, BGD87, GT87, Gly85, Gul87, GR89, HTL85, KD86, McG86, McG87, McG89b, MR83, OWZ89, PY86, SO87, TL87b, TL87a, Tan88, Tra88, Wro84, dB86, dB87b, van86b]. Morula [FUM84, Mor85a]. mosaic [DW87]. Motile [KMM88, Amb86a]. moult [Ang87]. moulting [Mor85b, WR84a, WR84b]. mouth [Whi89]. Movement [Cow81, BHP^{+}86, FF86a, Hoe88]. Movements [IR81, BP84b, HPV^{+}87, KB88, PG88, SP88]. muciferous [EM89]. Mucilage [LNB81, LNV81]. mucous [DD85]. mucus [Cof84, JKMR84]. mucus-net [JKMR84]. Mud [LN81, SJSO82, BL88, CPM89, KSM88, LB86, SVK83]. mud-crab [KSM88]. muddy [AH86, ID88, SMB85]. mudflat [LSD85, NT85]. mudsnail [Lev85]. mufria [MUC84]. Mugil [KOO80]. Mulinia [Luc84]. Mullet [KOO80]. multicalorimeter [GMQ^{+}88]. Multifactorial [MR81]. multifida [DBFM80]. Multiphasic [AH81, NNA84]. Multiple [CA85, TL87b]. multiscale [RB84, YY89]. Multispecies [FCW82, YY89]. Multivariate [Bro88, KLE85, OWP80, ID88]. Munida [Zel85]. muscular [Rii89]. Mussel [Gri80b, HL80, LNF81, SNL82, SFN82, Asm87, BHNI89, BS83b, BTFF85, BH88, CR84, Cha88, HSB^{+}84, HB85, HK84, Jac87, JLMR88, KLN88, LN89, LJIO84, LNS^{+}87, MKG88, SDRN84, Spr84c, Spr84d, Spr84a, Spr84b, Tho84]. mussel-raft [LJIO84]. Mussel-Rafts [LNF81]. Mussels [BW80, GW81, HS81, PY81, WS84, BBH^{+}88, CSH83, CH84, Fis86, Fis89, FT86, Lin89, Phi85, PR88, SDMN88, WDS^{+}84, ap 86]. Myctophum [SCF^{+}82]. myopsid [RSM88]. Myrtea [DSS^{+}85]. Mysid [WM82, LM87b, LL83, WW88b]. Mysidacea [O’B88a, WPW87, WPW88]. Mysidopsis [LS84a]. Mysids [Sch81b, Kaa89, LL83, O’B88a]. Mysis [WM82]. Mytiius [RYD84, Spr84a, Spr84b]. Mytilus [AAH87, Auf88, BW80, BT88, BHNI89, BTFF85, BTQS88, CL88, CL84a, CR84, CSH83, CSH89, Dah86, Fis86, Fis88, GT87, GW81, HC80, Ham80, HSB^{+}84, HB85, HL80, Joh88, KE87, Liv88, Low88, LNS^{+}87, MKG88, Oka86, Spr84c, Spr84d, Tho84, Tra88, TC87, TN85, TN86, WDS^{+}84, WJ88, ZJ83, ap 86].

N [BLK88, DL86a, GH83, LNF81, CS84, DL86a, Gar84, KP89, MT87, MC88, Thi87]. N-assimilation [DL86a]. N-mineralization [BLK88]. N-remineralization [GH83]. N. [CSB81, DW83]. N/C [DL86a]. N2 [SG87a]. Nabeta [Hir86, Hir87]. NaCI [RBA86]. NADPH [Liv88, Moo88c]. NADPH-cytochrome [Liv88]. NADPH-ferrihemoprotein [Moo88c]. Nahant [PM88]. Namibia [Mac85b, Mac88]. Namsu [Lat86]. nanoflagellates [CS86, HAA^{+}88, PST87, Wei89b]. Nanomolar [FF86b]. nanoplankton [CC86b, DCJS85, JOP86, KW88, MF88, Ver86a]. nanoplanktonic [WH88]. Naphthalene [LN81, SKS^{+}84]. Narragansett [JT85, KS84, Kel88, SM86, Tra88, Ver86a, Ver86b]. Nassarius [Tal80]. Natal [BS80c, BvdEH^{+}82, BS83b, Sch81a, Sch89a]. Natantia [RL83]. native [CH88]. ’natural [McG88, AP80, AH81, Bro82, Che82, Cho86, DW83, EC87a, Gri80a, GHCH81, Gul87, JK80a, KFPH88, McG86, McG87, OSM82, Pau81, Thi81, WH87, Wol82, ZW89a, CT84, CSH89, GKA88, HMH89, HF88, JP84, Kim85, Kri83a, KBU85, LlJC83, MHL85, McG89a, MF88, Moe85, MBP^{+}86, MR83, NM85, Sch88c, Sch88b, SRTM83, Ste86, Ver86b, WRKD84]. Naturally [WB80, WPW87]. Nature [JMSK86]. Nauplii [SJSO82, DCNS84, Mar84, MK85, PL89, Pau83, LS84a]. Navarino [SO83b, VCS84]. Near [Gun80, HBFA80, BTL86, BCCH84, HC86a, HH89, Sch89a, SW85, SCSD86, WY82, War84]. near-shore [Sch89a, SW85]. Nearshore [HL82, KC82, LRH82, OB88c, TW82, ABT89, BK85b, JL86, MSS86, MP88a, PS88, PM88, RDD88, SB85a, WF83]. nebulosa [GAA89]. need [PBP85]. neglecta [Sie82]. Nekton [Het89, WB83b]. Nematoda [VTHVd81]. Nematode [FT82, Jen81, Jen82, Alo87, EC87b, Lam86, LPV87, RBA83]. nematodes [Jen87, PF83, SLP83, SRTM83, TW83]. Neocalanus [DW83, Fro87]. Neomysis [LL83]. Neopanope [BL88, FWB89, LB86]. Neosarmatium [GLNR86, NGR86]. Nephtys [Beu87, SR86, ZW89a]. Nereis [BMRN84, EZ89, JK80a, JK80b, Kri83a, Kri83b, Kri83c, ZE89]. neritic [Hob88, Uch88, Ued87]. neritina [Keo89]. nerring [Lam87]. Net [BC81, JK80a, DR86, JKMR84, KM87a, LMO86, LT87b]. Netherlands [dB81b, LVdVdV85, LV86]. netplankton [KW88]. nettle [FK84]. networks [VP88]. neustonic [TÓ89]. Neutral [PL85, GL88b, LP88]. Neverita [RNK87]. Newfoundland [SHIP86, FHKR85, KSS85, KSF89, PH83, SGG87, SH85]. newly [Keo89, WB83a, ZWO89, vdVB86, vB87]. newnesi [RTK^{+}89]. NH [LBH89, ZCF89]. Niche [Bro80, Eva82, ET85, FR80, HL89, McC84, Pea85]. niches [Rei87b]. Nicolas [LWE87]. niger [Mag88]. night [AI88, Eva87]. nigrescens [MS89, OS84a, SO84]. nigrofuscus [FMM87, MP88b]. niloticus [BLR86]. Niño [And89a, Gly85, MH89]. Niño-associated [Gly85]. Nitrate [DM82, KZ85, ARGL87, BSC85, Col84, JS88, Le 86]. Nitrification [ES87]. nitrite [VGJ89]. Nitrogen [ABM81, ABB80, Cap82, FB86, Pro85, Ric82, Sei87, SG87a, SS87, Ten82, WC88, Abo84, ARGL87, BW83, BSC85, BR83, BBR88, Col84, DSW89, DL86a, ET84, ES87, FD89, FWE89, Gar84, HK86, HH85, HB85, HlWPH84, HHKA88, Kiø89, KM83, KHHB86, LH87a, LBH89, MM86, Mor87, NL84, PW87, PCH85, SDRN84, SN85, SAS87, SC84b, SLE^{+}89, SP84a, SH89, SPSS89, WWD^{+}88, ZIO84, ZSA87, VB85b, WS83]. Nitrogen-limited [WC88, SP84a]. Nitrogenase [OC89]. NO [ZCF89, FER85, OFGG82, SM87b, JS85, LT87a]. nobilis [CP87]. Noctiluca [Daa87]. Nocturnal [RK82, SB85a, Ste86]. nodosum [JS87a]. noise [RL83]. nomothetic [Ste89]. Non [HO80, JNCS89, Bri85, LDER83, LB86, VK85]. non-burrowing [LB86]. Non-occurrence [JNCS89]. non-pigmented [LDER83]. non-polar [VK85]. non-selective [Bri85]. Non-Tidal [HO80]. noninear [Gal89]. Nonlinear [GTG81]. nonstratified [BSSW89]. Nordasvannet [Was85]. nordmanni [PY86]. normal [Gra83]. normani [Nil87]. Norsminde [JS88]. North [BTL86, Bul82, Cus89, HJ84, Hov89, IBMD81, Lan83, HMM87, PPCR86, SSM^{+}84, BHP^{+}86, BPH87, BAH87, Bd89, BTQS88, BBH^{+}88, CSS82, CSS83, Daa87, DCB^{+}86, DK86, DCJS85, ESR^{+}88, FPRW82, HJ84, Het89, JD85, KKT81, Leb82, LP87, Mee85, Mit87, NR89a, RGLK87, Ric85a, SN89, Shu80, SK85, WR81, WWD^{+}88, Yan82, YLH86, ZW89c]. north-eastern [SSM^{+}84]. North-West [HJ84]. northeast [GL89a, KW88, BK80, Rii88b]. northeastern [Kin89, Wil89]. northern [BW83, Fev89, GOAYH86, GMF88, GFM89, GM89, Kro88, KK89, LBH89, McC84, PE89, SBR84, WH87, WC83, Wro84, GHC83, HM87, Her88, HPF89, IBMD81]. northwest [HL84, Rüg89, Gif88, Sch83b, TR89]. northwestern [MR88a, BCCH84]. norvegica [Nic84]. Norway [AM88, KAA88, LG88, SW86b, Was83, Was85]. Norwegian [BNVP87, BHJ^{+}89, JBAH83, LGS^{+}89, MFS83]. Nostocales [TBRBA85]. Note [JH80b]. Notostomum [SBR84]. Nova [Gif88, HPV^{+}87, HH87, JNCS89, KM88b, EC87a, JS87a, Sam88, STM86, ZW89b]. novaeangliae [BPH87]. novaeanglieae [BHP^{+}86]. novel [BB88, Hic88, PAF86]. November [HC86a]. Np [Dah86, PH81]. Nucella [LH85]. nucleosides [AV87a]. nucleotid [Tam89]. nucleotides [AV87a]. Nuclepore [LD85]. Nucula [BT88, LlJC82, LlJC83]. Nuculanidae [HH84c]. Nuculidae [LlJC82]. nudibranchs [FG83, Yos86a, Yos86b]. number [GG88]. numbers [Sch88c]. Numenius [ZE89]. numerical [AWSB83, ZSA87]. Nursery [KD84, LRH82, Sto88, WAH89b]. nutricional [SGSD86]. Nutrient [AA84, AGCD86, BK85a, GCD87, HL82, HW82, KHE^{+}83, SFHU82, SBK88, TKS^{+}85, BH83a, EMA88, FF89, HPC89, JWC83a, KZ89, KBNO85, LT87a, RDF^{+}88, SG88b, WE88, ZK85, ZM84]. Nutrients [AG82, KNL82b, Ovi81, And86, ALAH85, CHAS83, KE87, LVdVdV85, LP87]. Nutrition [BK81, Ham82b, KN83, BSC85, HVK85, Pol84]. Nutritional [Klu84, KSR80, Sv80b, Thu84a, BW83, CVF89, WH85, LS84a]. nuttallii [GG82]. NW [VP85, FL88, SRS89]. Nyctiphanes [O’B88b].

O [JS85, MC88, ORK^{+}86]. O. [NS87]. Oahu [Hod85]. Objective [GP82, PGJ83]. obligately [IEK86]. Observation [BDKK83]. Observations [CFH89, Cow81, DE89, Ric82, AI88, Gei88, O’B88a, UKTA83]. Observed [MHL85, ZH81, BM86a, FHS88]. observer [Lin88]. obsoleta [Bre82, LSD85]. Occurrence [Arm89, Fen82d, Jør82, NvB89, TW82, BSS89, CTG86, HSRM86, JNCS89, KS84, KMDS85]. Occurring [McD82, WB80, Kan84, VB81, WPW87]. Ocean [Bou81, BC81, Law86, PRS^{+}82, SFHU82, Smi81, Alb83, IM86, KZ85, MW83, VP88, Bla84, BK80, BCCH84, CSW84, Car87, DGHK82, Fev89, Fro87, IEK86, KZ89, KC82, KB80, Le 86, Mee85, SKM85, Sor85, SK85, WR81]. Oceanic [SST89, Bat86, BBSS89, CS84, DS84, DPJ87, HH85, Hei88, Mau88, Par89, RA87, TB89]. oceanica [CD88, LB87, SDY87]. oceanographic [LGS^{+}89]. oceanography [SR85]. oceans [PHL^{+}89, RB84]. octachlorostyrene [SKS^{+}84]. Octocoral [LC83, Gri88]. Octocorallia [BL87b, CLS^{+}82, LCS86]. octocorals [LC88]. octopus [Amb86a, Amb82, GHCH81, GF82, MG81]. Ocypoda [NS87]. ocypodid [NS87]. oersstedii [BL80]. oersted [Mor83]. Off [DM82, LH82, BCA89, BS80c, BSF84, EC87a, EMG89, GB84, HMM87, JJN85, KFPH88, Mac85b, Mac88, MP88a, Mor85b, O’B88b, Pet88, Sam88, SW86a, SSM^{+}84, SH85, WPS89]. offshore [BBH^{+}88, KWG85, Pag86, RLB89, SV85a]. Offspring [OTC82, KGV84]. Oikopleura [DD85, DT85, DP87b, KSF89]. Oil [FPRW82, Jac80, LL83, LR80, Nel81b, OFGG82, AE88, Bod88, CL84a, DLT^{+}83, FC83, FER85, Pag86, RG85, SM87b, TAK88]. Oiled [FKW81]. Oithona [Uch88]. old [Pet86a]. Oligochaete [Gie81, GLD82]. oligotrich [FJ88a]. Oligotrichida [Gif85]. Oligotrichina [FJ88a, Jon86]. oligotrichous [Jon86]. oligotrichs [Gif85]. oligotrophic [Bie85, HAA^{+}88, IEK86, NNA84]. omnivorous [DGK88]. Omnivory [CT85, Ohm84]. omorii [Uch88]. Oncholaimidae [LPV87]. One [Gun80, Thr82]. onset [Jay89]. onto [WWB86]. ontogenetic [Ued87]. Ontogeny [LD89, LN81, SH84]. Oosterschelde [LVdVdV85, LV86]. opacum [KD86]. Open [KAE81, Fro87, IM86, KEK^{+}84, LNS^{+}87]. Operated [BLP87]. Operation [Aik81]. operational [CP84]. operative [BB84]. Opheliidae [Tam85]. Ophiophragmus [TM80]. Ophiuroidea [TM80]. opisthobranch [FG83]. Opportunistic [LM87b, WZ85]. Optimal [FH89, AAH87, LH85]. Optimization [GL88b, BA83, CA86, HSB^{+}84]. Orbiniidae [RBR86]. order [PF86]. Oreaster [Sch80a]. Oregon [GB84, Pet88]. Organ [FV82]. Organic [Cam82, CSS82, EHF^{+}81, Ise81, LlJC82, LC82, LNV81, MRF80, NLVS80, Rei81b, Ric82, RMZ82, RD89, Ten82, Wol82, de 80, ABT89, BAC^{+}88a, BW88, BAN89, Cam85, CSS83, FST85a, Gar84, GT87, HAA^{+}88, HP89, HHN^{+}83, HF88, KN84, KJ88, KWW^{+}88, LTH87, MT87, MC87, MR86, Mø83, NJ83, OOF83, OWKG85, PT89, Pet89, PBP85, PCH85, RWK83, SW85, SW86b, Tam89, TIS85, TP85, Thi87, Was83, ZW89b, ZW89c, HAHA84]. organically [BV83, LPV87, Tsu87]. organics [LS88]. Organism [RG82]. Organism-Sediment [RG82]. organismic [Gol85]. Organisms [MRF80, MST81, NLL81, SNL82, DRCA83, EMHS86, Hir86, Hir87, LML83, McG88, McG89a, MR88b, SHPH86]. Organization [Liv82, MKHD85, SMO81, JS87b, TN86]. organizing [SP84b]. Organochlorine [Har86]. Organochlorines [Phi85]. organotrophic [LDH83]. organs [AS83]. oriented [DD84]. Origin [VS80, VCS81, Ebe86, LS84a, NvB89]. Origins [EHF^{+}81, SW85]. ornatus [TP89]. Oscillation [And89a]. Oscillations [CT84]. Oscillatoria [BSS89, Scr83]. Oslofjord [BV83]. Osmeridae [Ste89]. Osmerus [LD89]. osmolality [NR89c]. osmotic [AA89]. Osteichthyes [Ben84]. ostralegus [SHG89]. Ostrea [BNF81, NH82, Wil83, WAN84]. Otaria [GNBO85]. Other [TP81, And89a, Beu87, Che84, FG83, GGA85, LGS^{+}89, MAWM83, OWKG85, RDD88, SR86]. otolith [Hov89, KBv89, TRMF89]. otoliths [Kv88]. Otter [BCFS82]. outdoor [RO84]. Outer [LWG^{+}81, SB82, WH82, SGG87]. outershelf [Rus84b]. outfall [SCSD86]. outflow [MF88]. Outline [Thi87]. output [MT85b]. outwelling [DCB^{+}86]. Ovalipes [WH87]. ovarian [NS87]. overall [Her88]. Overgrowth [Lan89, Rub85, WW86]. overlying [Sch89a]. overview [BAC^{+}88e, CFP88]. Overwintering [Hir83, TRMF89]. Oviposition [NS80]. oxic [GG88]. oxic/anoxic [GG88]. Oxidants [RRB^{+}80]. Oxidase [SC82]. oxidation [BR85, JMG84, VGJ89]. oxidising [DSS^{+}85]. oxidizer [JP85, JM85a, JM85b]. Oxygen [AK88, BB80, BNF81, Dal85, HW82, KM83, MW87, SK82, AA89, AT86, AH87, ABAI87, Bau85, BK85a, Cam85, Fis86, JMSK86, KCS89, Kri83b, Kri83c, LV86, MFP87, NVHHH88, Thu84b, TW88]. Oxygen-Independence [BB80]. oxygenase [AE88]. Oxyrrhis [FF89]. Oyster [BNF81, DP81, NH82, RRB^{+}80, SK82, BBR88, Bro88, DSW89, LN89, NJ83, NL86, OWZ89, WAN84, WH85]. oystercatcher [SHG89, WLB84]. oystercatchers [HD89, LWE87]. Oysters [BP81, PY81, DCMS89, PWM85].

P [HMF85, Liv88, MT87, SWG88, Thi87, VA85b]. P-450 [Liv88, SWG88]. P-I [HMF85]. P. [DAC^{+}83, WB80]. pachyptila [CFH89, GH86]. Pacific [BK80, IEK86, Bai81, BHP^{+}86, BPH87, Bat86, Boa80, Bro88, ESR^{+}88, Fro87, GR89, KB80, KW88, LL81, McG89b, Rob82, Sch88a, Shu80, Sor85]. pacifica [Bat86, Ohm84, Yos86a, Yos86b]. pacificus [HSRM86]. pack [HC86a, KSS85]. pack-ice [HC86a, KSS85]. pagurus [LH85, Law89, MGT89, Thr86]. PAH [SDMN88]. PAH-metabolizing [SDMN88]. Palaemon [Ber84, Dal85, HO80, MT84, NR89b, NR89c, TS88a, WR82, WR84a, WR84b, WR86]. Palaemonetes [NR89b]. Palaemonidae [TS88a]. palinurid [Mac85a]. Palinuridae [Jay89]. pallasi [McG89b]. pallida [Cla85, LNB81, LNV81, NLVS80, SLN81]. palmata [ID88]. Panama [DE87, McG89a]. Pandalid [BLP87]. panicea [Bar86, Bar88b, CM87, HSdWM88]. Panopeus [McD82]. Panulirus [BS80c, HB86, Jay89, TP89]. Papua [MD86]. Paracalanus [HSRM86, Tur84]. paradox [BT85, SWH84]. paradoxus [GLB87]. Paralichthys [JH80c]. Paralithodes [SS89b]. Parallel [Rus83b, EB86]. parameter [DJ84, Pat88]. parameters [FDM86, FEHA85, HMF85, HPC89, HM87, LVdVdV85, LM87a, LGS^{+}89, OWKG85, Rei88, RGH89, SC84a, SPH87, VB85a]. Paramoeba [JNCS89]. Paramphithoe [OS85]. pararmatus [RBA83]. Parasitic [Shi80, Ans89]. parasitism [SBR84]. Parathemisto [WR81]. parelegantul [VTHVd81]. parental [BRS86]. Parhyale [AS87]. Parika [KM87c]. Parophrys [GB84]. part [WZ83a]. partheneia [Sch83b]. parthenogenetic [Ama83]. Partial [HS87, TS88a]. partial-emersion [TS88a]. partially [MKD^{+}86, MCPW88]. Participants [BAC^{+}88d]. Participation [ALP84]. Particle [Cam82, Car87, Jon86, KM81, LNS^{+}87, WY82, ASSF87, Bro86, DP87b, JF88, KJ88, PT84, Rii88b, RD89, SHCL87]. particle-attached [ASSF87]. Particles [Moo81, AN88, KM88b, MF88, NBR^{+}88, NBMM88, PV85, PCM86, TKP86]. particularly [Fro87]. Particulate [CSS82, Ise81, Las81, OS80, PW87, Rei81b, SN81, SLN81, BSSW89, EB86, Gar84, LTH87, LNS^{+}87, MCA89, PT89, Par87b, SP86b, SW86b, Was83, Was85]. Particulates [RMZ82]. Partitioning [Joh82, Mac81, SHPH86, Ben84, BB86, HD89, HLH87, HGD86, HHN^{+}83, Mac86a, SF89, SK84]. partly [HSB88]. parvus [HSRM86]. Passage [HHE89, SACC88, BBL84, KN83, PD88]. passive [Pal88]. Patagonian [Pat88]. Patch [Sch80a, SS86a, SS89a, SH84]. patches [TB86]. Patchiness [ACF^{+}80, Kah82, SGG87, AH84b, ABAI87, GLT^{+}86, McG86, McG87, SO85]. Patella [BS80a, NLBFD80]. Patelloida [Fle87, Jer85, MUC84]. Pathogen [SA81, Mil85]. pathology [SM84]. pathways [Rei87a, WF83]. Patinopecten [ELK84]. patronus [GHC83]. Pattern [GLD82, BY83, PB85b, VK85, YY89]. Patterns [DCMS89, DG82b, FCW82, FR80, HC86a, KC81, McD82, NK81, OKSB86, Rid88, AW86, BC88, BGR86, CC86a, CC88a, DB88, FAG89, GA83, HMF85, HM86, HH87, HGD86, Hoe88, JG89, JWC83a, Jos87, Kaa86, KZ85, KZ89, KS84, McC89c, McG88, MKHD85, PL85, RB84, Rus83b, Ryl83, SS89a, TO88, TB88, VCS84, WB86, WW86, Yos86b, dB87a, Rus84b]. patters [Kin89]. pauciradiatus [Sog84]. payability [NGR86]. Pb [FBCH83]. PCBs [ITBW81]. peaks [Wal85a]. pecdation [HD89]. Pecten [KAE87]. Pectinaria [DS86b]. Pectinariidae [DS86b]. Pectinidae [Fev89, SC84a]. Pedal [Cla85]. Pedro [KHW88]. Pedum [DE88]. Pelagic [AG81, Abr83, FC82, GMF88, GFM89, GM89, Kle80, OF82, AS86, AN88, ALH86, BP84b, ER87, GB84, HAA^{+}88, HHD^{+}84, Her88, HSB88, KBNO85, KB88, KK89, MN87, Mau88, McG86, Sha83, SF85, SRS89, SG84, WCGG86, Wei89b, Wil85]. Pelagic-benthic [GMF88, GFM89, GM89, ER87]. pellet [ER87]. Pellets [Ise81, BNVP87, DT85, GABM84, Mor87, NvB89, Tur84]. Penaeus [CS81, DAC^{+}83, Emm80, MF89a, SV85a, SV85b, Sto88]. penetration [BBSS89, KKKW86]. Penguin [Wil85, AB89]. penicillatus [KO87]. Penicillus [Sto85]. Peninsula [Hir86, Hir87, KS87b]. Pennellidae [Ans89]. peptization [PT84]. Perception [Sto80b]. perennial [Jer86]. Pérès [Gol85]. Performance [Joh81b, DR86, HMH89, KM85, Rei87b]. period [DA85b, KS84]. Periodic [MF89a]. Periodicities [HH84a, Hei88, TB86]. periodicity [DL86a, HMT83, HHA85, KWG85, Mac85b, Rum89, WK88]. Periods [DB82]. perlae [TL85a]. perleve [CM87]. permanent [BNVP87]. Perna [BS83b, Cha88, Phi85, Sha89]. Persistence [NT85]. perspective [Mar89b]. Perspectives [Ric82]. perturbation [PKMS85]. Peru [DM82]. Peruvian [DE87, Wol88]. peruvianum [PB89]. petroleum [MBP^{+}86]. pH [MP88b, dHS88]. Phacosoma [TO88]. Phaeocystis [LM87a, AV87a, Arm89, BF87, LMO86, SPK^{+}87, VA85a]. Phaeocystis-dominated [LM87a, BF87]. Phaeodactylum [PW87, SGB89]. Phaeophycae [dB87b]. Phaeophyceae [Ger89, RHvd81, dB87a, dKB88a, dKB88b, tBB83]. Phaeophyta [CB88, De 84, OS84a, SO84]. Phaeophyta. [RW86]. phaeopigment [LDKR88]. phagotrophic [PB87]. phagotrophy [CS86]. Phallodrilus [Gie81, GLD82]. phanerogam [LB87]. pharyngeal [DP87b]. phase [CS86, HH84a, KM87c]. phase-lagged [HH84a]. phases [HS85, MN87]. phenan [SKS^{+}84]. phenan-threne [SKS^{+}84]. phenolic [WBVS86]. phenomena [BLP87]. Pheophyta [Law86]. Pheromone [DBFM80]. philippinarum [RNK87]. Philippines [Ang86]. Phoca [Boa80]. phosphatase [AV87a]. Phosphate [GCK82, Tam89]. Phosphoglucose [BS80a]. Phosphorus [BW83, Har82, Sor85, DSW89, HH85, SGB89, Tam89, AGCD86]. photic [HH84a]. photo [NVHHH88]. photo-synthetic [NVHHH88]. photo-system [NVHHH88]. Photoadaptation [HMT83, MLDT88, GP86, LCP84]. photoadaptational [SDY87]. Photoadaptive [BGL^{+}88]. photoautotrophic [SPL^{+}84]. Photographic [AI88, MW84]. photography [KM88b]. photoinhibition [BA87, GD84a]. Photometry [JBS82]. photomhibition [NPMW87]. Photoperiod [DD80, Rie82b]. Photoperiodic [Lün88, tBB83]. Photosensitivity [BBSS89]. Photosynthate [HGD86]. Photosynthates [KSR80]. Photosynthesis [BS89b, DB82, HMF85, HlWPH84, RR82, BA87, BSP89, CP84, CMC89, DDR89, GPR86, HMT83, HH84a, LCF89, LS83, MHL85, MLDT88, PDP87, PGV^{+}89, RC86, SP86b, SPK^{+}87, Riv89]. Photosynthesis-irradiance [BS89b]. ’Photosynthetic [FV82, BA87, BAG80b, DL82, GKA88, GD84b, JH80a, JH80b, PPCR86, SJ88b, BP84a, BP85, BP86, DTLN85, Dou83, ET84, FDM86, HMH89, JOP86, KWG85, LDG87, LDDH88, LSP84, SP85, VA85a]. Photosynthetically [PRS^{+}82]. Photosynthetically-Competent [PRS^{+}82]. phototrophic [LDDH88, Ver86a]. phototrophs [CMC89]. phoxo [OOF83]. phoxo-cephalid [OOF83]. Phoxocephalid [OOS82]. phycoerythrin [BA87]. phycoerythrin-containing [BA87]. Phymatolithon [CSB81]. Physalia [Pur84]. physalis [Pur84]. Physical [Arm88b, Coh85, CH89, DK85, GLT^{+}86, Ken87a, Ken87b, Lon81b, Mit87, Möl86, Thi81, BB85, DM83, DW87, HKS86, HHB88, LT84, SPH87, dKB88a]. physico [HG86]. physico-chemical [HG86]. physiologic [DG88a]. Physiological [Bak88, BT88, Cap88a, Cap88b, DB82, Joh81a, Joh81b, Joh82, PDWF85, SED81, Spr84c, Spr84d, Spr84a, Spr84b, WJ88, WW83, WWL83, HR89, TD89, TP89, Vid87, WDS^{+}84]. phytal [Gib88]. Phyto [Jen81, LH82, FST85a, FST85b, GH83, LVdVdV85]. phyto- [FST85a, FST85b, GH83, LVdVdV85]. Phyto-Chemical [Jen81]. Phyto-Zooplankton [LH82]. phytodetritus [GL89a, Rud88]. phytopiankters [FBT84]. phytoplankter [Cra89]. Phytoplankton [AC88, ABB80, Bat86, BT85, Car82, Clo82, DL82, EMA88, GTG81, IBMD81, JBAH83, JH80a, JH80b, KAE81, Kah82, Kiø89, KS87a, LT87a, Mø83, Nel89, NLL81, Ovi81, PS86, PWMl89, PRS^{+}82, Red83, RS80, THSW80, VDM89, WBJA81, Wol82, ASMV88, BM86a, Bar84, BP84a, Bau85, BP88, Bid88, BH83a, BNF^{+}89, BA83, BS89b, CC88b, CC88c, CA86, CA87, CC84, CC87a, CS84, CP84, DTLN85, DL86a, EAM87, FDM86, Fro87, FEHA85, FHS88, GP85, Gal89, Gei88, GD84b, GGV89, HJ84, HMF85, HYT^{+}88, HGD86, Hob88, Hsi88, JH87, JS86, JP88, KWG85, KLE85, KS84, Kel88, KHHB86, Lan83, Le 86, LT84, LCP84, LSP84, LNL83, LP87, LT87b, MKD^{+}86, MCPW88, MHL85, MJPH^{+}88, NL84, NVLL88, PT89, PLM89, Par87b, PHT85, PCH85, PHL^{+}86, Pro85]. phytoplankton [RTLR88, RNH^{+}88, Riv89, Roy89, SC85, SN85, SN88, SP86b, SCH89b, SSM^{+}84, SP85, Spi87, TTH83, Van86a, VDL88, Vid87, WPW87, WJ87, ZCF89, GCD87]. Phytoplankton-bacteria [BT85]. phytoplankton-derived [BP88, Bid88]. Phytoplankton-microflagellate [GCD87]. Phytoplanktonic [LM87a]. phytoplanktor [WRKD84]. Pico [GPCW88]. Pico- [GPCW88]. Picocyanobacteria [GL89b]. picoplankton [DCJS85, Dou83, JP84, JOP86, LDDH88, RHS89, SPL^{+}84]. picoplanktonic [PDTH86]. picta [Sha89]. Pigment [JH80b, Dec88, KS87a, RGLK87, RLB89, WJ87]. pigmentation [PGV^{+}89, TC87]. pigmented [LDER83]. Pigments [JH80a, JH80b, HJ84, KM85, Nel89, PD88, PPCR86, Roy89, RHP89]. Pilayella [PM88, RW86]. pileatus [Tur84]. pileus [Sch87b]. Pinna [CP87]. pinniger [BK80]. Pinniped [WLB84]. Pinnotheridae [BS83a]. Pipefish [Rye88, BRS86, SWB89]. Pirenella [TP81]. Pisces [CL84b, FAG89, KC81, KM87c, McC89b, NC89, OTK80, SH84, SF89, Ste89, WB80]. pistillata [HGS89]. piston [Rii89]. pit [Thr86]. pit-digging [Thr86]. Pitfalls [Blo81]. placement [SS86a]. Placopecten [BGSS88, LRS87, MT85a, MT85b, Mac86a, RLB89]. Plaice [SC82, Cus89, Hov89, KBv89, vdV85, vdVB86, van86b, vB87]. plana [MGC82, SSN87, WWB86, WW83, WWL83]. planci [DRB86, DE88, Gly85, Ols85, Wil86a]. Planktivorous [Thr82, KM88a, Kin89]. Plankton [BDKK83, HAOC^{+}86, ITBW81, KEK^{+}84, KMR^{+}88, Kui81, Ric85a, FAG89, HlWPH84, HSW89, KHE^{+}83, KBv^{+}83a, KBv^{+}83b, LHF84, Mar84, NL84, NR89a, Pur88, SSPA89, SLE^{+}89]. Planktonic [ABB80, LOvM89, WC83, BSF84, DLT^{+}83, DRCA83, Gif85, Grö88, Hod85, Jon86, Jon89, KMH85, LDH83, LGS^{+}89, LvB87, PV85, PDWF85, SS86b, SA89, Sor85, STM88]. Planktotrophic [Sv80b, Mar89a]. plant [KZ89, KM89a, NL86, RO87b, SN85, STM86, TKS^{+}85, WRC83, tBB83]. plant-dominated [KM89a]. plantation [SB85b]. Plants [Ten82, Bar89, Gle86, TKS^{+}85]. planula [Grö89, Ric85b]. plastic [AV87b, Kri83a]. plate [HF87]. Plateau [CSB81]. platensis [TBRBA85]. plates [Hir86, Hir87]. platessa [Hov89, KBv89, vdV85, vdVB86, van86b, vB87]. platessoides [SC82]. Platform [FPRW82, Sv80a, Jer85, KM89a, Pag86]. Plathelminthes [NW84, Rei87b]. Platichthys [AE88, OA88, SM84, SWG88, VL88, vdV85]. platytarsis [NS87]. Plerogyra [FV82]. Pleurobrachia [Sch87b].

Pleuromamma [BBSS89]. Pleuronectes [Hov89, KBv89, vdV85, vdVB86, van86b, vB87, CLB88]. Pleurophycus [Ger89]. Plexaura [Ger84, PF89]. plexaurid [Las84]. plotting [Gra83]. Plovers [Sch81b]. plumchrus [DW83]. plume [KSVB89, MF88, SN89]. plumes [Alb83]. plunderfish [Dan83]. plunge [HS88b]. Plutonium [GF82, MGC82, BTFF85, MG85b]. Pmax [NM85]. Pocillopora [Col80, SB85b]. Podoclavella [Dav89]. Podon [JKJ88]. pods [BBSS89]. Pogonophora [SC86]. Polar [Cla82, VK85]. polaris [HGH87]. Pollicipes [Pag86]. pollock [Bai89, DCNS84, Pau83, Wal85b]. pollutant [Add88b, Add88a, Bak88, CL88, Cap88a, Cap88b, Moo88a, War88b, War88c]. pollutants [BCG88, HF88, Liv88]. polluted [AE88, Bod88, Joh88, LPV87, OPN^{+}84, Tsu87]. Pollution [GP82, Gri84, LR80, AV87b, FS84, Gra83, LL83, PB85a, PGJ83, Sch88c, Sch88b, SLP83, SDMN88, UP88, WCC^{+}88]. Pollution-Induced [GP82, PGJ83, SLP83]. Polonium [CHH83]. Polonium-210 [CHH83]. polutant [Moo88b]. polycerus [Mar89a]. Polychaeta [BK81, DS86b, Gas87, GMT89, JK80a, JK80b, Kri83b, Kri83c, PH83, RBR86, SWSL81, Tam85, TW80, Tsu87]. Polychaete [FR80, HT81, Rei81a, Alo85, Beu87, Cam85, CT84, GAA89, Kri83a, Mar89a, Oli84, Pet84, PT84, Rii89, Tam87, VAF83, ZW89a]. polychaetes [Beu87, DB88, PF83, SR86]. Polychaetous [ML80]. Polychlorinated [ITBW81]. polylepis [BHJ^{+}89]. polymers [HL85]. polymerus [Pag86]. Polynesia [LDDH88]. Polynesian [Gal87a, Gal87b]. Polyp [Sam82]. polyps [Grö88]. Pomatomus [OSB85]. Pomatoschistus [Mag88, WB80]. pompejana [GH86]. Ponds [MCGP82]. Pontes [DRB86, DE88]. Pontonema [LPV87]. pools [HG86]. Population [AS86, BS83a, CS89a, CF84, DM82, GGA85, GR89, Him86, HLH87, HK84, ITBW81, Leb82, NGV^{+}82, Nic84, Rum89, SO83b, She80, Tal80, TD81a, Tsu87, AM88, BHP^{+}86, BGSS88, BB88, BSS89, CA83, CSH83, Dah86, Dav89, GT87, Gar89, GMT89, Gri88, HH89, HGS89, HY88, HH84c, IEK86, JP84, KH88a, KM87a, LDH83, Lam87, Mac88, McC89c, MAWM83, MD87, MWM86, MR83, Pag86, PH87, SC84a, Sma83, Smi86, SM87b, Tam85, Tho84, Wei89a, Zel85, dB87b, vdVB86, van86b, vB87]. Populations [AP80, BS80a, BW80, Bro80, Bro82, FC82, Gun80, JH80a, She82, Wol82, de 80, AAK^{+}88, ASMV88, BTQS88, CT84, Daa87, Dav89, DO86, EC87b, FF89, GL89a, HMH89, HlWPH84, HF88, Ken89, KC89, Kim85, LNS^{+}87, MF88, Moe85, Ola89, PE89, PGV^{+}89, Pro85, RLB89, RA87, SO87, TB85, Tow84, TW88, Ver86b, WDS^{+}84, WW83, WWL83, dB87a, dKB88a]. pore [ZM84]. Porites [DE89, PDIF85]. Port [KHW88]. Portuguese [Pur84]. Portunidae [Cho86, Pau81, WH87]. POSER [BDKK83, KHE^{+}83, KBv^{+}83a, KBv^{+}83b]. Posidonia [CD88, LB87, You81]. Position [Spa80, Kee87]. position-dependent [Kee87]. Possible [NL81, OR80, PD88, Sma83, CG87, Dun86, Ger86, HAA^{+}88, LSD85, Mil85]. post [Gle86, Gly85, KD86, Luc84, Ov87, OWZ89, SVK83, TL87b, TL87a, TC87, TÓ89]. post-emergent [TL87b, TL87a]. post-larval [Gle86, SVK83, TC87, TÓ89]. post-settlement [KD86, Luc84, Ov87, OWZ89]. postlarvae [LOvM89]. postlarval [BL88, Bd89, HB86, LB86, SV85b, van86b]. potatorum [CH88]. Potential [BCF82, Blo81, Kem87, KB84, HMD88, Hig85, KMM88, LDKR88, SWH84, STM86, TB85, WC88]. potentials [HMLF88]. pouchetii [AV87a, LMO86, SPK^{+}87, VA85a]. poutassou [BCA89]. power [WRC83]. Practical [FCW82, BCG88, LP88]. practice [MGM86]. prasina [SCT87]. prasinomonad [PDTH86]. Prawn [CS81, Emm80, Ber84, MT84, TS88a]. Prawns [DB81a, SV85a, SV85b]. Precise [JBS82]. Predation [BL88, Bre82, Eva82, FG88, Lon81b, Nel81a, Pea82, Pur84, Sch81b, WAH89b, vB87, Amb86a, CYC83, Daa87, DS84, DRB86, DE88, FUM84, FF89, Fui89, Gib88, HS87, Hop88, Kim84, KM89b, LB86, LWE87, Mag88, ME89, Par84, Pie87, Rus87, SP84b, TKT86, vdV85]. Predator [Pea80, WW88b, DRB86, ET85, HL84, NPVW86, OS85, TL87a, TE86a, WH88, Wil85]. predator/prey [HL84]. Predators [OOS82, CYC83, Gib88, HHE89, Lar86, Pie83]. Predatory [Law89, Pau81, SR86, Amb84, Amb86b, Beu87, CH84, Fai88, HGH87, Mor85a, Wil86b]. predicted [MHL85]. Predicting [GC83, Wei89a, DL86a]. Predictions [MR83]. predictive [KSF89]. preference [AAN87, LWE87, RNK87]. Preferences [Rie82a, VB81, Gle86, Las85, TW83, WP88]. Preferential [NJ83]. Preliminary [FUM84, KAK88, LC82, Sch80b]. Prenatal [PY86]. Presence [Kle80, NL81, PH81, CTG86, FK84, Jen85]. present [CSH83]. preserved [Pur88]. Presettlement [KM87c]. pressure [CC86b, DRB86, FWB89, Las85, MW83, ME89, RA89]. Pretreatment [HL80]. Prey [Bar88a, CL84b, CL82, DD80, HL84, Las85, LWE87, Pea80, AP83, BBG^{+}87, BB84, BK85b, Col87, Eva87, Fai88, FL86b, HHE89, HSRM86, Kim84, Law89, NC89, Ohm84, Pau83, RNK87, Rye88, SH84, SRL^{+}86, WH88, WW88b, ZE89]. preyed [WWB86]. preying [RNK87, ZE89]. Priapulida [OS88]. Primary [Cd83, LT87b, LTH87, PSB^{+}81, VP85, Vid87, BSF84, Col86, DK86, GD84a, GKM^{+}87, Hsi88, IM88, KS87b, LMO86, WC88]. primolecta [FF89]. PRIZE [Gra86b, Gra86c]. prizes [Kin86]. probability [VDM89]. probably [NvB89]. Problem [Ric82, JH87, SF86]. procaryotes [MB87]. procedure [CA87]. Processes [EWM87, GCBM82, Bar88b, BGR86, Coh85, FHS88, Her88, KEK^{+}84, KZ85, KZ89, LH88, Mar84, TE86b, dKB88a]. Processing [CSDZ86, AH84b, BLR86, DSW89, DS86b]. Prochaetoderma [Sch87a]. Proctoeces [Shi80]. Produced [KSR80, RRB^{+}80, LTS86]. Production [BW80, BS80c, BK82, Car82, DCNS84, Eva82, FP82, HHG84, KCFS89, KM80, LNF81, Mac86a, PGBN81, Pie87, RDF^{+}88, Sie82, Tho84, VN82, dB81b, AM88, Alb83, AS87, AN88, Asm87, Bar88b, CA83, CR84, CFP88, Cd83, Col85, Col87, Col86, CH89, DK86, DM83, DL83, DNF87, Dua89, Eme89, Eml86, ES87, FNH83, GT87, GD84a, GMB87, GMT89, GKM^{+}87, Hsi88, HMLF88, IM88, JOP86, Jos87, Kem87, KMH85, KMAM84, KNH86, KH88b, KSVB89, KS87b, KGS87, LMO86, Lar86, LS88, LT87b, LTH87, MT85b, MKD^{+}86, MCPW88, MMS89, MP88a, MW87, NVLL88, PLM89, PS86, Pet86a, Pet88, PHL^{+}89, SHPH86, SO85, Sha88a, SSM^{+}84, SSPA89, SACC88, SS87, STW89, SM86, TGA89, TR87, TKS^{+}85, VB85a, VB85b, VP85, Vid87, WYG85]. production-related [Jos87]. production/biomass [Dua89]. Productivity [BCCH84, LDH83, Li82, MP81, MCGP82, PSB^{+}81, Alo85, ABT89, BSF84, CC84, CC87a, Cro84, GM87, GP85, Hol88, JP88, Kel88, KMD87, LV86, OKSB86, PGV^{+}89, Sha88b, TL85b, Ver89, WC88]. products [VA85a]. productus [Bai81]. Profile [RG82, Gra86a, RTK^{+}89]. Profiles [SN81, CS84]. program [GL88b]. programme [BGR86]. Progress [NH82]. prop [TCH87]. propagation [BH87]. Propallene [NS80]. Properties [Rii89, Spa80, JLMR88, Par84, Rii88a, Sch84b]. Proportional [KR82]. Proposal [ABB80]. Prorocentrum [HMT83, HMH89, KMDS85, RWK83]. Prosobranchia [FG87, FL86a, Sie82, Tal80, TE86b]. Prosobranchiata [BS80a]. Prostaglandin [Ger84]. Prostaglandins [RT85, PF89]. protect [DRB86, LB86]. protected [McC89a]. Protein [SA89, KNH86, LMO86, LT87b, LTH87, MSS86]. proteins [JCL86]. Proteolytic [RA87, HSB88, VBFS85]. protists [GFS89]. Protogonyaulax [GFG89]. Protogynous [Cha82]. Prototrophic [DCJS85]. Protozoa [SSA^{+}86, ASSF87]. protozoan [NvB89, SSPA89, SE85]. protozoans [LDMT86]. Province [Din82a]. provincialis [BTQS88]. Proximate [Law86]. Prydz [HH89]. psammodytes [WPW88]. Pseudalibrotus [BC88]. pseudo [TÓ89, CLB88]. pseudo-neustonic [TÓ89]. Pseudo-pleuronectes [CLB88]. Pseudobodo [PDTH86]. pseudonana [BP84a, SDY87, SP84a]. Pseudopleuronectes [Buc82, FKW81, FHKR85]. psychrophilic [Rüg89]. psychrotrophic [Rüg89]. Pterygophora [BGD87, De 84]. Ptychodera [DG88b]. Pu [CSH89, Dah86, HC80, MG85b]. puber [Cho86, aH85, ap 86]. Puerto [Sto88, SO83b, VCS84]. pufferfish [GR89]. pugilator [Chr89, CF84, WP87]. pulchra [Has81]. pullnet [HKB87]. pulses [SSM^{+}84]. pulsing [GD84b]. pump [DR86, JFK^{+}86, JLMR88, Rii88a, Rii89]. pumping [Wil83]. punctatum [SCF^{+}82]. purpuratus [TD81a, Wol88]. purpureum [BH87]. purse [WEHF85]. purse-seine [WEHF85]. pushnet [HKB87]. pusilla [PDTH86]. pusillum [Jer86]. pycnocline [HM87]. Pycnogonid [NS80]. pyrifera [JJN85, MJ84, OS84b, SO83a, SO83b, VCS84, ZK85]. Pyura [Klu84, SN88].

quahog [MW83]. Qualitative [BGR86]. Quality [CDL87, HR87, Klu84, NBR^{+}88, NBMM88, PV85, Riv89]. Quantification [Pur88, CC88c, CS86]. quantify [HA87]. Quantifying [WW86, CP84, MW84]. Quantitative [AH86, Bou81, Fen82d, LNL83, NLVS80, NW84, Pet86a, WW88a, Fis88, RHP89]. quantity [Klu84, NBR^{+}88, NBMM88, PV85]. quantum [BSP89, SPK^{+}87]. quasimodo [Tur84]. queen [SLMB88, Sto89]. Queensland [JG89, Bul82]. Quempillen [WAN84]. Question [AG82]. quinquecirrha [FK84]. quinquedens [Sv80b]. quinquiesperforata [BS83a]. Quoddy [GT87]. quota [BH83a].

R. [JNCS90]. racemosum [Bar88a]. Radiation [SA81]. Radioactive [GHCH81]. Radiocarbon [Li82]. radioisotope [Mon83]. Radiolabelling [LC82]. Radionuclides [HC80, CSH83]. raft [LJIO84]. rafting [Hic88, Hig85]. Rafts [LNF81]. Random [SS86a, Sch82a]. range [PW87]. Ranina [Ken89, KC89]. Rapid [BL80, Chr89, ZCF89, FF86b, PH87, Sch87a]. rapidly [dHS88]. rare [PH87]. Raritan [MHW88]. Rate [AP80, FITF80, Joh81a, NLL81, NH82, VCS81, BH83a, Bue87, CC86a, DO86, DRCA83, HSdWM88, Kiø89, KN83, LT87b, LlJC83, MT84, Pag86, PWMl89, Reg84, Rie89, Rii88b, SM87a, STW89, VA85a, Wil83]. Rates [BH81a, Ban82, BS80c, HL82, Pet88, ZH81, AC88, AH87, Bie85, BP86, CC88a, CC88b, CC88c, Cro84, DW83, Dah86, DeW87, Dec86, Dec88, DL86a, Dou83, DNF87, DJ84, FL86a, Gei88, GCSP88, Gle86, Har85, Hei88, IM88, JF88, JS86, Jon86, KM87a, KNH86, KH88b, KSF89, KA84, LV86, Lar86, MT85a, Mon84, MR83, NBT88, PB87, PW87, PGV^{+}89, Red83, RNJ^{+}84, SO85, SHCL87, SA89, Spi87, Tan88, Ver86b, WAH89b]. Ratio [BAG80b, Pea85, BTH^{+}84, FF89, HK86, KS83, LT87a, MC88]. Ratio-based [Pea85]. Ration [RTW81, Mac85b, SK85]. rationale [BCG88]. Ratios [de 80, DL86a, DRCA83, HL84, Hov89, SKIR89, SGSD86, STM86, VA85b, WH85]. re [PF89]. re-evaluation [PF89]. Reactive [Har82]. really [HS88b]. Reared [DB81a, JH80c, ALPS83, Ang84, CLB88, Chr89, GMS85, KA84, Nil87]. Rearing [SS82]. reassessment [Har86]. Rebuttal [JNCS90]. recently [FL86a]. reciprocal [PST87, RYD84]. recognition [SCH89b, VK85]. Recolonization [DL83, Gra81, Pea81b, ZW82a]. Record [DB81a, PH83]. recorded [Fis89]. recoveries [Bal84]. Recovery [OFGG82, OPN^{+}84, Pea81b, HY88, MAWM83, TB85, Wil88]. Recruitment [Bai81, CS80, Gri88, Gun80, HY88, Kee87, Mar84, SO84, SS80a, SLO82, TD81a, TRMF89, WWB86, Wil82, BL87b, Chr89, CJS87, DJTL88, ELR88, HF87, KB87, Luc84, Ola89, OWZ89, Pal88, PH87, Qui88, SO87, SV85a, Sto86, SB87, Wal85a, Wol88]. Recurrence [SMN86]. recurrent [LML83]. Recycling [Har82]. Red [BL87b, DM82, FITF80, SF89, SLO82, Sv80b, BH87, Han89, Jer86, NVHHH88, PB89, SS89b, TCH87, WW88a, BS89a, BSS89, FH83, MP88b, RO87a, Wei82, Wei89b, WF80]. Red-Tide [DM82]. Reduced [BGSS88, HSdWM88]. Reducing [SO85, SGSD86]. Reductase [DM82, Liv88, Moo88c]. Reduction [Cap82, HL82, SA81, JS85, JS88, SG87a, Tra88]. Reef [BS80c, BvdEH^{+}82, Bou81, BLR86, BY81, BR82, Col80, DP81, Ham82a, Ham82b, HSD87, KC81, LR80, McG89a, OE81, Rob82, Sch80b, Sch81a, Sch84b, She80, She82, Thr82, Wil82, dB81b, BL87a, BG84, CSDZ86, CHAS83, Cro84, DSW89, DE89, GABM84, HDFG88, HLS83, HGS89, Hop88, HSB88, JWC^{+}83b, JWC83a, Kar86, KM89a, KM87c, KM88a, Kin89, KMM88, KHW88, LDDH88, Lin88, Mar83a, MD86, McC89a, McC89b, McC89c, MR88b, PB85a, RB84, Rid88, RO87a, RA89, Sam83, SSN87, Sch84a, SF89, SO87, TB89, Van88a, WS83, Wil86a, WC88, dB86, dB87a, dB87b, AA84, AG82, CLS^{+}82, DCMS89, Din82a, Din82b, Rid88, Rus84a, Rus84b, SR87, She85, WH82, Wil86a]. reef-associated [McC89a, TB89]. Reef-building [Sch84b]. Reefs [Bul82, Pea81b, Rei81b, WH82, BCd84, BLR86, Gal87a, Gal87b, Gly85, Gri83, KMD87, McC89a, RO87a, Rus84b, Ryl83, SS86a, SS89a, SR87]. Reference [Jac80, NLBFD80, SJSO82, Sor80, FF86a, JKMR84, KHHB86, MCA89, SLP83, VL88, WPW88]. references [Gas85]. refractory [KLN88, NL86]. refuges [BB86]. Regard [BAG80b]. regenerated [SS87]. Regeneration [KNL82b, SFHU82, BK85a, HSD87, Hop88, KP89, Pro87, Sel88, Tam89, VB85a, VB85b, YD84]. regime [HH84a]. regimes [SBK88]. Region [AG81, GT87, WtD89, Abr83, AS87, FWE89, PMJH84, Rüg89, RM88, Sch88a, SM85, TR89, VDL88, Whi89, FC82]. Regional [Din82a, Fis89, JAS88, TB89]. regions [BBH^{+}88, Par89]. regularly [Het89]. regularly-flooded [Het89]. regulating [HB86, Möl86, TNR86, WH88]. Regulation [HNR^{+}88, MWM86, RNH^{+}88, RHvd81, WR82, WR84a, NR89b, SVK83, Sul84, TE86b]. Regulator [TC81]. Regulatory [Lon81b]. rejection [HSRM86]. Related [DG82b, AC85, Asm87, Auf88, Bro88, FEHA85, GB89, Jos87, KB88, KSM88, Lam87, LM87a, Pia89, STW89, TE86a, Tan88, Van83, Wal85b, dB86, vdVB86]. Relating [GCBM82]. Relation [Che84, KAE81, PSB^{+}81, Pea80, Sv80a, WU80, Wei82, dB81b, AHF87, Ang87, BSS89, CE89, Col87, DW87, Fis83, GG88, Ger89, HHB88, HMD88, JMSK86, KMH85, KMR87, LSD85, LH88, LTH87, LVP^{+}89, LNS^{+}87, MUC84, MR84, MK85, OWKG85, Par84, Pie83, PM88, RR87, SDRN87, SSM^{+}84, Ste86, WP88, dB87b, NBMM88]. Relations [GD81, MCGP82, Pol84, RG82, SPH87, MERW84]. Relationship [Din81, SP86b, BBSS89, MC87, PGV^{+}89, RA87]. Relationships [BW81, DL82, HHE89, LCP84, PT89, Sto80a, AB89, BTQS88, BS89b, Dua89, DSG^{+}89, HG86, JJN85, KSVB89, LNL83]. Relative [Eva87, SN85, WDS^{+}84, dB81b, CHH83, HB85, KP89, Pos87]. Relatively [LL81]. Release [ALAH85, RW86, Cam85, Fuh87, Lan83, Mac85a, Mø83]. Released [LNB81, LNV81, RMZ82, Wol82, Dun86, TE86b]. relict [HS88a]. remineralization [GH83]. Remote [Aik81, RG82, Smi81]. Remotely [BLP87]. Remots^{T M} [RG82]. Removal [Moo81, SO83a]. Renewed [DB82]. replete [BH83a]. replicate [HH84a]. reply [Amb86b, Gra83, McG87, Tod85]. Reproduction [Cla82, GG82, Hig82, KB87, OTC82, Sam82, Sch87a, Sha89, VS82, Bar86, BS83a, FW86, GMT89, HC86a, Jen85, Las84, ME88, Oli84, Pol84, Wah85, WB83a]. Reproductive [BPH87, Bar89, Ber84, Hei88, Jen82, KQ82, McC89b, TD81b, TMBG85, Van83, Wil89, BEM87, CFH89, Cla87, GB89, GFS89, GMK85, HS85, LRS87, ME88, RSM88, SCT87, Ste89, Tho84, Tra88, WLB84, MT85b, vDBS81]. requirement [SJ88a]. Requirements [Spa80, BB88, Ber84, Che89, HYT^{+}88, HB85, Thu84a, WtD89]. research [Mar89b]. resident [OWZ89]. residing [WB83b]. Residual [LSB85]. Resistance [Din81, HL80, SU87]. resistivity [Tam87]. Resolute [SACC88]. resolution [PHT85, VK85]. Resource [BH81a, Ben84, HC86b, Mac81, SB82, TW82, HSB^{+}84, Mac86a, SDRN84, SN88, SK84]. Resources [BH81b, Joh82, DG88b, FD89, Lev85, SN85]. respect [JG89]. Respiration [Ban82, Gri80b, HS88a, Has81, KMR87, AHMB89, Bau85, HlWPH84, JA85, KB84, KMH85, Klu84, SHPH86, Spr84a]. Respiratory [LN81, Sch88a, BBR88, MR88a]. responding [dHS88]. Response [BNF81, Sam82, SS80b, WE88, Bar89, BP84a, CL88, EMA88, FL86a, FA83, Gil85, GKA88, GL89a, HAOC^{+}86, HR89, HR87, KJ88, LSP84, Mag88, MW83, MCM84, MH84, MB85, MT84, PV85, PLM89, RW86, TS88a, TAK88, WK88, ZK85, tBB83]. Response-Surface [BNF81]. Responses [FER85, GP88b, LN81, Liv88, MR81, ZW82a, ZW82b, And89a, BA87, BAC^{+}88b, CL84a, DK85, ET84, FC83, FSZ89, FWB89, GD84b, GR89, HMF85, Mar89a, Moo88c, OOF83, PDWF85, SDY87, SJ88b, SAS87, SP85, SRL^{+}86]. resting [SM86]. restoration [FF86a]. restrained [Gil85]. restricted [PS88]. resulting [Low88]. Results [Bod88]. resupply [Col84]. Resuspension [BS80b]. Retention [Kaa89, SM85, Wil83, AS85, Hic84, LD89, Rii88b, SHCL87, SP88]. reticulatus [Sch80a, Tal80]. Reunion [Bou81]. Reutilization [ABB80]. revealed [KBv89, Lam86]. Reversed [BRS86]. Reversible [Ric85b]. review [Gol85, LP88, Pal88, RT85]. reworking [RBR86]. Rhithropanopeus [LN81]. rhizomes [HKd88]. Rhizophora [BH85]. Rhizosphere [Cap82, AK88]. Rhode [SM86, Tra88, Gal89]. Rhodochorton [BH87]. Rhodomela [D’A85a]. Rhodophyceae [RR82]. Rhodophyta [KH88a, RK81, Rie82b]. Rhodophyta. [HS85]. rhomboidalis [DD80]. Rhone [KSVB89, MR88a]. rhythmicity [CK89]. Rhythms [DB81a, Has81, HM86]. Ria [HAOC^{+}86, LNF81, VP85]. Ribbed [SFN82, KLN88, Lin89]. ribulose [SP85]. richardii [Boa80]. Rico [Sto88]. Riftia [CFH89, GH86]. Right [HBFA80]. rigidula [tBB83]. rings [DJ84]. risk [Mag88]. River [BA81, FW86, Gal89, Man88, NVLL88, Alb83, MCPW88, MR88a, ELR88, KSVB89]. rivulets [RMO88]. RNA [Bam83, BTH^{+}84, DRCA83, Hov89, WH85]. RNA/DNA [DRCA83, Hov89]. rock [BB88, HL89, HG86, Jer85, Mac85a, Mac86b, Van88b]. Rockall [CSB81]. Rockfish [BK80]. rocks [BL88]. Rocky [Boa80, Amb86a, BH86, CB88, GT87, Him86, Jer86, Kro88, Lin88, LML83, Mac86b, MB84b, MB85, MH89, OD89, Pag86]. Role [Amb84, AFF^{+}82, BSC85, D’A85a, DA85b, EHF^{+}81, KE87, Lon81b, Rus87, Sch87b, SRTM83, SLE^{+}89, SP84b, War84, AN88, Bro88, CL84b, FB86, Fro87, HD89, LT87a, LDKR88, McG86, NR89a, Ohm84, PAF86, PF89, Scr83, She85, TB86, TNR86, WZ83a]. roles [BRS86, HAA^{+}88, Lev85, SH84]. Romanian [BSF84]. Root [Thu84b, PB89, TCH87]. root-boring [PB89]. Root-exuded [Thu84b]. roots [HKd88]. Roscoff [Jac80, KS87a]. roscoffensis [DG82a]. rosea [BCS84, TMBG85]. Rosfjord [KBv^{+}83a].

Rosfjorden [BDKK83, KHE^{+}83]. rotula [EB86]. rotundata [HR89]. rotundatum [Coh85]. ROV [BLP87]. Royal [KHW88]. Ru [CSH89]. rubrum [Cra89]. Ruditapes [RNK87]. rugulosum [CSB81]. runoff [WS88]. rupestre [LN86]. rupicola [BB84, Bla84]. Ruppia [Thu84a, Thu84b].

S [Lee88, VL88]. S-transferase [Lee88, VL88]. S. [BK80, Øre87, TD81a]. S0 [JS85]. Saanich [JB86, Lar86]. saccarina [MKS80]. saccharina [Lün88]. Saduria [HS88a]. Sagartia [SBR87]. Sagitta [Kim84, Øre86, Øre87, Pea81a]. Sagittaria [PDP87]. Saint [SMB85]. saira [OTK80]. Salangid [IR81]. salar [Lar84]. salina [RBA86, RBA88]. Saline [SA82]. Salinities [LN81]. Salinity [BBBM80, Joh81b, Joh82, LH81, LHH81, MR81, NR89c, Reg84, SSK80, TW80, TM80, AT86, AAN87, BK85a, BMRN84, Fis86, Geo85, Kri83c, NR89b, PDP87, TAK88]. Salinity-induced [Reg84]. Salinlty [Joh81a]. salinus [CL84a]. Salmo [Lar84]. salmon [Lar84]. Salp [Ise81, HMD88]. salps [AN88, HMD88]. Salt [KPJ84, Kne81, EWM87, FC83, FCFP84, Gle86, HKd88, Het89, Lin89, MW85, OK89, RD89, SC84b, WBVS86, WWD^{+}88]. salt-marsh [Het89, SC84b]. saltatrix [OSB85]. salterns [Ama83]. saltmarsh [FJ88b]. saltwater [CFP88]. Sample [KR82, Lin88]. Samplers [BH83b]. samples [BR85, Dem85, DR86, HKB87, KS83, Pur88, Sch84a, SF86, WRKD84]. Sampling [BC81, FM88, GB86, Ign87, Ken89, KC89, Mac86c, PPCR86]. San [Clo82, NT85, AC88, CL86, CDL87, CC84, LWE87, Oka86]. Sand [Sch80a, BS83a, BTH^{+}84, CF84, D’A85a, Eml86, HP86b, ID88, Lew84, LML83, MP86, MKHD85, Sch89a, SHB^{+}83, Tam85, Tam87, VP85]. sand-dwelling [HP86b]. sandeels [FG88]. Sandflat [Gra81]. sandpiper [BFP87]. sands [NW84]. Sandy [KNL82a, JD85, MM86, MH84, Mil89, PPCR86, vM87]. Sanriku [SN81]. Santa [SLE^{+}89]. Santos [Jac87]. Sapelo [PAN89]. sapidus [HLH87, HS81, Ov87]. Sardine [WEHF85]. Sargasso [BP85, FSCP89, GPCW88, IM88, PGV^{+}89]. Sargassum [Ang86, McC84, SG84, TW81]. sarsi [SM87a]. satellite [BLR86, KB88]. Saturated [MST81]. saturation [CP84]. Saury [OTK80]. saxatilis [Che89]. sayi [BL88, FWB89]. scaber [KM87c]. scabra [HD89]. scabridum [LN86]. scaevola [AZ85]. Scale [KAE81, OE81, SD85, SP86a, AH84b, ABAI87, GCSP88, Hic88, JL86, Jen88, Jos87, LH88, MF89b, SH89, Ued87, Wah85, Wat88]. Scale-dependent [SD85, SP86a]. Scaled [Ban82]. Scales [BR82, EM89, Fai88, Gal87a, Gal87b, Mar83a]. scallop [BGSS88, DE88, ELK84, Fev89, MT85a, MT85b, Mac86a, RLB89, Wol88, WRC83]. scallops [BEM87, LRS87, WK88]. Scanner [BM86a]. Scanning [Fuh81]. Schedules [Ovi81]. Schlei [Gas85]. schlosseri [BBBM80]. School [DD84]. School-oriented [DD84]. schooling [O’B88a, O’B88b, SP86a]. sciaenids [HHA85]. Scleractinia [BL87a, KQ82, PDIF85, Sch84b, Sie88, Van83, Van88a]. Scleractinian [Bou81, Bul82, FV82, Har85, HF87, Wal85a]. Scoloplos [RBR86]. Scomber [PA84]. scombrus [PA84]. Scope [BNF81, WJ88]. scorpaenid [Hoe88]. Scotia [EC87a, JNCS89, JS87a, Sam88, STM86, ZW89b, Gif88, HPV^{+}87, HH87, KM88b]. Scotian [MN87]. Scotland [HMM87, TL84]. scour [Hei89, Ken88]. Scrobicularia [MGC82, WW83, WWL83]. SCUBA [JAN85]. scurra [MS89]. Scurria [MS89]. Scyllarides [STPA88]. scyphomedusa [CK89, SG88a, FK84]. Scyphozoa [Kee87]. scyphozoan [Grö89]. scyphozoans [Grö88]. SE [Wal85b]. Sea [AA81, AFF^{+}82, BL87b, Bra81, CS80, Car87, DD80, GKM^{+}87, Hic81, Ise81, KC82, KGS87, Lat86, MD87, Mil85, MG85b, OTC82, PKMS85, PSMS87, Ric85a, SED81, SF89, Sv80a, Sv80b, TD81a, TW82, Ama83, AC85, And89b, AI88, AH84b, BCd84, BGL^{+}88, BCS84, Bri85, BM86b, CC86b, CC87b, CMC89, DJTL88, Dem85, DJ84, Ebe86, FK84, FHC88, GH86, GFS89, Gar89, GNBO85, GLT^{+}86, GMB87, HGH87, HTC84, HVK85, Him86, HHKA88, Hsi88, HH84c, Ign87, KGV84, KN83, KS87b, LRS87, Lar84, LDG87, LGK87, MG85a, MB84b, MLDT88, MGT89, MS84, Paw89, PH87, Pet88, PE89, RLB89, Rum89, Sch87a, SCT87, SPK^{+}87, SHH^{+}88, TMBG85, VB85a, VB85b, VAF83, VCS84, WPS89, YALD88, dB86, dB87b, Bai89, BM86a, BP85, BNVP87]. Sea [Bd89, BSSW89, BFP87, BSF84, BBH^{+}88, BS89a, BSS89, BR83, BC88, BCFS82, CM82, CSW84, CD81, Cus89, Daa87, DCNS84, FPRW82, FH83, FSCP89, Gas85, GG88, GPCW88, GCBM82, GB83, HC86a, HM87, Her88, HPF89, Hov89, IEK86, IBMD81, IM88, JD85, JS86, JP84, JOP86, JP87, KAE81, KEK^{+}84, KKT81, KKKW86, KD84, KK89, Lan83, LN86, MB84a, MR88a, MP88b, NR89a, NVLL88, NvB89, PH81, PB85b, PGV^{+}89, RTLR88, RS87, RGH89, RO87a, SN81, Sch82b, SRS89, SS87, SPSS89, Spi87, SBK88, SR85, SRL^{+}86, SHG89, TS88b, TL85b, Wal85b, Wei82, Wei89b, WZ83a, WF80, Yan82, vdV85, vdVB86, van86b, vB87]. sea-ice [BGL^{+}88, GLT^{+}86, LDG87, MLDT88]. sea-side [Ama83]. sea-star [HGH87]. sea-surface [CMC89, WPS89]. Sea-temperature [MD87]. sea-water [FHC88, HHKA88]. Sea/Skagerrak [Ric85a]. Seabird [AG81, BH86, FC82, Han86, HS88b, Sch82b, SRL^{+}86, BBG^{+}87, BP84b, SD85]. seabirds [Abr83, AA84, And89a, SP86a]. Seafloor [RG82]. seafoam [Arm89]. Seagrass [GL82, Jac80, Liv82, MP81, Nel81a, WY82, You81, BMZZ88, CD88, Dua89, FF86a, GB86, Hic86, How85, KCFS89, KN83, MW87, OB88c, Ov87, PB88, SPH87, SWB89, STM86, SP84b, Vel86, VC86, WW88a, WB86, War84, WB83b, Wil88, ZM84]. Seagrass-Associated [GL82]. Seagrasses [RMZ82, Sto80b, RC86, Wil87]. Seal [Boa80, RM88]. Seals [SS82]. seamount [LDMT86]. seamounts [KB88]. Search [DB82]. Seas [GMF88, GFM89, GM89]. season [Dah86, HMLF88, OSB85]. Seasonal [AP80, BW88, BR83, BP81, Cro84, DS86c, EC87b, EZ89, FB87, HO80, HB85, JOP86, JS85, KC82, Lar86, MCA89, TW82, VN82, Wal85a, WR81, YLH86, Båm88, Bam83, Bro88, DCJS85, FWE89, GD84a, HC86a, Hsi88, JG89, JAS88, KS84, LSD85, LWE87, MR83, NR89a, OD89, PCW86, SM86, Ov87, RZ85]. Seasonality [McC88, Mor83, STPA88, TW83]. seasonally [CL86, GL89a]. seasons [GKOW87, Øre87]. seastar [AP83]. Seawater [MT80, MCGP82, AV87a, AFHA84, BP88, Bid88, Bjø86, BB87, BR85, Fis89, FB85, FF86b, Fuh87, FST85a, FST85b, HAHA84, KJ88, KS87b, LD85, Man89, PKMS85, SS86b]. Seaweed [PH86, DDR89, HDFG88, Smi86]. Seaweeds [DG82b, KB80, LNF81, RR82, TH84]. Sebastes [BK80]. Secondary [Asm87, FP82, GMT89, Eme89]. Sediment [ABM81, BNF^{+}89, Blo80, DGHK82, DS86b, Gra81, Gra84, HC80, Hic88, Hoc81, NK81, PB81, RG82, SED81, Sv80a, WY82, AH84a, And86, Arm88a, Arm88b, BAN89, CDL87, Chr89, DM83, DL83, Dem85, DL86b, Eme89, ES87, FF86a, FL86a, Gra86a, HL85, HA87, JAS88, JS88, KN88, Kem88, KA85, KDT^{+}88, Kri83a, LBA86, LS88, LBH89, LVP^{+}89, MB87, MG85b, Mon84, NK86, NM89, OOF83, PF83, RBR86, Ryg85, Sch88c, SC84b, SG88b, SJNL89, WW88a, WCC^{+}88, WBAN85, Wil86b, WZ83b, YD84, ZM84]. sediment-associated [DM83, DL83]. sediment-bound [HL85]. Sedimentary [LlJC82, LC82, ES84, FL86a]. Sedimentation [BSSW89, ITBW81, OS80, SW86b, Was83, Was85, AN88, BNVP87, FSZ89, Gra87, Rie89, TRL89]. sedimented [HP89]. sedimenting [RS87]. Sediments [Cam82, Cap82, FKW81, GCBM82, HL82, MGC82, MP81, PGBN81, PY81, ABT89, AH87, AH86, BV83, BLK88, BK85a, CJS87, EWM87, ES87, FNH83, FCFP84, Gar84, GMQ^{+}88, JNCS89, JD85, JS85, KCS89, MMS89, MSS86, MR86, Mil89, MW85, OC89, PPCR86, PM88, Rei88, Rei83, RGLK87, Rid88, SR86, SG87a, SG87b, SGSD86, VAF83, Wat88, WYG85, dHS88]. Seeded [Bro82]. seep [JS87b, MBP^{+}86]. seepage [WS88]. Segregation [Eva82, Han86, HL89]. seine [WEHF85]. selected [Bje88, ELK84]. Selection [CS80, GP82, Ham82a, HS81, KM81, NC89, NH82, Oli84, AP83, Bar88a, Bro86, DeW87, DP87b, HL84, Jon86, PGJ83, PR84, Pih85, SH84, STPA88, SWB89, ap 86]. Selective [AS85, Che82, GOAYH86, Hic81, Kim84, NK82, PE89, Sch81b, WR89, ALH86, Bri85, CC86b, DS86b, HD89, PA84, SSA^{+}86]. Selectivity [LlJC82, CL84b, GNBO85, JF88, LlJC83, BH81a]. selenium [Bje88, HYT^{+}88]. self [DR86]. self-contained [DR86]. Semi [Has81]. Semi-Lunar [Has81]. Semidiel [SP88]. Semihlanus [RSWJ80]. senile [Wah85]. sense [AS83]. Sensitive [GP82, PGJ83]. Sensitivity [GC85, Jen81, CA87]. Senslng [Smi81]. Sensor [Aik81]. separate [RN89]. separation [ET85]. September [ESR^{+}88]. sequence [DS86a]. sequencing [Sin86]. Seriatopora [HGS89]. series [Gei88, ID88, Par87b]. serpulid [Mar89a]. Serpulidae [SWSL81]. serra [Don87]. Serranidae [SF89]. Serve [SC82]. Sesarmidae [GLNR86]. sesarminid [NGR86]. Sessile [Sch82a, BB86, Hir86, Hir87, McG89a, SW84]. Seston [Hic84, BHNI89]. setiferus [MF89a]. setosa [Øre86, Øre87]. settle [Paw89]. settled [Lan89, ZWO89, vdVB86, vB87]. Settlement [CR84, GW85, Luc84, Nel81b, RZ85, SWSL81, Sch82a, TD81b, Tod85, DCMS89, Grö89, HF87, HB86, Hoe88, KD86, LOvM89, MKG88, Ov87, OWZ89, PT84, Van88a, van86b].

Settlement-Timing [TD81b, GW85, Tod85]. settling [Kee87]. Several [Pea80, Fai88, Ger86, SAS87, Tam87]. Severe [TP89]. Severn [CW82]. sewage [Lam86, OQM^{+}87, PB85a, SKIR89, SCSD86]. sewage-distorted [SKIR89]. sex [BRS86, HLH87, SWB89]. sexual [Jay89, ME88]. Shade [Din82a, Din82b, Ken88]. Shade-Dwelling [Din82a, Din82b]. shaded [TD89]. Shallow [BS80c, Eva82, Ham82a, Ham82b, Jør82, LWG^{+}81, CSW84, Car87, DR86, Eme89, ET85, GD84a, LV86, LDKR88, LM85, MPR85, Möl86, Par89, PR84, Pih85, SM87a, SC86, SJNL89, YY89]. Shallow-Water [LWG^{+}81, Par89, SM87a, SJNL89, YY89]. Shark [WW88a]. Shatt [AS87]. Shaw [Gra83]. shedding [RV84]. Sheldon [Pea85]. Shelf [CL82, FR80, LWG^{+}81, WH82, YLH86, BC88, Epi88, GMF88, GFM89, GM89, HJ84, HlWPH84, LNL83, Rus84a, Rus84b, SMB85, SdT86, STM88, Tur84, CSB81, LBH89, MN87, Sam88, SB82, SC86]. shell [BH88, FS84, LGP88, MT85a, TO88, TC87, Fis83, Spr84c, TS88b]. shells [BM86b]. Shelter [Amb82, Sv80a, SLO82, MR88b, STPA88]. shift [HK86]. shifts [LSD85]. Shine [Som80]. shoals [LM87b]. shoots [TD89]. Shore [BB86, KNL82b, AZ85, BV85, CH84, Gra84, Jer86, Qui88, RN89, Sch89a, SW85, aH85, ap 86]. shore-bird [Gra84]. shore-crabs [CH84]. shorebird [Pie83]. shorebirds [Pie87]. shores [Kro88]. shoreward [Sha83]. Short [GKOW87, NM85, SV85b, TL87b, TL87a, Zel88, CJS87, Gar84, HMF85, How85, LS83, OK89, TD89, Wil86a]. short- [HMF85]. short-incubation-time [LS83]. Short-term [GKOW87, NM85, SV85b, TL87b, TL87a, Zel88, CJS87, Gar84, How85, Wil86a]. Shrimp [Cla82, HO80, WM82, WR82, BLP87, Dal85, Gle86, HS86, KD84, LL83, MF89a, PR84, Reg84, TE86a]. Shrimps [DB81a, KM83, WW88b]. Siam [Lat86]. Sibling [BS80d]. side [Ama83]. Sightings [HBFA80]. Significance [AAK^{+}88, CC84, Hop83, Joh80, KSR80, NLVS80, NFG82, NL84, SW82b, Sv80b, vM87, AS83, GSC83, HSRM86, JS85, Kem88, LBH89, SN85, TH84, TS88a]. significant [IM86]. sihama [Lee81, LHH81]. silicate [HA87]. Siliceous [MS84]. Silicic [YD84, KZ85, RS87]. silicium [SW86b]. silicoflagellate [TL85a]. Silicon [OR80, DOB^{+}89]. Sillago [Lee81, LHH81]. silver [NCG^{+}83]. Silverside [GD81]. SIMCA [VK85]. Simeon [Cos82]. Similarity [Blo81, KR82, LCP84]. Simple [Hoc81, Ign87]. Simulated [OFGG82, BLR86, DK85, Gra87, MT84, Ste86, WRC83]. simulating [AN88]. Simulation [VN82, ZSA87]. Simultaneous [BDKK83, SSPA89]. Sinai [TP81]. since [KKKW86]. single [TL85a]. single-cell [TL85a]. Sinking [JS86, BH83a, Bie85, PWMl89, Rie89, TKP86]. sinuosa [FV82]. SiO [LT87a]. Sippewissett [KPJ84]. Sipuncula [Mur84]. Sirvival [DD80]. site [Bro88, KSS85]. site-related [Bro88]. sites [AE88, EC87b, OA88, RB84, VL88]. situ [BBR88, KM88b, ZH81, CJS87, Dah86, DTLN85, HW82, How85, JMG84, Mon84, MT84, Ols85, TL87b, TKP86, ZK85]. situations [Col84]. Size [ALH86, BAG80b, Bie85, Cam82, Fuh81, GPR86, GHS85, Ham82a, HSW89, HS81, Jay89, Joh81a, Kne81, KR82, Leb82, Pea80, RS80, Sch81b, WY82, ap 86, CG87, DP87b, ER87, FL86b, GG88, Gul87, HLH87, JF88, Jon86, Klu84, LDG87, LS88, Mac88, MC87, ME88, Øre86, Pea85, PT84, PCM86, Pro85, RTLR88, RNK87, RD89, Rum89, Rye88, SM87a, SHPH86, Sch88b, Tam85, TN85, WCGG86, Wat86, WR87]. size-class [GG88]. size-efficiency [Pea85]. size-fractionated [LDG87, Pro85]. Size-Selective [Sch81b, ALH86]. Size-Specific [Kne81]. size-structure [Rum89]. sized [MF88, NBT88]. sizes [Mau88]. Skagerrak [Ric85a, Jos87]. Skeletal [Hug87, DJ84, Las84]. Skeletonema [CSD88, EBH^{+}83, LDER83, Pet89, SF86]. skeletons [BL87a]. skills [CH84]. skin [RV84]. skin-shedding [RV84]. slabberi [WPW87]. slicks [Sha83]. slipper [STPA88]. slope [STM88, Wil86a]. Slopes [She80, She82, WH82]. Slowly [Kem81]. sludge [OQM^{+}87]. slurries [SC84b]. Small [AH84b, Ban82, SF80, Ued87, Wat88, Hic88, LS83, TN85, TN86, TNR86, Wah85, Wat86]. Small-scale [AH84b, Ued87, Wat88, Hic88]. smell [Van86a]. smelt [LD89]. Smelter [WY82, War84]. smithi [GLNR86, NGR86]. Snail [TP81]. Snow [KM80, DS86a, GKM^{+}87, Her88]. soak [Ken89]. soak-time [Ken89]. Soft [Blo80, CLS^{+}82, Eva82, Las81, OOS82, SS80a, SS80b, Thi81, Amb84, Amb86b, BL87b, BH88, Cof84, CA85, CJS87, Eme89, ET85, LCS86, MPR85, WCC^{+}88, Wil86b]. Soft-Bottom [Eva82, OOS82, SS80a, SS80b, Amb84, Amb86b, CA85]. soft-corals [BL87b]. Soft-Sediment [Blo80, Eme89, WCC^{+}88, Wil86b]. Solaster [MC87]. sole [AS83, GB84]. Solea [AS83]. solidum [BSv81, SvDB81, vDBS81]. solidungula [DS86c]. solitary [GSC83, Sma83, TH84]. Soluble [Har82, RSWJ80, LL83]. Solution [AG82]. somatic [BEM87, CF84, MT85a]. Some [HG86, KKT81, Sch82a, Bri85, CSH83, FG83, HL84, HKS86, Mac85b, McG89a, MPR85, Pie83, PR84, SG84, Wah89a, Fen82a]. sordidus [Rus87]. sorus [Lün88]. Sound [KGS87, PKMS85, Rum89, SPK^{+}87, CC82, CC88a, MP86]. Source [AG82, CM82, Gas85, GABM84, GMF88, HMD88, JM85b, Kem88, Sog84, VC86]. Sources [Ohm88, BB88, FB86, LN89, LvB87]. South [BS80c, DCB^{+}86, Don87, Jer86, Lat86, Law86, Mac88, Sch89a, WWD^{+}88, CSW84, CH88, KAK88, MJ84, TCH87, TL84, BS83b, Clo82, GNBO85, IEK86, MR84, MERW84, MW85, Mor85b, NT85, OK89, SO83a, SS80b, WY82, WS88]. south-east [KAK88]. south-western [CSW84, TL84]. Southeastern [GCBM82, Sch82b, BBH^{+}88, DCNS84, NvB89, STPA88]. Southern [And89a, SO83b, Abr83, BCCH84, BK85b, CH88, Daa87, GLB87, HHD^{+}84, HP86a, HHE89, JBAH83, JD85, KFPH88, KN83, OS84b, PAH89, PWMl89, Pro85, RM88, SO83a, SF85, SB85a, Stu86, TB85, TC88, Whi89, WAN84, Bla84, Car82, Cus89, EHF^{+}81, FEHA85, FC82, Hov89, KZ89, Lan83, RK82, SCSD86, VDL88, YALD88]. southwestern [Car87, EC87a, LM85]. Southwestward [DE87]. sp [BB87, MWM86, PDTH86, RR82, SN89, Tam85]. space [BM86b, Fai88, Gar89, SO85]. space-mediated [Gar89]. Spacing [LCS86]. Spain [FL88, HAOC^{+}86, LNF81, VP85]. spanner [Ken89, KC89]. Sparidae [DD80, KC81]. Spartina [AH84a, And86, DK86, HKd88, KA85, KLN88, NL86, OK89, WBVS86]. spat [RZ85]. Spatial [AW86, HSB88, Hsi88, Jac87, JAN85, KAE81, RB84, Rei87b, SB82, SW86a, ST82, Thr86, BB86, BY83, CF84, HM87, IM88, JS87b, KH88a, Ken89, LC88, MB89, McG86, McG87, PS86, Pia89, PPCR86, SB87, WW86, WC83, YY89, YLH86, Gal87a, ZW82a, ZW82b]. Spatio [McC89c]. Spatio-temporal [McC89c]. spawned [GMS85]. Spawning [CL82, SSM^{+}84, SHB^{+}83, Wol88, Cus89, Dun86, HHA85, Jay89, Lam87, SM85]. Special [Jac80, NLBFD80, Gas85, JKMR84, KHHB86, MCA89, SLP83, WPW88]. speciation [Ste89]. Species [BS80a, BK81, BK80, Bre82, BS80d, CC82, CC88b, CC88c, Col80, CD81, GP82, JH80a, JH80b, JK80a, JK80b, Las81, Lon81b, ML80, MGC82, Pea80, Rie82a, RS80, SG85, She80, TW82, VB81, WB80, WCGG86, ARGL87, Bam83, Ber84, BRS86, Col85, CC87b, DA84, DL86a, Epi88, FF86a, FJ88b, GD84b, GL89a, GHC83, Gri84, HJ84, HS85, HMH89, Hop88, HC86b, HH84c, KS84, KH88a, Ken87b, Ken88, KL84, Kri83b, Kri83c, Mag88, McC84, OWZ89, Paw89, PGJ83, PL85, Rii88b, SS89a, SR87, SCH89b, SBR87, SF89, Sin86, SBR84, SWB89, Sto88, TE86a, Tam87, TAK88, TN86, TC88, VAF83, VCS84, War88a, WZ85, WW88b]. Species-specific [CC88b, CC88c, HMH89]. Specific [Kne81, Lon81a, Pea82, CC88b, CC88c, Dou83, DNF87, FNH83, HMH89, IM88, KMH85, KMR87, LTS86, Red83, ZWO89]. specificity [Bar86]. spectabilis [McC89b]. spectra [GHS85, NVHHH88, PCM86, Sch83c, Sch88b]. Spectral [HMH89, PG88, SPK^{+}87, PSMS87, Riv89]. spectrophotometer [ABJS88]. spectrum [GG88, Pea85]. speed [Lin88]. sperm [Dun86, WPS89]. Sphacelaria [tBB83]. Sphaeroma [PB89]. Spheniscus [Wil85]. spider [ALPS83, Ang87, AHMB89, KA84]. Spill [Jac80, Bod88, FC83, RG85]. spines [Ebe86]. spinifera [DSS^{+}85]. spinulosus [OS88]. Spiny [BS80c, HB86]. Spionidae [BK81]. spiral [Don87]. Spirobranchus [DRB86, Mar89a]. Spirulina [TBRBA85]. splashpool [OFGL88]. spondyloideum [DE88]. sponge [Bar86, Bar88b, HSdWM88, OS85]. sponges [CP87, CM87, Hop88, ME88]. Spontaneous [AAN87]. Sporangia [BL80]. spot [GOAYH86, NC89, BLR86]. spotted [Sog84]. spp [BA87, BSS89, BCCH84, CC86a, CC86b, CC88a, Fro87, GPCW88, GGA85, IM88, Kri83a, RBR86, WBAN85]. sprat [FG88]. Spring [CC89, BF87, FSZ89, GMS85, HAOC^{+}86, HH87, JBAH83, JS85, KS87a, LVdVdV85, LM87a, NK89, Par87b, Rie89, Roy89, SW86b, ZW89b]. springtime [MLDT88]. squamata [ACF^{+}80]. squares [Pat88]. squid [Pat88, RSM88]. Sri [Jay89]. St [HHD^{+}84, HP86a, Hei89, PT89]. St. [HHD^{+}84, HP86a, Hei89, PLT87, PCM86, WYG85]. Stability [DL82]. Stable [CD88, SKIR89, STM86]. staff [Kin86]. stage [HLH87, LOvM89]. stages [DW87, JA85, OFGL88, Øre86, SVK83]. staging [BFP87]. stain [Ans89]. staining [How85]. stalked [Pag86]. stand [SW86a]. stand-off [SW86a]. Standard [VS80, LTS86]. Standing [Gun80, KAE81, NVLL88, DCNS84, Lar86, Law86, MH89, TR89]. stands [CH88]. Star [SED81, HGH87, Rum89]. starfish [Bla84, MC87, Ols85]. stars [ME88]. starvation [Cam85, Col84, SM84]. starvation-induced [SM84]. state [DG88a, HAHA84, LT87b, Vid87]. States [Sha88a, SSM^{+}84]. station [BNVP87, WAN84]. Statistical [CG88]. status [BW83, FF89, OK89]. steady [HAHA84, LT87b]. steady-state [HAHA84, LT87b]. steinbergae [Yos86a]. steinbergiae [Yos86b]. stenolepis [WM82]. step [LMO86]. Stephanoeca [GL88a]. sterols [MCA89]. stimpsoni [MC87]. stimulation [Gil85]. stipes [MS89]. stirring [EMA88]. stock [DCNS84, Fro87, GLB87, NVLL88, TR89, vdV85]. Stocked [MCGP82]. Stocks [Gun80, Lar86, MH89, SF85]. Stoichactis [HVK85]. stolonifera [Klu84, SN88]. Stomnformes [LHT88]. Stone [Sv80b, Wil89]. stony [Van83, Van88a]. storage [Asm87]. stores [RLB89]. storey [Ken87b]. storm [DV83]. storms [Bar89, OQ83, RD89]. Strait [NK89, Alb83, HHE89, LvB87, PSB^{+}81, PCH85, TP89]. Stranding [SCF^{+}82]. Strategies [RR82, TD81b, Bar88a, BGL^{+}88, CVF89, Epi88, GSC83, JS87b, LCS86, LN86, MG85a, PPCR86, SU87, SSM^{+}84, TNR86, Van83]. Strategy [BH81b, FCW82, vDBS81, Wah85]. stratified [ABAI87, BSSW89, BH83b, CS84, HM86, HM87, HHN^{+}83, HlWPH84, MKD^{+}86, MCPW88, PN88, PCH85, TL85b, Was83]. Stream [AW86, Han86]. Stress [Sam82, WU80, AA89, Bak88, CSH89, Eme89, FG87, RW86, RO84, SC85, TAK88, TP89, VMM^{+}88, WH85]. striae [CTG86]. striped [Che89]. Strombidium [FJ88a]. Strombus [Sto89]. strong [DD88]. Strongylocentrotus [KGV84, LGK87, TD81a]. Structural [BB85, Gie81]. Structure [AG81, BCFS82, DOB^{+}89, FVDBD87, Gal87a, Gal87b, GL82, HFK82, JS87a, KAE81, Kro88, LNF81, OE81, SW82a, SH85, Tar81, TD81a, WY82, WH82, Ang86, BHP^{+}86, BV83, Bie85, BH86, BSS89, BY83, FK86, GLB87, GFM89, Gri83, HP86b, Hei89, HH89, Hug84, HH84c, KH88a, KM89a, Lam87, Low88, McC89c, MB85, MR88a, ME88, Nic84, NR89a, OD89, Pag86, PLM89, PB88, PST87, Rum89, Sam83, SO83a, Sma83, TN85, TN86, TNR86, WG84, War88b, War88c, Wei89a, WB83b, Wil86a]. Structures [Leb82, She80]. structuring [Amb84, War84, dKB88a]. studied [FA83, Fuh87, KS87a]. Studies [Blo81, BP81, DG82a, FPRW82, Gie81, GF82, JH80a, JH80b, Joh81a, Joh81b, Joh82, LC82, Nel81a, WF80, Ans89, Bak88, CPM89, FSBOG84, FB85, KN84, KP89, MJ84, RBR86, Rüg89, TL85a, TR89, PST87]. Study [Bou81, CLS^{+}82, LTS86, OE81, SJSO82, SSK80, VS80, VCS81, BLP87, CG88, Eml86, Fis88, FL85, FL86b, Gar84, GMQ^{+}88, Grö88, Grö89, JP84, KS87a, LS84a, MG85b, NC89, PAN89, PR88, SG88b, SJNL89, TB85, VMM^{+}88, WW88a, Wat88, WWD^{+}88]. stutchburyi [MB89]. stylifera [Abo84]. Stylophora [HGS89]. Sub [Lam86, AB89, BAC^{+}88b, BB85, HLH87, RHS89, BB84]. sub-Antarctic [AB89, BB84]. Sub-catastrophic [Lam86]. sub-estuary [HLH87, RHS89]. sub-lethal [BAC^{+}88b]. sub-tidal [BB85]. subarctic [Bat86, DSG^{+}89, Fro87, Tho84]. subduction [JS87b]. Sublethal [ME89, TR87]. sublittoral [CA83, CJS87, Hei89, WCC^{+}88]. Submarine [Joh80, Rei81b, KCFS89, KCS89, WBAN85]. submerged [RO87b, Thu84a]. submersed [TKS^{+}85]. Submersion [DG82b]. subsequent [Bal84, Tam89]. Substances [WM82, Wol82]. substrata [BL87b, D’A85a, GSC83, JG89, Lew84, Lan89, McG89a, Van88a]. Substrate [CS80, TE86a, LML83, She85, Bar86]. Substrate-related [TE86a]. substrates [BL88, Hop83, MR86]. Substratum [Sto80b, MB84b, Sch89a]. Subsurface [ESR^{+}88, FA83, Kim85]. Subtidal [BCFS82, Sch80b, Sch81a, Amb86a, Bal84, BV83, DW87, Dav89, Fle87, Jos87, KSS85, MKHD85, OD89, RGLK87, Sch84a, Sto86, SB87, Thr86, Van88b]. subtropical [DPJ87, NK86, TL85a, UKTA83]. subulata [RSM88]. success [WLB84, WPS89]. Succession [GSC83, Hir86, Hir87, LNF81, VS84, ZW82b, BP88, Gri83, LT84, SM86, Van88b, Hir87]. successional [DW87]. suitable [MHW88]. sulcatum [FJ88a]. Sulfate [HL82]. Sulfide [PCR80, MFP87]. Sulfide-System [PCR80]. Sulphur [DSS^{+}85]. Sulphur-oxidising [DSS^{+}85]. summary [Add88a, Cap88b, Moo88a, War88c]. Summer [JH80c, Daa87, Par87a, RTLR88, SW86b, SBK88, JH80a]. Sunlight [SS86b]. Sunlight-induced [SS86b]. Sunrise [Mac85a]. superba [EM87, HHE89, MM81, Mor85b, SHH^{+}88, Wat86]. superciliosus [Gib88]. supplemented [VB85a, VB85b]. supply [GMF88, GM89, HGS89, HlWPH84, PW87, ZWO89]. Suprabenthic [SMB85, JL86, Kaa89]. surf [LM87b, PM88, WPW88, vM87]. surf-zone [LM87b, PM88]. Surface [BC81, BNF81, DGHK82, HC80, O’B88b, Pie83, Sha83, AS85, CMC89, EMG89, Gar84, GMB87, HM86, Ign87, KN88, Kee87, KJ88, Kim85, LTH87, MW84, Nic84, NK86, PDJ86, Pia89, WPS89]. Surfaces [SED81, Mac86b]. surfperch [SH84]. surfzone [TB86]. surgeon [MP88b]. surgeon-fish [MP88b]. surgeonfish [FMM87, WP88]. surplus [GFS89]. Survey [BAG80b, HYT^{+}88, BCS84]. Survival [BSv81, Joh81a, JM85b, LH81, LHH81, RRB^{+}80, SW82b, TNR86, VS80, AT86, BTH^{+}84, DCMS89, FL85, FL86b, Jen85, KMR^{+}88, LH87b, Lev85, Luc84, Pet84, RV84, SO84, SS89b]. survivorship [AC85, Lin89, Pet86b, Sto86, ZWO89]. susceptibility [CT89, PH86]. Suspended [Moo81, LNS^{+}87, Mac86a]. Suspension [KM81, BBL84, HNR^{+}88, Jør83, LN89, Ola89, Rii89, RR87]. suspension-feeders [LN89]. Suspension-Feeding [KM81, HNR^{+}88, Ola89, RR87]. suspensions [Wil83]. SW [BC88]. swamp [BH85, KAK88]. swarming [O’B88a]. swarms [Nic84, UKTA83, Wat86]. Sweden [Båm88, LH82, LH88, Øre87, PR84, Sel88, Tal80, TW82]. Swimming [Jen81, MW83, Epi88, Jon89, Lin88, MK85, SMB85, aH85]. switch [GB89]. sword [SK85]. sword-fish [SK85]. Syllidae [PH83]. Symbiosis [DG82a]. Symbiotic [KSR80, KM85, MWM86, SvDB81]. sympagic [AA89]. Sympatric [BS80a, Bro80, CC87b, HTC84, McC84]. Synchronous [SLMB88]. Synechococcus [BA87, BSP89, CC86a, CC86b, CC88a, GPCW88, IM88, PGV^{+}89]. syngnathid [TB88]. Syngnathidae [BRS86]. Synoptic [CDL87, LlJC83]. synthase [CA86]. Synthesis [Cla82, MP81, RMZ82, Bid84, CC88b, CC88c, Fis88, GPR86, KNH86, LMO86, LT87b, SA89]. synthetase [BA83, CA87]. synthetic [NVHHH88]. Syringodium [Wil87]. Systeilaspis [WR87]. System [Car82, DBFM80, Liv82, PCR80, SFHU82, SC82, AH87, BAN89, BT85, CFP88, EMHS86, FNH83, FHS88, LBA86, LTS86, LM85, Man88, Moo88c, NS87, NVHHH88, OPN^{+}84, ORK^{+}86, PWMl89, Pro85, Sch88a, SJNL89, Vel86, RG82]. Systems [ABB80, Din81, LL81, SA81, Alo87, KNH86, Sha88a, Sha88b].

T. [SDY87]. tables [WWB86]. Tachidius [HHG84]. tactics [HS88b]. tactile [Gil85]. tag [TC87]. Tahiti [DE87]. Talitridae [AS87]. tall [OK89]. tamarense [Han89]. tamarensis [GKA88, GFG89, MHW88]. tannins [NGR86]. Tapetum [Som80]. Tasman [Bra81, GB83]. Tasmania [O’B88b]. taste [Van86a]. Tautogolabrus [BL88, LB86]. taxa [FSZ89]. taxonomic [VK85, War88a]. Taxonomy [EM89]. TBT [LGP88]. Tc [Dah86]. Technique [LC82, SFN82, BBL84, CC86a, DL85, PB87, PHT85, Red83, TL85a, TRMF89]. Techniques [OWP80, CC86b, ES84, GB86, RGLK87]. tehuelchus [RZ85]. teleostean [FL88]. teleosts [LM87b]. Telinidae [Mee85]. Tellinacea [Leb82]. Temora [Abo84, RHP89]. temperate [CA83, Hob88, KM87c, KM88a, LH87a, Lin88, McC89b, McC89c, MW87, SP84b, UKTA83]. temperate-zone [SP84b]. Temperature [AP80, BBBM80, Buc82, Cla87, Joh81a, Joh81b, Joh82, KS84, LH81, LSP84, MR81, RHvd81, Rie82b, SW82b, SP85, TW80, WtD89, AT86, Ang87, BMRN84, BTH^{+}84, Dah86, Dal85, Fis86, FA83, Geo85, GSC83, JA85, JM85a, Kri83c, LS88, MT85a, MT85b, MW83, MB84b, MD87, Moe85, OSB85, Pau83, Pie83, Rei87a, SM86, TL84, WR84a, WPS89, WF80]. Temperature-Salinity [MR81]. Temperatures [BNF81, HL80, JH80c, LN81]. Temporal [BH88, Dua89, IM88, LC88, Øre86, SS89a, SB82, Wil86a, You81, ZW82a, ZW82b, AW86, HSB88, Jac87, JAN85, JS87b, KH88a, McC89c, PS86, WC83, ZK85, Gal87b, Ken87a]. Ten [THSW80, JH87]. Tenacity [HD89]. Tenacity-mediated [HD89]. tension [AT86]. Tentative [Yan82]. tenuis [BT88, Jen81, Jen82]. terebellid [DB88]. Term [DL82, OE81, BL87b, CJS87, ELK84, FER85, FHS88, Gar84, GKOW87, GC83, HMF85, How85, ID88, LH88, NM85, NCG^{+}83, PH87, Rei87b, RG85, SSN87, SV85b, TL87b, TL87a, Wil86a, Zel88]. ternatanus [PD86]. terrestrial [AZ85]. territorial [GMK85, Hoe88]. Territoriality [SW82a, Sam83, WS83, Zel88]. territories [KMD87]. tertiolecta [SGB89]. tertiotecta [BP84a]. Test [KOO80, VS80, BMGB88, Par84, Phi85, Sin86]. testing [Wei89a]. Tests [DBFM80, Ger86, SDMN88]. testudinum [TD89, Wil87, ZIO84]. Tetembia [BLR86]. tetrodotoxin [KDT^{+}88]. texani [LB86]. Texas [FR80]. TH [TC81]. Thaiassiosira [BP84a]. Thalassia [Lew84, TD89, Wil87, ZIO84]. Thalassinidea [DG88a]. Thalassiosira [EB86, SDY87, Sch83b, SP84a]. Thalia [HMD88]. Their [BP81, Fen82a, PSB^{+}81, She82, ABB80, AS83, BK85b, CT84, ES87, FF86a, Har86, HHE89, HSB88, IM88, KM85, PF83, Pie83, Pie87, PPCR86, Sto89, TB86, TE86a, TW80, TNR86, Wat88]. Thematic [BLR86]. theoretical [LP88]. Theory [Cha82, Ebe86, EM89, LH85, MGM86, MR83, TP85, Thi87, Wei89a]. Theragra [Bai89, DCNS84, Pau83]. there [RN89]. Thermal [WRC83, OSB85]. these [SC82]. thiebautii [OF82, Scr83]. Thiobacterium [SN89]. thorns [Ols85]. thread [Bd89]. thread-drifting [Bd89]. Three [JWC83a, JK80a, JK80b, Las81, MGC82, PCR80, VB81, Yon80, BL88, Bod88, BMRN84, Col85, DA84, EC87b, GHC83, HLS83, Hop88, KL84, Kri83b, Kri83c, Mar83b, MD87, RB84, SBR84, TC88, VAF83, WW83, WWL83, YALD88]. three-year [EC87b]. threne [SKS^{+}84]. thy [DNF87]. thy-midine-specific [DNF87]. thymidine [Hol88, KH88b, PAN89, RNJ^{+}84]. Tidal [BS80b, Bd89, HO80, Has81, KdC81, LV86, Leb82, PB81, WU80, BB85, CE89, Dec88, DK85, EZ89, FCFP84, JD85, KPJ84, KD84, LD89, LDKR88, Ov87, Reg84, RD89, RO87b, RMO88, SO84, SR86, Vid87, WB83b, WZ83a, WZ83b, WS88]. Tidal-current [Bd89]. tidally [BW88, HlWPH84, LT87a, Sha83, vdVB86]. tidally-dominated [BW88]. Tide [DM82, FITF80, Han89, Tra88]. tides [LG88]. tightly [WH88]. Tigriopus [OFGL88]. Time [Col86, GTG81, HR89, HS81, Li82, Par87b, PCH85, Rud88, BBL84, BR85, Fai88, Gei88, GMK85, HSB^{+}84, ID88, Ken89, LS83, MR84, SO85, VC86, WZ83b, BA87]. time-averaged [HSB^{+}84]. Time-Course [GTG81]. Time-lag [Col86]. time-series [Gei88, ID88]. times [HP89]. Timing [TD81b, GW85, SS89b, Tod85]. Tintinnid [CC82, Han89, Ver86b]. tintinnids [CTG86]. tintinnine [HC86a, Hei88]. Tisbe [VB81]. tissue [BH88, CL86, FT86, MT85a, WR84b]. tissues [BFP87, ELK84, MGT89, PD86]. Tolerance [BAG80a, LHH81, NL81, SSK80, TM80, WF80, Cha88, HS86, OFGL88, Sie88, TW88, WtD89, WRC83]. toluene [BR85]. Tom [Gra86c]. Tonda [FW86]. tonsa [HR87, KMH85, PS88, STW89, SM86, TC81, TR87]. tool [MC88, Mil85]. topography [BB86]. Toro [SO83b, VCS84]. Torres [TP89]. Total [EBH^{+}83, MT85b, PD86]. totteni [Lev85, LSD85, LlJC83, Pet84]. Toxic [PHL^{+}86, GFG89]. Toxicities [CLS^{+}82]. Toxicity [THSW80, BNF^{+}89, BMRN84, CDL87, EM89, FBT84, LGP88, RBA88, War84]. toxotes [Pau81]. Trace [BBH^{+}88, BP81, PY81, WY82, AS88, BAC^{+}88a, CL86, OOF83, PWM85, Phi85, PR88, RYD84, VMM^{+}88]. tracer [BBL84, FHC88, Gar84, Mon83, SP84a]. Traits [Lon81a, CLB88, LS84b]. Transect [ABM81]. Transfer [RTW81, SF80, VA85a, MG85b]. transferase [Lee88, TTH83, VL88]. Transient [Col84, ET84]. transition [NPVW86]. transitions [HR89, PDWF85]. Transplant [dKB88b, RYD84]. transplantation [WW83]. Transplanted [HFK82]. Transplants [BCFS82]. Transport [Bai81, CSS82, CSS83, EMG89, Gra81, Ise81, OS80, PGBN81, SC82, ZH81, AH87, Bd89, HMM87, HA87, LB87, MB87, Pal88, PAH89, Sch88a, Sha83, Tam87, WS88, JK80a]. Transuranics [GHCH81]. Transuranium [HC80]. trap [Dem85, KC89, LBA86]. Trapezia [GGA85, HC86b]. traps [KC89]. trawl [Abr83]. trawl-fisheries [Abr83]. trawling [RM88]. Tree [Thr82]. Trematoda [Shi80]. Trematodes [TP81]. Trematomus [RTK^{+}89]. trend [EZ89]. Triad [CDL87]. tributyltin [LGP88]. Trichodesmium [BSS89, Scr83, OF82]. tricornutum [PW87, SGB89]. tridens [KSM88]. Trididemnum [BSv81, KD86, SvDB81, vDBS81]. Trieste [HM87, HPF89]. trilobata [DG88a]. trimethylamine [SG87b]. tripton [PBP85]. tripunctulatus [BH81a, BH81b, HB81]. triquetra [CC88c]. triterpene [PAF86]. trochid [MUC84]. Trochus [BLR86]. troglodytes [SBR87]. Trophic [DSG^{+}89, Liv82, MB85, SSA^{+}86, Tar81, Yan82, AB89, AS86, CA85, DL86b, JS87b, KMM88, Pea85, PN88, WF83]. trophodynamic [KSF89]. Trophodynamics [BH81a, BH81b, HB81, CS87]. Tropical [BSv81, HJ84, Rei81b, RMZ82, SvDB81, VS82, vDBS81, Alo87, ABT89, AHMB89, Bal84, BH85, BW88, BAN89, Bue87, CT89, DL86b, HS87, LH87a, Mor83, MWM86, NGR86, OC89, Par89, PF86, RDD88, SP86b, SO87, SP85, SKM85, Sor85, SP89, Sto88, TGA89, WP88, YALD88, Zel88, dKB88a, BW81]. tropics [SV85a]. truncata [FL86a]. trypsin [Båm88]. Tubastraea [Sch84b]. tube [CFH89, TB85]. Tubes [GH86, Alo85, Kri83a, MFP87]. Tubificid [Gie81]. tubing [PHL^{+}86]. Tubularia [Sch83a]. Turbellaria [Mar83b, TW80]. Turbellarians [PCR80]. turbidity [Che89]. turbidostats [PW87]. turbulence [ASMV88, Che89, EAM87]. turbulent [Wro84]. turf [SB87]. turfs [WC88, Zel88]. Turnover [ABB80, Asm87, FF86b, How85, LBH89, NK89]. Twin [SS82, BBL84]. Two [BW80, BK81, Bul82, BS80d, JH80c, JS88, KQ82, Lon81a, Lon81b, SCF^{+}82, TD81a, WB80, You81, ARGL87, AT86, AS85, Ber84, BRS86, BCS84, BK80, CG87, CC87b, CYC83, Dav89, DP87a, Don87, Epi88, EC87b, FBCH83, FJ88b, GP88a, GB86, HS85, HH84c, Kar86, KM89b, KN83, LN89, Mag88, MW87, Nil87, RGLK87, SDY87, SBR87, STW89, SWB89, TE86a, TB88, TNR86, Van83, VAF83, WDS^{+}84, WW88b, dKB88b]. Tygon [PHL^{+}86]. Type [PT84, Som80, DD84, GMT89, LVP^{+}89, Roy89, RHP89, CT84]. types [HF87, KN83, OS88]. Typical [KKT81, BCS84]. typicus [WR89]. tyrannus [FAG89].

Uca [Chr89, CF84, WP87]. Ultilization [Ov87]. ultrafiltration [DL85]. ultraplankton [GPCW88, Joc89]. ultrasonication [ES84]. Ultrastructure [DD85, DP87b, DP87a]. Ulva [HR89, RR82, SJ88a, SJ88b, WAH89b]. ulvae [FJ88b, TE86b]. Unambiguous [CS86]. Uncinial [DB88]. undatum [HGH87]. under-ice [PSMS87]. under-storey [Ken87b]. underestimation [SHCL87]. Underice [CM82]. understorey [Ken88]. understory [SO83a].

Underwater [UKTA83]. undulates [GHC83]. undulatum [RHvd81]. undulatus [GOAYH86]. unfragmented [CSD88]. unguiculata [CK89]. Unialgal [OF82, EBH^{+}83]. Unicellular [BAG80b, BAG80a, KSR80, SvDB81, WU82, HBJ^{+}88]. uniformly [Cam85]. uniformly-labeled [Cam85]. unique [Kin88]. United [Sha88a, SSM^{+}84]. Unoccupied [She85]. unorthodox [BB88]. unpolluted [Sei87]. unprotected [McC89a]. unshaded [TD89]. Unstocked [MCGP82]. unusual [BNF^{+}89, BB84]. Upogebia [KM83]. upon [Alb83, BL88, HL85, WWB86, ZE89]. Upper [SA82, HH84b, KSS85, LL81, PT89]. Uptake [AA81, AH81, CSH89, DM82, GTG81, Har82, JK80a, JK80b, KW88, Li82, LB87, MRF80, MGC82, SHH81, STP84, TW80, VAF83, ALP84, Bje88, Cam85, Col84, CS84, Fuh87, FHC88, Gei88, HH85, HPC89, Kri83b, Kri83c, LV86, Le 86, MG85b, MGT89, NNA84, NR89b, NR89c, NBT88, PHT85, PAN89, PCH85, Pro85, RBA86, Spr84d, Tam89, ZJ83]. Upwelling [AG82, Bar84, FD89, FWE89, HAOC^{+}86, LT87a, MP88a, PWMl89, Pro85, Pro87, Rüg89, Sch83b, SdT86, TR89, WF83]. upwellings [DE87]. Uranium [Ham80]. Urchin [CS80, BCd84, CC87b, DJ84, Ebe86, KGV84, LGK87, Mil85, dB86, dB87b]. Urchins [TD81a, AC85, And89b, DJTL88, Gar89, Him86, PH87, VCS84]. urea [KW88, LBH89]. Urophycis [SLO82]. USA [BTL86, BGSS88, CE89, EMG89, Het89, LWE87, PAN89, Pet88, PM88, Sha88b, SMN86]. Usage [SLO82]. Use [BLP87, FHC88, Gra83, HBJ^{+}88, LC82, RG89, SB82, STM86, TKT86, TW82, WH85, CC86a, DR86, FF89, HB86, Het89, RO87b, TC87]. Used [DB81a, GM87, KM87a, aH85]. user [EMHS86]. user-friendly [EMHS86]. Ushant [PMJH84, Vid87]. Using [BCF82, OWP80, RG82, SFN82, BNF^{+}89, CC86b, CS86, DL85, Dec88, DO86, How85, JMG84, KB88, MB89, Mon83, MR83, Pat88, RHP89, Sch83c, TRMF89, VP88, VK85]. Utilisation [LNB81, LNV81, Sch80b, SDRN84, SLN81, Sch84a]. Utility [CP84]. Utilization [Amb82, BH81a, Gle86, JH80c, KLN88, LN89, MT87, Sog84, TCH87, Thi87, BWH87, BS83b, CC89, HAHA84, HB85, HHKA88, HC86b, KCFS89, LM87b]. UV [Sie88, SA81].

V [Col84]. Vagile [You81, GB86]. Valenciennellus [BH81a, BH81b, HB81]. Value [SJSO82, TE86a, Tho84]. values [BR85, WA87]. vanhoeffeni [DD85, DT85, DP87b]. vannus [CSD88, GH83]. varia [SC84a]. Variability [BS80a, BP81, DK86, GCK82, JCL86, WR81, Wil82, ABAI87, Col84, CD88, Dua89, EC87b, EM87, FDM86, GPCW88, IM88, LC88, MB89, Nel89, Ohm88, PS86, Rus84a, SS89a, SD85, SS89b, SB87]. variable [Fis83, HGD86]. variables [GKOW87, HG86]. varians [NR89b]. Variation [BK80, BP86, BP81, CLB88, DD88, Din82a, GL82, KA84, MC87, PB81, ZW82a, ZW82b, Alo87, Bam83, BH83a, BH88, BEM87, Bro88, CT89, EZ89, FG87, FB87, HB85, JAS88, Kar86, KH88a, KPJ84, MKD^{+}86, MJPH^{+}88, ME88, MFS83, NM85, SG84, TO88, WR84b, Wil86a, ZW89a]. Variations [AP80, HO80, Has81, Mac87, MH89, Wat86, Båm88, BW88, BR83, Cro84, EWM87, FEHA85, HM87, Hsi88, HMLF88, LV86, LH88, MD87, NR89a, Ov87, RGLK87]. variegata [dB87a, dB87b, dKB88a, dKB88b]. variegate [CT89]. variety [Mac87]. variopedatus [Rii89]. Various [Pea81a, Rie82a, Bie85, BS89a, OS88, SO84]. varying [HH84a]. Vascular [Ten82, TKS^{+}85]. vegetated [FCFP84, WS88]. vegetation [Ang86]. Vegetational [BW81]. vegetations [dB87b]. Vegetative [BH87]. Vehicle [BLP87]. veliger [Man89]. velvet [aH85]. Veneridae [TO88]. Venetian [BBBM80]. Venezuela [Bri85]. vent [CFH89, GH86]. ventilabrum [OS85]. Ventilation [Kri83b, Kri83c, HS88a, JMSK86, Kri83a]. ventrosa [Sie82]. vents [SN89]. verification [CC88c, Dec88]. vermetid [Sma83]. Verrill [BB84]. Versatility [ZE89]. versus [CSD88, KNH86, LCF89, LvB87, OWZ89, SS87, VC86]. vertebrate [KM89b]. Vertical [BSS89, DG82b, GP85, GG88, HPV^{+}87, HHB88, HHN^{+}83, Jon89, NBR^{+}88, NBMM88, OS80, PN88, PAH89, RTLR88, RGH89, SN81, Spa80, Wei82, WR81, AS85, Bai89, BK85b, CPM89, DCJS85, HKS86, KAE87, LD89, LSD85, LCP84, MF89b, MERW84, NvB89, OSB85, SP88, Ued87, WB86, HB81, KM87b, NBR^{+}88]. vertically [Kaa89]. Very [Ban82, SJ88b]. vesiculosus [HC80, JS87a, KKKW86]. Vestimentifera [GH86]. vetulus [GB84]. VI [GKM^{+}87]. via [CC88b, CC88c, WS88]. Viability [TC81]. Victoria [BK81]. VII [PSMS87]. VIII [KGS87]. villosus [MFS83, Ste89, TL87b, TL87a]. violacea [TMBG85]. Virgin [WYG85]. Virginia [Man88]. virginica [LN89, NJ83, NL86, OWZ89, PWM85, RRB^{+}80, SK82, WU82, WH85, ZJ83]. viridis [Cha88, Phi85]. vis [TL85a]. vital [Ans89]. Vitamin [DL85]. vitamin-binder [DL85]. vitro [CA87]. vitulina [Boa80]. viviparity [PH83]. vivo [DTLN85]. Volume [Hob88, BB87, FL86b, Kin88, LS83]. Volume-frequency [Hob88]. Volumetric [DCB82]. volutator [BMRN84, Jen85, MR81, MGC82, NK82]. vs [LL81, TH84]. vulgare [LPV87]. vulgaris [AS83, GHCH81, GF82, MG81]. vulgata [NLBFD80]. Vulnerability [Fui89, GFG89].

W [CSB81, LNF81, DJ84]. Wadden [CD81, Bd89, KD84, SHG89, WZ83a, vdV85, vdVB86, van86b, vB87]. wading [Eva87]. wailesii [RHP89]. Wales [KB87, Jer86]. wall [Van88b]. walleye [Bai89, DCNS84, Wal85b]. Wandering [KSM88]. war [Pur84]. Warm [Bra81, JH80a, JH80b, GB83, SC86]. Warm-Core [Bra81, GB83]. Washington [DD88]. Washington-Alaska [DD88]. wasn’t [Cra89]. waste [Lam86]. wasting [SMN86]. Water [AFF^{+}82, BS80d, DGHK82, Dec86, DL82, JB86, LWG^{+}81, Sch81a, TM80, And86, BGSS88, BSSW89, BAN89, BH83b, CC86b, CS89b, ES87, FHC88, GC85, GD84a, HS86, HS88b, HA87, HM87, HSD87, Hop83, HHKA88, JNCS89, KHE^{+}83, Kim85, KB88, LL83, MH84, MR88a, MG85b, MGT89, OA88, Par89, RBA83, SMB85, SM87a, Sch84a, Sch89a, SSA^{+}86, SSPA89, SG88b, SJNL89, SF86, TAK88, TRL89, VB85a, VB85b, VBFS85, Wil88, WZ83b, WRKD84, YY89, ZM84, dB87a, dB87b]. water- [SG88b]. water-borne [OA88]. Water-Column [AFF^{+}82, JB86]. Water-cover [Dec86]. waterflow [HSdWM88]. Waters [EHF^{+}81, LH82, PY81, Rei81b, SA82, ALP84, AS86, Bie85, BF87, BK85b, DE87, FA83, FHS88, GKOW87, HJ84, HM86, HHN^{+}83, HlWPH84, Joh88, KM88a, KSF89, LMO86, LM87a, LDDH88, LNL83, McC88, MFS83, MS84, OWKG85, PAN89, PN88, Phi85, Pia89, RA87, Sel88, SRS89, SC86, STM88, Tis89, TL85b, Tur84, UKTA83, WC83, YLH86, ZW89b, ZW89c]. waterway [BW88]. Wave [BY81, BGD87, KM89a, MB84b, MB85, Tam87]. wave-induced [BGD87]. wavelength [NPMW87]. Waves [Kah82, Sha83]. Weaning [SS82]. web [HHD^{+}84, HP86a, IM86, SKIR89, SE85, VP88, BCF82]. webs [SW85, SR85]. Weddell [HC86a, JS86, MB84a, NvB89, Spi87]. Weight [SC82, BR83, CL86, Cam85, DPJ87, Fis83, HL84, HMD88, Mø83]. well [AS85, LD89]. well-mixed [AS85, LD89]. West [HJ84, Mac88, KS87b, TÓ89, Vid87, IEK86, Sha88a, YLH86]. Western [CHAS83, Joc89, SW86b, WW88a, AM88, Båm88, CSW84, KAA88, Mee85, Øre86, PR84, PB85b, RTLR88, SK85, TL84, Was83, Was85, WZ83a, ZW89c, vdV85, vdVB86, van86b, vB87, ID88, LH82, SP89, TS88b, YLH86]. Westerschelde [HKd88]. wet [SV85a]. wet-dry [SV85a]. WH7803 [BSP89]. Whale [HBFA80]. whales [BHP^{+}86, BPH87, WPS89]. Whatman [Hic84]. whelk [FUM84, HGH87]. whelks [Fai88]. which [Blo80]. White [CS81, DJTL88, MF89a]. whiting [BCA89]. whole [FNH83, JM85a, SDMN88]. whole-cell [JM85a]. whole-system [FNH83]. wilkitzkii [AA89]. will [FNH83]. Williamson [Tod85]. Wilson [Amb86b]. Wind [Eme89, DK85]. Winner [Gra86b, Gra86c]. Winter [BM86a, Buc82, FKW81, Gra86a, KAA88, Sto89, Bar89, BS89b, CLB88, Coh85, FHKR85, KS87b, MB84a, Reg84, Rie89]. winter/spring [Rie89]. Wintering [HBFA80]. Within [Cam82, KC81, DG88a, ET85, How85, Kee87, Wal85b, WWD^{+}88]. without [Sch84b]. wlth [LNF81]. wood [BH85]. Workshop [AS88, BFMS88, BAC^{+}88d, BAC^{+}88e, FM88, KWW^{+}88, BCG88]. world [KZ85]. worm [CFH89, Oli84]. worms [GH86]. wrightii [PB88].

X [LB81]. X-Irradiated [LB81]. Xanthidae [McD82, Wil89]. xanthurus [GHC83, GOAYH86]. xenobiotics [Moo88c, SKS^{+}84]. Xenophyophorea [LDMT86]. Xiphias [SK85, BBSS89]. XIV [VS80]. XLI [LTS86]. XV [VCS81]. XXIX [LS84a]. XXVI [SJSO82].

Yangtse [NVLL88]. year [EC87b, KS84, Kin88, MN87]. yearling [Fui89]. Yearly [Wil82]. yellowfin [Pol84]. Yellowtail [LH81, Col87]. yessoensis [ELK84]. yield [BSP89, Bjø86, LNS^{+}87, SPK^{+}87]. Yolk [JH80c]. yongei [Sch87a]. York [KLE85]. Young [OTK80, HTL85].

Zealand [Kin89, WH87]. zeaxanthin [BSP89]. Zebrasoma [WP88]. Zinc [NLBFD80, NL81, WR82, WR84b, WU82, Cha88, FT86, NR89b, NR89c, RWK83, WR84a, WRKD84]. Zinc- [NL81]. Zn [Dah86, FBT84, HL85, JCL86, RSWJ80]. ZoBell [Mon89]. Zonation [DB82, She80, Tam85, Alo87, LSD85, Rus84b]. Zonations [Gol85]. Zone [BM86a, Boa80, Gun80, PRS^{+}82, CS84, DD88, FH83, HK84, JS86, JP87, Kro88, LM87b, MP88a, Ohm88, OD89, PM88, RS87, SBK88, SP84b, Wil89, vM87]. Zones [Wei82, BS89a, JS87b, KMD87]. zooflagellate [PDTH86]. zooflagellates [TKP86]. Zooplankton [CC82, Che82, CW82, DT85, EHF^{+}81, KM87a, KM87b, KM89b, LH82, RDD88, RS80, Tur84, Wei82, AW86, Båm88, Bam83, DR86, ER87, FST85a, FST85b, Har86, HKB87, HNR^{+}88, JG89, KM88a, KFPH88, LH88, MC88, PA84, PHL^{+}86, Sch88a, SSM^{+}84, SB85a, Tow84, BAH87, Ovi81]. zooxanthellae [CK89, HGS89, Sch84b]. Zostera [BTL86, Cap82, DA85b, GB86, HFK82, KMBSJ87, KMAM84, NG86, VN82, You81, ZSA87]. Zygogenetic [Ama83]. zygotes [BRS86].

References

Ammerman:1981:DCA

[AA81] J. W. Ammerman and F. Azam. Dissolved cycllc adenosine monophosphate (CAMP) in the sea and uptake of CAMP by marine bacteria. Marine Ecology Progress Series, 5:85–89, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

           tronic). URL https://www.int-res.com/articles/meps/5/
           m005p085.pdf.

Allaway:1984:NIS

[AA84] W. G. Allaway and A. E. Ashford. Nutrient input by seabirds to the forest on a coral island of the Great Barrier Reef. Marine Ecology Progress Series, 19:297–298, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

           www.int-res.com/articles/meps/19/m019p297.pdf.

Aarset:1989:EOS

[AA89] A. V. Aarset and T. Aunaas. Effects of osmotic stress on oxygen consumption and ammonia excretion of the Arctic sympagic amphipod Gammarus wilkitzkii. Marine Ecology Progress Series, 58: 217–224, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

           1599 (electronic). URL https://www.int-res.com/articles/
           meps/58/m058p217.pdf.

Ameyaw-Akumfi:1987:BCM

[AAH87] C. Ameyaw-Akumfi and R. N. Hughes. Behaviour of Carcinus maenas feeding on large Mytilus edulis. How do they assess the optimal diet? Marine Ecology Progress Series, 38:213–216, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
38/m038p213.pdf.

Aksnes:1988:SAC

[AAK^{+}88] D. L. Aksnes, J. Aure, S. Kaartvedt, T. Magnesen, and J. Richard. Significance of advection for the carrying capacities of fjord populations. Marine Ecology Progress Series, 50:263–274, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/50/
m050p263.pdf.

Ameyaw-Akumfi:1987:SIC

[AAN87] C. Ameyaw-Akumfi and E. Naylor. Spontaneous and induced components of salinity preference behaviour in Carcinus maenas. Marine Ecology Progress Series, 37:153–158, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/37/m037p153.pdf.

Adams:1989:DDT

[AB89] N. J. Adams and C. R. Brown. Dietary differentiation and trophic relationships in the sub-Antarctic penguin community at Marion Island. Marine Ecology Progress Series, 57:249–258, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/57/
m057p249.pdf.

Atkinson:1987:FSO

[ABAI87] M. J. Atkinson, T. Berman, B. R. Allanson, and J. Imberger. Fine-scale oxygen variability in a stratified estuary: patchiness in aquatic environments. Marine Ecology Progress Series, 36:1– 10, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
36/m036p001.pdf.

Antia:1980:PAN

[ABB80] N. J. Antia, B. R. Berland, and D. J. Bonin. Proposal for an abridged nitrogen turnover cycle in certain marine planktonic systems involving hypoxanthine–guanine excretion by ciliates and their reutilization by phytoplankton. Marine Ecology Progress Series, 2:97–103, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/2/m002p097.pdf.

Arneson:1988:CCD

[ABJS88] A. C. Arneson, S. Benyamin, B. Jones, and G. W. Schmidt. A charge-coupled device (CCD) spectrophotometer for measuring marine bioluminescence. Marine Ecology Progress Series, 43:277– 283, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
43/m043p277.pdf.

Anderson:1981:IBC

[ABM81] J. G. Anderson, P. Boonruang, and P. S. Meadows. Interrelationships between chlorophylls, carbon, nitrogen and heterotrophic bacteria in an intertidal sediment transect. Marine Ecology Progress Series, 6:277–283, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/6/m006p277.pdf.

AbouDebs:1984:CNB

[Abo84] C. Abou Debs. Carbon and nitrogen budget of the calanoid copepod Temora stylifera: effect of concentration and composition of food. Marine Ecology Progress Series, 15:213–223, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/15/m015p213.pdf.

Abrams:1983:PST

[Abr83] R. W. Abrams. Pelagic seabirds and trawl-fisheries in the southern Benguela Current region. Marine Ecology Progress Series, 11:151– 156, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
11/m011p151.pdf.

Alongi:1989:EEM

[ABT89] D. M. Alongi, K. G. Boto, and F. Tirendi. Effect of exported mangrove litter on bacterial productivity and dissolved organic carbon fluxes in adjacent tropical nearshore sediments. Marine Ecology Progress Series, 56:133–144, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/56/m056p133.pdf.

Andrew:1985:HRD

[AC85] N. L. Andrew and J. H. Choat. Habitat related differences in the survivorship and growth of juvenile sea urchins. Marine Ecology Progress Series, 27:155–161, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/27/m027p155.pdf.

Alpine:1988:PGR

[AC88] A. E. Alpine and J. E. Cloern. Phytoplankton growth rates in a light-limited environment, San Francisco Bay. Marine Ecology Progress Series, 44:167–173, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/44/m044p167.pdf.

Aldrich:1980:AEG

[ACF^{+}80] J. C. Aldrich, W. Crowe, M. Fitzgerald, M. Murphy, C. McManus, B. Magennis, and D. Murphy. Analysis of environmental gradients and patchiness in the distribution of the epiphytic marine hydroid Clava squamata. Marine Ecology Progress Series, 2:293– 301, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
2/m002p293.pdf.

Addison:1988:BEPb

[Add88a] R. F. Addison. Biochemical effects of a pollutant gradient — summary. Marine Ecology Progress Series, 46:75–77, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/46/m046p075.pdf.

Addison:1988:BEPa

[Add88b] R. F. Addison. Biochemical effects of a pollutant gradient introduction MEPS 46:31. Marine Ecology Progress Series, ??:??, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p031.pdf.

Addison:1988:HMM

[AE88] R. F. Addison and A. J. Edwards. Hepatic microsomal monooxygenase activity in flounder Platichthys flesus from polluted sites in langesundfjord and from mesocosms experimentally dosed with diesel oil and copper. Marine Ecology Progress Series, 46:51– 54, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p051.pdf.

Azam:1982:ERW

[AFF^{+}82] F. Azam, T. Fenchel, J. G. Field, J. S. Gray, L. A. Meyer-Reil, and F. Thingstad. The ecological role of water-column microbes in the sea. Marine Ecology Progress Series, 10:257–263, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/10/m010p257.pdf.

Ammerman:1984:BGS

[AFHA84] J. W. Ammerman, J. A. Fuhrman, Å Hagström, and F. Azam. Bacterioplankton growth in seawater: I. Growth kinetics and cellular characteristics in seawater cultures. Marine Ecology Progress Series, 18:31–39, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/18/m018p031.pdf.

Abrams:1981:ESP

[AG81] R. W. Abrams and A. M. Griffiths. Ecological structure of the pelagic seabird community in the Benguela Current region. Marine Ecology Progress Series, 5:269–277, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/5/m005p269.pdf.

Andrews:1982:USN

[AG82] J. C. Andrews and P. Gentien. Upwelling as a source of nutrients for the Great Barrier Reef ecosystems: a solution to Darwin’s question? Marine Ecology Progress Series, 8:257–269, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/8/m008p257.pdf.

Amouroux:1989:MCA

[AGA89] J. M. Amouroux, A. Gremare, and J. Amouroux. Modelling of consumption and assimilation in Abra alba (Mollusca, Bivalvia). Marine Ecology Progress Series, 51:87–97, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/51/m051p087.pdf.

Andersen:1986:NCM

[AGCD86] O. K. Andersen, J. C. Goldman, D. A. Caron, and M. R. Dennett. Nutrient cycling in a microflagellate food chain: III. Phosphorus dynamics. Marine Ecology Progress Series, 31:47–55, 1986.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/31/
m031p047.pdf.

Azam:1981:MKD

[AH81] F. Azam and R. E. Hodson. Multiphasic kinetics for D-glucose uptake by assemblages of natural marine bacteria. Marine Ecology Progress Series, 6:213–222, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/6/m006p213.pdf.

Andersen:1984:ESD

[AH84a] F. Ø. Andersen and B. T. Hargrave. Effects of Spartina detritus enrichment on aerobic/anaerobic benthic metabolism in an intertidal sediment. Marine Ecology Progress Series, 16:161–171, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
16/m016p161.pdf.

Astheimer:1984:SSP

[AH84b] H. Astheimer and H. Haardt. Small-scale patchiness of the chlorophyll-fluorescence in the sea: aspects of instrumentation, data processing, and interpretation. Marine Ecology Progress Series, 15:233–245, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/15/m015p233.pdf.

apRheinallt:1985:HMU

[aH85] T. ap Rheinallt and R. N. Hughes. Handling methods used by the velvet swimming crab Liocarcinus puber when feeding on molluscs and shore crabs. Marine Ecology Progress Series, 25:63–70, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
25/m025p063.pdf.

Armonies:1986:QEL

[AH86] W. Armonies and M. Hellwig. Quantitative extraction of living meiofauna from marine and brackish muddy sediments. Marine Ecology Progress Series, 29:37–43, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/29/m029p037.pdf.

Andersen:1987:COM

[AH87] F. Ø. Andersen and W. Helder. Comparison of oxygen microgradients, oxygen flux rates and electron transport system activity in coastal marine sediments. Marine Ecology Progress Series, 37:259– 264, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
37/m037p259.pdf.

Al-Hasan:1987:GCR

[AHF87] R. H. Al-Hasan and G. E. Fogg. Glycollate concentrations in relation to hydrography in Liverpool Bay. Marine Ecology Progress Series, 37:305–307, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/37/m037p305.pdf.

Anger:1989:GRD

[AHMB89] K. Anger, J. Harms, M. Montu, and C. Bakker. Growth and respiration during the larval development of a tropical spider crab, Libinla ferreirae (Decapoda: Majidae). Marine Ecology Progress Series, 54:43–50, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/54/m054p043.pdf.

Aoki:1988:POB

[AI88] I. Aoki and T. Inagaki. Photographic observations on the behaviour of Japanese anchovy Engraulis japonica at night in the sea. Marine Ecology Progress Series, 43:213–221, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/43/m043p213.pdf.

Aiken:1981:CSA

[Aik81] J. Aiken. A chlorophyll sensor for automatic, remote, operation in the marine environment. Marine Ecology Progress Series, 4:235– 239, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
4/m004p235.pdf.

Andersen:1988:OMR

[AK88] F. Ø. Andersen and E. Kristensen. Oxygen microgradients in the rhizosphere of the mangrove Avicennia marina. Marine Ecology Progress Series, 44:201–204, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/44/m044p201.pdf.

Andersson:1985:RAI

[ALAH85] A. Andersson, C. Lee, F. Azam, and A. Hagström. Release of aminoacids and inorganic nutrients by heterotrophic marine micro-flagellates. Marine Ecology Progress Series, 23:99–106, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
23/m023p099.pdf.

Albright:1983:IRO

[Alb83] L. J. Albright. Influence of river–ocean plumes upon bacterioplankton production of the Strait of Georgia, British Columbia. Marine Ecology Progress Series, 12:107–113, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/12/m012p107.pdf.

Andersson:1986:SSG

[ALH86] A. Andersson, U. Larsson, and Å. Hagström. Size-selective grazing by a microflagellate on pelagic bacteria. Marine Ecology Progress Series, 33:51–57, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/33/m033p051.pdf.

Alongi:1985:MMB

[Alo85] D. M. Alongi. Microbes, meiofauna, and bacterial productivity on tubes constructed by the polychaete Capitella capitata. Marine Ecology Progress Series, 23:207–208, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/23/m023p207.pdf.

Alongi:1987:IEV

[Alo87] D. M. Alongi. Inter-estuary variation and intertidal zonation of free-living nematode communities in tropical mangrove systems. Marine Ecology Progress Series, 40:103–114, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/40/m040p103.pdf.

Admiraal:1984:PDA

[ALP84] W. Admiraal, R. W. P. M. Laane, and H. Peletier. Participation of diatoms in the amino acid cycle of coastal waters; uptake and excretion in cultures. Marine Ecology Progress Series, 15:303– 306, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
15/m015p303.pdf.

Anger:1983:CBC

[ALPS83] Klaus Anger, Norbert Laasch, Cornelia Püschel, and Franziska Schorn. Changes in biomass and chemical composition of spider crab (Hyas araneus) larvae reared in the laboratory. Marine Ecology Progress Series, 12:91–101, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/12/m012p091.pdf.

Aksnes:1988:PDA

[AM88] D. L. Aksnes and T. Magnesen. A population dynamics approach to the estimation of production of four calanoid copepods in Lindåspollene, western Norway. Marine Ecology Progress Series, 45:57–68, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/45/m045p057.pdf.

Amat:1983:ZPA

[Ama83] F. Amat. Zygogenetic and parthenogenetic Artemia in Cadiz sea-side salterns. Marine Ecology Progress Series, 13:291–293, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p291.pdf.

Ambrose:1982:SUM

[Amb82] R. F. Ambrose. Shelter utilization by the molluscan cephalopod Octopus bimaculatus. Marine Ecology Progress Series, 7:67–73, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
7/m007p067.pdf.

Ambrose:1984:RPI

[Amb84] W. G. Ambrose, Jr. Role of predatory infauna in structuring marine soft-bottom communities. Marine Ecology Progress Series, 17: 109–115, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/17/m017p109.pdf.

Ambrose:1986:EOP

[Amb86a] R. F. Ambrose. Effects of octopus predation on motile invertebrates in a rocky subtidal community. Marine Ecology Progress Series, 30:261–273, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/30/m030p261.pdf.

Ambrose:1986:IPI

[Amb86b] W. G. Ambrose, Jr. Importance of predatory infauna in marine soft-bottom communities: reply to Wilson. Marine Ecology Progress Series, 32:41–45, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/32/m032p041.pdf.

Andersen:1988:PEM

[AN88] V. Andersen and P. Nival. A pelagic ecosystem model simulating production and sedimentation of biogenic particles: role of salps and copepods. Marine Ecology Progress Series, 44:37–50, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
44/m044p037.pdf.

Andersen:1986:ISD

[And86] F. Ø. Andersen. Influence of Spartina detritus enrichment on exchange of nutrients between sediment and water in an intertidal area of Bay of Fundy. Marine Ecology Progress Series, 29:7–14, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
29/m029p007.pdf.

Anderson:1989:DRB

[And89a] D. J. Anderson. Differential responses of boobies and other seabirds in the Galapagos to the 1986–87 El Niño–Southern Oscillation event. Marine Ecology Progress Series, 52:209–216, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/52/
m052p209.pdf.

Andrew:1989:CEI

[And89b] N. L. Andrew. Contrasting ecological implications of food limitation in sea urchins and herbivorous gastropods. Marine Ecology Progress Series, 51:189–193, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/51/m051p189.pdf.

Anger:1984:DGL

[Ang84] K. Anger. Development and growth in larval and juvenile Hyas coarctatus (Decapoda, Majidae) reared in the laboratory. Marine Ecology Progress Series, 19:115–123, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/19/m019p115.pdf.

Ang:1986:AVS

[Ang86] P. O. Ang, Jr. Analysis of the vegetation structure of a Sargassum community in the Philippines. Marine Ecology Progress Series, 28:9–19, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/28/m028p009.pdf.

Anger:1987:ESC

[Ang87] K. Anger. Energetics of spider crab Hyas araneus megalopa in relation to temperature and the moult cycle. Marine Ecology Progress Series, 36:115–122, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/36/m036p115.pdf.

Anstensrud:1989:VSS

[Ans89] M. Anstensrud. A vital stain for studies of behaviour and ecology of the parasitic copepod Lernaeocera branchialis (Pennellidae). Marine Ecology Progress Series, 53:47–50, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/53/m053p047.pdf.

apRheinallt:1986:SSC

[ap 86] T. ap Rheinallt. Size selection by the crab Liocarcinus puber feeding on mussels Mytilus edulis and on shore crabs Carcinus maenas: the importance of mechanical factors. Marine Ecology Progress Series, 29:45–53, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/29/m029p045.pdf.

Admiraal:1980:ISV

[AP80] W. Admiraal and H. Peletier. Influence of seasonal variations of temperature and light on the growth rate of cultures and natural populations of intertidal diatoms. Marine Ecology Progress Series, 2:35–43, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/2/m002p035.pdf.

Annett:1983:FBP

[AP83] C. Annett and R. Pierotti. Foraging behavior and prey selection of the leather seastar Dermasterias imbricate. Marine Ecology Progress Series, 14:197–206, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/14/m014p197.pdf.

Admiraal:1987:IAN

[ARGL87] W. Admiraal, C. Riaux-Gobin, and R. W. P. M. Laane. Interactions of ammonium, nitrate, and d- and l-amino acids in the nitrogen assimilation of two species of estuarine benthic diatoms. Marine Ecology Progress Series, 40:267–273, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/40/m040p267.pdf.

Armonies:1988:AEM

[Arm88a] W. Armonies. Active emergence of meiofauna from intertidal sediment. Marine Ecology Progress Series, 43:151–159, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/43/m043p151.pdf.

Armonies:1988:PFI

[Arm88b] W. Armonies. Physical factors influencing active emergence of meiofauna from boreal intertidal sediment. Marine Ecology Progress Series, 49:277–286, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/49/m049p277.pdf.

Armonies:1989:OMP

[Arm89] W. Armonies. Occurrence of meiofauna in Phaeocystis seafoam. Marine Ecology Progress Series, 53:305–309, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/53/m053p305.pdf.

Appelbaum:1983:DSO

[AS83] S. Appelbaum and Ch. Schemmel. Dermal sense organs and their significance in the feeding behaviour of the common sole Solea vulgaris. Marine Ecology Progress Series, 13:29–36, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/13/m013p029.pdf.

Anderson:1985:SRT

[AS85] D. M. Anderson and K. D. Stolzenbach. Selective retention of two dinoflagellates in a well-mixed estuarine embayment: the importance of diel vertical migration and surface avoidance. Marine Ecology Progress Series, 25:39–50, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/25/m025p039.pdf.

Andersen:1986:PDT

[AS86] P. Andersen and H. M. Sorensen. Population dynamics and trophic coupling in pelagic microorganisms in eutrophic coastal waters. Marine Ecology Progress Series, 33:99–109, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/33/m033p099.pdf.

Ali:1987:GPA

[AS87] M. H. Ali and S. D. Salman. Growth and production of the amphipod Parhyale basrensis (Talitridae) in the Shatt al-Arab region. Marine Ecology Progress Series, 40:231–238, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/40/m040p231.pdf.

Abdullah:1988:GWT

[AS88] M. I. Abdullah and I. Steffenak. The GEEP workshop: trace metal analyses. Marine Ecology Progress Series, 46:27–30, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/46/
m046p027.pdf.

Asmus:1987:SPI

[Asm87] H. Asmus. Secondary production of an intertidal mussel bed community related to its storage and turnover compartments. Marine Ecology Progress Series, 39:251–266, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/39/m039p251.pdf.

Alcaraz:1988:ETD

[ASMV88] Miguel Alcaraz, Enric Saiz, Celia Marrasé, and Dolors Vaqué. Effects of turbulence on the development of phytoplankton biomass and copepod populations in marine microcosms. Marine Ecology Progress Series, 49:117–125, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/49/m049p117.pdf.

Albright:1987:GCP

[ASSF87] L. J. Albright, E. B. Sherr, B. F. Sherr, and R. D. Fallon. Grazing of ciliated protozoa on free and particle-attached bacteria. Marine Ecology Progress Series, 38:125–129, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/38/m038p125.pdf.

Agnew:1986:EOT

[AT86] D. J. Agnew and A. C. Taylor. Effects of oxygen tension, temperature, salinity, and humidity on the survival of two intertidal gammarid amphipods. Marine Ecology Progress Series, 32:27– 33, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
32/m032p027.pdf.

Auffret:1988:HCR

[Auf88] M. Auffret. Histopathological changes related to chemical contamination in Mytilus edulis from field and experimental conditions. Marine Ecology Progress Series, 46:101–107, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/46/m046p101.pdf.

Admiraal:1987:DNN

[AV87a] W. Admiraal and M. J. W. Veldhuis. Determination of nucleosides and nucleotides in seawater by HPLC; application to phosphatase activity in cultures of the alga Phaeocystis pouchetii. Marine Ecology Progress Series, 36:277–285, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/36/m036p277.pdf.

Azzarello:1987:MBP

[AV87b] M. Y. Azzarello and E. S. Van Vleet. Marine birds and plastic pollution. Marine Ecology Progress Series, 37:295–303, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/37/
m037p295.pdf.

Allison:1986:STP

[AW86] S. K. Allison and K. F. Wishner. Spatial and temporal patterns of zooplankton biomass across the Gulf Stream. Marine Ecology Progress Series, 31:233–244, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/31/m031p233.pdf.

Abel:1983:NNM

[AWSB83] D. J. Abel, W. T. Williams, P. W. Sammarco, and J. S. Bunt. A new numerical model for coral distribution. Marine Ecology Progress Series, 12:257–265, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/12/m012p257.pdf.

Achituv:1985:ACS

[AZ85] Y. Achituv and M. Ziskind. Adaptation of Coenobita scaevola (Forskål) (Crustacea, Anemura) to terrestrial life in desertbordered shore line. Marine Ecology Progress Series, 25:189–197, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
25/m025p189.pdf.

Bell:1981:AFF

[BA81] C. R. Bell and L. J. Albright. Attached and free-floating bacteria in the Fraser River Estuary, British Columbia, Canada. Marine Ecology Progress Series, 6:317–327, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/6/m006p317.pdf.

Bressler:1983:DGS

[BA83] S. L. Bressler and S. I. Ahmed. Detection of glutamine synthetase activity in marine phytoplankton: optimization of the biosynthetic assay. Marine Ecology Progress Series, 14:207–217, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/14/m014p207.pdf.

Barlow:1987:PCP

[BA87] R. G. Barlow and R. S. Alberte. Photosynthetic characteristics of phycoerythrin-containing marine Synechococcus spp. II. Time course responses of photosynthesis to photoinhibition. Marine Ecology Progress Series, 39:191–196, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/39/m039p191.pdf.

Bayne:1988:ACO

[BAC^{+}88a] B. L. Bayne, R. F. Addison, J. M. Capuzzo, K. R. Clarke, J. S. Gray, M. N. Moore, and R. M. Warwick. Appendix 1: Chemistry: organic and trace metal data. Marine Ecology Progress Series, 46:

245–254, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/46/m046p245.pdf.

Bayne:1988:ABS

[BAC^{+}88b] B. L. Bayne, R. F. Addison, J. M. Capuzzo, K. R. Clarke, J. S. Gray, M. N. Moore, and R. M. Warwick. Appendix 2: Biology: sub-lethal responses. Marine Ecology Progress Series, 46:255–262, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p255.pdf.

Bayne:1988:ABB

[BAC^{+}88c] B. L. Bayne, R. F. Addison, J. M. Capuzzo, K. R. Clarke, J. S. Gray, M. N. Moore, and R. M. Warwick. Appendix 3: Biology: benthic community data. Marine Ecology Progress Series, 46:263– 276, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p263.pdf.

Bayne:1988:APG

[BAC^{+}88d] B. L. Bayne, R. F. Addison, J. M. Capuzzo, K. R. Clarke, J. S. Gray, M. N. Moore, and R. M. Warwick. Appendix 4: Participants in the GEEP workshop. Marine Ecology Progress Series, 46:277– 278, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p277.pdf.

Bayne:1988:OGW

[BAC^{+}88e] B. L. Bayne, R. F. Addison, J. M. Capuzzo, K. R. Clarke, J. S. Gray, M. N. Moore, and R. M. Warwick. An overview of the GEEP workshop. Marine Ecology Progress Series, 46:235–243, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p235.pdf.

Ben-Amotz:1980:CMUb

[BAG80a] A. Ben-Amotz and A. Gilboa. Cryopreservation of marine unicellular algae. II. Induction of freezing tolerance. Marine Ecology Progress Series, 2:221–224, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/2/m002p221.pdf.

Ben-Amotz:1980:CMUa

[BAG80b] B. Ben-Amotz and A. Gilboa. Cryopreservation of marine unicellular algae. I. A survey of algae with regard to size, culture, age, photosynthetic activity and chlorophyll-to-cell ratio. Marine Ecology Progress Series, 2:157–161, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/2/m002p157.pdf.

Beckmann:1987:CCC

[BAH87] W. Beckmann, A. Auras, and C. Hemleben. Cyclonic cold-core eddy in the eastern North Atlantic. III. Zooplankton. Marine Ecology Progress Series, 39:165–173, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/39/m039p165.pdf.

Bailey:1981:LTR

[Bai81] K. M. Bailey. Larval transport and recruitment of Pacific hake Merluccius productus. Marine Ecology Progress Series, 6:1–9, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
6/m006p001.pdf.

Bailey:1989:IBV

[Bai89] K. M. Bailey. Interaction between the vertical distribution of juvenile walleye pollock Theragra chalcogramma in the eastern Bering Sea, and cannibalism. Marine Ecology Progress Series, 53:205– 213, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
53/m053p205.pdf.

Bakke:1988:PEL

[Bak88] T. Bakke. Physiological energetics of Littorina littorea under combined pollutant stress in field and mesocosm studies. Marine Ecology Progress Series, 46:123–128, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/46/m046p123.pdf.

Ballantine:1984:HIM

[Bal84] D. L. Ballantine. Hurricane-induced mass mortalities to a tropical subtidal algal community and subsequent recoveries. Marine Ecology Progress Series, 20:75–83, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/20/m020p075.pdf.

Bamstedt:1983:RCZ

[Bam83] U. Bamstedt. RNA concentration in zooplankton: seasonal variation in boreal species. Marine Ecology Progress Series, 11:291– 297, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
11/m011p291.pdf.

Baamstedt:1988:ISD

[Båm88] U. Båmstedt. Interspecific, seasonal and diel variations in zooplankton trypsin and amylase activities in kosterfjorden, western Sweden. Marine Ecology Progress Series, 44:15–24, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/44/m044p015.pdf.

Banse:1982:MSR

[Ban82] K. Banse. Mass-scaled rates of respiration and intrinsic growth in very small invertebrates. Marine Ecology Progress Series, 9:281– 297, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
9/m009p281.pdf.

Boto:1989:DOC

[BAN89] K. G. Boto, D. M. Alongi, and A. L. J. Nott. Dissolved organic carbon-bacteria interactions at sediment–water interface in a tropical mangrove system. Marine Ecology Progress Series, 51:243– 251, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
51/m051p243.pdf.

Barlow:1984:DDP

[Bar84] R. G. Barlow. Dynamics of the decline of a phytoplankton bloom after an upwelling event. Marine Ecology Progress Series, 16:121– 126, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
16/m016p121.pdf.

Barthel:1986:ESH

[Bar86] D. Barthel. On the ecophysiology of the sponge Halichondria panicea in Kiel Bight. I. Substrate specificity, growth and reproduction. Marine Ecology Progress Series, 32:291–298, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/32/m032p291.pdf.

Barange:1988:PSC

[Bar88a] M. Barangé. Prey selection and capture strategies of the benthic hydroid Eudendrium racemosum. Marine Ecology Progress Series, 47:83–88, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/47/m047p083.pdf.

Barthel:1988:ESH

[Bar88b] D. Barthel. On the ecophysiology of the sponge Halichondria panicea in Kiel Bight. II. Biomass, production, energy budget and integration in environmental processes. Marine Ecology Progress Series, 43:87–93, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/43/m043p087.pdf.

Barry:1989:RRM

[Bar89] J. P. Barry. Reproductive response of a marine annelid to winter storms: an analog to fire adaptation in plants? Marine Ecology Progress Series, 54:99–107, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/54/m054p099.pdf.

Batchelder:1986:PBO

[Bat86] H. P. Batchelder. Phytoplankton balance in the oceanic subarctic Pacific: grazing impact of Metridia pacifica. Marine Ecology Progress Series, 34:213–225, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/34/m034p213.pdf.

Bauerfeind:1985:DPD

[Bau85] S. Bauerfeind. Degradation of phytoplankton detritus by bacteria, estimation of bacterial consumption and respiration in an oxygen chamber. Marine Ecology Progress Series, 21:27–36, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/21/
m021p027.pdf.

Bridges:1980:OCO

[BB80] C. R. Bridges and A. R. Brand. Oxygen consumption and oxygenindependence in marine crustaceans. Marine Ecology Progress Series, 2:133–141, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/2/m002p133.pdf.

Blankley:1984:COP

[BB84] W. O. Blankley and G. M. Branch. Co-operative prey capture and unusual brooding habits of Anasterias rupicola (Verrill) (Asteroidea) at sub-Antarctic Marion Island. Marine Ecology Progress Series, 20:171–176, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/20/m020p171.pdf.

Brault:1985:SCE

[BB85] S. Brault and E. Bourget. Structural changes in an estuarine sub-tidal epibenthic community biotic and physical causes. Marine Ecology Progress Series, 21:63–73, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/21/m021p063.pdf.

Bergeron:1986:STS

[BB86] P. Bergeron and E. Bourget. Shore topography and spatial partitioning of crevice refuges by sessile epibenthos in an ice disturbed environment. Marine Ecology Progress Series, 28:129–145, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
28/m028p129.pdf.

Borsheim:1987:CVC

[BB87] K. Y. Borsheim and G. Bratbak. Cell volume to cell carbon conversion factors for a bacterivorous Monas sp. enriched from seawater. Marine Ecology Progress Series, 36:171–175, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/36/m036p171.pdf.

Barkai:1988:ERD

[BB88] A. Barkai and G. M. Branch. Energy requirements for a dense population of rock lobsters Jasus lalandii: novel importance of unorthodox food sources. Marine Ecology Progress Series, 50:83– 96, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
50/m050p083.pdf.

Brunetti:1980:CET

[BBBM80] R. Brunetti, L. Beghi, M. Bressan, and M. G. Marin. Combined effects of temperature and salinity on colonies of Botryllus schlosseri and Botfylloides leachi (Ascidiacea) from the Venetian Lagoon.

Marine Ecology Progress Series, 2:303–314, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/2/m002p303.pdf.

Birt:1987:AHD

[BBG^{+}87] V. L. Birt, T. P. Birt, D. Goulet, D. K. Cairns, and W. A. Montevecchi. Ashmole’s halo: direct evidence for prey depletion by a seabird. Marine Ecology Progress Series, 40:205–208, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
40/m040p205.pdf.

Borchardt:1988:TMC

[BBH^{+}88] T. Borchardt, S. Burchert, H. Hablizel, L. Karbe, and R. Zeitner. Trace metal concentrations in mussels: comparison between estuarine, coastal and offshore regions in the southeastern North Sea from 1983 to 1986. Marine Ecology Progress Series, 42:17– 31, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
42/m042p017.pdf.

Bricelj:1984:AGP

[BBL84] V. M. Bricelj, A. E. Bass, and G. R. Lopez. Absorption and gut passage time of microalgae in a suspension feeder — an evaluation of the ^{51}Cr : ^{14}C twin tracer technique. Marine Ecology Progress Series, 17:57–63, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/17/m017p057.pdf.

Boucher:1988:SMR

[BBR88] G. Boucher and R. Boucher-Rodoni. In situ measurement of respiratory metabolism and nitrogen fluxes at the interface of oyster beds. Marine Ecology Progress Series, 44:229–238, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/44/m044p229.pdf.

Buskey:1989:POC

[BBSS89] Edward J. Buskey, Karen S. Baker, Raymond C. Smith, and Elijah Swift. Photosensitivity of the oceanic cope-pods Pleuromamma gracilis and Pleuromamma xiphias and its relationship to light penetration and daytime depth distribution. Marine Ecology Progress Series, 55:207–216, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/55/m055p207.pdf.

Brown:1981:NNS

[BC81] D. M. Brown and L. Cheng. New net for sampling the ocean surface. Marine Ecology Progress Series, 5:225–227, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/5/m005p225.pdf.

Boudrias:1988:LHP

[BC88] M. A. Boudrias and A. G. Carey, Jr. Life history patterns of Pseudalibrotus litoralis (Crustacea: Amphipoda) on the inner continental shelf, SW Beaufort Sea. Marine Ecology Progress Series, 49: 249–257, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/49/m049p249.pdf.

Bas:1989:EAE

[BCA89] C. Bas and L. E. Calderon-Aguilera. Effect of anthropogenic and environmental factors on the blue whiting Micromesistius poutassou off the Catalonian coast, 1950–1982. Marine Ecology Progress Series, 54:221–228, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/54/m054p221.pdf.

Brady-Champbell:1984:PKL

[BCCH84] M. M. Brady-Champbell, D. B. Campbell, and M. M. Harlin. Productivity of kelp (Laminaria spp.) near the southern limit in the Northwestern Atlantic Ocean. Marine Ecology Progress Series, 18: 79–88, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/18/m018p079.pdf.

Bak:1984:DSU

[BCd84] R. P. M. Bak, M. J. E. Carpay, and E. D. de Ruyter van Steveninck. Densities of the sea urchin Diadema antillarum before and after mass mortalities on the coral reefs of curaçao. Marine Ecology Progress Series, 17:105–108, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/17/m017p105.pdf.

Beviss-Challinor:1982:ABC

[BCF82] M. H. Beviss-Challinor and J. G. Field. Analysis of a benthic community food Web using isotopically labelled potential food. Marine Ecology Progress Series, 9:223–230, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/9/m009p223.pdf.

Breen:1982:CSC

[BCFS82] P. A. Breen, T. A. Carson, J. B. Foster, and E. A. Stewart. Changes in subtidal community structure associated with British Columbia Sea otter transplants. Marine Ecology Progress Series, 7:13–20, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/7/m007p013.pdf.

Bayne:1988:BRP

[BCG88] B. L. Bayne, K. R. Clarke, and J. S. Gray. Background and rationale to a practical workshop on biological effects of pollutants. Marine Ecology Progress Series, 46:1–5, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/46/m046p001.pdf.

Bisol:1984:EGS

[BCS84] P. M. Bisol, R. Costa, and M. Sibuet. Ecological and genetical survey on two deep-sea holothurians: Benthogone rosea and Benthodytes typical. Marine Ecology Progress Series, 15:275–281, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
15/m015p275.pdf.

Beukema:1989:TCT

[Bd89] J. J. Beukema and J. de Vlas. Tidal-current transport of threaddrifting postlarval juveniles of the bivalve Macoma balthica from the Wadden Sea to the North Sea. Marine Ecology Progress Series, 52:193–200, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/52/m052p193.pdf.

Brockmann:1983:CPP

[BDKK83] U. H. Brockmann, E. Dahl, J. Kuiper, and G. Kattner. The concept of POSER (plankton observation with simultaneous enclosures in rosfjorden). Marine Ecology Progress Series, 14:1–8, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
14/m014p001.pdf.

Bricelj:1987:IVR

[BEM87] V. M. Bricelj, J. Epp, and R. E. Malouf. Intraspecific variation in reproductive and somatic growth cycles of bay scallops Argopecten irradians. Marine Ecology Progress Series, 36:123– 137, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
36/m036p123.pdf.

Bengtson:1984:RPM

[Ben84] D. A. Bengtson. Resource partitioning by Menidia menidia and Menidia beryllina (Osteichthyes: Atherinidae). Marine Ecology Progress Series, 18:21–30, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/18/m018p021.pdf.

Berglund:1984:RAT

[Ber84] A. Berglund. Reproductive adaptation in two Palaemon prawn species with differing habitat requirements. Marine Ecology Progress Series, 17:77–83, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/17/m017p077.pdf.

Beukema:1987:IPP

[Beu87] J. J. Beukema. Influence of the predatory polychaete Nephtys hombergii on the abundance of other polychaetes. Marine Ecology Progress Series, 40:95–101, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/40/m040p095.pdf.

Billen:1987:DPD

[BF87] G. Billen and A. Fontigny. Dynamics of a Phaeocystis-dominated spring bloom in Belgian coastal waters. II. Bacterioplankton dynamics. Marine Ecology Progress Series, 37:249–257, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/37/
m037p249.pdf.

Bakke:1988:GWM

[BFMS88] T. Bakke, O. A. Folium, K. A. Moe, and K. Sørensen. The GEEP workshop: mesocosm exposures. Marine Ecology Progress Series, 46:13–18, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/46/m046p013.pdf.

Blomqvist:1987:MLK

[BFP87] S. Blomqvist, A. Frank, and L. R. Petersson. Metals in liver and kidney tissues of autumn-migrating dunlin Calidrisalpina and curlew sandpiper Calidris ferruginea staging at the Baltic Sea. Marine Ecology Progress Series, 35:1–13, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/35/m035p001.pdf.

Bell:1984:ILC

[BG84] J. O. Bell and R. Galzin. Influence of live coral cover on coral-reef fish communities. Marine Ecology Progress Series, 15:265–274, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
15/m015p265.pdf.

Biedka:1987:BAW

[BGD87] R. F. Biedka, J. M. Gosline, and R. E. De Wreede. Biomechanical analysis of wave-induced mortality in the marine alga Pterygophora californica. Marine Ecology Progress Series, 36:163–170, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
36/m036p163.pdf.

Barlow:1988:PSS

[BGL^{+}88] R. G. Barlow, M. Gosselin, L. Legendre, J.-C Therriault, S. Demers, R. F. C. Mantoura, and C. A. Llewellyn. Photoadaptive strategies in sea-ice microalgae. Marine Ecology Progress Series, 45:145–152, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/45/m045p145.pdf.

Bradbury:1986:QPP

[BGR86] R. H. Bradbury, D. G. Green, and R. E. Reichelt. Qualitative patterns and processes in marine ecology: a conceptual programme. Marine Ecology Progress Series, 29:299–304, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/29/m029p299.pdf.

Barber:1988:RFD

[BGSS88] B. J. Barber, R. Getchell, S. Shumway, and D. Schick. Reduced fecundity in a deep-water population of the giant scallop Placopecten magellanicus in the Gulf of Maine, USA. Marine Ecology Progress Series, 42:207–212, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/42/m042p207.pdf.

Baird:1981:TFVa

[BH81a] R. C. Baird and T. L. Hopkins. Trophodynamics of the fish Valenciennellus tripunctulatus. II. Selectivity, grazlng rates and resource utilization. Marine Ecology Progress Series, 5:11–19, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
5/m005p011.pdf.

Baird:1981:TFVb

[BH81b] R. C. Baird and T. L. Hopkins. Trophodynamics of the fish Valenciennellus tripunctulatus. III. Energetics, resources and feeding strategy. Marine Ecology Progress Series, 5:21–28, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/5/m005p021.pdf.

Bienfang:1983:CVS

[BH83a] P. K. Bienfang and P. J. Harrison. Co-variation of sinking rate and cell quota among nutrient replete marine phytoplankton. Marine Ecology Progress Series, 14:297–300, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/14/m014p297.pdf.

Brockmann:1983:SES

[BH83b] U. H. Brockmann and G. Hentzschel. Samplers for enclosed stratified water columns. Marine Ecology Progress Series, 14:107–109, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
14/m014p107.pdf.

Benner:1985:MDL

[BH85] R. Benner and R. E. Hodson. Microbial degradation of the leachable and lignocellulosic components of leaves and wood from Rhizophora mangle in a tropical mangrove swamp. Marine Ecology Progress Series, 23:221–230, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/23/m023p221.pdf.

Bosman:1986:SGD

[BH86] A. L. Bosman and P. A. R. Hockey. Seabird guano as a determinant of rocky intertidal community structure. Marine Ecology Progress Series, 32:247–257, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/32/m032p247.pdf.

Breeman:1987:VPR

[BH87] A. M. Breeman and B. W. Hoeksema. Vegetative propagation of the red alga Rhodochorton purpureum by means of fragments that escape digestion by herbivores. Marine Ecology Progress Series, 35:197–201, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/35/m035p197.pdf.

Borrero:1988:TVS

[BH88] F. J. Borrero and T. J. Hilbish. Temporal variation in shell and soft tissue growth of the mussel Geukensia demissa. Marine Ecology Progress Series, 42:9–15, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/42/m042p009.pdf.

Borsheim:1989:FCC

[BHJ^{+}89] K. Y. Børsheim, T. Harboe, T. Johnsen, S. Norland, and K. Nygaard. Flow cytometric characterization and enumeration of Chrysochromulina polylepis during a bloom along the Norwegian coast. Marine Ecology Progress Series, 54:307–309, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/54/m054p307.pdf.

Bayne:1989:ESC

[BHNI89] B. L. Bayne, A. J. S. Hawkins, E. Navarro, and I. P. Iglesias. Effects of seston concentration on feeding, digestion and growth in the mussel Mytilus edulis. Marine Ecology Progress Series, 55:47– 54, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
55/m055p047.pdf.

Baker:1986:MMP

[BHP^{+}86] C. S. Baker, L. M. Herman, A. Perry, W. S. Lawton, J. M. Straley, A. A. Wolman, G. D. Kaufman, H. E. Winn, J. D. Hall, J. M. Reinke, and J. Ostman. Migratory movement and population structure of humpback whales (Megaptera novaeanglieae) in the central and eastern North Pacific. Marine Ecology Progress Series, 31:105–119, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/31/m031p105.pdf.

Biddanda:1984:MSM

[Bid84] B. A. Biddanda. Microbial synthesis of macroparticulate matter. Marine Ecology Progress Series, 20:241–251, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/20/m020p241.pdf.

Biddanda:1988:MADb

[Bid88] B. A. Biddanda. Microbial aggregation and degradation of phytoplankton-derived detritus in seawater. II. Microbial metabolism. Marine Ecology Progress Series, 42:89–95, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/42/
m042p089.pdf.

Bienfang:1985:SSS

[Bie85] P. K. Bienfang. Size structure and sinking rates of various microparticulate constituents in oligotrophic Hawaiian waters. Marine Ecology Progress Series, 23:143–151, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/23/m023p143.pdf.

Bjerregaard:1988:ESC

[Bje88] P. Bjerregaard. Effect of selenium on cadmium uptake in selected benthic invertebrates. Marine Ecology Progress Series, 48:17–28, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
48/m048p017.pdf.

Bjornsen:1986:BGY

[Bjø86] P. K. Bjørnsen. Bacterioplankton growth yield in continuous seawater cultures. Marine Ecology Progress Series, 30:191–196, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
30/m030p191.pdf.

Boehlert:1980:VGL

[BK80] G. W. Boehlert and R. F. Kappenman. Variation of growth with latitude in two species of rockfish (Sebastes pinniger and S. diploproa) from the Northeast Pacific Ocean. Marine Ecology Progress Series, 3:1–10, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/3/m003p001.pdf.

Blake:1981:LDL

[BK81] J. A. Blake and J. D. Kudenov. Larval development, larval nutrition and growth for two Boccardia species (Polychaeta: Spionidae) from Victoria, Australia. Marine Ecology Progress Series, 6:175– 182, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
6/m006p175.pdf.

Burkill:1982:PCE

[BK82] P. H. Burkill and T. F. Kendall. Production of the copepod Eurytemora affinis in the Bristol Channel. Marine Ecology Progress Series, 7:21–31, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/7/m007p021.pdf.

Boynton:1985:NRO

[BK85a] W. R. Boynton and W. M. Kemp. Nutrient regeneration and oxygen consumption by sediments along an estuarine salinity gradient. Marine Ecology Progress Series, 23:45–55, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/23/m023p045.pdf.

Brewer:1985:DVD

[BK85b] G. D. Brewer and G. S. Kleppel. Diel vertical distribution of fish larvae their prey in nearshore waters of southern California. Marine Ecology Progress Series, 27:217–226, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/27/m027p217.pdf.

Bonotto:1980:MRC

[BL80] S. Bonotto and A. Lüttke. Method for rapid counting of sporangia in the green alga Batophora oersstedii. Marine Ecology Progress Series, 3:75–77, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/3/m003p075.pdf.

Bak:1987:ABB

[BL87a] R. P. M. Bak and R. W. P. M. Laane. Annual black bands in skeletons of reef corals (Scleractinia). Marine Ecology Progress Series, 38:169–175, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/38/m038p169.pdf.

Benayahu:1987:LTR

[BL87b] Y. Benayahu and Y. Loya. Long-term recruitment of soft-corals (Octocorallia: Alcyonacea) on artificial substrata at Eilat (Red Sea). Marine Ecology Progress Series, 38:161–167, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/38/m038p161.pdf.

Barshaw:1988:PUP

[BL88] D. E. Barshaw and K. L. Lavalli. Predation upon postlarval lobsters Homarus americanus by cunners Tautogolabrus adspersus and mud crabs Neopanope sayi on three different substrates: eelgrass, mud and rocks. Marine Ecology Progress Series, 48:119– 123, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
48/m048p119.pdf.

Blankley:1984:ESA

[Bla84] W. O. Blankley. Ecology of the starfish Anasterias rupicola at Marion Island (Southern Ocean). Marine Ecology Progress Series, 18:131–137, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/18/m018p131.pdf.

Blackburn:1988:CMS

[BLK88] T. H. Blackburn, B. Aa Lund, and M. D. Krom. C- and n-mineralization in the sediments of earthen marine fishponds. Marine Ecology Progress Series, 44:221–227, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/44/m044p221.pdf.

Bloom:1980:ISS

[Blo80] S. A. Bloom. An intertidal soft-sediment avian exclosure which minimizes sediment alterations. Marine Ecology Progress Series, 3:79–81, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/3/m003p079.pdf.

Bloom:1981:SIC

[Blo81] S. A. Bloom. Similarity indices in community studies: Potential pitfalls. Marine Ecology Progress Series, 5:125–128, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/5/m005p125.pdf.

Bergstrom:1987:URO

[BLP87] B. I. Bergström, J. Larsson, and J.-O. Pettersson. Use of a remotely operated vehicle (ROV) to study marine phenomena: I. Pandalid shrimp densities. Marine Ecology Progress Series, 37:97– 101, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
37/m037p097.pdf.

Bour:1986:TMR

[BLR86] W. Bour, L. Loubersac, and P. Rual. Thematic mapping of reefs by processing of simulated SPOT satellite data: application to the Trochus niloticus biotope on Tetembia Reef (New Caledonia). Marine Ecology Progress Series, 34:243–249, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/34/m034p243.pdf.

Banse:1986:WBP

[BM86a] K. Banse and C. R. McClain. Winter blooms of phytoplankton in the Arabian Sea as observed by the coastal zone color scanner. Marine Ecology Progress Series, 34:201–211, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/34/m034p201.pdf.

Brooks:1986:ICS

[BM86b] W. R. Brooks and R. N. Mariscal. Interspecific competition for space by hydroids and a sea anemone living on gastropod shells inhabited by hermit crabs. Marine Ecology Progress Series, 28: 241–244, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/28/m028p241.pdf.

Brunetti:1988:MTE

[BMGB88] R. Brunetti, F. Majone, I. Gola, and C. Beltrame. The micronucleus test: examples of application to marine ecology. Marine Ecology Progress Series, 44:65–68, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/44/m044p065.pdf.

Bryant:1984:ETS

[BMRN84] V. Bryant, D. S. McLusky, K. Roddie, and D. M. Newbery. Effect of temperature and salinity on the toxicity of chromium to three estuarine invertebrates (Corophium volutator, Macoma balthica, Nereis diversicolor). Marine Ecology Progress Series, 20:137–149, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
20/m020p137.pdf.

Blum:1988:ABF

[BMZZ88] L. K. Blum, A. L. Mills, J. C. Zieman, and R. T. Zieman. Abundance of bacteria and fungi in seagrass and mangrove detritus. Marine Ecology Progress Series, 42:73–78, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/42/m042p073.pdf.

Buxton:1981:RSA

[BNF81] C. D. Buxton, R. C. Newell, and J. G. Field. Response-surface analysis of the combined effects of exposure and acclimation temperatures on filtration, oxygen consumption and scope for growth in the oyster Ostrea edulis. Marine Ecology Progress Series, 6:73– 82, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
6/m006p073.pdf.

Bihari:1989:STA

[BNF^{+}89] N. Bihari, M. Najdek, R. Floris, R. Batel, and R. K. Zahn. Sediment toxicity assessment using bacterial bioluminescence: effect of an unusual phytoplankton bloom. Marine Ecology Progress Series, 57:307–310, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/57/m057p307.pdf.

Bathmann:1987:CFP

[BNVP87] U. V. Bathmann, T. T. Noji, M. Voss, and R. Peinert. Copepod fecal pellets: abundance, sedimentation and content at a permanent station in the Norwegian Sea in May/June 1986. Marine Ecology Progress Series, 38:45–51, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/38/m038p045.pdf.

Boal:1980:PHS

[Boa80] J. Boal. Pacific harbor seal (Phoca vitulina richardii). Haul out impact on the rocky midtidal zone. Marine Ecology Progress Series, 2:265–269, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/2/m002p265.pdf.

Bodin:1988:REM

[Bod88] P. Bodin. Results of ecological monitoring of three beaches polluted by the ‘Amoco Cadiz’ oil spill: development of meiofauna from 1978 to 1984. Marine Ecology Progress Series, 42:105–123, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
42/m042p105.pdf.

Bouchon:1981:QSS

[Bou81] C. Bouchon. Quantitative study of the scleractinian coral communities of a fringing reef of Reunion Island (Indian Ocean). Marine Ecology Progress Series, 4:273–288, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p273.pdf.

Boyden:1981:SVI

[BP81] C. R. Boyden and D. J. H. Phillips. Seasonal variation and inherent variability of trace elements in oysters and their lmplications for indicator studies. Marine Ecology Progress Series, 5:29–40, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
5/m005p029.pdf.

Bates:1984:FIM

[BP84a] S. S. Bates and T. Piatt. Fluorescence induction as a measure of photosynthetic capacity in marine phytoplankton: response of Thaiassiosira pseudonana (Baciliariophyceae) and Dunalielta tertiotecta (Chlorophyceae). Marine Ecology Progress Series, 18:67– 77, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
18/m018p067.pdf.

Blomqvist:1984:CPS

[BP84b] S. Blomqvist and M. Peterz. Cylones and pelagic seabird movements. Marine Ecology Progress Series, 20:85–92, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/20/m020p085.pdf.

Bates:1985:FIC

[BP85] S. S. Bates and T. Piatt. Fluorescence induction of chlorophyll-a in the Sargasso Sea and on the Grand Banks: correlation with photosynthetic capacity. Marine Ecology Progress Series, 27:29– 38, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
27/m027p029.pdf.

Boulding:1986:VPR

[BP86] E. G. Boulding and T. Piatt. Variation in photosynthetic rates among individual cells of a marine dinoflagellate. Marine Ecology Progress Series, 29:199–203, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/29/m029p199.pdf.

Biddanda:1988:MADa

[BP88] B. A. Biddanda and L. R. Pomeroy. Microbial aggregation and degradation of phytoplankton-derived detritus in seawater. I. Microbial succession. Marine Ecology Progress Series, 42:79–88, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
42/m042p079.pdf.

Baker:1987:RHF

[BPH87] C. S. Baker, A. Perry, and L. M. Herman. Reproductive histories of female humpback whales Megaptera novaeangliae in the North Pacific. Marine Ecology Progress Series, 41:103–114, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/41/m041p103.pdf.

Bradbury:1982:FDC

[BR82] R. H. Bradbury and R. E. Reichelt. Fractal dimension of a coral reef at ecological scales. Marine Ecology Progress Series, 10:169– 171, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/

10/m010p169.pdf. See discussion [Mar83a].

Bottrell:1983:SVL

[BR83] H. H. Bottrell and D. B. Robins. Seasonal variations in length, dry weight, carbon and nitrogen of Calanus helgolandicus from the Celtic Sea. Marine Ecology Progress Series, 14:259–268, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/14/
m014p259.pdf.

Button:1985:ETE

[BR85] D. K. Button and B. R. Robertson. Effect of toluene exposure time and concentration on induction of high affinity values for toluene oxidation by bacteria of estuarine seawater samples. Marine Ecology Progress Series, 26:187–193, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/26/m026p187.pdf.

Brandt:1981:EWC

[Bra81] S. B. Brandt. Effects of a warm-core eddy on fish distributions in the Tasman Sea of East Australia. Marine Ecology Progress Series, 6:19–33, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/6/m006p019.pdf.

Bratbak:1987:CFE

[Bra87] G. Bratbak. Carbon flow in an experimental microbial ecosystem. Marine Ecology Progress Series, 36:267–276, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/36/m036p267.pdf.

Brenchley:1982:PEL

[Bre82] G. A. Brenchley. Predation on encapsulated larvae by adults: Effects of introduced species on the gastropod Ilyanassa obsoleta. Marine Ecology Progress Series, 9:255–262, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/9/m009p255.pdf.

Brey:1986:IMA

[Bre86] T. Brey. Increase in macrozoobenthos above the halocline in Kiel Bay comparing the 1960s with the 1980s. Marine Ecology Progress Series, 28:299–302, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/28/m028p299.pdf.

Bradbury:1983:FEM

[BRG83] R. H. Bradbury, R. E. Reichelt, and D. G. Green. Fractals in ecology: methods and interpretation. Marine Ecology Progress Series, 14:295–296, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/14/m014p295.pdf.

Briggs:1985:DFS

[Bri85] K. B. Briggs. Deposit feeding by some deep-sea megabenthos from the Venezuela Basin. selective or non-selective. Marine Ecology Progress Series, 21:127–134, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/21/m021p127.pdf.

Brock:1980:END

[Bro80] V. Brock. Evidence for niche differences in sympatric populations of Cerastoderma edule and C. lamarcki. Marine Ecology Progress Series, 2:75–80, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/2/m002p075.pdf.

Broom:1982:AGA

[Bro82] M. J. Broom. Analysis of the growth of Anadara granosa (Bivalvia: Arcidae) in natural, artificially seeded and experimental populations. Marine Ecology Progress Series, 9:69–79, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/9/m009p069.pdf.

Brown:1986:FIB

[Bro86] S. L. Brown. Feces of intertidal benthic invertebrates: influence of particle selection in feeding ace element concentration. Marine Ecology Progress Series, 28:219–231, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/28/m028p219.pdf.

Brown:1988:MAR

[Bro88] J. R. Brown. Multivariate analyses of the role of environmental factors in seasonal and site-related growth variation in the Pacific oyster Crassostrea gigas. Marine Ecology Progress Series, 45:225– 236, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
45/m045p225.pdf.

Berglund:1986:RSR

[BRS86] A. Berglund, G. Rosenqvist, and I. Svensson. Reversed sex roles and parental energy investment in zygotes of two pipefish (Syngnathidae) species. Marine Ecology Progress Series, 29:209–215, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
29/m029p209.pdf.

Badino:1980:PIV

[BS80a] G. Badino and G. Sella. Phosphoglucose isomerase variability in sympatric populations of Mediterranean species of Patella (Gastropoda, Prosobranchiata). Marine Ecology Progress Series, 2: 315–320, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/2/m002p315.pdf.

Bell:1980:FIM

[BS80b] S. S. Bell and K. M. Sherman. A field investigation of meiofaunal dispersal: Tidal resuspension and implications. Marine Ecology Progress Series, 3:245–249, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/3/m003p245.pdf.

Berry:1980:EPC

[BS80c] P. F. Berry and M. J. Smale. An estimate of production and consumption rates in the spiny lobster Panulirus homarus on a shallow littoral reef off the Natal Coast, South Africa. Marine Ecology Progress Series, 2:337–343, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/2/m002p337.pdf.

Bulnheim:1980:EGD

[BS80d] H.-P. Bulnheim and A. Scholl. Evidence of genetic divergence between two brackish water gammaridean sibling species. Marine Ecology Progress Series, 3:163–165, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/3/m003p163.pdf.

Bell:1983:PDR

[BS83a] J. L. Bell and S. E. Stancyk. Population dynamics and reproduction of Dissodactylus mellitae (Brachyura: Pinnotheridae) on its sand dollar host Mellita quinquiesperforata (Echinodermata).

Marine Ecology Progress Series, 13:141–149, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/13/m013p141.pdf.

Berry:1983:BMP

[BS83b] P. F. Berry and M. H. Schleyer. The brown mussel Perna perna on the Natal coast, South Africa: utilization of available food and energy budget. Marine Ecology Progress Series, 13:201–210, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p201.pdf.

Bottger-Schnack:1989:BLF

[BS89a] R. Böttger-Schnack. Body length of female Macrosetella gracilis (Copepoda: Harpacticoida) from various depth zones in the Red Sea. Marine Ecology Progress Series, 52:33–37, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/52/m052p033.pdf.

Brightman:1989:PIR

[BS89b] R. I. Brightman and W. O. Smith, Jr. Photosynthesis-irradiance relationships of Antarctic phytoplankton during austral winter. Marine Ecology Progress Series, 53:143–151, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/53/m053p143.pdf.

Boto:1985:RNN

[BSC85] K. Boto, P. Saffigna, and B. Clough. Role of nitrate in nitrogen nutrition of the mangrove Avicennia marina. Marine Ecology Progress Series, 21:259–265, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/21/m021p259.pdf.

Bologa:1984:ACP

[BSF84] A. S. Bologa, H. V. Skolka, and P. T. Frangopol. Annual cycle of planktonic primary productivity off the Romanian Black Sea coast. Marine Ecology Progress Series, 19:25–32, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/19/m019p025.pdf.

Bidigare:1989:IZQ

[BSP89] R. R. Bidigare, O. Schofield, and B. B. Prezelin. Influence of zeaxanthin on quantum yield of photosynthesis of Synechococcus clone WH7803 (DC2). Marine Ecology Progress Series, 56:177– 188, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
56/m056p177.pdf.

Bottger-Schnack:1989:VDP

[BSS89] R. Böttger-Schnack and D. Schnack. Vertical distribution and population structure of Macrosetella gracilis (Copepoda: Harpacticoida) in the Red Sea in relation to the occurrence of Oscillatoria (Trichodesmium) spp. (Cyanobacteria). Marine Ecology Progress Series, 52:17–31, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/52/m052p017.pdf.

Bhosle:1989:SPM

[BSSW89] N. B. Bhosle, S. S. Sawant, P. D. Sankaran, and A. B. Wagh. Sedimentation of particulate material in stratified and nonstratified water columns in the Bombay High area of the Arabian Sea. Marine Ecology Progress Series, 57:225–236, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/57/m057p225.pdf.

Bak:1981:ETC

[BSv81] R. P. M. Bak, J. Sybesma, and F. C. van Duyl. The ecology of the tropical compound ascidian Trididemnum solidum. II. Abundance, growth and survival. Marine Ecology Progress Series, 6:43–52, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
6/m006p043.pdf.

Bratbak:1985:PBI

[BT85] G. Bratbak and T. F. Thingstad. Phytoplankton-bacteria interactions: an apparant paradox? Analysis of a model system with both competition and commensalism. Marine Ecology Progress Series, 25:23–30, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/25/m025p023.pdf.

Bayne:1988:PMN

[BT88] B. L. Bayne and F. P. Thurberg. Physiological measurements on Nucula tenuis and on isolated gills of Mytilus edulis and Carcinus maenas. Marine Ecology Progress Series, 46:129–134, 1988.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/46/
m046p129.pdf.

Bjerregaard:1985:BAP

[BTFF85] P. Bjerregaard, S. Topçuolu, N. S. Fisher, and S. W. Fowler. Biokinetics of americium and plutonium in the mussel Mytilus edulis. Marine Ecology Progress Series, 21:99–111, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/21/m021p099.pdf.

Buckley:1984:ETF

[BTH^{+}84] L. J. Buckley, S. I. Turner, T. A. Halavik, A. S. Smigielski, S. M. Drew, and G. C. Laurence. Effects of temperature and food availability on growth, survival, and RNA–DNA ratio of larval sand lance (Ammodytes americanus). Marine Ecology Progress Series, 15:91–97, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/15/m015p091.pdf.

Bach:1986:EDE

[BTL86] S. D. Bach, G. W. Thayer, and M. W. LaCroix. Export of detritus from eelgrass (Zostera marina) beds near Beaufort, North Carolina, USA. Marine Ecology Progress Series, 28:265–278, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
28/m028p265.pdf.

Blot:1988:GRA

[BTQS88] M. Blot, C. Thiriot-Quiévreux, and J. Soyer. Genetic relationships among populations of Mytilus desolationis from Kerguelen, M. edulis from the North Atlantic and M. gallo-provincialis from the Mediterranean. Marine Ecology Progress Series, 44:239–247, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
44/m044p239.pdf.

Buckley:1982:ETG

[Buc82] L. J. Buckley. Effects of temperature on growth and biochemical composition of larval winter flounder Pseudopleuronectes americanus. Marine Ecology Progress Series, 8:181–186, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/8/m008p181.pdf.

Buesa:1987:GRT

[Bue87] R. J. Buesa. Growth rate of tropical demersal fishes. Marine Ecology Progress Series, 36:191–199, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/36/m036p191.pdf.

Bull:1982:SCC

[Bul82] G. D. Bull. Scleractinian coral communities of two inshore high island fringing reefs at Magnetic Island, North Queensland. Marine Ecology Progress Series, 7:267–272, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/7/m007p267.pdf.

Berge:1983:EEM

[BV83] J. A. Berge and V. A. Valderhaug. Effect of epibenthic macropredators on community structure in subtidal organically enriched sediments in the inner Oslofjord. Marine Ecology Progress Series, 11:15–22, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/11/m011p015.pdf.

Bjerregaard:1985:ECH

[BV85] P. Bjerregaard and T. Vislie. Effects of cadmium on hemolymph composition in the shore crab Carcinus maenas. Marine Ecology Progress Series, 27:135–142, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/27/m027p135.pdf.

Berry:1982:DBI

[BvdEH^{+}82] P. F. Berry, R. P. van der Elst, P. Hanekom, C. S. W. Joubert, and M. J. Smale. Density and biomass of the ichthyofauna of a Natal littoral reef. Marine Ecology Progress Series, 10:49–55, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
10/m010p049.pdf.

Bayne:1980:GPM

[BW80] B. L. Bayne and C. M. Worrall. Growth and production of mussels Mytilus edulis from two populations. Marine Ecology Progress Series, 3:317–328, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/3/m003p317.pdf.

Bunt:1981:VRM

[BW81] J. S. Bunt and W. T. Williams. Vegetational relationships in the mangroves of Tropical Australia. Marine Ecology Progress Series, 4:349–359, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/4/m004p349.pdf.

Boto:1983:PNN

[BW83] K. G. Boto and J. T. Wellington. Phosphorus and nitrogen nutritional status of a northern Australian mangrove forest. Marine Ecology Progress Series, 11:63–69, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/11/m011p063.pdf.

Boto:1988:SVC

[BW88] K. G. Boto and J. T. Wellington. Seasonal variations in concentrations and fluxes of dissolved organic and inorganic materials in a tropical, tidally-dominated, mangrove waterway. Marine Ecology Progress Series, 50:151–160, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/50/m050p151.pdf.

Bell:1987:HUH

[BWH87] S. S. Bell, K. Walters, and M. O. Hall. Habitat utilization by harpacticoid copepods: a morphometric approach. Marine Ecology Progress Series, 35:59–64, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/35/m035p059.pdf.

Bradbury:1981:EMF

[BY81] R. H. Bradbury and P. C. Young. The effects of a major forcing function, wave energy, on a coral reef ecosystem. Marine Ecology Progress Series, 5:229–241, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/5/m005p229.pdf.

Bradbury:1983:CIC

[BY83] R. H. Bradbury and P. C. Young. Coral interactions and community structure: an analysis of spatial pattern. Marine Ecology Progress Series, 11:265–271, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/11/m011p265.pdf.

Carrasco:1983:LHP

[CA83] F. D. Carrasco and D. F. Arcos. Life history and production of a cold-temperate population of the sublittoral amphipod Ampelisca araucana. Marine Ecology Progress Series, 14:245–252, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
14/m014p245.pdf.

Commito:1985:MTL

[CA85] J. A. Commito and W. G. Ambrose, Jr. Multiple trophic levels in soft-bottom communities. Marine Ecology Progress Series, 26: 289–293, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/26/m026p289.pdf.

Clayton:1986:DGS

[CA86] J. R. Clayton, Jr. and S. I. Ahmed. Detection of glutamate synthase (GOGAT) activity in phytoplankton: evaluation of cofactors and assay optimization. Marine Ecology Progress Series, 32:115– 122, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
32/m032p115.pdf.

Clayton:1987:MAP

[CA87] J. R. Clayton, Jr. and S. I. Ahmed. Modified assay procedure for enhanced sensitivity of in vitro glutamine synthetase (GS) activity measurements in marine phytoplankton. Marine Ecology Progress Series, 36:177–180, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/36/m036p177.pdf.

Cammen:1982:EPS

[Cam82] L. M. Cammen. Effect of particle size on organic content and microbial abundance within four marine sediments. Marine Ecology Progress Series, 9:273–280, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/9/m009p273.pdf.

Cammen:1985:MLO

[Cam85] L. M. Cammen. Metabolic loss of organic carbon by the polychaete Capitella capitata (Fabricius) estimated from initial weight decrease during starvation, oxygen uptake, and release of ^{14}C by uniformly-labeled animals. Marine Ecology Progress Series, 21: 163–167, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/21/m021p163.pdf.

Capone:1982:NFA

[Cap82] D. G. Capone. Nitrogen fixation (acetylene reduction) by rhizosphere sediments of the eelgrass Zostera marina. Marine Ecology Progress Series, 10:67–75, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/10/m010p067.pdf.

Capuzzo:1988:PEPa

[Cap88a] J. M. Capuzzo. Physiological effects of a pollutant gradient introduction MEPS 46:111. Marine Ecology Progress Series, ??: ??, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p111.pdf.

Capuzzo:1988:PEPb

[Cap88b] J. M. Capuzzo. Physiological effects of a pollutant gradient summary. Marine Ecology Progress Series, 46:147–148, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/46/m046p147.pdf.

Carter:1982:PBP

[Car82] R. A. Carter. Phytoplankton biomass and production in a Southern Benguela kelp bed system. Marine Ecology Progress Series, 8:9–14, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/8/m008p009.pdf.

Carey:1987:PFB

[Car87] A. G. Carey, Jr. Particle flux beneath fast ice in the shallow southwestern Beaufort Sea, Arctic Ocean. Marine Ecology Progress Series, 40:247–257, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/40/m040p247.pdf.

Castilla:1988:HER

[CB88] Juan C. Castilla and Rodrigo H. Bustamante. Human exclusion from rocky intertidal of Las Cruces, central Chile: effects on Durvillaea antarctica (Phaeophyta, Durvilleales). Marine Ecology Progress Series, 50:203–214, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/50/m050p203.pdf.

Capriulo:1982:ASC

[CC82] G. M. Capriulo and E. J. Carpenter. Abundance, species composition and feeding impact of tintinnid micro-zooplankton in Central Long Island Sound. Marine Ecology Progress Series, 10:277–288, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
10/m010p277.pdf.

Cole:1984:SBL

[CC84] B. E. Cole and J. E. Cloern. Significance of biomass and light availability to phytoplankton productivity in San Francisco Bay. Marine Ecology Progress Series, 17:15–24, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/17/m017p015.pdf.

Campbell:1986:DPC

[CC86a] L. Campbell and E. J. Carpenter. Diel patterns of cell division in marine Synechococcus spp. (Cyanobacteria): use of the frequency of dividing cells technique to measure growth rate. Marine Ecology Progress Series, 32:139–148, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/32/m032p139.pdf.

Campbell:1986:EGP

[CC86b] L. Campbell and E. J. Carpenter. Estimating the grazing pressure of heterotrophic nanoplankton on Synechococcus spp. using the sea water dilution and selective inhibitor techniques. Marine Ecology Progress Series, 33:121–129, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/33/m033p121.pdf.

Cole:1987:EME

[CC87a] B. E. Cole and J. E. Cloern. An empirical model for estimating phytoplankton productivity in estuaries. Marine Ecology Progress Series, 36:299–305, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/36/m036p299.pdf.

Contreras:1987:FBM

[CC87b] S. Contreras and J. C. Castilla. Feeding behavior and morphological adaptations in two sympatric sea urchin species in central Chile. Marine Ecology Progress Series, 38:217–224, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/38/m038p217.pdf.

Carpenter:1988:DPC

[CC88a] E. J. Carpenter and L. Campbell. Diel patterns of cell division and growth rates of Synechococcus spp. in Long Island Sound. Marine Ecology Progress Series, 47:179–183, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/47/m047p179.pdf.

Carpenter:1988:SSP

[CC88b] E. J. Carpenter and J. Chang. Species-specific phytoplankton growth rates via diel DNA synthesis cycles. I. Concept of the method. Marine Ecology Progress Series, 43:105–111, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/43/
m043p105.pdf.

Chang:1988:SSP

[CC88c] J. Chang and E. J. Carpenter. Species-specific phytoplankton growth rates via diel DNA synthesis cycles. II DNA quantification and model verification in the dino-flagellate Heterocapsa triquetra. Marine Ecology Progress Series, 44:287–296, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/44/m044p287.pdf.

Craven:1989:SFM

[CC89] D. B. Craven and A. F. Carlucci. Spring and fall microheterotrophic utilization of dissolved free amino acids in a California borderland basin. Marine Ecology Progress Series, 51:229– 235, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
51/m051p229.pdf.

Colijn:1981:SCB

[CD81] F. Colijn and K. S. Dijkema. Species composition of benthic diatoms and distribution of chlorophyll-a on an lntertidal flat in the Dutch Wadden Sea. Marine Ecology Progress Series, 4:9–21, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
4/m004p009.pdf.

Colijn:1983:PPM

[Cd83] F. Colijn and V. N. de Jonge. Primary production of microphytobenthos in the Ems–Dollard Estuary. Marine Ecology Progress Series, 14:185–196, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/14/m014p185.pdf.

Cooper:1988:SCI

[CD88] L. W. Cooper and M. J. DeNiro. Stable carbon isotope variability in the seagrass Posidonia oceanica: evidence for light intensity effects. Marine Ecology Progress Series, 50:225–229, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/50/m050p225.pdf.

Chapman:1987:SMS

[CDL87] P. M. Chapman, R. N. Dexter, and E. R. Long. Synoptic measures of sediment contamination, toxicity and infaunal community composition (the sediment quality triad) in San Francisco Bay. Marine Ecology Progress Series, 37:75–96, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/37/m037p075.pdf.

Clancy:1989:DCL

[CE89] M. Clancy and C. E. Epifanio. Distribution of crab larvae in relation to tidal fronts in Delaware Bay, USA. Marine Ecology Progress Series, 57:77–82, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/57/m057p077.pdf.

Colby:1984:PDS

[CF84] D. R. Colby and M. S. Fonseca. Population dynamics, spatial dispersion and somatic growth of the sand fiddler crab Uca pugilator. Marine Ecology Progress Series, 16:269–279, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/16/m016p269.pdf.

Cary:1989:ORB

[CFH89] S. Craig Cary, Horst Felbeck, and Nicholas D. Holland. Observations on the reproductive biology of the hydrothermal vent tube worm Riftia pachyptila. Marine Ecology Progress Series, 52:89– 94, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
52/m052p089.pdf.

Cole:1988:BPF

[CFP88] J. J. Cole, S. Findlay, and M. L. Pace. Bacterial production in fresh and saltwater ecosystems: a cross-system overview. Marine Ecology Progress Series, 43:1–10, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/43/m043p001.pdf.

Claustre:1987:ABC

[CG87] H. Claustre and J. Gostan. Adaptation of biochemical composition and cell size to irradiance in two microalgae: possible ecological implications. Marine Ecology Progress Series, 40:167–174, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
40/m040p167.pdf.

Clarke:1988:SDA

[CG88] K. R. Clarke and R. H. Green. Statistical design and analysis for a ’biological effects’ study. Marine Ecology Progress Series, 46:213– 226, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p213.pdf.

Cunningham:1984:LPS

[CH84] P. N. Cunningham and R. N. Hughes. Learning of predatory skills by shore-crabs Carcinus maenas feeding on mussels and dogwhelks. Marine Ecology Progress Series, 16:21–26, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/16/m016p021.pdf.

Cheshire:1988:BDN

[CH88] Anthony C. Cheshire and Neil D. Hallam. Biomass and density of native stands of Durvillaea potatorum (southern bullkelp) in south eastern Australia. Marine Ecology Progress Series, 48:277– 283, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
48/m048p277.pdf.

Cota:1989:PCA

[CH89] G. F. Cota and E. P. W. Home. Physical control of Arctic ice algal production. Marine Ecology Progress Series, 52:111–121, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
52/m052p111.pdf.

Charnov:1982:ALH

[Cha82] E. L. Charnov. Alternative life-histories in protogynous fishes: a general evolutinonary theory. Marine Ecology Progress Series, 9: 305–307, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/9/m009p305.pdf.

Chan:1988:ATC

[Cha88] H. M. Chan. Accumulation and tolerance to cadmium, copper, lead and zinc by the green mussel Perna viridis. Marine Ecology Progress Series, 48:295–303, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/48/m048p295.pdf.

Crossland:1983:DNH

[CHAS83] C. J. Crossland, B. G. Hatcher, M. J. Atkinson, and S. V. Smith. Dissolved nutrients of a high-latitude coral reef, Houtman Abrolhos Islands, Western Australia. Marine Ecology Progress Series, 14:159–163, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/14/m014p159.pdf.

Checkley:1982:SFA

[Che82] D. M. Checkley, Jr. Selective feeding by Atlantic herring (Clupea harengus) larvae on zooplankton in natural assemblages. Marine Ecology Progress Series, 9:245–253, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/9/m009p245.pdf.

Checkley:1984:RGI

[Che84] D. M. Checkley, Jr. Relation of growth to ingestion for larvae of Atlantic herring Clupea harengus and other fish. Marine Ecology Progress Series, 18:215–224, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/18/m018p215.pdf.

Chesney:1989:EFR

[Che89] E. J. Chesney, Jr. Estimating the food requirements of striped bass larvae Morone saxatilis: effects of light, turbidity and turbulence. Marine Ecology Progress Series, 53:191–200, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/53/m053p191.pdf.

Cherry:1983:PRE

[CHH83] R. D. Cherry, M. Heyraud, and J. J. W. Higgo. Polonium-210: its relative enrichment in the hepatopancreas of marine invertebrates. Marine Ecology Progress Series, 13:229–236, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/13/
m013p229.pdf.

Choy:1986:NDF

[Cho86] S. C. Choy. Natural diet and feeding habits of the crabs Liocarcinus puber and L. holsatus (Decapoda, Brachyura, Portunidae). Marine Ecology Progress Series, 31:87–99, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/31/m031p087.pdf.

Christy:1989:RDM

[Chr89] J. H. Christy. Rapid development of megalopae of the fiddler crab Uca pugilator reared over sediment: implications for models of larval recruitment. Marine Ecology Progress Series, 57:259–265, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
57/m057p259.pdf.

Crowe:1987:IAD

[CJS87] W. A. Crowe, A. B. Josefson, and I. Svane. Influence of adult density on recruitment into soft sediments: a short-term in situ sublittoral experiment. Marine Ecology Progress Series, 41:61– 69, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
41/m041p061.pdf.

Costello:1989:CRL

[CK89] J. H. Costello and P. M. Kremer. Circadian rhythmicity in the location of zooxanthellae of the scyphomedusa Linuche unguiculata. Marine Ecology Progress Series, 57:279–286, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/57/m057p279.pdf.

Cohen:1982:PFL

[CL82] R. E. Cohen and R. G. Lough. Prey field of larval Herring Clupea harengus on a continental shelf spawning area. Marine Ecology Progress Series, 10:211–222, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/10/m010p211.pdf.

Carr:1984:BRG

[CL84a] R. S. Carr and O. Linden. Bioenergetic responses of Gammarus salinus and Mytilus edulis to oil and oil dispersants in a model ecosystem. Marine Ecology Progress Series, 19:285–291, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/19/
m019p285.pdf.

Clements:1984:PSF

[CL84b] W. H. Clements and R. J. Livingston. Prey selectivity of the fringed filefish Monacanthus ciliatus (Pisces: Monacanthidae): role of prey accessibility. Marine Ecology Progress Series, 16:291– 295, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
16/m016p291.pdf.

Cain:1986:ESC

[CL86] D. J. Cain and S. N. Luoma. Effect of seasonally changing tissue weight on trace metal concentrations in the bivalve Macoma balthica in San Francisco Bay. Marine Ecology Progress Series, 28:209– 217, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
28/m028p209.pdf.

Capuzzo:1988:LCD

[CL88] J. M. Capuzzo and D. F. Leavitt. Lipid composition of the digestive glands of Mytilus edulis and Carcinus maenas in response to pollutant gradients. Marine Ecology Progress Series, 46:139– 145, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p139.pdf.

Clarke:1982:LSR

[Cla82] A. Clarke. Lipid synthesis and reproduction in the polar shrimp Chorismus antarcticus. Marine Ecology Progress Series, 9:81–90, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
9/m009p081.pdf.

Clayton:1985:PLA

[Cla85] W. S. Clayton, Jr. Pedal laceration by the anemone Aiptasia pallida. Marine Ecology Progress Series, 21:75–80, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/21/m021p075.pdf.

Clarke:1987:TLR

[Cla87] A. Clarke. Temperature, latitude and reproductive effort. Marine Ecology Progress Series, 38:89–99, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/38/m038p089.pdf.

Chambers:1988:VAE

[CLB88] R. Christopher Chambers, William C. Legett, and Joseph A. Brown. Variation in and among early life history traits of laboratory-reared winter flounder Pseudo-pleuronectes americanus. Marine Ecology Progress Series, 47:1–15, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/47/m047p001.pdf.

Cloern:1982:DBC

[Clo82] J. E. Cloern. Does the benthos control phytoplankton biomass in South San Francisco Bay? Marine Ecology Progress Series, 9:191– 202, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
9/m009p191.pdf.

Coll:1982:CDS

[CLS^{+}82] J. C. Coll, S. La Barre, P. W. Sammarco, W. T. Williams, and G. J. Bakus. Chemical defences in soft corals (Coelenterata: Octocorallia) of the Great Barrier Reef: a study of comparative toxicities. Marine Ecology Progress Series, 8:271–278, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/8/m008p271.pdf.

Carey:1982:ASI

[CM82] A. G. Carey, Jr. and P. A. Montagna. Arctic Sea ice faunal assemblage: First approach to description and source of the underice meiofauna. Marine Ecology Progress Series, 8:1–8, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/8/m008p001.pdf.

Costello:1987:AFS

[CM87] M. J. Costello and A. A. Myers. Amphipod fauna of the sponges Halichondria panicea and Hymeniaci-don perleve in Lough Hyne, Ireland. Marine Ecology Progress Series, 41:115–121, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/41/
m041p115.pdf.

Cullen:1989:DPP

[CMC89] J. J. Cullen, H. L. Maclntyre, and D. J. Carlson. Distributions and photosynthesis of phototrophs in sea-surface films. Marine Ecology Progress Series, 55:271–278, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/55/m055p271.pdf.

Coffroth:1984:IIC

[Cof84] M. A. Coffroth. Ingestion and incorporation of coral mucus aggregates by a gorgonian soft coral. Marine Ecology Progress Series, 17:193–199, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/17/m017p193.pdf.

Cohen:1985:PPE

[Coh85] R. R. H. Cohen. Physical processes and the ecology of a winter dinoflagellate bloom of Katodinium rotundatum. Marine Ecology Progress Series, 26:135–144, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/26/m026p135.pdf.

Coles:1980:SDD

[Col80] S. L. Coles. Species diversity of decapods associated with living and dead reef coral Pocillopora meandrina. Marine Ecology Progress Series, 2:281–291, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/2/m002p281.pdf.

Collos:1984:TSN

[Col84] Y. Collos. Transient situations in nitrate assimilation by marine diatoms. V. Interspecific variability in biomass and uptake during nitrogen starvation and resupply. Marine Ecology Progress Series, 17:25–31, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/17/m017p025.pdf.

Collie:1985:LHP

[Col85] J. S. Collie. Life history and production of three amphipod species on Georges Bank. Marine Ecology Progress Series, 22:229–238, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
22/m022p229.pdf.

Collos:1986:TLA

[Col86] Y. Collos. Time-lag algal growth dynamics: biological constraints on primary production in aquatic environments. Marine Ecology Progress Series, 33:193–206, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/33/m033p193.pdf.

Collie:1987:FCY

[Col87] J. S. Collie. Food consumption by yellowtail flounder in relation to production of its benthic prey. Marine Ecology Progress Series, 36: 205–213, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/36/m036p205.pdf.

Costlow:1982:MSA

[Cos82] J. D. Costlow, Jr. In memory of Simeon Abramovitch Mileikovsky, 1935 to 1980. Marine Ecology Progress Series, 9:309–310, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/9/
m009p309.pdf.

Cowan:1981:FOC

[Cow81] M. E. Cowan. Field observations of colony movement and division of the ascidian Didemnum molle. Marine Ecology Progress Series, 6:335–337, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/6/m006p335.pdf.

Cote:1984:ULS

[CP84] B. Cote and T. Piatt. Utility of the light-saturation curve as an operational model for quantifying the effects of environmental conditions on phytoplankton photosynthesis. Marine Ecology Progress Series, 18:57–66, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/18/m018p057.pdf.

Corriero:1987:ESB

[CP87] G. Corriero and R. Pronzato. Epibiontic sponges on the bivalve Pinna nobilis. Marine Ecology Progress Series, 35:75–82, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
35/m035p075.pdf.

Coull:1989:CVD

[CPM89] B. C. Coull, M. A. Palmer, and P. E. Myers. Controls on the vertical distribution of meiobenthos in mud: field and flume studies with juvenile fish. Marine Ecology Progress Series, 55:133–139, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
55/m055p133.pdf.

Ceccherelli:1984:SGP

[CR84] V. U. Ceccherelli and R. Rossi. Settlement, growth and production of the mussel Mytilus galloprovincialis. Marine Ecology Progress Series, 16:173–184, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/16/m016p173.pdf.

Crawford:1989:MRP

[Cra89] D. W. Crawford. Mesodinium rubrum: the phytoplankter that wasn’t. Marine Ecology Progress Series, 58:161–174, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/58/
m058p161.pdf.

Crossland:1984:SVR

[Cro84] C. J. Crossland. Seasonal variations in the rates of calcification and productivity in the coral Acropora formosa on a high-latitude reef. Marine Ecology Progress Series, 15:135–140, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/15/m015p135.pdf.

Cameron:1980:SUR

[CS80] R. A. Cameron and S. C. Schroeter. Sea urchin recruitment: Effect of substrate selection on juvenile distribution. Marine Ecology Progress Series, 2:243–247, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/2/m002p243.pdf.

Chong:1981:FFH

[CS81] V. C. Chong and A. Sasekumar. Food and feeding habits of the white prawn Penaeus merguiensis. Marine Ecology Progress Series, 5:185–191, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/5/m005p185.pdf.

Collos:1984:UMP

[CS84] Y. Collos and G. Slawyk. ^{13}C and ^{15}N uptake by marine phytoplankton. III. Interactions in euphotic zone profiles of stratified oceanic areas. Marine Ecology Progress Series, 19:223–231, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
19/m019p223.pdf.

Cynar:1986:UDI

[CS86] F. J. Cynar and J. McN. Sieburth. Unambiguous detection and improved quantification of phagotrophy in apochlorotic nanoflagellates using fluorescent microspheres and concomitant phase contrast and epifluorescence microscopy. Marine Ecology Progress Series, 32:61–70, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/32/m032p061.pdf.

Coffin:1987:MTD

[CS87] R. B. Coffin and J. H. Sharp. Microbial trophodynamics in the Delaware Estuary. Marine Ecology Progress Series, 41:253–266, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
41/m041p253.pdf.

Carlotti:1989:PDM

[CS89a] F. Carlotti and A. Sciandra. Population dynamics model of Euterpina acutifrons (Copepoda. Harpacticoida) coupling individual growth and larval development. Marine Ecology Progress Series, 56:225–242, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/56/m056p225.pdf.

Cartes:1989:FED

[CS89b] J. E. Cartes and F. Sarda. Feeding ecology of the deep-water aristeid crustacean Aristeus antennatus. Marine Ecology Progress Series, 54:229–238, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/54/m054p229.pdf.

Clokie:1981:DMC

[CSB81] J. J. P. Clokie, T. P. Scoffin, and A. D. Boney. Depth maxima of Conchocelis and Phymatolithon rugulosum on the N. W. Shelf and Rockall Plateau. Marine Ecology Progress Series, 4:131–133, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
4/m004p131.pdf.

Capriulo:1988:DGE

[CSD88] G. M. Capriulo, R. A. Schreiner, and B. L. Dexter. Differential growth of Euplotes vannus fed fragmented versus unfragmented chains of Skeletonema costatum. Marine Ecology Progress Series, 47:205–209, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/47/m047p205.pdf.

Chrzanowski:1986:PMB

[CSDZ86] T. H. Chrzanowski, J. D. Spurrier, R. F. Dame, and R. G. Zingmark. Processing of microbial biomass by an intertidal reef community. Marine Ecology Progress Series, 30:181–189, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/30/
m030p181.pdf.

Clifton:1983:CDS

[CSH83] R. J. Clifton, H. E. Stevens, and E. I. Hamilton. Concentration and depuration of some radionuclides present in a chronically exposed population of mussels (Mytilus edulis). Marine Ecology Progress Series, 11:245–256, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/11/m011p245.pdf.

Clifton:1989:UDM

[CSH89] R. J. Clifton, H. E. Stevens, and E. I. Hamilton. Uptake and depuration of^{241}Am,^{239+240}Pu^{238}Pu,^{137}Cs and^{106}Ru by Mytilus edulis under natural stress. Marine Ecology Progress Series, 54:91– 98, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
54/m054p091.pdf.

Chrzanowski:1982:TPO

[CSS82] T. H. Chrzanowski, L. H. Stevenson, and J. D. Spurrier. Transport of particulate organic carbon through the North Inlet Ecosystem. Marine Ecology Progress Series, 7:231–245, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/7/m007p231.pdf.

Chrzanowski:1983:TDO

[CSS83] T. H. Chrzanowski, L. H. Stevenson, and J. D. Spurrier. Transport of dissolved organic carbon through a major creek of the North Inlet Ecosystem. Marine Ecology Progress Series, 13:167– 174, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p167.pdf.

Carey:1984:DES

[CSW84] A. G. Carey, Jr., P. H. Scott, and K. R. Walters. Distributional ecology of shallow south-western Beaufort Sea (Arctic Ocean) bivalve Mollusca. Marine Ecology Progress Series, 17:125–134, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
17/m017p125.pdf.

Chesney:1984:OLP

[CT84] E. J. Chesney, Jr. and K. R. Tenore. Oscillations of laboratory populations of the polychaete Capitella capitata (Type I): their cause and implications for natural populations. Marine Ecology Progress Series, 20:289–296, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/20/m020p289.pdf.

Conley:1985:OCM

[CT85] W. J. Conley and J. T. Turner. Omnivory by the coastal marine copepods Centropages hamatus and Labidocera aestiva. Marine Ecology Progress Series, 21:113–120, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/21/m021p113.pdf.

Coen:1989:MVD

[CT89] L. D. Coen and C. E. Tanner. Morphological variation and differential susceptibility to herbivory in the tropical brown alga Lobophora variegate. Marine Ecology Progress Series, 54:287–298, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
54/m054p287.pdf.

Capriulo:1986:CFE

[CTG86] G. M. Capriulo, J. Taveras, and K. Gold. Ciliate feeding: effect of food presence or absence on occurrence of striae in tintinnids. Marine Ecology Progress Series, 30:145–158, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/30/m030p145.pdf.

Cushing:1989:HCS

[Cus89] D. H. Cushing. Hydrographic containment of a spawning group of plaice in the Southern Bight of the North Sea. Marine Ecology Progress Series, 58:287–297, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/58/m058p287.pdf.

Cary:1989:HCN

[CVF89] S. C. Cary, R. D. Vetter, and H. Felbeck. Habitat characterization and nutritional strategies of the endosymbiont-bearing bivalve Lucinoma aequizonata. Marine Ecology Progress Series, 55:31– 45, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
55/m055p031.pdf.

Collins:1982:ZCB

[CW82] N. R. Collins and R. Williams. Zooplankton communities in the Bristol Channel and Severn Estuary. Marine Ecology Progress Series, 9:1–11, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/9/m009p001.pdf.

Cowden:1983:DPM

[CYC83] C. Cowden, C. M. Young, and F. S. Chia. Differential predation on marine invertebrate larvae by two benthic predators. Marine Ecology Progress Series, 14:145–149, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/14/m014p145.pdf.

DeJong:1984:CBT

[DA84] L. De Jong and W. Admiraal. Competition between three estuarine benthic diatom species in mixed cultures. Marine Ecology Progress Series, 18:269–275, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/18/m018p269.pdf.

DAntonio:1985:RSD

[D’A85a] C. M. D’Antonio. Role of sand in the domination of hard substrata by the intertidal alga Rhodomela larix. Marine Ecology Progress Series, 27:263–275, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/27/m027p263.pdf.

Dennison:1985:RDL

[DA85b] William C. Dennison and Randall S. Alberte. Role of daily light period in the depth distribution of Zostera marina (eelgrass). Marine Ecology Progress Series, 25:51–61, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/25/m025p051.pdf.

Daan:1987:IEP

[Daa87] R. Daan. Impact of egg predation by Noctiluca miliaris on the summer development of copepod populations in the southern North Sea. Marine Ecology Progress Series, 37:9–17, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/37/
m037p009.pdf.

DeMatthaeis:1983:GDB

[DAC^{+}83] E. De Matthaeis, G. Allegrucci, A. Caccone, D. Cesaroni, M. Cobolli Sbordoni, and V. Sbordoni. Genetic differentation between Penaeus kerathurus and P. japonicus (Crustacea, Decapoda). Marine Ecology Progress Series, 12:191–197, 1983.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/12/
m012p191.pdf.

Dahlgaard:1986:EST

[Dah86] H. Dahlgaard. Effects of season and temperature on longterm in situ loss rates of Pu, Am, Np, Eu, Ce, Ag, Tc, Zn, Co and Mn in a Baltic Mytilus edulis population. Marine Ecology Progress Series, 33:157–165, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/33/m033p157.pdf.

DallaVia:1985:OCT

[Dal85] J. Dalla Via. Oxygen consumption and temperature change in the shrimp Palaemon elegans. Marine Ecology Progress Series, 26: 199–202, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/26/m026p199.pdf.

Daniels:1983:DCA

[Dan83] R. A. Daniels. Demographic characteristics of an Antarctic plunderfish, Harpagifer bispinis antarcticus. Marine Ecology Progress Series, 13:181–187, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/13/m013p181.pdf.

Davis:1989:CPD

[Dav89] A. R. Davis. Contrasting population dynamics and life histories in two populations of the colonial subtidal ascidian Podoclavella moluccensis. Marine Ecology Progress Series, 51:107–119, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/51/
m051p107.pdf.

Dalley:1981:NAU

[DB81a] R. Dalley and H. Bailey. A new apparatus used to record the locomotor rhythms of laboratory reared prawns and shrimps. Marine Ecology Progress Series, 4:229–234, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p229.pdf.

deRuytervanSteveninck:1981:BRC

[dB81b] E. D. de Ruyter van Steveninck and A. M. Breeman. Biomass and relative coverage of benthic algae in the fore-reef of Curaçao (Netherlands Antilles) in relation to production. Marine Ecology Progress Series, 6:257–265, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/6/m006p257.pdf.

Dring:1982:PIB

[DB82] M. J. Dring and F. A. Brown. Photosynthesis of intertidal brown algae during and after periods of emersion: a renewed search for physiological causes of zonation. Marine Ecology Progress Series, 8:301–308, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/8/m008p301.pdf.

deRuytervanSteveninck:1986:CAC

[dB86] E. D. de Ruyter van Steveninck and R. P. M. Bak. Changes in abundance of coral-reef bottom components related to mass mortality of the sea urchin Diadema antillarum. Marine Ecology Progress Series, 34:87–94, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/34/m034p087.pdf.

deRuytervanSteveninck:1987:DWP

[dB87a] E. D. de Ruyter van Steveninck and A. M. Breeman. Deep water populations of Lobophora variegata (Phaeophyceae) on the coral reef of Curaçao: influence of grazing and dispersal on distribution patterns. Marine Ecology Progress Series, 38:241–250, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
38/m038p241.pdf.

deRuytervanSteveninck:1987:DWV

[dB87b] E. D. de Ruyter van Steveninck and A. M. Breeman. Deep water vegetations of Lobophora variegata (Phaeophycae) in the coral reef of Curaçao: population dynamics in relation to mass mortality of the sea urchin Diadema antillarum. Marine Ecology Progress Series, 36:81–90, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/36/m036p081.pdf.

Duchene:1988:UPA

[DB88] J. C. Duchene and M. Bhaud. Uncinial patterns and age determination in terebellid polychaetes. Marine Ecology Progress Series, 49:267–275, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/49/m049p267.pdf.

Derenbach:1980:ITP

[DBFM80] J. B. Derenbach, W. Boland, E. Fölster, and D. G. Müller. Interference tests with the pheromone system of the brown alga Cutleria multifida. Marine Ecology Progress Series, 3:357–361, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
3/m003p357.pdf.

Diaz:1982:VME

[DCB82] H. Diaz, J. E. Conde, and M. Bevilacqua. A volumetric method for estimating fecundity in Decapoda. Marine Ecology Progress Series, 10:203–206, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/10/m010p203.pdf.

Dame:1986:OHN

[DCB^{+}86] R. Dame, T. Chrzanowski, K. Bildstein, B. Kjerfve, H. McKellar, D. Nelson, J. Spurrier, S. Stancyk, H. Stevenson, J. Vernberg, and R. Zingmark. The outwelling hypothesis and North Inlet, South Carolina. Marine Ecology Progress Series, 33:217– 229, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
33/m033p217.pdf.

Davis:1985:PAC

[DCJS85] P. G. Davis, D. A. Caron, P. W. Johnson, and J. McN. Sieburth. Prototrophic and apochlorotic components of picoplankton and nanoplankton in the North Atlantic: geographic, vertical, seasonal and diel distributions. Marine Ecology Progress Series, 21:15–26, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
21/m021p015.pdf.

Dayton:1989:PSS

[DCMS89] P. K. Dayton, J. H. Carleton, A. G. Mackley, and P. W. Sammarco. Patterns of settlement, survival and growth of oysters across the Great Barrier Reef. Marine Ecology Progress Series, 54:75–90, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
54/m054p075.pdf.

Dagg:1984:PSS

[DCNS84] M. J. Dagg, M. E. Clarke, T. Nishiyama, and S. L. Smith. Production and standing stock of copepod nauplii, food items for larvae of the walleye pollock Theragra chalcogramma in the southeastern Bering Sea. Marine Ecology Progress Series, 19:7–16, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
19/m019p007.pdf.

Dowd:1980:CEP

[DD80] C. E. Dowd and Houde E. D. Combined effects of prey concentration and photoperiod on sirvival and growth of larval sea bream Archosargus rhomboidalis (Sparidae). Marine Ecology Progress Series, 3:181–185, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/3/m003p181.pdf.

Dafni:1984:SOM

[DD84] J. Dafni and A. Diamant. School-oriented mimicry, a new type of mimicry in fishes. Marine Ecology Progress Series, 20:45–50, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
20/m020p045.pdf.

Deibel:1985:UMF

[DD85] D. Deibel and C. V. L. Dickson, M.-L, Powell. Ultrastructure of the mucous feeding filter of the house of the appendicularian Oikopleura vanhoeffeni. Marine Ecology Progress Series, 27:79– 86, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
27/m027p079.pdf.

Dethier:1988:VSI

[DD88] M. N. Dethier and D. O. Duggins. Variation in strong interactions in the intertidal zone along a geographical gradient: a Washington-Alaska comparison. Marine Ecology Progress Series, 50:97–105, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
50/m050p097.pdf.

Davison:1989:EFS

[DDR89] I. R. Davison, S. R. Dudgeon, and H.-M. Ruan. Effect of freezing on seaweed photosynthesis. Marine Ecology Progress Series, 58: 123–131, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/58/m058p123.pdf.

deJonge:1980:FOC

[de 80] V. N. de Jonge. Fluctuations in the organic carbon to chlorophylla ratios for estuarine benthic diatom populations. Marine Ecology Progress Series, 2:345–353, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/2/m002p345.pdf.

DeWreede:1984:GAC

[De 84] R. E. De Wreede. Growth and age class distribution of Pterygophora california (Phaeophyta). Marine Ecology Progress Series, 19:93–100, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/19/m019p093.pdf.

Dandonneau:1987:SEC

[DE87] Y. Dandonneau and G. Eldin. Southwestward extent of chlorophyll-enriched waters from the Peruvian and equatorial upwellings between Tahiti and Panama. Marine Ecology Progress Series, 38:283–294, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/38/m038p283.pdf.

DeVantier:1988:SPS

[DE88] L. M. DeVantier and R. Endean. The scallop Pedum spondyloideum mitigates the effects of Acanthaster planci predation on the host coral Pontes: host defence facilitated by exaptation? Marine Ecology Progress Series, 47:293–301, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/47/m047p293.pdf.

DeVantier:1989:OCF

[DE89] L. M. DeVantier and R. Endean. Observations of colony fission following ledge formation in massive reef corals of the genus Porites.

Marine Ecology Progress Series, 58:191–195, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/58/m058p191.pdf.

Decho:1986:WCI

[Dec86] A. W. Decho. Water-cover influences on diatom ingestion rates by meiobenthic copepods. Marine Ecology Progress Series, 33:139– 146, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
33/m033p139.pdf.

Decho:1988:HDH

[Dec88] A. W. Decho. How do harpacticoid grazing rates differ over a tidal cycle? Field verification using chlorophyll-pigment analyses. Marine Ecology Progress Series, 45:263–270, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/45/m045p263.pdf.

Deming:1985:BGD

[Dem85] J. W. Deming. Bacterial growth in deep-sea sediment trap and boxcore samples. Marine Ecology Progress Series, 25:305–312, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
25/m025p305.pdf.

DeWitt:1987:MSC

[DeW87] T. H. DeWitt. Microhabitat selection and colonization rates of a benthic amphipod. Marine Ecology Progress Series, 36:237–250, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
36/m036p237.pdf.

Doonan:1982:ESS

[DG82a] S. A. Doonan and G. W. Gooday. Ecological studies of symbiosis in Convoluta roscoffensis. Marine Ecology Progress Series, 8:69– 73, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
8/m008p069.pdf.

Druehl:1982:VDI

[DG82b] L. D. Druehl and J. M. Green. Vertical distribution of intertidal seaweeds as related to patterns of submersion and emersion. Marine Ecology Progress Series, 9:163–170, 1982. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/9/m009p163.pdf.

Dobbs:1988:CTC

[DG88a] Fred C. Dobbs and James B. Guckert. Callianassa trilobata (Crustacea: Thalassinidea) influences abundance of meiofauna and biomass, composition, and physiologic state of microbial communities within its burrow. Marine Ecology Progress Series, 45:69– 79, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
45/m045p069.pdf.

Dobbs:1988:MFR

[DG88b] Fred C. Dobbs and James B. Guckert. Microbial food resources of the macrofaunal-deposit feeder Ptychodera bahamensis (Hemichordata: Enteropneusta). Marine Ecology Progress Series, 45: 127–136, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/45/m045p127.pdf.

Dahlback:1982:MIS

[DGHK82] B. Dahlbäck, L. Å. H. Gunnarsson, M. Hermansson, and S. Kjelleberg. Microbial investigations of surface microlayers, water column, ice and sediment in the Arctic Ocean. Marine Ecology Progress Series, 9:101–109, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/9/m009p101.pdf.

Daan:1988:COC

[DGK88] R. Daan, S. R. Gonzalez, and W. C. M. Klein Breteler. Cannibalism in omnivorous calanoid copepods. Marine Ecology Progress Series, 47:45–54, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/47/m047p045.pdf.

deJong:1988:CCR

[dHS88] S. A. de Jong, P. A. G. Hofman, and A. J. J. Sandee. Construction and calibration of a rapidly responding pH mini-electrode: application to intertidal sediments. Marine Ecology Progress Series, 45: 187–192, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/45/m045p187.pdf.

Dini:1981:RBB

[Din81] F. Dini. Relationship between breeding systems and resistance to mercury in Euplotes crassus (Ciliophora Hypotrichida). Marine Ecology Progress Series, 4:195–202, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p195.pdf.

Dinesen:1982:RVS

[Din82a] Z. D. Dinesen. Regional variation in shade-dwelling coral assemblages of the Great Barrier Reef province. Marine Ecology Progress Series, 7:117–123, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/7/m007p117.pdf.

Dinesen:1982:SDC

[Din82b] Z. D. Dinesen. Shade-dwelling corals of the Great Barrier Reef. Marine Ecology Progress Series, 10:173–185, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/10/m010p173.pdf.

Duineveld:1984:DGR

[DJ84] G. C. A. Duineveld and M. I. Jenness. Differences in growth rates of the sea urchin Echinocardium cordatum as estimated by the parameter w of the von Bertalanffy equation applied to skeletal rings. Marine Ecology Progress Series, 19:65–72, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/19/m019p065.pdf.

Dean:1988:DEG

[DJTL88] T. A. Dean, F. R. Jacobsen, K. Thies, and S. L. Lagos. Differential effects of grazing by white sea urchins on recruitment of brown algae. Marine Ecology Progress Series, 48:99–102, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/48/m048p099.pdf.

Donaghay:1985:PCB

[DK85] P. L. Donaghay and E. Klos. Physical, chemical and biological responses to simulated wind and tidal mixing in experimental marine ecosystems. Marine Ecology Progress Series, 26:35–45, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
26/m026p035.pdf.

Dame:1986:VSA

[DK86] R. F. Dame and P. D. Kenny. Variability of Spartina alterniflora primary production in the euhaline North Inlet estuary. Marine Ecology Progress Series, 32:71–80, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/32/m032p071.pdf.

deRuytervanSteveninck:1988:IPB

[dKB88a] E. D. de Ruyter van Steveninck, P. Kamermans, and A. M. Breeman. Importance of physical and biological processes in structuring tropical intertidal populations of Lobophora variegata (Phaeophyceae). Marine Ecology Progress Series, 44:77–84, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
44/m044p077.pdf.

deRuytervanSteveninck:1988:TET

[dKB88b] E. D. de Ruyter van Steveninck, P. Kamermans, and A. M. Breeman. Transplant experiments with two morphological growth forms of Lobophora variegata (Phaeophyceae). Marine Ecology Progress Series, 49:191–194, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/49/m049p191.pdf.

Davis:1983:BDA

[DKMR83] C. L. Davis, K. Koop, D. G. Muir, and F. T. Robb. Bacterial diversity in adjacent kelp-dominated ecosystems. Marine Ecology Progress Series, 13:115–119, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/13/m013p115.pdf.

Demers:1982:WCS

[DL82] S. Demers and L. Legendre. Water column stability and photosynthetic capacity of estuarine phytoplankton: Long-term relationships. Marine Ecology Progress Series, 7:337–340, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/7/m007p337.pdf.

Davis:1983:RSA

[DL83] M. W. Davis and H. Lee II. Recolonization of sediment-associated microalgae and effects of estuarine infauna on microalgal production. Marine Ecology Progress Series, 11:227–232, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/11/m011p227.pdf.

Davies:1985:VBB

[DL85] A. G. Davies and J. W. Leftley. Vitamin b,2 binding by microalgal exocrines: dissociation constant of the vitamin-binder complex determined using an ultrafiltration technique. Marine Ecology Progress Series, 21:267–273, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/21/m021p267.pdf.

DiTullio:1986:DPN

[DL86a] G. R. DiTullio and E. A. Laws. Diel periodicity of nitrogen and carbon assimilation in five species of marine phytoplankton: accuracy of methodology for predicting n-assimilation rates and N/C composition ratios. Marine Ecology Progress Series, 32:123–132, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
32/m032p123.pdf.

Dye:1986:MMF

[DL86b] A. H. Dye and T. A. Lasiak. Microbenthos, meiobenthos and fiddler crabs: trophic interactions in a tropical mangrove sediment. Marine Ecology Progress Series, 32:259–264, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/32/m032p259.pdf.

Dahl:1983:EEC

[DLT^{+}83] E. Dahl, M. Laake, K. Tjessem, K. Eberlein, and B. Bohle. Effects of ekofisk crude oil on an enclosed planktonic ecosystem. Marine Ecology Progress Series, 14:81–91, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/14/m014p081.pdf.

Dortch:1982:DUN

[DM82] Q. Dortch and H. Maske. Dark uptake of nitrate and nitrate reductase activity of a red-tide population off Peru. Marine Ecology Progress Series, 9:299–303, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/9/m009p299.pdf.

Davis:1983:EPG

[DM83] M. W. Davis and C. D. Mclntire. Effects of physical gradients on the production dynamics of sediment-associated algae. Marine Ecology Progress Series, 13:103–114, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/13/m013p103.pdf.

Douglas:1987:MBE

[DNF87] Donald J. Douglas, James A. Novitsky, and Robert O. Fournier. Microautoradiography-based enumeration of bacteria with estimates of thy-midine-specific growth and production rates. Marine Ecology Progress Series, 36:91–99, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/36/m036p091.pdf.

Doering:1986:AFR

[DO86] P. H. Doering and C. A. Oviatt. Application of filtration rate models to field populations of bivalves: an assessment using experimental mesocosms. Marine Ecology Progress Series, 31:265– 275, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
31/m031p265.pdf.

Dobbs:1983:MDA

[Dob83] F. C. Dobbs. Monitoring defecation activity of infaunal deposit feeders. Marine Ecology Progress Series, 12:47–50, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/12/m012p047.pdf.

Doering:1989:SFM

[DOB^{+}89] P. H. Doering, C. A. Oviatt, L. L. Beatty, V. F. Banzon, R. Rice, S. P. Kelly, B. K. Sullivan, and J. B. Frithsen. Structure and function in a model coastal ecosystem: silicon, the benthos and eutrophication. Marine Ecology Progress Series, 52:287–299, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/52/
m052p287.pdf.

Donn:1987:LDD

[Don87] T. E. Donn, Jr. Longshore distribution of Donax serra in two log-spiral bays in the eastern Cape, South Africa. Marine Ecology Progress Series, 35:217–222, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/35/m035p217.pdf.

Douglas:1983:MBE

[Dou83] D. J. Douglas. Microautoradiography-based enumeration of photosynthetic picoplankton with estimates of carbon-specific growth rates. Marine Ecology Progress Series, 14:223–228, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/14/m014p223.pdf.

Dame:1981:AEF

[DP81] R. F. Dame and B. C. Patten. Analysis of energy flows in an intertidal oyster reef. Marine Ecology Progress Series, 5:115–124, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
5/m005p115.pdf.

Deibel:1987:CUF

[DP87a] D. Deibel and C. V. L. Powell. Comparison of the ultrastructure of the food-concentrating filter of two appendicularians. Marine Ecology Progress Series, 39:81–85, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/39/m039p081.pdf.

Deibel:1987:UPF

[DP87b] D. Deibel and C. V. L. Powell. Ultrastructure of the pharyngeal filter of the appendicularian Oikopleura vanhoeffeni: implications for particle size selection and fluid mechanics. Marine Ecology Progress Series, 35:243–250, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/35/m035p243.pdf.

DeFlaun:1987:DMW

[DPJ87] M. F. DeFlaun, J. H. Paul, and W. H. Jeffrey. Distribution and molecular weight of dissolved DNA in subtropical estuarine and oceanic environments. Marine Ecology Progress Series, 38:65–73, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
38/m038p065.pdf.

Dixon:1986:CSC

[DR86] P. Dixon and A. I. Robertson. A compact, self-contained zooplankton pump for use in shallow coastal habitats: design and performance compared to net samples. Marine Ecology Progress Series, 32:97–100, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/32/m032p097.pdf.

DeVantier:1986:DSG

[DRB86] L. M. DeVantier, R. E. Reichelt, and R. H. Bradbury. Does Spirobranchus giganteus protect host Pontes from predation by Acanthaster planci: predator pressure as a mechanism of coevolution? Marine Ecology Progress Series, 32:307–310, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/32/m032p307.pdf.

Dortch:1983:RDR

[DRCA83] Q. Dortch, T. L. Roberts, J. R. Clayton, Jr., and S. I. Ahmed. RNA/DNA ratios and DNA concentrations as indicators of growth rate and biomass in planktonic marine organisms. Marine Ecology Progress Series, 13:61–71, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/13/m013p061.pdf.

Davis:1984:EOM

[DS84] P. G. Davis and J. McN. Sieburth. Estuarine and oceanic microflagellate predation of actively growing bacteria estimation by frequency of dividing-divided bacteria. Marine Ecology Progress Series, 19:237–246, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/19/m019p237.pdf.

Davoll:1986:MSA

[DS86a] P. J. Davoll and M. W. Silver. Marine snow aggregates: life history sequence and microbial community of abandoned larvacean houses from Monterey Bay, California. Marine Ecology Progress Series, 33:111–120, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/33/m033p111.pdf.

Dobbs:1986:SPS

[DS86b] Fred C. Dobbs and Teresa A. Scholly. Sediment processing and selective feeding by Pectinaria koreni (Polychaeta: Pectinariidae). Marine Ecology Progress Series, 29:165–176, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/29/m029p165.pdf.

Dunton:1986:SCB

[DS86c] K. H. Dunton and D. M. Schell. Seasonal carbon budget and growth of Laminaria solidungula in the Alaskan high Arctic. Marine Ecology Progress Series, 31:57–66, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/31/m031p057.pdf.

Dunton:1989:TRI

[DSG^{+}89] K. H. Dunton, S. M. Saupe, A. N. Golikov, D. M. Schell, and S. V. Schonberg. Trophic relationships and isotopic gradients among Arctic and subarctic marine fauna. Marine Ecology Progress Series, 56:89–97, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/56/m056p089.pdf.

Dando:1985:SOB

[DSS^{+}85] P. R. Dando, A. J. Southward, E. C. Southward, N. B. Terwilliger, and R. C. Terwilliger. Sulphur-oxidising bacteria and haemoglobin in gills of the bivalve mollusc Myrtea spinifera. Marine Ecology Progress Series, 23:85–98, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/23/m023p085.pdf.

Dame:1989:CNP

[DSW89] R. F. Dame, J. D. Spurrier, and T. G. Wolaver. Carbon, nitrogen and phosphorus processing by an oyster reef. Marine Ecology Progress Series, 54:249–256, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/54/m054p249.pdf.

Deibel:1985:ZFE

[DT85] D. Deibel and J. T. Turner. Zooplankton feeding ecology: contents of fecal pellets of the appendicularian Oikopleura vanhoeffeni. Marine Ecology Progress Series, 27:67–78, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/27/m027p067.pdf.

Demers:1985:LVF

[DTLN85] S. Demers, J.-C. Therriault, L. Legendre, and J. Neveux. An in vivo fluorescence method for the continuous in situ estimation of phytoplankton photosynthetic characteristics. Marine Ecology Progress Series, 27:21–27, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/27/m027p021.pdf.

Duarte:1989:TBV

[Dua89] C. M. Duarte. Temporal biomass variability and production/ biomass relationships of seagrass communities. Marine Ecology Progress Series, 51:269–276, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/51/m051p269.pdf.

Dundas:1986:FPE

[Dun86] I. E. D. Dundas. Fate and possible effects of excessive sperm released during spawning. Marine Ecology Progress Series, 30:287– 290, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
30/m030p287.pdf.

Dobbs:1983:IES

[DV83] Fred C. Dobbs and Joseph M. Vozarik. Immediate effects of a storm on coastal infauna. Marine Ecology Progress Series, 11:273– 279, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
11/m011p273.pdf.

Dagg:1983:NIR

[DW83] M. J. Dagg and K. D. Wyman. Natural ingestion rates of the copepods Neocalanus plumchrus and N. cristatus calculated from gut contents. Marine Ecology Progress Series, 13:37–46, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/13/
m013p037.pdf.

Davis:1987:ADR

[DW87] A. N. Davis and R. T. Wilce. Algal diversity in relation to physical disturbance: a mosaic of successional stages in a subtidal cobble habitat. Marine Ecology Progress Series, 37:229–237, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
37/m037p229.pdf.

Estrada:1987:ETC

[EAM87] M. Estrada, M. Alcaraz, and C. Marrasé. Effects of turbulence on the composition of phytoplankton assemblages in marine microcosms. Marine Ecology Progress Series, 38:267–281, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
38/m038p267.pdf.

Eberlein:1986:DPD

[EB86] K. Eberlein and U. H. Brockmann. Development of particulate and dissolved carbohydrates in parallel enclosure experiments with monocultures of Thalassiosira rotula. Marine Ecology Progress Series, 32:133–138, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/32/m032p133.pdf.

Ebert:1986:NTE

[Ebe86] T. A. Ebert. A new theory to explain the origin of growth lines in sea urchin spines. Marine Ecology Progress Series, 34:197–199, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
34/m034p197.pdf.

Eberlein:1983:TDC

[EBH^{+}83] K. Eberlein, U. H. Brockmann, K. D. Hammer, G. Kattner, and M. Laake. Total dissolved carbohydrates in an enclosure experiment with unialgal Skeletonema costatum culture. Marine Ecology Progress Series, 14:45–58, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/14/m014p045.pdf.

Elner:1987:NDL

[EC87a] R. W. Elner and A. Campbell. Natural diets of lobster Homarus americanus from barren ground and macroalgal habitats off southwestern Nova Scotia, Canada. Marine Ecology Progress Series, 37: 131–140, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/37/m037p131.pdf.

Eskin:1987:STY

[EC87b] R. A. Eskin and B. C. Coull. Seasonal and three-year variability of meiobenthic nematode populations at two estuarine sites. Marine Ecology Progress Series, 41:295–303, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/41/m041p295.pdf.

Eppley:1981:ODO

[EHF^{+}81] R. W. Eppley, S. G. Horrigan, J. A. Fuhrman, E. R. Brooks, C. C. Price, and K. Sellner. Origins of dissolved organic matter in Southern California coastal waters: Experiments on the role of zooplankton. Marine Ecology Progress Series, 6:149–159, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/6/m006p149.pdf.

Evtushenko:1984:BCS

[ELK84] Z. S. Evtushenko, O. N. Lukyanova, and N. K. Khristoforova. Biochemical changes in selected body tissues of the scallop Patinopecten yessoensis under long-term exposure to low Cd concentrations. Marine Ecology Progress Series, 20:165–170, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/20/
m020p165.pdf.

Epifanio:1988:DRF

[ELR88] C. E. Epifanio, K. T. Little, and P. M. Rowe. Dispersal and recruitment of fiddler crab larvae in the Delaware River estuary. Marine Ecology Progress Series, 43:181–188, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/43/m043p181.pdf.

Everson:1987:MVD

[EM87] I. Everson and E. Murphy. Mesoscale variability in the distribution of krill Euphausia superba. Marine Ecology Progress Series, 40:53– 60, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
40/m040p053.pdf.

Estep:1989:TLC

[EM89] K. W. Estep and F. Maclntyre. Taxonomy, life cycle, distribution and dasmotrophy of Chrysochromulina: a theory accounting for scales, haptonema, muciferous bodies and toxicity. Marine Ecology Progress Series, 57:11–21, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/57/m057p011.pdf.

Estrada:1988:PRI

[EMA88] M. Estrada, C. Marrasé, and M. Alcaraz. Phytoplankton response to intermittent stirring and nutrient addition in marine microcosms. Marine Ecology Progress Series, 48:225–234, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/48/
m048p225.pdf.

Emerson:1989:WSL

[Eme89] C. W. Emerson. Wind stress limitation of benthic secondary production in shallow, soft-sediment communities. Marine Ecology Progress Series, 53:65–77, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/53/m053p065.pdf.

Epifanio:1989:TBC

[EMG89] C. E. Epifanio, A. K. Masse, and R. W. Garvine. Transport of blue crab larvae by surface currents off Delaware Bay, USA. Marine Ecology Progress Series, 54:35–41, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/54/m054p035.pdf.

Estep:1986:MUF

[EMHS86] K. W. Estep, F. Maclntyre, E. Hjörleifsson, and J. McN. Sieburth. Maclmage: a user-friendly image-analysis system for the accurate mensuration of marine organisms. Marine Ecology Progress Series, 33:243–253, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/33/m033p243.pdf.

Emlet:1986:LPD

[Eml86] R. B. Emlet. Larval production, dispersal, and growth in a fjord: a case study on larvae of the sand dollar Dendraster excentricus. Marine Ecology Progress Series, 31:245–254, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/31/m031p245.pdf.

Emmerson:1980:IMP

[Emm80] W. D. Emmerson. Induced maturation of prawn Penaeus indicus. Marine Ecology Progress Series, 2:121–131, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/2/m002p121.pdf.

Epifanio:1988:DST

[Epi88] C. E. Epifanio. Dispersal strategies of two species of swimming crab on the continental shelf adjacent to Delaware Bay. Marine Ecology Progress Series, 49:243–248, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/49/m049p243.pdf.

Emerson:1987:IFP

[ER87] C. W. Emerson and J. C. Roff. Implications of fecal pellet size and zooplankton behaviour to estimates of pelagic-benthic carbon flux. Marine Ecology Progress Series, 35:251–257, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/35/m035p251.pdf.

Ellery:1984:CHU

[ES84] W. N. Ellery and M. H. Schleyer. Comparison of homogenization and ultrasonication as techniques in extracting attached sedimentary bacteria. Marine Ecology Progress Series, 15:247–250, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
15/m015p247.pdf.

Enoksson:1987:NDA

[ES87] V. Enoksson and M.-O. Samuelsson. Nitrification and dissimilatory ammonium production and their effects on nitrogen flux over the sediment–water interface in bioturbated coastal sediments. Marine Ecology Progress Series, 36:181–189, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/36/m036p181.pdf.

Eppley:1988:SCM

[ESR^{+}88] R. W. Eppley, E. Swift, D. G. Redalje, M. R. Landry, and L. W. Haas. Subsurface chlorophyll maximum in August–September 1985 in the CLIMAX area of the North Pacific. Marine Ecology Progress Series, 42:289–301, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/42/m042p289.pdf.

Elrifi:1984:TPR

[ET84] I. R. Elrifi and D. H. Turpin. Transient photosynthetic responses of nitrogen limited macroalgae to nitrogen addition. Marine Ecology Progress Series, 20:253–258, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/20/m020p253.pdf.

Evans:1985:NSW

[ET85] S. Evans and B. Tallmark. Niche separation within the mobile predator guild on marine shallow soft bottoms. Marine Ecology Progress Series, 23:279–286, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/23/m023p279.pdf.

Evans:1982:PPF

[Eva82] S. Evans. Production, predation and food niche segregation in a marine shallow soft-bottom community. Marine Ecology Progress Series, 10:147–157, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/10/m010p147.pdf.

Evans:1987:RAP

[Eva87] A. Evans. Relative availability of the prey of wading birds by day and by night. Marine Ecology Progress Series, 37:103–107, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
37/m037p103.pdf.

Ember:1987:PIC

[EWM87] L. M. Ember, D. F. Williams, and J. T. Morris. Processes that influence carbon isotope variations in salt marsh sediments. Marine Ecology Progress Series, 36:33–42, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/36/m036p033.pdf.

Esselink:1989:STB

[EZ89] P. Esselink and L. Zwarts. Seasonal trend in burrow depth and tidal variation in feeding activity of Nereis diversicolor. Marine Ecology Progress Series, 56:243–254, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/56/m056p243.pdf.

Fuhrman:1983:ABM

[FA83] J. A. Fuhrman and F. Azam. Adaptations of bacteria to marine subsurface waters studied by temperature response. Marine Ecology Progress Series, 13:95–98, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/13/m013p095.pdf.

Friedland:1989:IPD

[FAG89] K. D. Friedland, D. W. Ahrenholz, and J. F. Guthrie. Influence of plankton on distribution patterns of the filter-feeder Brevoortia tyrannus (Pisces: Clupeidae). Marine Ecology Progress Series, 54:1–11, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/54/m054p001.pdf.

Fairweather:1988:CPW

[Fai88] P. G. Fairweather. Correlations of predatory whelks with intertidal prey at several scales of space and time. Marine Ecology Progress Series, 45:237–243, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/45/m045p237.pdf.

Fuhrman:1985:BCM

[FB85] J. A. Fuhrman and T. M. Bell. Biological considerations in the measurement of dissolved free amino acids in seawater and implications for chemical and microbiological studies. Marine Ecology Progress Series, 25:13–21, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/25/m025p013.pdf.

Flynn:1986:NSG

[FB86] K. J. Flynn and I. Butler. Nitrogen sources for the growth of marine microalgae: role of dissolved free amino acids. Marine Ecology Progress Series, 34:281–304, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/34/m034p281.pdf.

Frankignoulle:1987:SVD

[FB87] M. Frankignoulle and J. M. Bouquegneau. Seasonal variation of the diel carbon budget of a marine macrophyte ecosystem. Marine Ecology Progress Series, 38:197–199, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/38/m038p197.pdf.

Fisher:1983:ACD

[FBCH83] N. S. Fisher, K. A. Burns, R. D. Cherry, and M. Heyraud. Accumulation and cellular distribution of ^{241}Am, ^{210}Po, and ^{210}Pb in two marine algae. Marine Ecology Progress Series, 11:233–237, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
11/m011p233.pdf.

Fisher:1984:ATC

[FBT84] N. S. Fisher, M. Bone, and J.-L. Teyssie. Accumulation and toxicity of Cd, Zn, Ag, and Hg in four marine phytopiankters. Marine Ecology Progress Series, 18:201–213, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/18/m018p201.pdf.

Furness:1982:IBB

[FC82] R. W. Furness and J. Cooper. Interactions between breeding seabird and pelagic fish populations in the Southern Benguela Region. Marine Ecology Progress Series, 8:243–250, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/8/m008p243.pdf.

Fleeger:1983:MRE

[FC83] J. W. Fleeger and G. T. Chandler. Meiofauna responses to an experimental oil spill in a Louisiana salt marsh. Marine Ecology Progress Series, 11:257–264, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/11/m011p257.pdf.

Fleeger:1984:ETC

[FCFP84] J. W. Fleeger, G. T. Chandler, G. R. Fitzhugh, and F. E. Phillips. Effects of tidal currents on meiofauna densities in vegetated salt marsh sediments. Marine Ecology Progress Series, 19:49–53, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/19/
m019p049.pdf.

Field:1982:PSA

[FCW82] J. G. Field, K. R. Clarke, and R. M. Warwick. A practical strategy for analysing multispecies distribution patterns. Marine Ecology Progress Series, 8:37–52, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/8/m008p037.pdf.

Fielding:1989:CNR

[FD89] P. J. Fielding and C. L. Davis. Carbon and nitrogen resources available to kelp bed filter feeders in an upwelling environment. Marine Ecology Progress Series, 55:181–189, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/55/m055p181.pdf.

Forbes:1986:DPC

[FDM86] J. R. Forbes, K. L. Denman, and D. L. Mackas. Determination of photosynthetic capacity in coastal marine phytoplankton: effects of assay irradiance and variability of photosynthetic parameters. Marine Ecology Progress Series, 32:181–191, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/32/m032p181.pdf.

Fuhrman:1985:DVB

[FEHA85] J. A. Fuhrman, R. W. Eppley, Å. Hägström, and F. Azam. Diel variations in bacterioplankton, phytoplankton, and related parameters in the Southern California Bight. Marine Ecology Progress Series, 27:9–20, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/27/m027p009.pdf.

Fenchel:1982:EHMa

[Fen82a] T. Fenchel. Ecology of heterotrophic microflagellates. I. Some important forms and their functional morphology. Marine Ecology Progress Series, 8:211–223, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/8/m008p211.pdf.

Fenchel:1982:EHMb

[Fen82b] T. Fenchel. Ecology of heterotrophic microflagellates. II. Bioenergetics and growth. Marine Ecology Progress Series, 8:225–231, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
8/m008p225.pdf.

Fenchel:1982:EHMc

[Fen82c] T. Fenchel. Ecology of heterotrophic microflagellates. III. Adaptations to heterogeneous environments. Marine Ecology Progress Series, 9:25–33, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/9/m009p025.pdf.

Fenchel:1982:EHMd

[Fen82d] T. Fenchel. Ecology of heterotrophic microflagellates. IV quantitative occurrence and importance as bacterial consumers. Marine Ecology Progress Series, 9:35–42, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/9/m009p035.pdf.

Frithsen:1985:RBM

[FER85] J. B. Frithsen, R. Elmgren, and D. T. Rudnick. Responses of benthic meiofauna to long-term, low-level additions of no. 2 fuel oil. Marine Ecology Progress Series, 23:1–14, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/23/m023p001.pdf.

Fevolden:1989:GDI

[Fev89] S. E. Fevolden. Genetic differentiation of the Iceland scallop Chlamys islandica (Pectinidae) in the northern Atlantic Ocean. Marine Ecology Progress Series, 51:77–85, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/51/m051p077.pdf.

Fonseca:1986:CCF

[FF86a] M. S. Fonseca and J. S. Fisher. A comparison of canopy friction and sediment movement between four species of seagrass with reference to their ecology and restoration. Marine Ecology Progress Series, 29:15–22, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/29/m029p015.pdf.

Fuhrman:1986:NCR

[FF86b] J. A. Fuhrman and R. L. Ferguson. Nanomolar concentrations and rapid turnover of dissolved free amino acids in seawater: agreement between chemical and microbiological measurements. Marine Ecology Progress Series, 33:237–242, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/33/m033p237.pdf.

Flynn:1989:CIE

[FF89] K. J. Flynn and J. Fielder. Changes in intracellular and extracellular amino acids during the predation of the chlorophyte Dunaliella primolecta by the heterotrophic dinoflagellate Oxyrrhis marina and the use of the glutamine/glutamate ratio as an indicator of nutrient status in mixed populations. Marine Ecology Progress Series, 53:117–127, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/53/m053p117.pdf.

Faulkner:1983:CDE

[FG83] D. J. Faulkner and M. T. Ghiselin. Chemical defense and evolutionary ecology of dorid nudibranchs and some other opisthobranch gastropods. Marine Ecology Progress Series, 13:295–301, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p295.pdf.

Fevolden:1987:ESA

[FG87] S. E. Fevolden and S. P. Garner. Environmental stress and allozyme variation in Littorina littorea (Prosobranchia). Marine Ecology Progress Series, 39:129–136, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/39/m039p129.pdf.

Fuiman:1988:PAH

[FG88] L. A. Fuiman and J. C. Gamble. Predation by Atlantic herring, sprat, and sandeels on herring larvae in large enclosures. Marine Ecology Progress Series, 44:1–6, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/44/m044p001.pdf.

Fricke:1983:CBB

[FH83] H. W. Fricke and L. Hottinger. Coral bioherms below the euphotic zone in the Red Sea. Marine Ecology Progress Series, 11:113–117, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
11/m011p113.pdf.

Fortier:1989:OFD

[FH89] L. Fortier and R. P. Harris. Optimal foraging and densitydependent competition in marine fish larvae. Marine Ecology Progress Series, 51:19–33, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/51/m051p019.pdf.

Fuhrman:1988:UTB

[FHC88] J. A. Fuhrman, S. G. Horrigan, and D. G. Capone. Use of ^{13}N as tracer for bacterial and algal uptake of ammonium from seawater. Marine Ecology Progress Series, 45:271–278, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/45/m045p271.pdf.

Fletcher:1985:AAC

[FHKR85] G. L. Fletcher, K. Haya, M. J. King, and H. M. Reisman. Annual antifreeze cycles in Newfoundland, New Brunswick and Long Island winter flounder Pseudopleuronectes americanus. Marine Ecology Progress Series, 21:205–212, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/21/m021p205.pdf.

Fukuchi:1988:PBA

[FHS88] M. Fukuchi, H. Hattori, and T. Sasaki, H.. Hoshiai. A phytoplankton bloom and associated processes observed with a longterm moored system in Antarctic waters. Marine Ecology Progress Series, 45:279–288, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/45/m045p279.pdf.

Fischer:1983:SWI

[Fis83] H. Fischer. Shell weight as an independent variable in relation to cadmium content of mollusks. Marine Ecology Progress Series, 12: 59–75, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/12/m012p059.pdf.

Fischer:1986:ITS

[Fis86] H. Fischer. Influence of temperature, salinity, and oxygen on the cadmium balance of mussels Mytilus edulis. Marine Ecology Progress Series, 32:265–278, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/32/m032p265.pdf.

Fischer:1988:MEQ

[Fis88] H. Fischer. Mytilus edulis as a quantitative indicator of dissolved cadmium. final study and synthesis. Marine Ecology Progress Series, 48:163–174, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/48/m048p163.pdf.

Fischer:1989:CSR

[Fis89] H. Fischer. Cadmium in seawater recorded by mussels: regional decline established. Marine Ecology Progress Series, 55:159–169, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
55/m055p159.pdf.

Fukazawa:1980:MRT

[FITF80] N. Fukazawa, T. Ishimaru, M. Takahashi, and Y. Fujita. A mechanism of red tide formation. I. Growth rate estimate by DCMUinduced fluorescence increase. Marine Ecology Progress Series, 3: 217–222, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/3/m003p217.pdf.

Fenchel:1988:FBS

[FJ88a] T. Fenchel and P. R. Jonsson. The functional biology of Strombidium sulcatum, a marine oligotrich ciliate (Ciliophora, Oligotrichina). Marine Ecology Progress Series, 48:1–15, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/48/
m048p001.pdf.

Frid:1988:IBT

[FJ88b] C. L. J. Frid and R. James. Interactions between two species of saltmarsh gastropod, Hydrobia ulvae and Littorina littorea. Marine Ecology Progress Series, 43:173–179, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/43/m043p173.pdf.

Feigenbaum:1984:CLC

[FK84] D. Feigenbaum and M. Kelly. Changes in the lower Chesapeake Bay food chain in presence of the sea nettle Chrysaora quinquecirrha (Scyphomedusa). Marine Ecology Progress Series, 19:39–47, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
19/m019p039.pdf.

Flint:1986:BEE

[FK86] R. W. Flint and R. D. Kalke. Biological enhancement of estuarine benthic community structure. Marine Ecology Progress Series, 31: 23–33, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/31/m031p023.pdf.

Fletcher:1981:EOS

[FKW81] G. L. Fletcher, J. W. Kiceniuk, and U. P. Williams. Effects of oiled sediments on mortality, feeding and growth of winter flounder Pseudopleuronectes americanus. Marine Ecology Progress Series, 4:91–96, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/4/m004p091.pdf.

Fortier:1985:DSL

[FL85] L. Fortier and W. C. Leggett. A drift study of larval fish survival. Marine Ecology Progress Series, 25:245–257, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/25/m025p245.pdf.

Forbes:1986:CFC

[FL86a] V. E. Forbes and G. R. Lopez. Changes in feeding and crawling rates of Hydrobia truncata (Prosobranchia: Hydrobiidae) in response to sedimentary chlorophyll-a and recently egested sediment. Marine Ecology Progress Series, 33:287–294, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/33/m033p287.pdf.

Frank:1986:EPA

[FL86b] K. T. Frank and W. C. Leggett. Effect of prey abundance and size on the growth and survival of larval fish: an experimental study employing large volume enclosures. Marine Ecology Progress Series, 34:11–22, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/34/m034p011.pdf.

Ferreiro:1988:DAT

[FL88] M. J. Ferreiro and U. Labarta. Distribution and abundance of teleostean eggs and larvae on the NW coast of Spain. Marine Ecology Progress Series, 43:189–199, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/43/m043p189.pdf.

Fletcher:1987:LHD

[Fle87] W. J. Fletcher. Life history dynamics of the limpet Patelloida alticostata in intertidal and subtidal environments. Marine Ecology Progress Series, 39:115–127, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/39/m039p115.pdf.

Folium:1988:GWF

[FM88] O. A. Folium and K. A. Moe. The GEEP workshop: field sampling. Marine Ecology Progress Series, 46:7–12, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/46/m046p007.pdf.

Fishelson:1987:BSA

[FMM87] L. Fishelson, L. W. Montgomery, and A. H. Myrberg, Jr. Biology of surgeonfish Acanthurus nigrofuscus with emphasis on changeover in diet and annual gonadal cycles. Marine Ecology Progress Series, 39:37–47, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/39/m039p037.pdf.

Fallon:1983:BPM

[FNH83] R. D. Fallon, S. Y. Newell, and C. S. Hopkinson. Bacterial production in marine sediments: will cell-specific measures agree with whole-system metabolism? Marine Ecology Progress Series, 11: 119–127, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/11/m011p119.pdf.

Fleeger:1982:SPE

[FP82] J. W. Fleeger and M. A. Palmer. Secondary production of the estuarine, meiobenthic copepod Microarthridion littorale. Marine Ecology Progress Series, 7:157–162, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/7/m007p157.pdf.

Forteath:1982:MGS

[FPRW82] G. N. R. Forteath, G. B. Picken, R. Ralph, and J. Williams. Marine growth studies on the North Sea oil platform montrose alpha. Marine Ecology Progress Series, 8:61–68, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/8/m008p061.pdf.

Flint:1980:PEN

[FR80] R. W. Flint and N. N. Rabalais. Polychaete ecology and niche patterns: Texas continental shelf. Marine Ecology Progress Series, 3: 193–202, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/3/m003p193.pdf.

Frost:1987:GCP

[Fro87] B. W. Frost. Grazing control of phytoplankton stock in the open subarctic Pacific Ocean: a model assessing the role of mesozooplankton, particularly the large calanoid copepods Neocalanus spp. Marine Ecology Progress Series, 39:49–68, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/39/m039p049.pdf.

Fang:1984:FEC

[FS84] L.-S. Fang and P. Shen. Foreign elements in a clam shell: a clue to the history of marine pollution events. Marine Ecology Progress Series, 18:187–189, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/18/m018p187.pdf.

Farke:1984:BCI

[FSBOG84] H. Farke, M. Schulz-Baldes, K. Ohm, and S. A. Gerlach. Bremerhaven caisson for intertidal field studies. Marine Ecology Progress Series, 16:193–197, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/16/m016p193.pdf.

Fuhrman:1989:DBB

[FSCP89] J. A. Fuhrman, T. D. Sleeter, C. A. Carlson, and L. M. Proctor. Dominance of bacterial biomass in the Sargasso Sea and its ecological implications. Marine Ecology Progress Series, 57:207– 217, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
57/m057p207.pdf.

Fukami:1985:MDPa

[FST85a] K. Fukami, U. Simidu, and N. Taga. Microbial decomposition of phyto- and zooplankton in seawater I changes in organic matter. Marine Ecology Progress Series, 21:1–5, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/21/m021p001.pdf.

Fukami:1985:MDPb

[FST85b] K. Fukami, U. Simidu, and N. Taga. Microbial decomposition of phyto- and zooplankton in seawater II. Changes in the bacterial community. Marine Ecology Progress Series, 21:7–13, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/21/
m021p007.pdf.

Fleeger:1989:MRS

[FSZ89] John W. Fleeger, Thomas C. Shirley, and David A. Ziemann. Meiofaunal responses to sedimentation from an Alaskan spring bloom. I major taxa. Marine Ecology Progress Series, 57:137– 145, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
57/m057p137.pdf.

Findlay:1982:EFL

[FT82] S. Findlay and K. R. Tenore. Effect of a free-living marine nematode (Diplolaimella chitwoodi) on detrital carbon mineralization. Marine Ecology Progress Series, 8:161–166, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/8/m008p161.pdf.

Fisher:1986:IFC

[FT86] N. S. Fisher and J.-L. Teyssie. Influence of food composition on the biokinetics and tissue distribution of zinc and americium in mussels. Marine Ecology Progress Series, 28:197–207, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/28/m028p197.pdf.

Fuhrman:1981:IMA

[Fuh81] J. A. Fuhrman. Influence of method on the apparent size distribution of bacterioplankton cells: Epifluorescence microscopy compared to scanning electron microscopy. Marine Ecology Progress Series, 5:103–106, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/5/m005p103.pdf.

Fuhrman:1987:CCB

[Fuh87] J. Fuhrman. Close coupling between release and uptake of dissolved free amino acids in seawater studied by an isotope dilution approach. Marine Ecology Progress Series, 37:45–52, 1987.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/37/
m037p045.pdf.

Fuiman:1989:VAH

[Fui89] L. A. Fuiman. Vulnerability of Atlantic herring larvae to predation by yearling herring. Marine Ecology Progress Series, 51:291–299, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
51/m051p291.pdf.

Fairweather:1984:PIP

[FUM84] P. G. Fairweather, A. J. Underwood, and M. J. Moran. Preliminary investigations of predation by the whelk Morula marginalba. Marine Ecology Progress Series, 17:143–156, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/17/m017p143.pdf.

Fricke:1982:SCP

[FV82] H. Fricke and E. Vareschi. A scleractinian coral (Plerogyra sinuosa) with ‘’photosynthetic organs’. Marine Ecology Progress Series, 7:273–278, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/7/m007p273.pdf.

Flores-Verdugo:1987:SLF

[FVDBD87] F. J. Flores-Verdugo, J. W. Day, Jr., and R. Briseno-Duenas. Structure, litter fall, decomposition, and detritus dynamics of mangroves in a Mexican coastal lagoon with an ephemeral inlet. Marine Ecology Progress Series, 35:83–90, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/35/m035p083.pdf.

Fukui:1986:DRF

[FW86] Y. Fukui and K. Wada. Distribution and reproduction of four intertidal crabs (Crustacea, Brachyura) in the Tonda River Estuary, Japan. Marine Ecology Progress Series, 30:229–241, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/30/
m030p229.pdf.

Forward:1989:BRL

[FWB89] R. B. Forward, Jr., C. A. Wellins, and C. U. Buswell. Behavioral responses of larvae of the crab Neopanope sayi to hydrostatic pressure. Marine Ecology Progress Series, 57:267–277, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/57/m057p267.pdf.

Fujita:1989:AMN

[FWE89] R. M. Fujita, P. A. Wheeler, and R. L. Edwards. Assessment of macroalgal nitrogen limitation in a seasonal upwelling region. Marine Ecology Progress Series, 53:293–303, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/53/m053p293.pdf.

Gotelli:1983:CPC

[GA83] N. J. Gotelli and L. G. Abele. Community patterns of coralassociated decapods. Marine Ecology Progress Series, 13:131–139, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p131.pdf.

Gremare:1989:MCA

[GAA89] A. Grémare, J. M. Amouroux, and J. Amouroux. Modelling of consumption and assimilation in the deposit-feeding polychaete Eupolymnia nebulosa. Marine Ecology Progress Series, 54:239– 248, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
54/m054p239.pdf.

Geesey:1984:FFP

[GABM84] G. G. Geesey, G. V. Alexander, R. N. Bray, and A. C. Miller. Fish fecal pellets are a source of minerals for inshore reef communities. Marine Ecology Progress Series, 15:19–25, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/15/m015p019.pdf.

Gray:1988:ACA

[GAC^{+}88] J. S. Gray, M. Aschan, M. R. Carr, K. R. Clarke, R. H. Green, T. H. Pearson, R. Rosenberg, and R. M. Warwick. Analysis of community attributes of the benthic macrofauna of Frierfjord/Langesundfjord and in a mesocosm experiment. Marine Ecology Progress Series, 46:151–165, 1988. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/46/m046p151.pdf.

Galzin:1987:SFCa

[Gal87a] R. Galzin. Structure of fish communities of French Polynesian coral reefs. I. Spatial scales. Marine Ecology Progress Series, 41: 129–136, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/41/m041p129.pdf.

Galzin:1987:SFCb

[Gal87b] R. Galzin. Structure of fish communities of French Polynesian coral reefs. II. Temporal scales. Marine Ecology Progress Series, 41:137–145, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/41/m041p137.pdf.

Gallegos:1989:MGP

[Gal89] C. L. Gallegos. Microzooplankton grazing on phytoplankton in the Rhode River, Maryland: noninear feeding kinetics. Marine Ecology Progress Series, 57:23–33, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/57/m057p023.pdf.

Garber:1984:TSS

[Gar84] J. H. Garber. ^{15}N tracer study of the short-term fate of particulate organic nitrogen at the surface of coastal marine sediments. Marine Ecology Progress Series, 16:89–104, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/16/m016p089.pdf.

Garnick:1989:LHA

[Gar89] E. Garnick. Lobster (Homarus americanus) population declines, sea urchins, and “barren grounds” a space-mediated competition hypothesis. Marine Ecology Progress Series, 58:23–28, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/58/
m058p023.pdf.

Gast:1985:BFS

[Gas85] V. Gast. Bacteria as a food source for microzooplankton in the Schlei Fjord and Baltic Sea with special references to ciliates. Marine Ecology Progress Series, 22:107–120, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/22/m022p107.pdf.

Gaston:1987:BPM

[Gas87] G. R. Gaston. Benthic Polychaeta of the Middle Atlantic Bight: feeding and distribution. Marine Ecology Progress Series, 36:251– 262, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
36/m036p251.pdf.

Griffiths:1983:DMD

[GB83] F. B. Griffiths and S. B. Brandt. Distribution of mesopelagic decapod Crustacea in and around a warm-core eddy in the Tasman Sea. Marine Ecology Progress Series, 12:175–184, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/12/m012p175.pdf.

Gadomski:1984:FEP

[GB84] D. M. Gadomski and G. W. Boehlert. Feeding ecology of pelagic larvae of English sole Parophrys vetulus and butter sole Isopsetta isolepis off the Oregon coast. Marine Ecology Progress Series, 20: 1–12, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/20/m020p001.pdf.

Gray:1986:CTC

[GB86] C. A. Gray and J. D. Bell. Consequences of two common techniques for sampling vagile macrofauna associated with the seagrass Zostera capricorni. Marine Ecology Progress Series, 28:43– 48, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
28/m028p043.pdf.

Gajardo:1989:ASR

[GB89] G. M. Gajardo and J. A. Beardmore. Ability to switch reproductive mode in Artemia is related to maternal heterozygosity. Marine Ecology Progress Series, 55:191–195, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/55/m055p191.pdf.

Gray:1983:PLT

[GC83] J. S. Gray and H. Christie. Predicting long-term changes in marine benthic communities. Marine Ecology Progress Series, 13:87–94, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p087.pdf.

Gill:1985:SIA

[GC85] C. W. Gill and D. J. Crips. Sensitivity of intact and antennule amputated copepods to water disturbance. Marine Ecology Progress Series, 21:221–227, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/21/m021p221.pdf.

Griffiths:1982:MPR

[GCBM82] R. P. Griffiths, B. A. Caldwell, W. A. Broich, and R. Y. Morita. Microbial processes relating to carbon cycling in southeastern Bering Sea sediments. Marine Ecology Progress Series, 10:265– 275, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
10/m010p265.pdf.

Goldman:1987:NCM

[GCD87] J. G. Goldman, D. A. Caron, and M. R. Dennett. Nutrient cycling in a microflagellate food chain: IV. Phytoplankton-microflagellate interactions. Marine Ecology Progress Series, 38:75–87, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/38/
m038p075.pdf.

Gold:1982:VCP

[GCK82] K. Gold, G. Capriulo, and K. Keeling. Variability in the calcium phosphate concretion load in the kidney of Mercenaria mercenaria. Marine Ecology Progress Series, 10:97–99, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/10/m010p097.pdf.

Giguere:1988:MRS

[GCSP88] L. A. Giguère, B. Coté, and J.-F. St-Pierre. Metabolic rates scale isometrically in larval fishes. Marine Ecology Progress Series, 50: 13–19, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/50/m050p013.pdf.

Gilmurray:1981:FRA

[GD81] M. C. Gilmurray and G. R. Daborn. Feeding relations of the Atlantic silverside Menidia menidia in the Minas Basin, Bay of Fundy. Marine Ecology Progress Series, 6:231–235, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/6/m006p231.pdf.

Goldman:1984:EPD

[GD84a] Joel C. Goldman and Mark R. Dennett. Effect of photoinhibition during bottle incubations on the measurement of seasonal primary production in a shallow coastal water. Marine Ecology Progress Series, 15:169–180, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/15/m015p169.pdf.

Goldman:1984:PRP

[GD84b] Joel C. Goldman and Mark R. Dennett. Photosynthetic responses of 15 phytoplankton species to ammonium pulsing. Marine Ecology Progress Series, 20:259–264, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/20/m020p259.pdf.

Geider:1988:EGL

[Gei88] R. J. Geider. Estimating the growth and loss rates of phytoplankton from time-series observations of ^{14}C-bicarbonate uptake. Marine Ecology Progress Series, 43:125–138, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/43/m043p125.pdf.

George:1985:DEI

[Geo85] V. S. George. Demographic evaluation of the influence of temperature and salinity on the copepod Eurytemora herdmani. Marine Ecology Progress Series, 21:145–152, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/21/m021p145.pdf.

Gerhart:1984:PAA

[Ger84] D. J. Gerhart. Prostaglandin A2: an agent of chemical defense in the Caribbean gorgonian Plexaura homomalla. Marine Ecology Progress Series, 19:181–187, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/19/m019p181.pdf.

Gerhart:1986:GGE

[Ger86] D. J. Gerhart. Gregariousness in the gorgonian-eating gastropod Cyphoma gibbosum: tests of several possible causes. Marine Ecology Progress Series, 31:255–263, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/31/m031p255.pdf.

Germann:1989:ACM

[Ger89] I. Germann. Aspects of carbon metabolism in relation to autumnal blade abscission in the kelp Pleurophycus gardneri (Phaeophyceae, Laminariales). Marine Ecology Progress Series, 54:179– 188, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
54/m054p179.pdf.

Guary:1982:ESB

[GF82] Jean-Claude Guary and Scott W. Fowler. Experimental studies on the biokinetics of plutonium and americium in the cephalopod Octopus vulgaris. Marine Ecology Progress Series, 7:327–335, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
7/m007p327.pdf.

Gosselin:1989:VMF

[GFG89] S. Gosselin, L. Fortier, and J. A. Gagne. Vulnerability of marine fish larvae to the toxic dinoflagellate Protogonyaulax tamarensis. Marine Ecology Progress Series, 57:1–10, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/57/m057p001.pdf.

Grebmeier:1989:PBCa

[GFM89] J. M. Grebmeier, H. M. Feder, and C. P. McRoy. Pelagic-benthic coupling on the shelf of the northern Bering and Chukchi Seas. II. Benthic community structure. Marine Ecology Progress Series, 51: 253–268, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/51/m051p253.pdf.

Galvao:1989:IGP

[GFS89] H. M. Galvao, A. T. Fritz, and R. Schmaljohann. Ingestion of gametes by protists: fate of surplus reproductive energy in the sea. Marine Ecology Progress Series, 51:215–220, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/51/m051p215.pdf.

Gallucci:1982:REH

[GG82] V. F. Gallucci and B. B. Gallucci. Reproduction and ecology of the hermaphroditic cockle Clinocardium nuttallii (Bivalvia: Cardiidae) in Garrison Bay. Marine Ecology Progress Series, 7:137– 145, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
7/m007p137.pdf.

Gast:1988:VDN

[GG88] V. Gast and K. Gocke. Vertical distribution of number, biomass and size-class spectrum of bacteria in relation to oxic/anoxic conditions in the Central Baltic Sea. Marine Ecology Progress Series, 45:179–186, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/45/m045p179.pdf.

Gotelli:1985:PBT

[GGA85] Nicholas J. Gotelli, Sandra L. Gilchrist, and Lawrence G. Abele. Population biology of Trapezia spp. and other coral-associated decapods. Marine Ecology Progress Series, 21:89–98, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/21/m021p089.pdf.

Gorsky:1989:AIA

[GGV89] G. Gorsky, P. Guilbert, and E. Valenta. The Autonomous Image Analyzer — enumeration, measurement and identification of marine phytoplankton. Marine Ecology Progress Series, 58:133– 142, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
58/m058p133.pdf.

Gast:1983:RPB

[GH83] V. Gast and U. Horstmann. N-remineralization of phyto- and bacterioplankton by the marine ciliate Euplotes vannus. Marine Ecology Progress Series, 13:55–60, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/13/m013p055.pdf.

Gaill:1986:TDS

[GH86] F. Gaill and S. Hunt. Tubes of deep sea hydrothermal vent worms Riftia pachyptila (Vestimentifera) and Alvinella pompejana (Annelida). Marine Ecology Progress Series, 34:267–274, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
34/m034p267.pdf.

Govoni:1983:CFT

[GHC83] J. J. Govoni, D. E. Hoss, and A. J. Chester. Comparative feeding of three species of larval fishes in the Northern Gulf of Mexico: Brevoortia patronus, Leiostomus xanthurus, and Micropogonias undulates. Marine Ecology Progress Series, 13:189–199, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p189.pdf.

Guary:1981:HCT

[GHCH81] J.-C. Guary, J. J. W. Higgo, R. D. Cherry, and M. Heyraud. High concentrations of transuranics and natural radioactive elements in the branchial hearts of the cephalopod Octopus vulgaris. Marine Ecology Progress Series, 4:123–126, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p123.pdf.

Gerlach:1985:SSB

[GHS85] S. A. Gerlach, A. E. Hahn, and M. Schrage. Size spectra of benthic biomass and metabolism. Marine Ecology Progress Series, 26:161– 173, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
26/m026p161.pdf.

Gibbons:1988:IPJ

[Gib88] M. J. Gibbons. Impact of predation by juvenile Clinus superciliosus on phytal meiofauna. are fish important as predators? Marine Ecology Progress Series, 45:13–22, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/45/m045p013.pdf.

Giere:1981:GMO

[Gie81] O. Giere. The gutless marine oligochaete Phallodrilus leukodermatus. Structural studies on an aberrant tubificid associated with bacteria. Marine Ecology Progress Series, 5:353–357, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
5/m005p353.pdf.

Gifford:1985:LCM

[Gif85] D. J. Gifford. Laboratory culture of marine planktonic oligotrichs (Ciliophora, Oligotrichida). Marine Ecology Progress Series, 23: 257–267, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/23/m023p257.pdf.

Gifford:1988:IGM

[Gif88] D. J. Gifford. Impact of grazing by microzooplankton in the Northwest Arm of Halifax Harbour, Nova Scotia. Marine Ecology Progress Series, 47:249–258, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/47/m047p249.pdf.

Gill:1985:RRC

[Gil85] C. W. Gill. The response of a restrained copepod to tactile stimulation. Marine Ecology Progress Series, 21:121–125, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/21/m021p121.pdf.

Glibert:1988:PRG

[GKA88] P. M. Glibert, T. M. Kana, and D. M. Anderson. Photosynthetic response of Gonyaulax tamarensis during growth in a natural bloom and in batch culture. Marine Ecology Progress Series, 42: 303–309, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/42/m042p303.pdf.

Grossi:1987:SIM

[GKM^{+}87] S. M. Grossi, S. T. Kottmeier, R. L. Moe, G. T. Taylor, and C. W. Sullivan. Sea ice microbial communities. VI. Growth and primary production in bottom ice under graded snow cover. Marine Ecology Progress Series, 35:153–164, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/35/m035p153.pdf.

Gocke:1987:STF

[GKOW87] K. Gocke, K. Kremling, C. Osterroht, and A. Wenck. Shortterm fluctuations of microbial and chemical variables during different seasons in coastal Baltic waters. Marine Ecology Progress Series, 40:137–144, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/40/m040p137.pdf.

Greening:1982:DVS

[GL82] H. S. Greening and R. J. Livingston. Diel variation in the structure of seagrass-associated epibenthic macroinvertebrate communities. Marine Ecology Progress Series, 7:147–156, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/7/m007p147.pdf.

Geider:1988:KEG

[GL88a] R. J. Geider and B. S. C. Leadbeater. Kinetics and energetics of growth of the marine choanoflagellate Stephanoeca diplocostata. Marine Ecology Progress Series, 47:169–177, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/47/m047p169.pdf.

Goldman:1988:OEC

[GL88b] N. Goldman and P. J. D. Lambshead. Optimization of the Ewens/Caswell neutral model program for community diversity analysis. Marine Ecology Progress Series, 50:255–261, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/50/
m050p255.pdf.

Gooday:1989:ISD

[GL89a] A. J. Gooday and P. J. D. Lambshead. Influence of seasonally deposited phytodetritus on benthic Foraminiferal populations in the bathyal northeast Atlantic. the species response. Marine Ecology Progress Series, 58:53–67, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/58/m058p053.pdf.

Gradinger:1989:PHA

[GL89b] R. Gradinger and J. Lenz. Picocyanobacteria in the high Arctic. Marine Ecology Progress Series, 52:99–101, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/52/m052p099.pdf.

Grant:1987:GSS

[GLB87] W. S. Grant, R. W. Leslie, and I. I. Becker. Genetic stock structure of the southern African hakes Merluccius capensis and M. paradoxus. Marine Ecology Progress Series, 41:9–20, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/41/m041p009.pdf.

Giere:1982:HCD

[GLD82] O. Giere, G. Liebezeit, and R. Dawson. Habitat conditions and distribution pattern of the gutless oligochaete Phallodrilus leukodermatus. Marine Ecology Progress Series, 8:291–299, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/8/m008p291.pdf.

Gleason:1986:USM

[Gle86] D. F. Gleason. Utilization of salt marsh plants by post-larval brown shrimp: carbon assimilation rates and food preferences. Marine Ecology Progress Series, 31:151–158, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/31/m031p151.pdf.

Giddins:1986:FEM

[GLNR86] R. L. Giddins, J. S. Lucas, M. J. Neilson, and G. N. Richards. Feeding ecology of the mangrove crab Neosarmatium smithi (Crustacea: Decapoda: Sesarmidae). Marine Ecology Progress Series, 33:147–155, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/33/m033p147.pdf.

Gosselin:1986:PCH

[GLT^{+}86] M. Gosselin, L. Legendre, J.-C. Therriault, S. Demers, and M. Rochet. Physical control of the horizontal patchiness of sea-ice microalgae. Marine Ecology Progress Series, 29:289–298, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/29/
m029p289.pdf.

Glynn:1985:NAD

[Gly85] P. W. Glynn. El Niño-associated disturbance to coral reefs and post disturbance mortality by Acanthaster planci. Marine Ecology Progress Series, 26:295–300, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/26/m026p295.pdf.

Gagne:1987:EFM

[GM87] J. A. Gagné and K. H. Mann. Evaluation of four models used to estimate kelp productivity from growth measurements. Marine Ecology Progress Series, 37:35–44, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/37/m037p035.pdf.

Grebmeier:1989:PBCb

[GM89] J. M. Grebmeier and C. P. McRoy. Pelagic-benthic coupling on the shelf of the northern Bering and Chukchi Seas. III. Benthic food supply and carbon cycling. Marine Ecology Progress Series, 53:79– 91, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
53/m053p079.pdf.

Goutx:1987:LEP

[GMB87] M. Goutx, S. Mutaftshiev, and J.-C. Bertrand. Lipid and exopolysaccharide production during hydrocarbon growth of a marine bacterium from the sea surface. Marine Ecology Progress Series, 40:259–265, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/40/m040p259.pdf.

Grebmeier:1988:PBC

[GMF88] J. M. Grebmeier, C. P. McRoy, and H. M. Feder. Pelagic-benthic coupling on the shelf of the northern Bering and Chukchi Seas. I. Food supply source and benthic bio-mass. Marine Ecology Progress Series, 48:57–67, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/48/m048p057.pdf.

Green:1985:FTA

[GMK85] J. M. Green, G. Martel, and E. A. Kingsland. Foraging time allocation in a territorial fish: influence of reproductive activities. Marine Ecology Progress Series, 24:23–26, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/24/m024p023.pdf.

Graf:1988:MSB

[GMQ^{+}88] G. Graf, V. Martens, W. Queisser, P. Weinholz, and A. Altenbach. A multicalorimeter for the study of biological activity in marine sediments. Marine Ecology Progress Series, 45:201–204, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
45/m045p201.pdf.

Gamble:1985:CGD

[GMS85] J. C. Gamble, P. MacLachlan, and D. D. Seaton. Comparative growth and development of autumn and spring spawned Atlantic herring larvae reared in large enclosed ecosystems. Marine Ecology Progress Series, 26:19–33, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/26/m026p019.pdf.

Gremare:1989:SPR

[GMT89] A. Grémare, A. G. Marsh, and K. R. Tenore. Secondary production and reproduction of Capitella capitata type I (Annelida. Polychaeta) during a population cycle. Marine Ecology Progress Series, 51:99–105, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/51/m051p099.pdf.

George-Nascimento:1985:FES

[GNBO85] M. George-Nascimento, R. Bustamante, and C. Oyarzun. Feeding ecology of the South American sea lion Otaria flavescens: food contents and food selectivity. Marine Ecology Progress Series, 21: 135–143, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/21/m021p135.pdf.

Govoni:1986:SFS

[GOAYH86] J. J. Govoni, P. B. Ortner, F. Al-Yamani, and L. C. Hill. Selective feeding of spot, Leiostomus xanthurus, and Atlantic croaker, Micropogonias undulatus, larvae in the northern Gulf of Mexico. Marine Ecology Progress Series, 28:175–183, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/28/m028p175.pdf.

Golikov:1985:ZOA

[Gol85] A. N. Golikov. Zonations and organismic assemblages: comments on the comprehensive review by Pérès (1982). Marine Ecology Progress Series, 23:203–206, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/23/m023p203.pdf.

Gray:1982:OSS

[GP82] J. S. Gray and T. H. Pearson. Objective selection of sensitive species indicative of pollution-induced change in benthic communities. I. Comparative methodology. Marine Ecology Progress Series, 9:111–119, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/9/m009p111.pdf.

Gallegos:1985:VAP

[GP85] C. L. Gallegos and T. Piatt. Vertical advection of phytoplankton and productivity estimates: a dimensional analysis. Marine Ecology Progress Series, 26:125–134, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/26/m026p125.pdf.

Geider:1986:MMP

[GP86] J. R. Geider and T. Piatt. A mechanistic model of photoadaptation in microalgae. Marine Ecology Progress Series, 30:85–92, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
30/m030p085.pdf.

Gadomski:1988:EFD

[GP88a] D. M. Gadomski and J. H. Petersen. Effects of food deprivation on the larvae of two flatfishes. Marine Ecology Progress Series, 44: 103–111, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/44/m044p103.pdf.

Gill:1988:RCD

[GP88b] C. W. Gill and S. A. Poulet. Responses of copepods to dissolved free amino acids. Marine Ecology Progress Series, 43:269–276, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
43/m043p269.pdf.

Glover:1988:PUS

[GPCW88] H. E. Glover, B. B. Prezelin, L. Campbell, and M. Wyman. Picoand ultraplankton Sargasso Sea communities: variability and comparative distributions of Synechococcus spp. and algae. Marine Ecology Progress Series, 49:127–139, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/49/m049p127.pdf.

Geider:1986:SDG

[GPR86] R. J. Geider, T. Piatt, and J. A. Raven. Size dependence of growth and photosynthesis in diatoms: a synthesis. Marine Ecology Progress Series, 30:93–104, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/30/m030p093.pdf.

Guzman:1989:PFR

[GR89] H. M. Guzmán and D. R. Robertson. Population and feeding responses of the corallivorous pufferfish Arothron meleagris to coral mortality in the eastern Pacific. Marine Ecology Progress Series, 55:121–131, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/55/m055p121.pdf.

Grant:1981:STD

[Gra81] J. Grant. Sediment transport and disturbance on an intertidal sandflat: infaunal distribution and recolonization. Marine Ecology Progress Series, 6:249–255, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/6/m006p249.pdf.

Gray:1983:UML

[Gra83] J. S. Gray. Use and misuse of the log-normal plotting method for detection of effects of pollution — a reply to Shaw et al. (1983). Marine Ecology Progress Series, 11:203–204, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/11/m011p203.pdf.

Grant:1984:SMS

[Gra84] J. Grant. Sediment microtopography and shore-bird foraging. Marine Ecology Progress Series, 19:293–296, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/19/m019p293.pdf.

Graf:1986:WIB

[Gra86a] G. Graf. Winter inversion of biomass and activity profile in a marine sediment. Marine Ecology Progress Series, 33:231–235, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
33/m033p231.pdf.

Gray:1986:CCW

[Gra86b] J. Gray. Colleen Cavanaugh: Winner of the Ecology Institute’s IRPE PRIZE 1986 in marine ecology. Marine Ecology Progress Series, 34:9–10, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/34/m034p009.pdf.

Gray:1986:TFW

[Gra86c] J. Gray. Tom Fenchel: Winner of the Ecology Institute Prize 1986 in marine ecology. Marine Ecology Progress Series, 34:7– 8, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
34/m034p007.pdf.

Graf:1987:BEF

[Gra87] G. Graf. Benthic energy flow during a simulated autumn bloom sedimentation. Marine Ecology Progress Series, 39:23–29, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/39/
m039p023.pdf.

Griffith:1980:NFA

[Gri80a] R. J. Griffith. Natural food availability and assimilation in the bivalve Choromytilus meridionalis. Marine Ecology Progress Series, 3:151–156, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/3/m003p151.pdf.

Griffiths:1980:FRA

[Gri80b] R. J. Griffiths. Filtration, respiration and assimilation in the black mussel Choromytilus meridionalis. Marine Ecology Progress Series, 3:63–70, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/3/m003p063.pdf.

Grigg:1983:CSS

[Gri83] R. W. Grigg. Community structure, succession and development of coral reefs in Hawaii. Marine Ecology Progress Series, 11:1– 14, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
11/m011p001.pdf.

Grizzle:1984:PIS

[Gri84] R. E. Grizzle. Pollution indicator species of macrobenthos in a coastal lagoon. Marine Ecology Progress Series, 18:191–200, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
18/m018p191.pdf.

Grigg:1988:RLD

[Gri88] R. W. Grigg. Recruitment limitation of a deep benthic hardbottom octocoral population in the Hawaiian Islands. Marine Ecology Progress Series, 45:121–126, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/45/m045p121.pdf.

Grondahl:1988:IBP

[Grö88] F. Gröndahl. Interactions between polyps of Aurelia aurita and planktonic larvae of scyphozoans: an experimental study. Marine Ecology Progress Series, 45:87–93, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/45/m045p087.pdf.

Grondahl:1989:EGS

[Grö89] F. Gröndahl. Evidence of gregarious settlement of planula larvae of the scyphozoan Aurelia aurita: an experimental study. Marine Ecology Progress Series, 56:119–125, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/56/m056p119.pdf.

Greene:1983:SMH

[GSC83] C. H. Greene, A. Schoener, and E. Corets. Succession on marine hard substrata: the adaptive significance of solitary and colonial strategies in temperature fouling communities. Marine Ecology Progress Series, 13:121–129, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/13/m013p121.pdf.

Gardner:1987:GMP

[GT87] J. P. A. Gardner and M. L. H. Thomas. Growth, mortality and production of organic matter by a rocky intertidal population of Mytilus edulis in the quoddy region of the Bay of Fundy. Marine Ecology Progress Series, 39:31–36, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/39/m039p031.pdf.

Goldman:1981:NTC

[GTG81] J. C. Goldman, C. D. Taylor, and P. M. Glibert. Nonlinear timecourse uptake of carbon and ammonium by marine phytoplankton. Marine Ecology Progress Series, 6:137–148, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/6/m006p137.pdf.

Gulland:1987:NMS

[Gul87] J. A. Gulland. Natural mortality and size. Marine Ecology Progress Series, 39:197–199, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/39/m039p197.pdf.

Gunnill:1980:RSS

[Gun80] F. C. Gunnill. Recruitment and standing stocks in populations of one green alga and five brown algae in the intertidal zone near La Jolla, California, during 1973–1977. Marine Ecology Progress Series, 3:231–243, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/3/m003p231.pdf.

Gosling:1981:EGM

[GW81] E. M. Gosling and N. P. Wilkina. Ecological genetics of the mussels Mytilus edulis and M. galloprovincialis on Irish coasts. Marine Ecology Progress Series, 4:221–227, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p221.pdf.

Grant:1985:STH

[GW85] A. Grant and P. Williamson. Settlement-timing hypothesis: a critique. Marine Ecology Progress Series, 23:193–196, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/23/m023p193.pdf.

Helder:1987:EAQ

[HA87] W. Helder and F. Ø. Andersen. An experimental approach to quantify biologically mediated dissolved silicate transport at the sediment–water interface. Marine Ecology Progress Series, 39:305– 311, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
39/m039p305.pdf.

Hagstrom:1988:MLO

[HAA^{+}88] Å Hagström, F. Azam, A. Andersson, J. Wikner, and F. Rassoulzadegan. Microbial loop in an oligotrophic pelagic marine ecosystem: possible roles of cyanobacteria and nanoflagellates in the organic fluxes. Marine Ecology Progress Series, 49:171–178, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
49/m049p171.pdf.

Hagstrom:1984:BGS

[HAHA84] Å. Hagström, J. W. Ammerman, S. Henrichs, and F. Azam. Bacterioplankton growth in seawater: II. Organic matter utilization during steady-state growth in seawater cultures. Marine Ecology Progress Series, 18:41–48, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/18/m018p041.pdf.

Hamilton:1980:CDU

[Ham80] E. I. Hamilton. Concentration and distribution of uranium in Mytilus edolis and associated materials. Marine Ecology Progress Series, 2:61–73, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/2/m002p061.pdf.

Hammond:1982:AGS

[Ham82a] L. S. Hammond. Analysis of grain-size selection by deposit-feeding holothurians and echinoids (Echinodermata) from a shallow reef lagoon, Discovery Bay, Jamaica. Marine Ecology Progress Series, 8:25–36, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/8/m008p025.pdf.

Hammond:1982:NDF

[Ham82b] L. S. Hammond. Nutrition of deposit-feeding holothuroids and echinoids (Echinodermata) from a shallow reef lagoon, Discovery Bay, Jamaica. Marine Ecology Progress Series, 10:297–305, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
10/m010p297.pdf.

Haney:1986:SSG

[Han86] J. C. Haney. Seabird segregation at Gulf Stream frontal eddies. Marine Ecology Progress Series, 28:279–285, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/28/m028p279.pdf.

Hansen:1989:RTD

[Han89] P. J. Hansen. The red tide dinoflagellate Alexandrium tamarense: effects on behaviour and growth of a tintinnid ciliate. Marine Ecology Progress Series, 53:105–116, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/53/m053p105.pdf.

Hanson:1986:PRF

[HAOC^{+}86] R. B. Hanson, M. T. Alvarez-Ossorio, R. Cal, M. J. Campos, M. Roman, G. Santiago, M. Varela, and J. A. Yoder. Plankton response following a spring upwelling event in the Ria de Arosa, Spain. Marine Ecology Progress Series, 32:101–113, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/32/m032p101.pdf.

Harrison:1982:URS

[Har82] W. G. Harrison. Uptake and recycling of soluble reactive phosphorus by marine microplankton. Marine Ecology Progress Series, 10: 127–135, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/10/m010p127.pdf.

Harriott:1985:MRS

[Har85] V. J. Harriott. Mortality rates of scleractinian corals before and during a mass bleaching event. Marine Ecology Progress Series, 21: 81–88, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/21/m021p081.pdf.

Harding:1986:ODB

[Har86] G. C. Harding. Organochlorine dynamics between zooplankton and their environment, a reassessment. Marine Ecology Progress Series, 33:167–191, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/33/m033p167.pdf.

Hastings:1981:SLV

[Has81] M. H. Hastings. Semi-lunar variations of endogenous circa-tidal rhythms of activity and respiration in the lsopod Eurydice pulchra. Marine Ecology Progress Series, 4:85–90, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/4/m004p085.pdf.

Hopkins:1981:TFV

[HB81] T. L. Hopkins and R. C. Baird. Trophodynamics of the fish Valenciennellus tripunctulatus. I. Vertical distribution, diet and feeding chronology. Marine Ecology Progress Series, 5:1–10, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
5/m005p001.pdf.

Hawkins:1985:SVR

[HB85] A. J. S. Hawkins and B. L. Bayne. Seasonal variation in the relative utilization of carbon and nitrogen by the mussel Mytilus edulis: budgets, conversion efficiencies and maintenance requirements. Marine Ecology Progress Series, 25:181–188, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/25/
m025p181.pdf.

Herrnkind:1986:FRP

[HB86] W. F. Herrnkind and M. J. Butler IV. Factors regulating postlarval settlement and juvenile microhabitat use by spiny lobsters Panulirus argus. Marine Ecology Progress Series, 34:23–30, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
34/m034p023.pdf.

Herman:1980:RWB

[HBFA80] L. M. Herman, C. S. Baker, P. H. Forestell, and R. C. Antinoja. Right whale Balaena glacialis sightings near Hawaii: a clue to the wintering grounds? Marine Ecology Progress Series, 2:271– 275, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
2/m002p271.pdf.

Heuillet:1988:UFE

[HBJ^{+}88] E. Heuillet, A. Bermond, N. Jeanne, S. Puiseux-Dao, and C. Ducauze. Use of factorial experimental design to evaluate how biotic and abiotic factors influence cadmium accumulation in the unicellular alga Dunaliella bioculata. Marine Ecology Progress Series, 44:69–75, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/44/m044p069.pdf.

Hamilton:1980:CDT

[HC80] E. I. Hamilton and R. J. Clifton. Concentration and distribution of the transuranium radionuclides ^{239+240}Pu, ^{238}Pu and ^{241}Am in Mytilus edulis, Fucus vesiculosus and surface sediment of Esk Estuary. Marine Ecology Progress Series, 3:267–277, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/3/m003p267.pdf.

Heinbokel:1986:PTA

[HC86a] J. F. Heinbokel and D. W. Coats. Patterns of tintinnine abundance and reproduction near the edge of seasonal pack-ice in the Weddell Sea, November 1983. Marine Ecology Progress Series, 33:71–80, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
33/m033p071.pdf.

Huber:1986:RUC

[HC86b] M. E. Huber and S. L. Coles. Resource utilization and competition among the five Hawaiian species of Trapezia (Crustacea, Brachyura). Marine Ecology Progress Series, 30:21–31, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/30/
m030p021.pdf.

Hahn:1989:TMS

[HD89] T. Hahn and M. Denny. Tenacity-mediated selective pecdation by oystercatchers on intertidal limpets and its role in maintaining habitat partitioning by ’Collisella’ scabra and Lottia digitalis. Marine Ecology Progress Series, 53:1–10, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/53/m053p001.pdf.

Hay:1988:CDS

[HDFG88] M. E. Hay, J. E. Duffy, W. Fenical, and K. Gustafson. Chemical defense in the seaweed Dictyopteris delicatula: differential effects against reef fishes and amphipods. Marine Ecology Progress Series, 48:185–192, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/48/m048p185.pdf.

Heinbokel:1988:RRP

[Hei88] J. F. Heinbokel. Reproductive rates and periodicities of oceanic tintinnine ciliates. Marine Ecology Progress Series, 47:239–248, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
47/m047p239.pdf.

Heine:1989:EIS

[Hei89] J. N. Heine. Effects of ice scour on the structure of sublittoral marine algal assemblages of St. Lawrence and St. Matthew Islands, Alaska. Marine Ecology Progress Series, 52:253–260, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/52/
m052p253.pdf.

Herndl:1988:EAA

[Her88] G. J. Herndl. Ecology of amorphous aggregations (marine snow) in the Northern Adriatic Sea. II. Microbial density and activity in marine snow and its implication to overall pelagic processes. Marine Ecology Progress Series, 48:265–275, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/48/m048p265.pdf.

Hettler:1989:NUR

[Het89] W. F. Hettler, Jr. Nekton use of regularly-flooded salt-marsh cordgrass habitat in North Carolina, USA. Marine Ecology Progress Series, 56:111–118, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/56/m056p111.pdf.

Harriott:1987:CSP

[HF87] V. J. Harriott and D. A. Fisk. A comparison of settlement plate types for experiments on the recruitment of scleractinian corals. Marine Ecology Progress Series, 37:201–208, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/37/m037p201.pdf.

Hudak:1988:EFO

[HF88] J. P. Hudak and J. A. Fuhrman. Effects of four organic pollutants on the growth of natural marine bacterioplankton populations. Marine Ecology Progress Series, 47:185–194, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/47/m047p185.pdf.

Homziak:1982:MCS

[HFK82] J. Homziak, M. S. Fonseca, and W. J. Kenworthy. Macrobenthic community structure in a transplanted eelgrass (Zostera marina) meadow. Marine Ecology Progress Series, 9:211–221, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
9/m009p211.pdf.

Huggett:1986:SRB

[HG86] J. Huggett and C. L. Griffiths. Some relationships between elevation, physico-chemical variables and biota of intertidal rock pools. Marine Ecology Progress Series, 29:189–197, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/29/m029p189.pdf.

Hitchcock:1986:PPC

[HGD86] Gary L. Hitchcock, Joel C. Goldman, and Mark R. Dennett. Photosynthate partitioning in cultured marine phytoplankton: metabolic patterns in a marine diatom under constant and variable light intensities. Marine Ecology Progress Series, 30:77–84, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
30/m030p077.pdf.

Harvey:1987:CPS

[HGH87] C. Harvey, F.-X Garneau, and J. H. Himmelman. Chemodetection of the predatory sea-star Leptasterias polaris by the whelk Buccinum undatum. Marine Ecology Progress Series, 40:79–86, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
40/m040p079.pdf.

Hoegh-Guldberg:1989:IPD

[HGS89] O. Hoegh-Guldberg and G. J. Smith. Influence of the population density of zooxanthellae and supply of ammonium on the biomass and metabolic characteristics of the reef corals Seriatopora hystrix and Stylophora pistillata. Marine Ecology Progress Series, 57:173– 186, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
57/m057p173.pdf.

Harding:1984:PPC

[HH84a] Lawrence W. Harding, Jr. and J. F. Heinbokel. Periodicities of photosynthesis and cell division: behavior of phase-lagged replicate cultures of Ditylum brightwellii in a diurnally varying photic regime. Marine Ecology Progress Series, 15:225–232, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
15/m015p225.pdf.

Hawkins:1984:FDU

[HH84b] S. J. Hawkins and R. G. Hartnoll. Factors determining the upper limits of intertidal canopy-forming algae. Marine Ecology Progress Series, 20:265–271, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/20/m020p265.pdf.

Hutchings:1984:GPS

[HH84c] J. A. Hutchings and R. L. Haedrich. Growth and population structure in two species of bivalves (Nuculanidae) from the deep sea. Marine Ecology Progress Series, 17:135–142, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/17/m017p135.pdf.

Harrison:1985:IDE

[HH85] W. G. Harrison and L. R. Harris. Isotope-dilution and its effects on measurements of nitrogen and phosphorus uptake by oceanic microplankton. Marine Ecology Progress Series, 27:253– 261, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
27/m027p253.pdf.

Head:1987:CFP

[HH87] E. J. H. Head and L. R. Harris. Copepod feeding patterns before and during a spring bloom in Bedford Basin, Nova Scotia. Marine Ecology Progress Series, 40:221–230, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/40/m040p221.pdf.

Higginbottom:1989:BPS

[HH89] I. R. Higginbottom and G. W. Hosie. Biomass and population structure of a large aggregation of krill near Prydz Bay, Antarctica. Marine Ecology Progress Series, 58:197–203, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/58/m058p197.pdf.

Holt:1985:DPS

[HHA85] G. J. Holt, S. A. Holt, and C. R. Arnold. Diel periodicity of spawning in sciaenids. Marine Ecology Progress Series, 27:1–7, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
27/m027p001.pdf.

Heath:1988:VDH

[HHB88] M. R. Heath, E. W. Henderson, and D. L. Baird. Vertical distribution of herring larvae in relation to physical mixing and illumination. Marine Ecology Progress Series, 47:211–228, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/47/m047p211.pdf.

Hargrave:1984:DPF

[HHD^{+}84] B. T. Hargrave, G. C. Harding, K. F. Drinkwater, T. C. Lambert, and W. G. Harrison. Dynamics of the pelagic food web in St. Georges Bay, southern Gulf of St. Lawrence. Marine Ecology Progress Series, 20:221–240, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/20/m020p221.pdf.

Heinemann:1989:RBD

[HHE89] D. Heinemann, G. Hunt, and I. Everson. Relationships between the distributions of marine avian predators and their prey, Euphausia superba, in Bransfield Strait and southern Drake Passage, Antarctica. Marine Ecology Progress Series, 58:3–16, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/58/
m058p003.pdf.

Herman:1984:PTD

[HHG84] Peter M. J. Herman, Carlo Heip, and Bernadette Guillemijn. Production of Tachidius discipes (Copepoda. Harpacticoida). Marine Ecology Progress Series, 17:271–278, 1984. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/17/m017p271.pdf.

Horrigan:1988:INU

[HHKA88] S. G. Horrigan, Å. Hägström, I. Koike, and F. Azam. Inorganic nitrogen utilization by assemblages of marine bacteria in sea-water culture. Marine Ecology Progress Series, 50:147–150, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
50/m050p147.pdf.

Holligan:1983:VDP

[HHN^{+}83] P. M. Holligan, R. P. Harris, R. C. Newell, D. S. Harbour, R. N. Head, E. A. S. Linley, M. I. Lucas, P. R. G. Tranter, and C. M. Weekley. Vertical distribution and partitioning of organic carbon in mixed, frontal and stratified waters of the English Channel. Marine Ecology Progress Series, 14:111–127, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/14/m014p111.pdf.

Hickman:1981:SDF

[Hic81] C. S. Hickman. Selective deposit feeding by the deep-sea archaeogastropod Bathybembix aeola. Marine Ecology Progress Series, 6:339–342, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/6/m006p339.pdf.

Hickel:1984:SRW

[Hic84] W. Hickel. Seston retention by whatman GF/C glass-fiber filters. Marine Ecology Progress Series, 16:185–191, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/16/m016p185.pdf.

Hicks:1986:DBM

[Hic86] G. R. F. Hicks. Distribution and behaviour of meiofaunal copepods inside and outside seagrass beds. Marine Ecology Progress Series, 31:159–170, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/31/m031p159.pdf.

Hicks:1988:SRN

[Hic88] G. R. F. Hicks. Sediment rafting: a novel mechanism for the smallscale dispersal of intertidal estuarine meiofauna. Marine Ecology Progress Series, 48:69–80, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/48/m048p069.pdf.

Highsmith:1982:RFC

[Hig82] R. C. Highsmith. Reproduction by fragmentation in corals. Marine Ecology Progress Series, 7:207–226, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/7/m007p207.pdf.

Highsmith:1985:FAR

[Hig85] R. C. Highsmith. Floating and algal rafting as potential dispersal mechanisms in brooding invertebrates. Marine Ecology Progress Series, 25:169–179, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/25/m025p169.pdf.

Himmelman:1986:PBG

[Him86] J. H. Himmelman. Population biology of green sea urchins on rocky barrens. Marine Ecology Progress Series, 33:295–306, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/33/
m033p295.pdf.

Hirche:1983:OCF

[Hir83] H.-J. Hirche. Overwintering of Calanus finmarchicus and Calanus helgolandicus. Marine Ecology Progress Series, 11:281–290, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/11/
m011p281.pdf.

Hirata:1986:SSO

[Hir86] T. Hirata. Succession of sessile organisms on experimental plates immersed in Nabeta Bay, Izu Peninsula, Japan. I. Algal succession. Marine Ecology Progress Series, 34:51–61, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/34/m034p051.pdf.

Hirata:1987:SSO

[Hir87] T. Hirata. Succession of sessile organisms on experimental plates immersed in Nabeta Bay, Izu Peninsula, Japan. II. Succession of invertebrates. Marine Ecology Progress Series, 38:25–35, 1987.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/38/
m038p025.pdf.

Hallegraeff:1984:TPS

[HJ84] G. M. Hallegraeff and S. W. Jeffrey. Tropical phytoplankton species and pigments of continental shelf waters of North and North-West Australia. Marine Ecology Progress Series, 20:59– 74, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
20/m020p059.pdf.

Hiwatari:1984:PDL

[HK84] T. Hiwatari and T. Kajihara. Population dynamics and life cycle of Hyale barbicornis (Amphipoda, Crustacea) in a blue mussel zone. Marine Ecology Progress Series, 20:177–183, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/20/m020p177.pdf.

Hammer:1986:DFA

[HK86] K. D. Hammer and G. Kattner. Dissolved free amino acids in the marine environment: a carbon to nitrogen ratio shift during diatom blooms. Marine Ecology Progress Series, 31:35–45, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
31/m031p035.pdf.

Hillmann-Kitalong:1987:CBP

[HKB87] A. Hillmann-Kitalong and C. Birkeland. Comparative biases from pushnet and pullnet zooplankton samples. Marine Ecology Progress Series, 38:131–135, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/38/m038p131.pdf.

Hemminga:1988:DSA

[HKd88] M. A. Hemminga, C. J. Kok, and W. de Munck. Decomposition of Spartina anglica roots and rhizomes in a salt marsh of the Westerschelde Estuary. Marine Ecology Progress Series, 48:175– 184, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
48/m048p175.pdf.

Haug:1986:ISP

[HKS86] T. Haug, E. Kjørsvik, and P. Solemdal. Influence of some physical and biological factors on the density and vertical distribution of Atlantic halibut Hippoglossus hippoglossus eggs. Marine Ecology Progress Series, 33:207–216, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/33/m033p207.pdf.

Huppert:1980:IPC

[HL80] H.-W. Huppert and H. Laudien. Influence of pretreatment with constant and changing temperatures on heat and freezing resistance in gill-epithelium of the mussel Mytilus edulis. Marine Ecology Progress Series, 3:113–120, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/3/m003p113.pdf.

Hines:1982:BNB

[HL82] M. E. Hines and Wm. B. Lyons. Biogeochemistry of nearshore Bermuda sediments. I. Sulfate reduction rates and nutrient generation. Marine Ecology Progress Series, 8:87–94, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/8/m008p087.pdf.

Hahm:1984:PSB

[HL84] W. Hahm and H. Langton. Prey selection based on predator/prey weight ratios for some northwest Atlantic fish. Marine Ecology Progress Series, 19:1–5, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/19/m019p001.pdf.

Harvey:1985:EAB

[HL85] R. W. Harvey and S. N. Luoma. Effect of adherent bacteria and bacterial extracellular polymers upon assimilation by Macoma balthica of sediment-bound Cd, Zn and Ag. Marine Ecology Progress Series, 22:281–289, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/22/m022p281.pdf.

Hudon:1989:NSB

[HL89] C. Hudon and G. Lamarche. Niche segregation between American lobster Homarus americanus and rock crab Cancer irroratus. Marine Ecology Progress Series, 52:155–168, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/52/m052p155.pdf.

Hines:1987:PDH

[HLH87] A. H. Hines, R. N. Lipcius, and A. M. Haddon. Population dynamics and habitat partitioning by size, sex, and molt stage of blue crabs Callinectes sapidus in a sub-estuary of central Chesapeake Bay. Marine Ecology Progress Series, 36:55–64, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/36/m036p055.pdf.

Highsmith:1983:GBT

[HLS83] Raymond C. Highsmith, Rebekka L. Lueptow, and Sandra C. Schonberg. Growth and bioerosion of three massive corals on the Belize barrier reef. Marine Ecology Progress Series, 13:261–271, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p261.pdf.

Holligan:1984:PRN

[HlWPH84] P. M. Holligan, P. J. leB. Williams, D. Purdie, and R. P. Harris. Photosynthesis, respiration and nitrogen supply of plankton populations in stratified, frontal and tidally mixed shelf waters. Marine Ecology Progress Series, 17:201–213, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/17/m017p201.pdf.

Harris:1986:DPA

[HM86] R. P. Harris and A. Malej. Diel patterns of ammonium excretion and grazing rhythms in Calanus helgolandicus in surface stratified waters. Marine Ecology Progress Series, 31:75–85, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/31/m031p075.pdf.

Herndl:1987:IPL

[HM87] G. J. Herndl and V. Malacic. Impact of the pycnocline layer on bacterioplankton: diel and spatial variations in microbial parameters in the stratified water column of the Gulf of Trieste (Northern Adriatic Sea). Marine Ecology Progress Series, 38:295–303, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
38/m038p295.pdf.

Heron:1988:LWR

[HMD88] A. C. Heron, P. S. McWilliam, and G. Dal Pont. Length–weight relation in the salp Thalia democratica and potential of salps as a source of food. Marine Ecology Progress Series, 42:125–132, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
42/m042p125.pdf.

Harding:1985:PPC

[HMF85] Lawrence W. Harding, Jr., Blanche W. Meeson, and Thomes R. Fisher, Jr. Photosynthesis patterns in Chesapeake Bay phytoplankton: short- and long-term responses of P-I curve parameters to light. Marine Ecology Progress Series, 26:99–111, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/26/
m026p099.pdf.

Harding:1989:SDS

[HMH89] L. W. Harding, M. E. Mallonee, Jr., and K. W. Henderson. Spectral distribution and species-specific photosynthetic performance of natural populations of Prorocentrum mariae-lebouriae (Dinophyceae) in the Chesapeake Bay. Marine Ecology Progress Series, 52:261–272, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/52/m052p261.pdf.

Hudak:1988:MPA

[HMLF88] J. P. Hudak, J. McDaniel, S. Lee, and J. A. Fuhrman. Mineralization potentials of aromatic hydrocarbons by estuarine microorganisms: variations with season, location, and bacterioplankton production. Marine Ecology Progress Series, 47:97–102, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/47/
m047p097.pdf.

Heath:1987:ICT

[HMM87] M. R. Heath, P. M. MacLachlan, and J. H. A. Martin. Inshore circulation and transport of herring larvae off the north coast of Scotland. Marine Ecology Progress Series, 40:11–23, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/40/m040p011.pdf.

Harding:1983:PDP

[HMT83] Lawrence W. Harding, Jr., B. W. Meeson, and M. A. Tyler. Photoadaptation and diel periodicity of photosynthesis in the dinoflagellate Prorocentrum mariaelebouriae. Marine Ecology Progress Series, 13:73–85, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/13/m013p073.pdf.

Horsted:1988:RZS

[HNR^{+}88] S. J. Horsted, T. G. Nielsen, B. Riemann, J. Pock-Steen, and P. K. Bjornsen. Regulation of zooplankton by suspension-feeding bivalves and fish in estuarine enclosures. Marine Ecology Progress Series, 48:217–224, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/48/m048p217.pdf.

Hagerman:1980:SDA

[HO80] L. Hagerman and J. Ostrup. Seasonal and diel activity variations in the shrimp Palaemon adspersus from a brackish, non-tidal area. Marine Ecology Progress Series, 2:329–335, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/2/m002p329.pdf.

Hobson:1988:VFD

[Hob88] L. A. Hobson. Volume-frequency distributions of marine phytoplankton in a temperate neritic environment. Marine Ecology Progress Series, 48:47–55, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/48/m048p047.pdf.

Hockin:1981:SEE

[Hoc81] D. C. Hockin. A simple elutriator for extracting meiofauna from sediment matrices. Marine Ecology Progress Series, 4:241–242, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
4/m004p241.pdf.

Hodgson:1985:ADP

[Hod85] G. Hodgson. Abundance and distribution of planktonic coral larvae in Kaneohe Bay, Oahu, Hawaii. Marine Ecology Progress Series, 26:61–71, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/26/m026p061.pdf.

Hoelzer:1988:JMP

[Hoe88] G. A. Hoelzer. Juvenile movement patterns in a territorial scorpaenid fish before and during settlement. Marine Ecology Progress Series, 45:193–195, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/45/m045p193.pdf.

Hollibaugh:1988:LTM

[Hol88] J. T. Hollibaugh. Limitations of the [^{3}H]thymidine method for estimating bacterial productivity due to thymidine metabolism. Marine Ecology Progress Series, 43:19–30, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/43/m043p019.pdf.

Hoppe:1983:SEA

[Hop83] H.-G. Hoppe. Significance of exoenzymatic activities in the ecology of brackish water: measurements by means of methylumbelliferylsubstrates. Marine Ecology Progress Series, 11:299–308, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/11/
m011p299.pdf.

Hoppe:1988:GRP

[Hop88] W. F. Hoppe. Growth, regeneration and predation in three species of large coral reef sponges. Marine Ecology Progress Series, 50: 117–125, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/50/m050p117.pdf.

Hovenkamp:1989:GDL

[Hov89] F. Hovenkamp. Growth differences in larval plaice Pleuronectes platessa in the Southern Bight of the North Sea as indicated by otolith increments and RNA/DNA ratios. Marine Ecology Progress Series, 58:205–215, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/58/m058p205.pdf.

Howard:1985:MST

[How85] R. K. Howard. Measurements of short-term turnover of epifauna within seagrass beds using an in situ staining method. Marine Ecology Progress Series, 22:163–168, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/22/m022p163.pdf.

Hargrave:1986:DBF

[HP86a] B. T. Hargrave and G. A. Phillips. Dynamics of the benthic food web in St. Georges Bay, southern Gulf of St. Lawrence. Marine Ecology Progress Series, 31:277–294, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/31/m031p277.pdf.

Harkantra:1986:CSS

[HP86b] S. N. Harkantra and A. H. Parulekar. Community structure of sand-dwelling macrofauna of an estuarine beach in Goa, India. Marine Ecology Progress Series, 30:291–294, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/30/m030p291.pdf.

Hargrave:1989:DTO

[HP89] B. T. Hargrave and G. A. Phillips. Decay times of organic carbon in sedimented detritus in a macrodidal estuary. Marine Ecology Progress Series, 56:271–279, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/56/m056p271.pdf.

Harrison:1989:DNU

[HPC89] P. J. Harrison, J. S. Parslow, and H. L. Conway. Determination of nutrient uptake kinetic parameters: a comparison of methods. Marine Ecology Progress Series, 52:301–312, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/52/m052p301.pdf.

Herndl:1989:BCM

[HPF89] G. J. Herndl, P. Peduzzi, and N. Fanuko. Benthic community metabolism and microbial dynamics in the Gulf of Trieste (Northern Adriatic Sea). Marine Ecology Progress Series, 53:169–178, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
53/m053p169.pdf.

Harding:1987:VDD

[HPV^{+}87] G. C. Harding, J. D. Pringle, W. P. Vass, S. Pearre, Jr., and S. J. Smith. Vertical distribution and daily movements of larval lobsters Homarus americanus over Browns Bank, Nova Scotia. Marine Ecology Progress Series, 41:29–41, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/41/m041p029.pdf.

Houde:1987:EFQ

[HR87] S. E. L. Houde and M. R. Roman. Effects of food quality on the functional ingestion response of the copepod Acartia tonsa. Marine Ecology Progress Series, 40:69–77, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/40/m040p069.pdf.

Henley:1989:TCP

[HR89] W. J. Henley and J. Ramus. Time course of physiological response of Ulva rotundata to growth irradiance transitions. Marine Ecology Progress Series, 54:171–177, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/54/m054p171.pdf.

Hughes:1981:SSM

[HS81] R. N. Hughes and R. Seed. Size selection of mussels by the blue crab Callinectes sapidus: Energy maximizer or time minimizer? Marine Ecology Progress Series, 6:83–89, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/6/m006p083.pdf.

Hannach:1985:EDB

[HS85] G. Hannach and B. Santelices. Ecological differences between the isomorphic reproductive phases of two species of Indaea (Rhodophyta. Gigartinales). Marine Ecology Progress Series, 22: 291–303, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/22/m022p291.pdf.

Hagerman:1986:BTA

[HS86] L. Hagerman and A. Szaniawska. Behaviour, tolerance and anaerobic metabolism under hypoxia in the brackish-water shrimp Crangon crangon. Marine Ecology Progress Series, 34:125–132, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/34/
m034p125.pdf.

Harvell:1987:PPT

[HS87] C. D. Harvell and T. H. Suchanek. Partial predation on tropical gorgonians by Cyphoma gibbosum (Gastropoda). Marine Ecology Progress Series, 38:37–44, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/38/m038p037.pdf.

Hagerman:1988:RVC

[HS88a] L. Hagerman and A. Szaniawska. Respiration, ventilation and circulation under hypoxia in the glacial relict Saduria (Mesidotea) entomon. Marine Ecology Progress Series, 47:55–63, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/47/
m047p055.pdf.

Haney:1988:SFT

[HS88b] J. Christopher Haney and Amy E. Stone. Seabird foraging tactics and water clarity are plunge divers really in the clear? Marine Ecology Progress Series, 49:1–9, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/49/m049p001.pdf.

Hawkins:1984:FRA

[HSB^{+}84] A. J. S. Hawkins, P. N. Salkeld, B. L. Bayne, E. Gnaiger, and D. M. Lowe. Feeding and resource allocation in the mussel Mytilus edulis: evidence for time-averaged optimization. Marine Ecology Progress Series, 20:273–287, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/20/m020p273.pdf.

Hoppe:1988:STD

[HSB88] H.-G. Hoppe, W. Schramm, and P. Bacolod. Spatial and temporal distribution of pelagic microorganisms and their proteolytic activity over a partly destroyed coral reef. Marine Ecology Progress Series, 44:95–102, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/44/m044p095.pdf.

Hopkinson:1987:MRA

[HSD87] C. S. Hopkinson, Jr., B. F. Sherr, and H. W. Ducklow. Microbial regeneration of ammonium in the water column of Davies Reef, Australia. Marine Ecology Progress Series, 41:147–153, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
41/m041p147.pdf.

Hummel:1988:BGM

[HSdWM88] H. Hummel, A. B. J. Sepers, L. de Wolf, and F. W. Melissen. Bacterial growth of the marine sponge Halichondria panicea induced by reduced waterflow rate. Marine Ecology Progress Series, 42: 195–198, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/42/m042p195.pdf.

Hsiao:1988:SSV

[Hsi88] S. I. C. Hsiao. Spatial and seasonal variations in primary production of sea ice microalgae and phytoplankton in Frobisher Bay, Arctic Canada. Marine Ecology Progress Series, 44:275–285, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
44/m044p275.pdf.

Huntley:1986:CMR

[HSRM86] M. Huntley, P. Sykes, S. Rohan, and V. Marin. Chemicallymediated rejection of dinoflagellate prey by the copepods Calanus pacificus and Paracalanus parvus: mechanism, occurrence and significance. Marine Ecology Progress Series, 28:105–120, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/28/
m028p105.pdf.

Hopkinson:1989:SFM

[HSW89] C. S. Hopkinson, Jr., B. Sherr, and W. J. Wiebe. Size fractionated metabolism of coastal microbial plankton. Marine Ecology Progress Series, 51:155–166, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/51/m051p155.pdf.

Hanson:1981:MMI

[HT81] R. B. Hanson and K. R. Tenore. Microbial metabolism and incorporation by the polychaete Capitella capitata of aerobically and anaerobically decomposed detritus. Marine Ecology Progress Series, 6:299–307, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/6/m006p299.pdf.

Haylor:1984:GED

[HTC84] G. S. Haylor, J. P. Thorpe, and M. A. Carter. Genetic and ecological differentiation between sympatric colour morphs of the common intertidal sea anemone Actinia equine. Marine Ecology Progress Series, 16:281–289, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/16/m016p281.pdf.

Hewitt:1985:CMY

[HTL85] R. P. Hewitt, G. H. Theilacker, and N. C. H. Lo. Causes of mortality in young jack mackerel. Marine Ecology Progress Series, 26:1– 10, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
26/m026p001.pdf.

Hughes:1984:MSD

[Hug84] R. G. Hughes. A model of the structure and dynamics of benthic marine invertebrate communities. Marine Ecology Progress Series, 15:1–11, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/15/m015p001.pdf.

Hughes:1987:SDG

[Hug87] T. P. Hughes. Skeletal density and growth form of corals. Marine Ecology Progress Series, 35:259–266, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/35/m035p259.pdf.

Herndl:1985:HNC

[HVK85] G. J. Herndl, B. Velimirov, and R. E. Krauss. Heterotrophic nutrition and control of bacterial density in the coelenteron of the giant sea anemone Stoichactis giganteum. Marine Ecology Progress Series, 22:101–105, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/22/m022p101.pdf.

Hopkinson:1982:LSM

[HW82] C. S. Hopkinson and R. L. Wetzel. In situ measurements of nutrient and oxygen fluxes in a coastal marine benthic community. Marine Ecology Progress Series, 10:29–35, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/10/m010p029.pdf.

Help:1988:ACA

[HWC^{+}88] C. Help, R. M. Warwick, M. R. Carr, P. M. J. Herman, R. Huys, N. Smol, and K. Van Holsbeke. Analysis of community attributes of the benthic meiofauna of Frierfjord/Langesundfjord. Marine Ecology Progress Series, 46:171–180, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/46/m046p171.pdf.

Hunte:1988:RPR

[HY88] W. Hunte and D. Younglao. Recruitment and population recovery of Diadema antillarum (Echinodermata; Echinoidea) in Barbados. Marine Ecology Progress Series, 45:109–119, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/45/m045p109.pdf.

Harrison:1988:SSR

[HYT^{+}88] P. J. Harrison, P. W. Yu, P. A. Thompson, N. M. Price, and D. J. Phillips. Survey of selenium requirements in marine phytoplankton. Marine Ecology Progress Series, 47:89–96, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/47/m047p089.pdf.

Ittekkot:1981:DFC

[IBMD81] V. Ittekkot, U. Brockmann, W. Michaelis, and E. T. Degens. Dissolved free and combined carbohydrates during a phytoplankton Bloom in the Northern North Sea. Marine Ecology Progress Series, 4:299–305, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/4/m004p299.pdf.

Ibanez:1988:LTC

[ID88] Frédéric Ibanez and Jean-Claude Dauvin. Long-term changes (1977 to 1987) in a muddy fine sand Abra alba–Melinna palmata community from the Western English Channel: multivariate time-series analysis. Marine Ecology Progress Series, 49:65– 81, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
49/m049p065.pdf.

Ishida:1986:EOO

[IEK86] Y. Ishida, M. Eguchi, and H. Kadota. Existence of obligately oligotrophic bacteria as a dominant population in the South China Sea and West Pacific Ocean. Marine Ecology Progress Series, 30: 197–203, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/30/m030p197.pdf.

Ignatiades:1987:SAS

[Ign87] L. Ignatiades. A simple apparatus for sampling sea surface microlayers. Marine Ecology Progress Series, 39:207–208, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/39/m039p207.pdf.

Iturriaga:1986:CCS

[IM86] R. Iturriaga and B. G. Mitchell. Chroococcoid cyanobacteria: a significant component in the food web dynamics of the open ocean. Marine Ecology Progress Series, 28:291–297, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/28/m028p291.pdf.

Iturriaga:1988:TSV

[IM88] R. Iturriaga and J. Marra. Temporal and spatial variability of chroococcoid cyanobacteria Synechococcus spp. specific growth rates and their contribution to primary production in the Sargasso Sea. Marine Ecology Progress Series, 44:175–181, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/44/m044p175.pdf.

Iwai:1981:CMG

[IR81] T. Iwai and H. Rosenthal. Ciliary movements in guts of early clupeoid and salangid larvae. Marine Ecology Progress Series, 4: 365–367, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/4/m004p365.pdf.

Iseki:1981:POM

[Ise81] K. Iseki. Particulate organic matter transport to the deep sea by salp fecal pellets. Marine Ecology Progress Series, 5:55–60, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
5/m005p055.pdf.

Iseki:1981:EPB

[ITBW81] K. Iseki, M. Takahashi, E. Bauerfeind, and C. S. Wong. Effects of polychlorinated biphenyls (PCBs) on a marine plankton population and sedimentation in controlled ecosystem enclosures. Marine Ecology Progress Series, 5:207–214, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/5/m005p207.pdf.

Jacobi:1985:ETR

[JA85] C. C. Jacobi and K. Anger. Effect of temperature on respiration of larval stages of Hyas araneus and H. coarctatus (Decapoda, Majidae). Marine Ecology Progress Series, 26:181–186, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/26/m026p181.pdf.

Jacobs:1980:EVA

[Jac80] R. P. W. M. Jacobs. Effects of the ’Amoco Cadiz’ oil spill on the seagrass community at Roscoff with special reference to the benthic lnfauna. Marine Ecology Progress Series, 2:207–212, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
2/m002p207.pdf.

Jacobi:1987:STD

[Jac87] C. M. Jacobi. Spatial and temporal distribution of Amphipoda associated with mussel beds from the Bay of Santos (Brazil). Marine Ecology Progress Series, 35:51–58, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/35/m035p051.pdf.

Jansson:1985:STD

[JAN85] B.-O. Jansson, G. Aneer, and S. Nellbring. Spatial and temporal distribution of the demersal fish fauna in a Baltic archipelago as estimated by SCUBA census. Marine Ecology Progress Series, 23: 31–43, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/23/m023p031.pdf.

Jensen:1988:DCB

[JAS88] M. H. Jensen, T. K. Andersen, and J. Sørensen. Denitrification in coastal bay sediment: regional and seasonal variation in Aarhus Bight, Denmark. Marine Ecology Progress Series, 48:155–162, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
48/m048p155.pdf.

Jayakody:1989:SOS

[Jay89] D. S. Jayakody. Size at onset of sexual maturity and onset of spawning in female Panulirus homarus (Crustacea: Decapoda: Palinuridae) in Sri Lanka. Marine Ecology Progress Series, 57:83– 87, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
57/m057p083.pdf.

Juniper:1986:WCD

[JB86] S. Kim Juniper and Ralph O. Brinkhurst. Water-column dark CO_{2} fixation and bacterial-mat growth in intermittently anoxic Saanich Inlet, British Columbia. Marine Ecology Progress Series, 33:41– 50, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
33/m033p041.pdf.

Jahnke:1983:PAE

[JBAH83] J. Jahnke, U. H. Brockmann, L. Aletsee, and K. D. Hammer. Phytoplankton activity in enclosed and free marine ecosystems in a southern Norwegian fjord during spring 1979. Marine Ecology Progress Series, 14:19–28, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/14/m014p019.pdf.

Johnson:1982:PAD

[JBS82] K. M. Johnson, C. M. Burney, and J. McN. Sieburth. Precise and accurate determination by infrared photometry of CO_{2} dynamics in marine ecosystems. Marine Ecology Progress Series, 10:251– 256, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
10/m010p251.pdf.

Johansson:1986:VFC

[JCL86] C. Johansson, D. J. Cain, and S. N. Luoma. Variability in the fractionation of Cu, Ag, and Zn among cytosolic proteins in the bivalve Macoma balthica. Marine Ecology Progress Series, 28:87– 97, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
28/m028p087.pdf.

Jennes:1985:ETC

[JD85] M. I. Jennes and G. C. A. Duineveld. Effects of tidal currents on chlorophyll-a content of sandy sediments in the southern North Sea. Marine Ecology Progress Series, 21:283–287, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/21/m021p283.pdf.

Jensen:1981:PCS

[Jen81] P. Jensen. Phyto-chemical sensitivity and swimming behaviour of the free-living marine nematode Chromadorita tenuis. Marine Ecology Progress Series, 4:203–206, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p203.pdf.

Jensen:1982:RBF

[Jen82] P. Jensen. Reproductive behaviour of the free-living marine nematode Chromadorita tenuis. Marine Ecology Progress Series, 10:89– 95, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
10/m010p089.pdf.

Jensen:1985:PBC

[Jen85] K. T. Jensen. The presence of the bivalve Cerastoderma edule affects migration, survival and reproduction of the amphipod Corophium volutator. Marine Ecology Progress Series, 25:269– 277, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
25/m025p269.pdf.

Jensen:1987:FEF

[Jen87] P. Jensen. Feeding ecology of free-living aquatic nematodes. Marine Ecology Progress Series, 35:187–196, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/35/m035p187.pdf.

Jenkins:1988:MFS

[Jen88] G. P. Jenkins. Micro- and fine-scale distribution of microplankton in the feeding environment of larval flounder. Marine Ecology Progress Series, 43:233–244, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/43/m043p233.pdf.

Jernakoff:1985:IBL

[Jer85] P. Jernakoff. Interactions between the limpet Patelloida latistrigata and algae on an intertidal rock platform. Marine Ecology Progress Series, 23:71–78, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/23/m023p071.pdf.

Jernakoff:1986:EII

[Jer86] P. Jernakoff. Experimental investigation of interactions between the perennial red alga Gelidium pusillum and barnacles on a New South Wales rocky shore. Marine Ecology Progress Series, 28:259– 263, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
28/m028p259.pdf.

James:1988:EPS

[JF88] A. G. James and K. P. Findlay. Effect of particle size and concentration on feeding behaviour, selectivity and rates of food ingestion by the Cape anchovy Engraulis capensis. Marine Ecology Progress Series, 50:275–294, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/50/m050p275.pdf.

Jorgensen:1986:BP

[JFK^{+}86] C. B. Jørgensen, P. Famme, H. S. Kristensen, P. S. Larsen, F. Møhlenberg, and H. U. Riisgård. The bivalve pump. Marine Ecology Progress Series, 34:69–77, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/34/m034p069.pdf.

Jacoby:1989:EZM

[JG89] C. A. Jacoby and J. G. Greenwood. Emergent zooplankton in Moreton Bay, Queensland, Australia: seasonal, lunar, and diel patterns in emergence and distribution with respect to substrata. Marine Ecology Progress Series, 51:131–154, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/51/m051p131.pdf.

Jeffrey:1980:SPSa

[JH80a] S. W. Jeffrey and G. M. Hallegraeff. Studies of phytoplankton species and photosynthetic pigments in a warm core eddy of the East Australian current. I. Summer populations. Marine Ecology Progress Series, 3:285–294, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/3/m003p285.pdf.

Jeffrey:1980:SPSb

[JH80b] S. W. Jeffrey and G. M. Hallegraeff. Studies of phytoplankton species and photosynthetic pigments in a warm core eddy of the East Australian current. II. A note on pigment methodology. Marine Ecology Progress Series, 3:295–301, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/3/m003p295.pdf.

Johns:1980:YUS

[JH80c] D. M. Johns and W. H. Howell. Yolk utilization in summer flounder (Paralichthys dentatus) embryos and larvae reared at two temperatures. Marine Ecology Progress Series, 2:1–8, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/2/m002p001.pdf.

Jeffrey:1987:CDT

[JH87] S. W. Jeffrey and G. M. Hallegraeff. Chlorophyllase distribution in ten classes of phytoplankton: a problem for chlorophyll analysis. Marine Ecology Progress Series, 35:293–304, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/35/m035p293.pdf.

Jackson:1985:MRA

[JJN85] G. A. Jackson, D. E. James, and W. J. North. Morphological relationships among fronds of giant kelp Macrocystis pyrifera off La Jolla, California. Marine Ecology Progress Series, 26:261–270, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
26/m026p261.pdf.

Jorgensen:1980:UAAa

[JK80a] N. O. G. Jørgensen and E. Kristensen. Uptake of amino acids by three species of Nereis (Annelida: Polychaeta). I. Transport kinetics and net uptake from natural concentrations. Marine Ecology Progress Series, 3:329–340, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/3/m003p329.pdf.

Jorgensen:1980:UAAb

[JK80b] N. O. G. Jørgensen and E. Kristensen. Uptake of amino acids by three species of Nereis (Annelida: Polychaeta). II. Effects of anaerobiosis. Marine Ecology Progress Series, 3:341–346, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/3/m003p341.pdf.

Jagger:1988:FCP

[JKJ88] R. A. Jagger, W. J. Kimmerer, and G. P. Jenkins. Food of the cladoceran Podon intermedius in a marine embayment. Marine Ecology Progress Series, 43:245–250, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/43/m043p245.pdf.

Jorgensen:1984:CMN

[JKMR84] C. B. Jørgensen, T. Kiørboe, F. Møhlenberg, and H. U. Riisgård. Ciliary and mucus-net filter feeding, with special reference to fluid mechanical characteristics. Marine Ecology Progress Series, 15: 283–292, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/15/m015p283.pdf.

Jahn:1986:FSD

[JL86] A. E. Jahn and R. J. Lavenberg. Fine-scale distribution of nearshore, suprabenthic fish larvae. Marine Ecology Progress Series, 31:223–231, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/31/m031p223.pdf.

Jorgensen:1988:MPP

[JLMR88] C. B. Jørgensen, P. S. Larsen, F. Møhlenberg, and H. U. Riisgård. The mussel pump: properties and modeling. Marine Ecology Progress Series, 45:205–216, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/45/m045p205.pdf.

Jones:1985:LTG

[JM85a] R. D. Jones and R. Y. Morita. Low-temperature growth and whole-cell kinetics of a marine ammonium oxidizer. Marine Ecology Progress Series, 21:239–243, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/21/m021p239.pdf.

Jones:1985:SMA

[JM85b] R. D. Jones and R. Y. Morita. Survival of a marine ammonium oxidizer under energy-source deprivation. Marine Ecology Progress Series, 26:175–179, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/26/m026p175.pdf.

Jones:1984:MEL

[JMG84] R. D. Jones, R. Y. Morita, and R. P. Griffiths. Method for estimating in situ chemo-lithotrophic ammonium oxidation using carbon monoxide oxidation. Marine Ecology Progress Series, 17:259–269, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
17/m017p259.pdf.

Jorgensen:1986:NRB

[JMSK86] C. Barker Jørgensen, F. Møhlenberg, and O. Sten-Knudsen. Nature of relation between ventilation and oxygen consumption in filter feeders. Marine Ecology Progress Series, 29:73–88, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/29/
m029p073.pdf.

Jellett:1989:NOF

[JNCS89] J. F. Jellett, J. A. Novitsky, J. A. Cantley, and R. E. Scheibling. Non-occurrence of free-living Paramoeba invadens in water and sediments of Halifax Harbour, Nova Scotia, Canada. Marine Ecology Progress Series, 56:205–209, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/56/m056p205.pdf. See com-

ment [Mil90] and rebuttal [JNCS90].

Jellett:1990:RRJ

[JNCS90] J. F. Jellett, J. A. Novitsky, J. A. Cantley, and R. E. Scheibling. Rebuttal to R. J. Miller’s comment on Jellett et al. (1989). Marine Ecology Progress Series, 63:307–308, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/63/

m063p307.pdf. See [JNCS89, Mil90].

Jochem:1989:DIA

[Joc89] F. Jochem. Distribution and importance of autotrophic ultraplankton in a boreal inshore area (Kiel Bight, Western Baltic). Marine Ecology Progress Series, 53:153–168, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/53/m053p153.pdf.

Johannes:1980:ESS

[Joh80] R. E. Johannes. The ecological significance of the submarine discharge of groundwater. Marine Ecology Progress Series, 3:365– 373, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
3/m003p365.pdf.

Johns:1981:PSCa

[Joh81a] D. M. Johns. Physiological studies on Cancer irroratus larvae. I. Effects of temperature and salinlty on survival, development rate and size. Marine Ecology Progress Series, 5:75–83, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/5/m005p075.pdf.

Johns:1981:PSCb

[Joh81b] D. M. Johns. Physiological studies on Cancer irroratus larvae II. Effects of temperature and salinity on physiological performance. Marine Ecology Progress Series, 6:309–315, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/6/m006p309.pdf.

Johns:1982:PSC

[Joh82] D. M. Johns. Physiological studies on Cancer irroratus larvae. III. Effects of temperature and salinity on the partitioning of energy resources during development. Marine Ecology Progress Series, 8: 75–85, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/8/m008p075.pdf.

Johnson:1988:DME

[Joh88] D. Johnson. Development of Mytilus edulis embryos: a bio-assay for polluted waters. Marine Ecology Progress Series, 46:135–138, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p135.pdf.

Jonsson:1986:PSS

[Jon86] P. R. Jonsson. Particle size selection, feeding rates and growth dynamics of marine planktonic oligotrichous ciliates (Ciliophora: Oligotrichina). Marine Ecology Progress Series, 33:265–277, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/33/
m033p265.pdf.

Jonsson:1989:VDP

[Jon89] P. R. Jonsson. Vertical distribution of planktonic ciliates — an experimental analysis of swimming behaviour. Marine Ecology Progress Series, 52:39–53, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/52/m052p039.pdf.

Joint:1986:SPP

[JOP86] I. R. Joint, N. J. P. Owens, and A. J. Pomeroy. Seasonal production of photosynthetic picoplankton and nanoplankton in the Celtic Sea. Marine Ecology Progress Series, 28:251–258, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/28/
m028p251.pdf.

Jorgensen:1982:HAO

[Jør82] N. O. G. Jørgensen. Heterotrophic assimilation and occurrence of dissolved free amino acids in a shallow estuary. Marine Ecology Progress Series, 8:145–159, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/8/m008p145.pdf.

Jorgensen:1983:FMA

[Jør83] C. B. Jørgensen. Fluid mechanical aspects of suspension feeding. Marine Ecology Progress Series, 11:89–103, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/11/m011p089.pdf.

Josefson:1987:LSP

[Jos87] A. B. Josefson. Large-scale patterns of dynamics in subtidal macrozoobenthic assemblages in the Skagerrak: effects of a production-related factor? Marine Ecology Progress Series, 38:13– 23, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
38/m038p013.pdf.

Joint:1984:EMS

[JP84] I. R. Joint and R. K. Pipe. An electron microscope study of a natural population of picoplankton from the Celtic Sea. Marine Ecology Progress Series, 20:113–118, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/20/m020p113.pdf.

Jones:1985:LCM

[JP85] R. D. Jones and F. G. Prahl. Lipid composition of a marine ammonium oxidizer grown at 5 and 25 ^{∘}C. Marine Ecology Progress Series, 26:157–159, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/26/m026p157.pdf.

Joint:1987:AHB

[JP87] I. R. Joint and A. J. Pomroy. Activity of heterotrophic bacteria in the euphotic zone of the Celtic Sea. Marine Ecology Progress Series, 41:155–165, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/41/m041p155.pdf.

Joint:1988:AEP

[JP88] I. R. Joint and A. J. Pomroy. Allometric estimation of the productivity of phytoplankton assemblages. Marine Ecology Progress Series, 47:161–168, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/47/m047p161.pdf.

Jorgensen:1985:SCN

[JS85] B. B. Jørgensen and J. Sørensen. Seasonal cycles of O_{2}, NO^{3−} and S0_{4}^{2−} reduction in estuarine sediments: the significance of an NO^{3−} reduction maximum in spring. Marine Ecology Progress Series, 24: 65–74, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/24/m024p065.pdf.

Johnson:1986:SRP

[JS86] T. O. Johnson and W. O. Smith, Jr. Sinking rates of phytoplankton assemblages in the Weddell Sea marginal ice zone. Marine Ecology Progress Series, 33:131–137, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/33/m033p131.pdf.

Johnson:1987:SDE

[JS87a] S. C. Johnson and R. E. Scheibling. Structure and dynamics of epifaunal assemblages on intertidal macroalgae Ascophyllum nodosum and Fucus vesiculosus in Nova Scotia, Canada. Marine Ecology Progress Series, 37:209–227, 1987. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/37/m037p209.pdf.

Juniper:1987:CSB

[JS87b] S. K. Juniper and M. Sibuet. Cold seep benthic communities in the Japan subduction zones: spatial organization, trophic strategies and evidence for temporal evolution. Marine Ecology Progress Series, 40:115–126, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/40/m040p115.pdf.

Jorgensen:1988:TAM

[JS88] K. S. Jørgensen and J. Sørensen. Two annual maxima of nitrate reduction and denitrification in estuarine sediment (Norsminde Fjord, Denmark). Marine Ecology Progress Series, 48:147–154, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
48/m048p147.pdf.

Jeffries:1985:CCA

[JT85] H. P. Jeffries and M. Terceiro. Cycle of changing abundances in the fishes of the Narragansett Bay area. Marine Ecology Progress Series, 25:239–244, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/25/m025p239.pdf.

Johannes:1983:TPN

[JWC83a] R. E. Johannes, W. J. Wiebe, and C. J. Crossland. Three patterns of nutrient flux in a coral reef community. Marine Ecology Progress Series, 12:131–136, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/12/m012p131.pdf.

Johannes:1983:LLC

[JWC^{+}83b] R. E. Johannes, W. J. Wiebe, C. J. Crossland, D. W. Rimmer, and S. V. Smith. Latitudinal limits of coral reef growth. Marine Ecology Progress Series, 11:105–111, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/11/m011p105.pdf.

Kunisch:1984:VDG

[KA84] M. Kunisch and K. Anger. Variation in development and growth rates of larval and juvenile spider crabs Hyas araneus reared in the laboratory. Marine Ecology Progress Series, 15:293–301, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
15/m015p293.pdf.

Kepkay:1985:AAM

[KA85] P. E. Kepkay and F. Ø. Andersen. Aerobic and anaerobic metabolism of a sediment enriched with Spartina detritus. Marine Ecology Progress Series, 21:153–161, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/21/m021p153.pdf.

Kaartvedt:1986:DAP

[Kaa86] S. Kaartvedt. Diel activity patterns in deep-living cumaceans and amphipods. Marine Ecology Progress Series, 30:243–249, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
30/m030p243.pdf.

Kaartvedt:1988:WDM

[KAA88] S. Kaartvedt, D. L. Aksnes, and A. Aadnesen. Winter distribution of macroplankton and micronekton in masfjorden, western Norway. Marine Ecology Progress Series, 45:45–55, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/45/m045p045.pdf.

Kaartvedt:1989:RVM

[Kaa89] S. Kaartvedt. Retention of vertically migrating suprabenthic mysids in fjords. Marine Ecology Progress Series, 57:119–128, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
57/m057p119.pdf.

Kahru:1981:CSS

[KAE81] M. Kahru, A. Aitsam, and J. Elken. Coarse-scale spatial structure of phytoplankton standing crop in relation to hydrography in the open Baltic Sea. Marine Ecology Progress Series, 5:311– 318, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
5/m005p311.pdf.

Kaartveldt:1987:ELV

[KAE87] S. Kaartveldt, D. L. Aksnes, and J. K. Egge. Effect of light on the vertical distribution of Pecten maximus larvae. Marine Ecology Progress Series, 40:195–197, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/40/m040p195.pdf.

Kahru:1982:PPG

[Kah82] M. Kahru. Phytoplankton patchiness generated by long internal waves: a model. Marine Ecology Progress Series, 10:111–117, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
10/m010p111.pdf.

Kristensen:1988:PAB

[KAK88] E. Kristensen, F. O. Andersen, and L. H. Kofoed. Preliminary assessment of benthic community metabolism in a south-east Asian mangrove swamp. Marine Ecology Progress Series, 48:137–145, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
48/m048p137.pdf.

Kane:1984:FHC

[Kan84] J. Kane. The feeding habits of co-occurring cod and haddock larvae from Georges Bank. Marine Ecology Progress Series, 16:9– 20, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
16/m016p009.pdf.

Karlson:1986:DCF

[Kar86] R. H. Karlson. Disturbance, colonial fragmentation, and dependent life history variation in two coral reef cnidarians. Marine Ecology Progress Series, 28:245–249, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/28/m028p245.pdf.

Khristoforova:1980:MCS

[KB80] N. K. Khristoforova and N. N. Bogdanova. Mineral composition of seaweeds from coral islands of the Pacific Ocean as a function of environmental conditions. Marine Ecology Progress Series, 3:25– 29, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
3/m003p025.pdf.

King:1984:PEI

[KB84] G. M. King and T. Berman. Potential effects of isotopic dilution on apparent respiration in ^{14}C heterotrophy experiments. Marine Ecology Progress Series, 19:175–180, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/19/m019p175.pdf.

Kendall:1987:RRB

[KB87] M. A. Kendall and M. L. Bedford. Reproduction and recruitment of the barnacle Chthamalus montagui at Aberystwyth (mid-Wales). Marine Ecology Progress Series, 38:305–308, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/38/
m038p305.pdf.

Klimley:1988:IEP

[KB88] A. P. Klimley and S. B. Butler. Immigration and emigration of a pelagic fish assemblage to seamounts in the Gulf of California related to water mass movements using satellite imagery. Marine Ecology Progress Series, 49:11–20, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/49/m049p011.pdf.

Kelly:1985:BPC

[KBNO85] J. R. Kelly, V. M. Berounsky, S. W. Nixon, and C. A. Oviatt. Benthic-pelagic coupling and nutrient cycling across an experimental eutrophication gradient. Marine Ecology Progress Series, 26:207–219, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/26/m026p207.pdf.

Kumagai:1985:GJM

[KBU85] S. Kumagai, T. Bagarinao, and A. Unggui. Growth of juvenile milkfish Chanos chanos in a natural habitat. Marine Ecology Progress Series, 22:1–6, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/22/m022p001.pdf.

Kuiper:1983:EME

[KBv^{+}83a] J. Kuiper, U. H. Brockmann, H. van het Groenewoud, G. Hoornsman, and P. Roele. Effects of mercury on enclosed plankton communities in the rosfjord during POSER. Marine Ecology Progress Series, 14:93–105, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/14/m014p093.pdf.

Kuiper:1983:IBD

[KBv^{+}83b] J. Kuiper, U. H. Brockmann, H. van net Groenewoud, G. Hoornsman, and K. D. Hammer. Influences of bag dimensions on the development of enclosed plankton communities during POSER. Marine Ecology Progress Series, 14:9–17, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/14/m014p009.pdf.

Karakiri:1989:GDG

[KBv89] M. Karakiri, R. Berghahn, and H. von Westernhagen. Growth differences in 0-group plaice Pleuronectes platessa as revealed by otolith microstructure analysis. Marine Ecology Progress Series, 55:15–22, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/55/m055p015.pdf.

Kingett:1981:ADD

[KC81] P. D. Kingett and J. H. Choat. Analysls of denslty and distribution patterns in Chrysophrys auratus (Pisces: Sparidae) within a reef environment: an experimental approach. Marine Ecology Progress Series, 5:283–290, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/5/m005p283.pdf.

Kern:1982:FAI

[KC82] John C. Kern and Andrew G. Carey, Jr. The faunal assemblage inhabiting seasonal sea ice in the nearshore Arctic Ocean with emphasis on copepods. Marine Ecology Progress Series, 10:159– 167, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
10/m010p159.pdf.

Kennelly:1989:ETD

[KC89] S. J. Kennelly and J. R. Craig. Effects of trap design, independence of traps and bait on sampling populations of spanner crabs Ranina ranina. Marine Ecology Progress Series, 51:49–56, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/51/m051p049.pdf.

Kenworthy:1989:PDH

[KCFS89] W. J. Kenworthy, C. A. Currin, M. S. Fonseca, and G. Smith. Production, decomposition, and heterotrophic utilization of the seagrass Halophila decipiens in a submarine canyon. Marine Ecology Progress Series, 51:277–290, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/51/m051p277.pdf.

King:1989:AMO

[KCS89] G. M. King, R. G. Carlton, and T. E. Sawyer. Anaerobic metabolism and oxygen distribution in the carbonate sediments of a submarine canyon. Marine Ecology Progress Series, 58:275– 285, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
58/m058p275.pdf.

Kuipers:1984:NFW

[KD84] B. R. Kuipers and R. Dapper. Nursery function of Wadden Sea tidal flats for the brown shrimp Crangon crangon. Marine Ecology Progress Series, 17:171–181, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/17/m017p171.pdf.

Keough:1986:ESP

[KD86] M. J. Keough and B. J. Downes. Effects of settlement and post-settlement mortality on the distribution of the ascidian Trididemnum opacum. Marine Ecology Progress Series, 33:279–285, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
33/m033p279.pdf.

Kuipers:1981:EFT

[KdC81] B. R. Kuipers, P. A. W. J. de Wilde, and F. Creutzberg. Energy flow in a tidal flat ecosystem. Marine Ecology Progress Series, 5: 215–221, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/5/m005p215.pdf.

Kogure:1988:ATM

[KDT^{+}88] K. Kogure, H. K. Do, E. V. Thuesen, K. Nanba, K. Ohwada, and U. Simidu. Accumulation of tetrodotoxin in marine sediment. Marine Ecology Progress Series, 45:303–305, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/45/m045p303.pdf.

Kautsky:1987:RBM

[KE87] N. Kautsky and S. Evans. Role of biodeposition by Mytilus edulis in the circulation of matter and nutrients in a Baltic coastal ecosystem. Marine Ecology Progress Series, 38:201–212, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/38/m038p201.pdf.

Keen:1987:RAA

[Kee87] S. L. Keen. Recruitment of Aurelia aurita (Cnidaria: Scyphozoa) larvae is position-dependent, and independent of conspecific density, within a settling surface. Marine Ecology Progress Series, 38: 151–160, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/38/m038p151.pdf.

Kahru:1984:PDP

[KEK^{+}84] M. Kahru, J. Elken, I. Kotta, M. Simm, and K. Vilbaste. Plankton distributions and processes across a front in the open Baltic Sea. Marine Ecology Progress Series, 20:101–111, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/20/m020p101.pdf.

Keller:1988:EPP

[Kel88] A. A. Keller. Estimating phytoplankton productivity from light availability and biomass in the MERL mesocosms and Narragansett Bay. Marine Ecology Progress Series, 45:159–168, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/45/
m045p159.pdf.

Kempf:1981:LLL

[Kem81] S. C. Kempf. Long-lived larvae of the gastropod Aplysia juliana: Do they disperse and metamorphose or just slowly fade away. Marine Ecology Progress Series, 6:61–65, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/6/m006p061.pdf.

Kemp:1987:PIB

[Kem87] P. F. Kemp. Potential impact on bacteria of grazing by a macrofaunal deposit-feeder, and the fate of bacterial production. Marine Ecology Progress Series, 36:151–161, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/36/m036p151.pdf.

Kemp:1988:BBC

[Kem88] P. F. Kemp. Bacterivory by benthic ciliates: significance as a carbon source and impact on sediment bacteria. Marine Ecology Progress Series, 49:163–169, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/49/m049p163.pdf.

Kennelly:1987:PDAa

[Ken87a] S. J. Kennelly. Physical disturbances in an Australian kelp community. I. Temporal effects. Marine Ecology Progress Series, 40: 145–153, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/40/m040p145.pdf.

Kennelly:1987:PDAb

[Ken87b] S. J. Kennelly. Physical disturbances in an Australian kelp community. II. Effects on under-storey species due to differences in kelp cover. Marine Ecology Progress Series, 40:155–165, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/40/m040p155.pdf.

Kennelly:1988:EKC

[Ken88] S. J. Kennelly. Effects of kelp canopies on understorey species due to shade and scour. Marine Ecology Progress Series, 50:215– 224, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
50/m050p215.pdf.

Kennelly:1989:EST

[Ken89] S. J. Kennelly. Effects of soak-time and spatial heterogeneity on sampling populations of spanner crabs Ranina ranina. Marine Ecology Progress Series, 55:141–147, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/55/m055p141.pdf.

Keough:1989:DBB

[Keo89] M. J. Keough. Dispersal of the bryozoan Bugula neritina and effects of adults on newly metamorphosed juveniles. Marine Ecology Progress Series, 57:163–171, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/57/m057p163.pdf.

Kleppel:1988:NDZ

[KFPH88] G. S. Kleppel, D. Frazel, R. E. Pieper, and D. V. Holliday. Natural diets of zooplankton off southern California. Marine Ecology Progress Series, 49:231–241, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/49/m049p231.pdf.

Kottmeier:1987:SIM

[KGS87] S. T. Kottmeier, S. M. Grossi, and C. W. Sullivan. Sea ice microbial communities. VIII. Bacterial production in annual sea ice of McMurdo Sound, Antarctica. Marine Ecology Progress Series, 35: 175–186, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/35/m035p175.pdf.

Khristoforova:1984:ECG

[KGV84] N. K. Khristoforova, S. M. Gnezdilova, and G. A. Vlasova. Effect of cadmium on gametogenesis and offspring of the sea urchin Strongylocentrotus intermedius. Marine Ecology Progress Series, 17:9–14, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/17/m017p009.pdf.

Kendrick:1988:EMC

[KH88a] G. A. Kendrick and M. W. Hawkes. The epiphyte Microcladia coulteri (Rhodophyta): changes in population structure with spatial and temporal variation in availability of host species. Marine Ecology Progress Series, 43:79–86, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/43/m043p079.pdf.

Kirchman:1988:BPD

[KH88b] D. L. Kirchman and M. P. Hoch. Bacterial production in the Delaware Bay estuary estimated from thymidine and leucine incorporation rates. Marine Ecology Progress Series, 45:169–178, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
45/m045p169.pdf.

Kattner:1983:NPD

[KHE^{+}83] G. Kattner, K. D. Hammer, K. Eberlein, U. H. Brockmann, J. Jahnke, and M. Krause. Nutrient and plankton development in rosfjorden and enclosed ecosystems captured from changing water bodies during POSER. Marine Ecology Progress Series, 14:29– 43, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
14/m014p029.pdf.

Koike:1986:INM

[KHHB86] I. Koike, O. Holm-Hansen, and D. C. Biggs. Inorganic nitrogen metabolism by Antarctic phytoplankton with special reference to ammonium cycling. Marine Ecology Progress Series, 30:105–116, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
30/m030p105.pdf.

Koslow:1988:EFR

[KHW88] J. A. Koslow, F. Hanley, and R. Wicklund. Effects of fishing on reef fish communities at Pedro Bank and Port Royal Cays, Jamaica. Marine Ecology Progress Series, 43:201–212, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/43/m043p201.pdf.

Kimmerer:1984:SPI

[Kim84] W. J. Kimmerer. Selective predation and its impact on prey of Sagitta enflata (Chaetognatha). Marine Ecology Progress Series, 15:55–62, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/15/m015p055.pdf.

Kim:1985:EHM

[Kim85] S.-J. Kim. Effect of heavy metals on natural populations of bacteria from surface micro-layers and subsurface water. Marine Ecology Progress Series, 26:203–206, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/26/m026p203.pdf.

Kinne:1986:EIA

[Kin86] O. Kinne. The Ecology Institute — its activities, prizes, funding, and staff. Marine Ecology Progress Series, 34:1–5, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/34/m034p001.pdf.

Kinne:1988:MUJ

[Kin88] O. Kinne. MEPS: a unique journal celebrates its 10th year of existence and the appearance of its 50th volume. Marine Ecology Progress Series, 50:1–2, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/50/m050p001.pdf.

Kingsford:1989:DPP

[Kin89] M. J. Kingsford. Distribution patters of planktivorous reef fish along the coast of northeastern New Zealand. Marine Ecology Progress Series, 54:13–24, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/54/m054p013.pdf.

Kiorboe:1989:PGR

[Kiø89] T. Kiørboe. Phytoplankton growth rate and nitrogen content: implications for feeding and fecundity in a herbivorous copepod. Marine Ecology Progress Series, 55:229–234, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/55/m055p229.pdf.

Kepkay:1988:MRO

[KJ88] P. E. Kepkay and B. D. Johnson. Microbial response to organic particle generation by surface coagulation in seawater. Marine Ecology Progress Series, 48:193–198, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/48/m048p193.pdf.

Kuosa:1989:BHF

[KK89] H. Kuosa and K. Kivi. Bacteria and heterotrophic flagellates in the pelagic carbon cycle in the northern Baltic Sea. Marine Ecology Progress Series, 53:93–100, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/53/m053p093.pdf.

Kautsky:1986:DDP

[KKKW86] N. Kautsky, H. Kautsky, U. Kautsky, and M. Waern. Decreased depth penetration of Fucus vesiculosus (L.) since the 1940’s indicates eutrophication of the Baltic Sea. Marine Ecology Progress Series, 28:1–8, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/28/m028p001.pdf.

Kattner:1981:LCS

[KKT81] G. Kattner, M. Krause, and J. Trahms. Lipid composition of some typical North Sea copepods. Marine Ecology Progress Series, 4:69– 74, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
4/m004p069.pdf.

Keser:1984:CGD

[KL84] M. Keser and B. R. Larson. Colonization and growth dynamics of three species of Fucus. Marine Ecology Progress Series, 15:125– 134, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
15/m015p125.pdf.

KleinBreteler:1980:CBM

[Kle80] W. C. M. Klein Breteler. Continuous breeding of marine pelagic copepods in the presence of heterotrophic dinoflagellates. Marine Ecology Progress Series, 2:229–233, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/2/m002p229.pdf.

Kaneta:1985:MAP

[KLE85] P. J. Kaneta, M. Levandowsky, and W. Esaias. Multivariate analysis of the phytoplankton community in the New York Bight. Marine Ecology Progress Series, 23:231–239, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/23/m023p231.pdf.

Kreeger:1988:URC

[KLN88] D. A. Kreeger, C. J. Langdon, and R. I. E. Newell. Utilization of refractory cellulosic carbon derived from Spartina alterniflora by the ribbed mussel Geukensia demissa. Marine Ecology Progress Series, 42:171–179, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/42/m042p171.pdf.

Klumpp:1984:NEA

[Klu84] D. W. Klumpp. Nutritional ecology of the ascidian Pyura stolonifera: influence of body size, food quantity and quality on filter-feeding, respiration, assimilation efficiency and energy balance. Marine Ecology Progress Series, 19:269–284, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/19/m019p269.pdf.

Kranck:1980:MPM

[KM80] K. Kranck and T. Milligan. Macroflocs: Production of marine snow in the laboratory. Marine Ecology Progress Series, 3:19– 24, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
3/m003p019.pdf.

Kiorboe:1981:PSS

[KM81] T. Kiørboe and F. Møhlenberg. Particle selection in suspensionfeeding bivalves. Marine Ecology Progress Series, 5:291–296, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
5/m005p291.pdf.

Koike:1983:OIN

[KM83] I. Koike and H. Mukai. Oxygen and inorganic nitrogen contents and fluxes in burrows of the shrimps Callianassa japonica and Upogebia major. Marine Ecology Progress Series, 12:185–190, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
12/m012p185.pdf.

Knight:1985:CCP

[KM85] R. Knight and R. F. C. Mantoura. Chlorophyll and carotenoid pigments in Foraminifera and their symbiotic algae: analysis by high performance liquid chromatography. Marine Ecology Progress Series, 23:241–249, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/23/m023p241.pdf.

Kimmerer:1987:ZMBa

[KM87a] W. J. Kimmerer and A. D. McKinnon. Zooplankton in a marine bay I. Horizontal distributions used to estimate net population growth rates. Marine Ecology Progress Series, 41:43–52, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
41/m041p043.pdf.

Kimmerer:1987:ZMBb

[KM87b] W. J. Kimmerer and A. D. McKinnon. Zooplankton in a marine bay II. Vertical migration to maintain horizontal distributions. Marine Ecology Progress Series, 41:53–60, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/41/m041p053.pdf.

Kingsford:1987:PPP

[KM87c] M. J. Kingsford and M. J. Milicich. Presettlement phase of Parika scaber (Pisces: Monacanthidae): a temperate reef fish. Marine Ecology Progress Series, 36:65–79, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/36/m036p065.pdf.

Kingsford:1988:IBP

[KM88a] M. J. Kingsford and A. B. MacDiarmid. Interrelations between planktivorous reef fish and zooplankton in temperate waters. Marine Ecology Progress Series, 48:103–117, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/48/m048p103.pdf.

Kranck:1988:MDS

[KM88b] K. Kranck and T. G. Milligan. Macroflocs from diatoms: in situ photography of particles in Bedford Basin, Nova Scotia. Marine Ecology Progress Series, 44:183–189, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/44/m044p183.pdf.

Kilar:1989:EWE

[KM89a] J. A. Kilar and J. McLachlan. Effects of wave exposure on the community structure of a plant-dominated, fringing-reef platform: intermediate disturbance and disturbance-mediated competition. Marine Ecology Progress Series, 54:265–276, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/54/m054p265.pdf.

Kimmerer:1989:ZMB

[KM89b] W. J. Kimmerer and A. D. McKinnon. Zooplankton in a marine bay. III. Evidence for influence of vertebrate predation on distributions of two common copepods. Marine Ecology Progress Series, 53:21–35, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/53/m053p021.pdf.

Kirchman:1984:EBP

[KMAM84] David L. Kirchman, Lucia Mazzella, Randall S. Alberte, and Ralph Mitchell. Epiphytic bacterial production on Zostera marina. Marine Ecology Progress Series, 15:117–123, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/15/m015p117.pdf.

Kemp:1987:DGE

[KMBSJ87] W. M. Kemp, L. Murray, J. Borum, and K. Sand-Jensen. Diel growth in eelgrass Zostera marina. Marine Ecology Progress Series, 41:79–86, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/41/m041p079.pdf.

Klumpp:1987:DTZ

[KMD87] D. W. Klumpp, D. McKinnon, and P. Daniel. Damselfish territories: zones of high productivity on coral reefs. Marine Ecology Progress Series, 40:41–51, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/40/m040p041.pdf.

Kimor:1985:CMO

[KMDS85] B. Kimor, A. G. Moigis, V. Dohms, and C. Stienen. A case of mass occurrence of Prorocentrum minimum in the Kiel fjord. Marine Ecology Progress Series, 27:209–215, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/27/m027p209.pdf.

Kiorboe:1985:BPC

[KMH85] T. Kiørboe, F. Møhlenberg, and K. Hamburger. Bioenergetics of the planktonic copepod Acartia tonsa: relation between feeding, egg production and respiration, and composition of specific dynamic action. Marine Ecology Progress Series, 26:85–97, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
26/m026p085.pdf.

Klumpp:1988:MCC

[KMM88] D. W. Klumpp, A. D. McKinnon, and C. N. Mundy. Motile cryptofauna of a coral reef: abundance, distribution and trophic potential. Marine Ecology Progress Series, 45:95–108, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/45/m045p095.pdf.

Kiorboe:1987:RGL

[KMR87] T. Kiorbøe, P. Munk, and K. Richardson. Respiration and growth of larval herring Clupea harengus: relation between specific dynamic action and growth efficiency. Marine Ecology Progress Series, 40:1–10, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/40/m040p001.pdf.

Kiorboe:1988:PDL

[KMR^{+}88] T. Kiørboe, P. Munk, K. Richardson, V. Christensen, and H. Paulsen. Plankton dynamics and larval herring growth, drift and survival in a frontal area. Marine Ecology Progress Series, 44: 205–219, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/44/m044p205.pdf.

Klumpp:1983:NSS

[KN83] D. W. Klumpp and P. D. Nichols. Nutrition of the southern sea garfish Hyporhamphus melanochir: gut passage rate and daily consumption of two food types and assimilation of seagrass components. Marine Ecology Progress Series, 12:207–216, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/12/m012p207.pdf.

Kelly:1984:ESE

[KN84] J. R. Kelly and S. W. Nixon. Experimental studies of the effect of organic deposition on the metabolism of a coastal marine bottom community. Marine Ecology Progress Series, 17:157–169, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
17/m017p157.pdf.

Karl:1988:DMG

[KN88] D. M. Karl and J. A. Novitsky. Dynamics of microbial growth in surface layers of a coastal marine sediment ecosystem. Marine Ecology Progress Series, 50:169–176, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/50/m050p169.pdf.

Kneib:1981:SSE

[Kne81] R. T. Kneib. Size-specific effects of density on the growth fecundity and mortality of the fish Fundulus heteroclitus in an intertidal salt Marsh. Marine Ecology Progress Series, 6:203–212, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/6/m006p203.pdf.

Kirchman:1986:IVB

[KNH86] D. L. Kirchman, S. Y. Newell, and R. E. Hodson. Incorporation versus biosynthesis of leucine: implications for measuring rates of protein synthesis and biomass production by bacteria in marine systems. Marine Ecology Progress Series, 32:47–59, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/32/m032p047.pdf.

Koop:1982:BCF

[KNL82a] K. Koop, R. C. Newell, and M. I. Lucas. Biodegradation and carbon flow based on kelp (Ecklonia maxima) debris in a sandy beach microcosm. Marine Ecology Progress Series, 7:315–326, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
7/m007p315.pdf.

Koop:1982:MRN

[KNL82b] K. Koop, R. C. Newell, and M. I. Lucas. Microbial regeneration of nutrients from the decomposition of macrophyte debris on the shore. Marine Ecology Progress Series, 9:91–96, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/9/m009p091.pdf.

Kurihara:1987:CGC

[KO87] Y. Kurihara and K. Okamoto. Cannibalism in a grapsid crab, Hemi-grapsus penicillatus. Marine Ecology Progress Series, 41: 123–127, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/41/m041p123.pdf.

Kraynovic-Ozretic:1980:ADA

[KOO80] M. Kraynovic-Ozretic and B. Ozretic. The ALA-D activity test in lead-exposed grey-mullet Mugil auratus. Marine Ecology Progress Series, 3:187–191, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/3/m003p187.pdf.

Kristiansen:1989:IMD

[KP89] S. Kristiansen and E. Paasche. An improved method for determining relative ^{15}N abundance in ammonium regeneration studies by direct diffusion. Marine Ecology Progress Series, 54:203–207, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
54/m054p203.pdf.

Kirchman:1984:BGT

[KPJ84] D. Kirchman, B. Peterson, and D. Juers. Bacterial growth and tidal variation in bacterial abundance in the great sippewissett salt Marsh. Marine Ecology Progress Series, 19:247–259, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/19/
m019p247.pdf.

Kojis:1982:RET

[KQ82] B. L. Kojis and N. J. Quinn. Reproductive ecology of two faviid corals (Coelenterata: Scleractinia). Marine Ecology Progress Series, 8:251–255, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/8/m008p251.pdf.

Kohn:1982:SSD

[KR82] A. J. Kohn and A. C. Riggs. Sample size dependence in measures of proportional similarity. Marine Ecology Progress Series, 9:147– 151, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
9/m009p147.pdf.

Kristensen:1983:CPN

[Kri83a] E. Kristensen. Comparison of polychaete (Nereis spp.) ventilation in plastic tubes and natural sediment. Marine Ecology Progress Series, 12:307–309, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/12/m012p307.pdf.

Kristensen:1983:VOUa

[Kri83b] E. Kristensen. Ventilation and oxygen uptake by three species of Nereis (Annelida: Polychaeta). I. Effects of hypoxia. Marine Ecology Progress Series, 12:289–297, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/12/m012p289.pdf.

Kristensen:1983:VOUb

[Kri83c] E. Kristensen. Ventilation and oxygen uptake by three species of Nereis (Annelida: Polychaeta). II. Effects of temperature and salinity changes. Marine Ecology Progress Series, 12:299–306, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
12/m012p299.pdf.

Kronberg:1988:SAF

[Kro88] I. Kronberg. Structure and adaptation of the fauna in the black zone (littoral fringe) along rocky shores in northern Europe. Marine Ecology Progress Series, 49:95–106, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/49/m049p095.pdf.

Klinken:1983:ILL

[KS83] J. Klinken and H. R. Skjoldal. Improvements of luciferin-luciferase methodology for determination of adenylate energy charge ratio of marine samples. Marine Ecology Progress Series, 13:305–309, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p305.pdf.

Karentz:1984:TSO

[KS84] D. Karentz and T. J. Smayda. Temperature and seasonal occurrence patterns of 30 dominant phytoplankton species in Narragansett Bay over a 22-year period (1959–1980). Marine Ecology Progress Series, 18:277–293, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/18/m018p277.pdf.

Klein:1987:DSD

[KS87a] B. Klein and A. Sournia. A daily study of the diatom spring bloom at Roscoff (France) in 1985. II phytoplankton pigment composition studied by HPLC analysis. Marine Ecology Progress Series, 37: 265–275, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/37/m037p265.pdf.

Kottmeier:1987:LWP

[KS87b] S. T. Kottmeier and C. W. Sullivan. Late winter primary production and bacterial production in sea ice and seawater west of the Antarctic Peninsula. Marine Ecology Progress Series, 36:287– 298, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
36/m036p287.pdf.

Knoechel:1989:CRO

[KSF89] R. Knoechel and D. Steel-Flynn. Clearance rates of Oikopleura in cold coastal Newfoundland waters: a predictive model and its trophodynamic implications. Marine Ecology Progress Series, 53: 257–266, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/53/m053p257.pdf.

Kurihara:1988:WBM

[KSM88] Y. Kurihara, K. Sekimoto, and M. Miyata. Wandering behaviour of the mud-crab Helice tridens related to evasion of cannibalism. Marine Ecology Progress Series, 49:41–50, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/49/m049p041.pdf.

Kremer:1980:FNS

[KSR80] B. P. Kremer, R. Schmaljohann, and R. Riittger. Features and nutritional significance of photosynthates produced by unicellular algae symbiotic with larger Foraminifera. Marine Ecology Progress Series, 2:225–228, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/2/m002p225.pdf.

Keats:1985:ABD

[KSS85] D. W. Keats, G. R. South, and D. H. Steele. Algal biomass and diversity in the upper subtidal at a pack-ice disturbed site in eastern Newfoundland. Marine Ecology Progress Series, 25:151–158, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
25/m025p151.pdf.

Kirchman:1989:BPR

[KSVB89] D. Kirchman, Y. Soto, F. Van Wambeck, and M. Bianchi. Bacterial production in the Rhone River plume: effect of mixing on relationships among microbial assemblages. Marine Ecology Progress Series, 53:267–275, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/53/m053p267.pdf.

Kuiper:1981:FEC

[Kui81] J. Kuiper. Fate and effects of cadmium in marine plankton communities in experimental enclosures. Marine Ecology Progress Series, 6:161–174, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/6/m006p161.pdf.

Karakiri:1988:AGO

[Kv88] M. Karakiri and H. von Westernhagen. Apparatus for grinding otoliths of larval and juvenile fish for microstructure analysis. Marine Ecology Progress Series, 49:195–198, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/49/m049p195.pdf.

Kokkinakis:1988:UAU

[KW88] S. A. Kokkinakis and P. A. Wheeler. Uptake of ammonium and urea in the northeast Pacific: comparison between netplankton and nanoplankton. Marine Ecology Progress Series, 43:113–124, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
43/m043p113.pdf.

Kana:1985:DPP

[KWG85] Todd M. Kana, Jennifer L. Watts, and Patricia M. Glibert. Diel periodicity in the photosynthetic capacity of coastal and offshore phytoplankton assemblages. Marine Ecology Progress Series, 25: 131–139, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/25/m025p131.pdf.

Klungsoyr:1988:GWO

[KWW^{+}88] J. Klungsoyr, S. Wilhelmsen, K. Westrheim, E. Saetvedt, and K. H. Palmork. The GEEP workshop: organic chemical analyses. Marine Ecology Progress Series, 46:19–26, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/46/m046p019.pdf.

Kamykowski:1985:NSA

[KZ85] D. Kamykowski and S.-J. Zentara. Nitrate and silicic acid in the world ocean: patterns and processes. Marine Ecology Progress Series, 26:47–59, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/26/m026p047.pdf.

Kamykowski:1989:CPN

[KZ89] D. Kamykowski and S.-J. Zentara. Circumpolar plant nutrient covariation in the Southern Ocean: patterns and processes. Marine Ecology Progress Series, 58:101–111, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/58/m058p101.pdf.

LopezGappa:1989:OCA

[Lan89] J. J. López Gappa. Overgrowth competition in an assemblage of encrusting bryozoans settled on artificial substrata. Marine Ecology Progress Series, 51:121–130, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/51/m051p121.pdf.

Lambshead:1986:SCS

[Lam86] P. J. D. Lambshead. Sub-catastrophic sewage and industrial waste contamination as revealed by marine nematode faunal analysis. Marine Ecology Progress Series, 29:247–259, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/29/m029p247.pdf.

Lambert:1987:DIS

[Lam87] T. C. Lambert. Duration and intensity of spawning in nerring Clupea harengus as related to the age structure of the mature population. Marine Ecology Progress Series, 39:209–220, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/39/m039p209.pdf.

Lancelot:1983:FAP

[Lan83] C. Lancelot. Factors affecting phytoplankton extracellular release in the Southern Bight of the North Sea. Marine Ecology Progress Series, 12:115–121, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/12/m012p115.pdf.

Larsson:1984:GBS

[Lar84] P.-O. Larsson. Growth of Baltic salmon Salmo salar in the sea. Marine Ecology Progress Series, 17:215–226, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/17/m017p215.pdf.

Larson:1986:SCS

[Lar86] R. J. Larson. Seasonal changes in the standing stocks, growth rates, and production rates of gelatinous predators in Saanich Inlet, British Columbia. Marine Ecology Progress Series, 33:89–98, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
33/m033p089.pdf.

Lasker:1981:CPF

[Las81] H. R. A. Lasker. Comparison of the particulate feeding abilities of three species of gorgonian soft coral. Marine Ecology Progress Series, 5:61–67, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/5/m005p061.pdf.

Lasker:1984:ARF

[Las84] H. R. Lasker. Asexual reproduction, fragmentation, and skeletal morphology of a plexaurid gorgonian. Marine Ecology Progress Series, 19:261–268, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/19/m019p261.pdf.

Lasker:1985:PPB

[Las85] H. R. Lasker. Prey preferences and browsing pressure of the butterflyfish Chaetodon capistratus on Caribbean gorgonians. Marine Ecology Progress Series, 21:213–220, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/21/m021p213.pdf.

Latypov:1986:CCN

[Lat86] Yu. Ya. Latypov. Coral communities of the Namsu Islands (Gulf of Siam, South China Sea). Marine Ecology Progress Series, 29:261– 270, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
29/m029p261.pdf.

Lawrence:1986:PCS

[Law86] J. M. Lawrence. Proximate composition and standing crop of Durvillaea antarctica (Pheophyta) in the Bay of Morbihan, Kerguelen (South Indian Ocean). Marine Ecology Progress Series, 33: 1–5, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
33/m033p001.pdf.

Lawton:1989:PIB

[Law89] P. Lawton. Predatory interaction between the brachyuran crab Cancer pagurus and decapod crustacean prey. Marine Ecology Progress Series, 52:169–179, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/52/m052p169.pdf.

Luttke:1981:LFX

[LB81] A. Lüttke and S. Bonotto. Loss of fertility in X-irradiated Acetabularia. Marine Ecology Progress Series, 5:107–109, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/5/m005p107.pdf.

Lavalli:1986:BPP

[LB86] Kari L. Lavalli and Diana E. Barshaw. Burrows protect postlarval lobsters Homarus americanus from predation by the nonburrowing cunner Tautogolabrus adspersus, but not from the burrowing mud crab Neopanope texani. Marine Ecology Progress Series, 32:13–16, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/32/m032p013.pdf.

Libes:1987:ULD

[LB87] M. Libes and C.-F. Boudouresque. Uptake and long-distance transport of carbon in the marine phanerogam Posidonia oceanica. Marine Ecology Progress Series, 38:177–186, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/38/m038p177.pdf.

Larsson:1986:NST

[LBA86] U. Larsson, S. Blomqvist, and B. Abrahamsson. A new sediment trap system. Marine Ecology Progress Series, 31:205–207, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
31/m031p205.pdf.

Lomstein:1989:ANC

[LBH89] B. Aa Lomstein, T. H. Blackburn, and K. Henriksen. Aspects of nitrogen and carbon cycling in the northern Bering Shelf sediment. I. The significance of urea turnover in the mineralization of NH_{4}^{+}. Marine Ecology Progress Series, 57:237–247, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/57/m057p237.pdf.

Lopez:1982:RSO

[LC82] G. R. Lopez and M. A. Crenshaw. Radiolabelling of sedimentary organic matter with ^{14}C-formaldehyde: Preliminary evaluation of a new technique for use in deposit-feeding studies. Marine Ecology Progress Series, 8:283–289, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/8/m008p283.pdf.

Lasker:1983:ODC

[LC83] H. R. Lasker and M. A. Coffroth. Octocoral distributions at Carrie Bow Cay, Belize. Marine Ecology Progress Series, 13:21–28, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p021.pdf.

Lasker:1988:TSV

[LC88] H. R. Lasker and M. A. Coffroth. Temporal and spatial variability among grazers: variability in the distribution of the gastropod Cyphoma gibbosum on octocorals. Marine Ecology Progress Series, 43:285–295, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/43/m043p285.pdf.

Lamontagne:1989:EFV

[LCF89] I. Lamontagne, A. Cardinal, and L. Fortier. Environmental forcing versus endogenous control of photosynthesis in intertidal epilithic microalgae. Marine Ecology Progress Series, 51:177–187, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/51/
m051p177.pdf.

Lewis:1984:RBV

[LCP84] M. R. Lewis, J. J. Cullen, and T. Piatt. Relationships between vertical mixing and photoadaptation of phytoplankton: similarity criteria. Marine Ecology Progress Series, 15:141–149, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/15/m015p141.pdf.

LaBarre:1986:CSS

[LCS86] S. C. La Barre, J. C. Coll, and P. W. Sammarco. Competitive strategies of soft corals (Coelenterata: Octocorallia): III. Spacing and aggressive interactions between alcyonaceans. Marine Ecology Progress Series, 28:147–156, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/28/m028p147.pdf.

Li:1985:GBS

[LD85] W. K. W. Li and P. M. Dickie. Growth of bacteria in seawater filtered through 0.2 μ m nuclepore membranes. implications for dilution experiments. Marine Ecology Progress Series, 26:245– 252, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
26/m026p245.pdf.

Laprise:1989:OIT

[LD89] R. Laprise and J. J. Dodson. Ontogeny and importance of tidal vertical migrations in the retention of larval smelt Osmerus mordax in a well-mixed estuary. Marine Ecology Progress Series, 55: 101–111, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/55/m055p101.pdf.

Legendre:1988:BPA

[LDDH88] L. Legendre, S. Demers, B. Delesalle, and C. Harnois. Biomass and photosynthetic activity of phototrophic picoplankton in coral reef waters (Moorea Island, French Polynesia). Marine Ecology Progress Series, 47:153–160, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/47/m047p153.pdf.

Laake:1983:MAI

[LDER83] M. Laake, A. B. Dahle, K. Eberlein, and K. Rein. A modelling approach to the interplay of carbohydrates, bacteria and nonpigmented flagellates in a controlled ecosystem experiment with Skeletonema costatum. Marine Ecology Progress Series, 14:71– 79, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
14/m014p071.pdf.

Legendre:1987:CPE

[LDG87] L. Legendre, S. Demers, and M. Gosselin. Chlorophyll and photosynthetic efficiency of size-fractionated sea-ice microalgae (Hudson Bay, Canadian Arctic). Marine Ecology Progress Series, 40:199– 203, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
40/m040p199.pdf.

Laake:1983:PPD

[LDH83] M. Laake, A. B. Dahle, and G. Hentzschel. Productivity and population diversity of marine organotrophic bacteria in enclosed planktonic ecosystems. Marine Ecology Progress Series, 14:59– 69, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
14/m014p059.pdf.

Litaker:1988:DCP

[LDKR88] W. Litaker, C. S. Duke, B. E. Kenney, and J. Ramus. Diel chl-a and phaeopigment cycles in a shallow tidal estuary: potential role of microzooplankton grazing. Marine Ecology Progress Series, 47: 259–270, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/47/m047p259.pdf.

Levin:1986:EGP

[LDMT86] L. A. Levin, D. J. DeMaster, L. D. McCann, and C. L. Thomas. Effects of giant protozoans (class: Xenophyophorea) on deepseamount benthos. Marine Ecology Progress Series, 29:99–104, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
29/m029p099.pdf.

LeBouteiller:1986:ECN

[Le 86] A. Le Bouteiller. Environmental control of nitrate and ammonium uptake by phytoplankton in the Equatorial Atlantic Ocean. Marine Ecology Progress Series, 30:167–179, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/30/m030p167.pdf.

Leber:1982:BTD

[Leb82] K. M. Leber. Bivalves (Tellinacea: Donacidae) on a North Carolina beach: Contrasting population size structures and tidal migrations. Marine Ecology Progress Series, 7:297–301, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/7/
m007p297.pdf.

Lee:1981:FAE

[Lee81] C.-S. Lee. Factors affecting egg characteristics in the fish Sillago sihama. Marine Ecology Progress Series, 4:361–363, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/4/m004p361.pdf.

Lee:1988:GTM

[Lee88] R. F. Lee. Glutathione s-transferase in marine invertebrates from langesundfjord. Marine Ecology Progress Series, 46:33–36, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p033.pdf.

Levinton:1985:CID

[Lev85] J. S. Levinton. Complex interactions of a deposit feeder with its resources, roles of density, a competitor, and detrital addition in the growth and survival of the mudsnail Hydrobia totteni. Marine Ecology Progress Series, 22:31–40, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/22/m022p031.pdf.

Lewis:1984:DMC

[Lew84] F. G. Lewis III. Distribution of macrobenthic crustaceans associated with Thalassia, Halodule and bare sand substrata. Marine Ecology Progress Series, 19:101–113, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/19/m019p101.pdf.

Lonne:1988:ITM

[LG88] O. J. Lønne and J. S. Gray. Influence of tides on microgrowth bands in Cerastoderma edule from Norway. Marine Ecology Progress Series, 42:1–7, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/42/m042p001.pdf.

Lipina:1987:CDG

[LGK87] I. G. Lipina, S. M. Gnezdilova, and N. K. Khristoforova. Cadmium distribution in gonads of the sea urchin Strongylocentrotus intermedius. Marine Ecology Progress Series, 36:263–266, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
36/m036p263.pdf.

Laughlin:1988:CEL

[LGP88] R. B. Laughlin, Jr., R. Gustafson, and P. Pendoley. Chronic embryo-larval toxicity of tributyltin (TBT) to the hard shell clam Mercenaria mercenaria. Marine Ecology Progress Series, 48:29– 36, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
48/m048p029.pdf.

Lapota:1989:CPB

[LGS^{+}89] D. Lapota, M. L. Geiger, A. V. Stiffey, D. E. Rosenberger, and D. K. Young. Correlations of planktonic bioluminescence with other oceanographic parameters from a Norwegian fjord. Marine Ecology Progress Series, 55:217–227, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/55/m055p217.pdf.

Laurence:1981:EIT

[LH81] Geoffrey C. Laurence and W. Huntting Howell. Embryology and influence of temperature and salinity on early development and survival of yellowtail flounder Limanda lerruginea. Marine Ecology Progress Series, 6:11–18, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/6/m006p011.pdf.

Lindahl:1982:PZC

[LH82] O. Lindahl and L. Hernroth. Phyto-zooplankton community in coastal waters of Western Sweden — an ecosystem off balance? Marine Ecology Progress Series, 10:119–126, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/10/m010p119.pdf.

Lawton:1985:FBC

[LH85] P. Lawton and R. N. Hughes. Foraging behaviour of the crab Cancer pagurus feeding on the gastropods Nucella lapillus and Littorina littorea: comparisons with optimal foraging theory. Marine Ecology Progress Series, 27:143–154, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/27/m027p143.pdf.

LaRoche:1987:CMN

[LH87a] J. LaRoche and W. G. Harrison. Compartmental models of nitrogen cycling in tropical and temperate marine environments.

Marine Ecology Progress Series, 38:137–149, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/38/m038p137.pdf.

Leak:1987:CGS

[LH87b] James C. Leak and Edward D. Houde. Cohort growth and survival of bay anchovy Anchoa mitchilli larvae in Biscayne Bay, Florida. Marine Ecology Progress Series, 37:109–122, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/37/m037p109.pdf.

Lindahl:1988:LSL

[LH88] O. Lindahl and L. Hernroth. Large-scale and long-term variations in the zooplankton community of the Gullmar fjord, Sweden, in relation to advective processes. Marine Ecology Progress Series, 43: 161–171, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/43/m043p161.pdf.

Landry:1984:DMP

[LHF84] M. R. Landry, L. W. Haas, and V. L. Fagerness. Dynamics of microbial plankton communities: experiments in Kaneohe Bay, Hawaii. Marine Ecology Progress Series, 16:127–133, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/16/
m016p127.pdf.

Lee:1981:STF

[LHH81] C.-S. Lee, F. Hu, and R. Hirano. Salinity tolerance of fertilized eggs and larval survival in the fish Sillago sihama. Marine Ecology Progress Series, 4:169–174, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/4/m004p169.pdf.

Lancraft:1988:AEM

[LHT88] T. M. Lancraft, T. L. Hopkins, and J. J. Torres. Aspects of the ecology of the mesopelagic fish Gonostoma elongatum (Gonostomatidae, Stomnformes) in the eastern Gulf of Mexico. Marine Ecology Progress Series, 49:27–40, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/49/m049p027.pdf.

Li:1982:EHB

[Li82] W. K. W. Li. Estimating heterotrophic bacterial productivity by inorganic radiocarbon uptake: Importance of establishing time courses of uptake. Marine Ecology Progress Series, 8:167–172, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
8/m008p167.pdf.

LincolnSmith:1988:EOS

[Lin88] M. P. Lincoln Smith. Effects of observer swimming speed on sample counts of temperate rocky reef fish assemblages. Marine Ecology Progress Series, 43:223–231, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/43/m043p223.pdf.

Lin:1989:ILS

[Lin89] J. Lin. Influence of location in a salt marsh on survivorship of ribbed mussels. Marine Ecology Progress Series, 56:105–110, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
56/m056p105.pdf.

Livingston:1982:TOF

[Liv82] R. J. Livingston. Trophic organization of fishes in a coastal seagrass system. Marine Ecology Progress Series, 7:1–12, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/7/
m007p001.pdf.

Livingstone:1988:RMN

[Liv88] D. R. Livingstone. Responses of microsomal NADPH-cytochrome c reductase activity and cytochrome P-450 in digestive glands of Mytilus edulis and Littorina littorea to environmental and experimental exposure to pollutants. Marine Ecology Progress Series, 46: 37–43, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/46/m046p037.pdf.

Lopez-Jamar:1984:CIM

[LJIO84] E. Lopez-Jamar, J. Iglesias, and J. J. Otero. Contribution of infauna and mussel-raft epifauna to demersal fish diets. Marine Ecology Progress Series, 15:13–18, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/15/m015p013.pdf.

Littler:1981:LMC

[LL81] M. M. Littler and D. S. Littler. lntertidal macrophyte comrnunities from Pacific Baja California and the Upper Gulf of California relatively constant vs. environmentally fluctuating systems. Marine Ecology Progress Series, 4:145–158, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p145.pdf.

Laughlin:1983:OPB

[LL83] Ft Laughlin and O. Linden. Oil pollution and Baltic mysids: acute and chronic effects of the water soluble fractions of light fuel oil on the mysid shrimp Neomysis integer. Marine Ecology Progress Series, 12:29–41, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/12/m012p029.pdf.

Lopez:1982:ISS

[LlJC82] G. R. Lopez and l J. Cheng. Ingestion selectivity of sedimentary organic matter by the deposit-feeder Nucula annulata (Bivalvia: Nuculidae). Marine Ecology Progress Series, 8:279–282, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
8/m008p279.pdf.

Lopez:1983:SMI

[LlJC83] G. R. Lopez and l. J. Cheng. Synoptic measurements of ingestion rate, ingestion selectivity, and absorption efficiency of natural foods in the deposit-feeding molluscs Nucula annulata (Bivalvia) and Hydrobia totteni (Gastropoda). Marine Ecology Progress Series, 11:55–62, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/11/m011p055.pdf.

Lukatelich:1985:DAB

[LM85] R. J. Lukatelich and A. J. McComb. Distribution and abundance of benthic microalgae in a shallow southwestern Australian estuarine system. Marine Ecology Progress Series, 27:287–297, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
27/m027p287.pdf.

Lancelot:1987:DPD

[LM87a] C. Lancelot and S. Mathot. Dynamics of a phaeocystis-dominated spring bloom in Belgian coastal waters. I. Phytoplanktonic activities and related parameters. Marine Ecology Progress Series, 37: 239–248, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/37/m037p239.pdf.

Lasiak:1987:OUM

[LM87b] T. Lasiak and A. McLachlan. Opportunistic utilization of mysid shoals by surf-zone teleosts. Marine Ecology Progress Series, 37: 1–7, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
37/m037p001.pdf.

Littler:1983:ERS

[LML83] M. M. Littler, D. R. Martz, and D. S. Littler. Effects of recurrent sand deposition on rocky intertidal organisms: importance of substrate heterogeneity in a fluctuating environment. Marine Ecology Progress Series, 11:129–139, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/11/m011p129.pdf.

Lancelot:1986:MPS

[LMO86] C. Lancelot, S. Mathot, and N. J. P. Owens. Modelling protein synthesis, a step to an accurate estimate of net primary production: Phaeocystis pouchetii colonies in Belgian coastal waters. Marine Ecology Progress Series, 32:193–202, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/32/m032p193.pdf.

Laughlin:1981:ORG

[LN81] R. B. Laughlin, Jr. and J. M. Neff. Ontogeny of respiratory and growth responses of larval mud crabs Rhithropanopeus harrisii exposed to different temperatures, salinities and naphthalene concentrations. Marine Ecology Progress Series, 5:319–332, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
5/m005p319.pdf.

Lavie:1986:GDM

[LN86] B. Lavie and E. Nevo. Genetic diversity of marine gastropods: contrasting strategies of Cerithium rupestre and C. scabridum in the Mediterranean Sea. Marine Ecology Progress Series, 28:99– 103, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
28/m028p099.pdf.

Langdon:1989:UDB

[LN89] C. J. Langdon and R. I. E. Newell. Utilization of detritus and bacteria as food sources by two bivalve suspension-feeders, the oyster Crassostrea virginica and the mussel Geukensia demissa. Marine Ecology Progress Series, 58:299–310, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/58/m058p299.pdf.

Linley:1981:HUM

[LNB81] E. A. S. Linley, R. C. Newell, and S. A. Bosma. Heterotrophic utilisation of mucilage released during fragmentation of kelp (Ecklonia maxima and Laminana pallida). I. Development of microbial communities associated with the degradation of kelp mucilage. Marine Ecology Progress Series, 4:31–41, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p031.pdf.

Lapointe:1981:CSS

[LNF81] B. E. Lapointe, F. X. Niell, and J. M. Fuentes. Community structure, succession, and production of seaweeds associated wlth mussel-rafts in the Ria de Arosa, N. W. Spain. Marine Ecology Progress Series, 5:243–253, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/5/m005p243.pdf.

Linley:1983:QRB

[LNL83] E. A. S. Linley, R. C. Newell, and M. I. Lucas. Quantitative relationships between phytoplankton, bacteria and heterotrophic microflagellates in shelf waters. Marine Ecology Progress Series, 12:77–89, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/12/m012p077.pdf.

Lucas:1987:PCY

[LNS^{+}87] M. I. Lucas, R. C. Newell, S. E. Shumway, L. J. Seiderer, and R. Bally. Particle clearance and yield in relation to bacterioplankton and suspended particulate availability in estuarine and open coast populations of the mussel Mytilus edulis. Marine Ecology Progress Series, 36:215–224, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/36/m036p215.pdf.

Lucas:1981:HUM

[LNV81] M. I. Lucas, R. C. Newell, and B. Velimirov. Heterotrophic utilisation of mucilage released during fragmentation of kelp (Ecklonia maxima and Laminana pallida) II. Differential utilisation of dissolved organic components from kelp mucilage. Marine Ecology Progress Series, 4:43–55, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/4/m004p043.pdf.

Lonsdale:1981:IAS

[Lon81a] D. J. Lonsdale. Influence of age-specific mortality on the life history traits of two estuarine copepods. Marine Ecology Progress Series, 5:333–340, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/5/m005p333.pdf.

Lonsdale:1981:RRP

[Lon81b] D. J. Lonsdale. Regulatory role of physical factors and predation for two Chesapeake Bay copepod species. Marine Ecology Progress Series, 5:341–351, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/5/m005p341.pdf.

Lipcius:1989:PAM

[LOvM89] Romuald N. Lipcius, Eugene J. Olmi III, and Jacques van Montîrans. Planktonic availability, molt stage and settlement of blue crab postlarvae. Marine Ecology Progress Series, 58:235– 242, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
58/m058p235.pdf.

Lowe:1988:ACS

[Low88] D. M. Lowe. Alterations in cellular structure of Mytilus edulis resulting from exposure to environmental contaminants under field and experimental conditions. Marine Ecology Progress Series, 46: 91–100, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/46/m046p091.pdf.

Lochte:1987:CCC

[LP87] K. Lochte and O. Pfannkuche. Cyclonic cold-core eddy in the eastern North Atlantic. II. Nutrients, phytoplankton and bacterioplankton. Marine Ecology Progress Series, 39:153–164, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/39/
m039p153.pdf.

Lambshead:1988:ADE

[LP88] P. J. D. Lambshead and H. M. Piatt. Analysing disturbance with the Ewens/Caswell neutral model: theoretical review and practical assessment. Marine Ecology Progress Series, 43:31–41, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
43/m043p031.pdf.

Lorenzen:1987:MAF

[LPV87] S. Lorenzen, M. Prein, and C. Valentin. Mass aggregations of the free-living marine nematode Pontonema vulgare (Oncholaimidae) in organically polluted fjords. Marine Ecology Progress Series, 37: 27–34, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/37/m037p027.pdf.

Loya:1980:EOP

[LR80] Y. Loya and B. Rinkevich. Effects of oil pollution on coral reef communities. Marine Ecology Progress Series, 3:167–180, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
3/m003p167.pdf.

Lenanton:1982:NAD

[LRH82] R. C. J. Lenanton, A. I. Robertson, and J. A. Hansen. Nearshore accumulations of detached macrophytes as nursery areas for fish. Marine Ecology Progress Series, 9:51–57, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/9/m009p051.pdf.

Langton:1987:FRE

[LRS87] R. W. Langton, W. E. Robinson, and D. Schick. Fecundity and reproductive effort of sea scallops Placopecten magellanicus from the Gulf of Maine. Marine Ecology Progress Series, 37:19–25, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
37/m037p019.pdf.

Lewis:1983:SVS

[LS83] M. R. Lewis and J. C. Smith. A small volume, short-incubation-time method for measurement of photosynthesis as a function of incident irradiance. Marine Ecology Progress Series, 13:99–102, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p099.pdf.

Leger:1984:ISA

[LS84a] P. Leger and P. Sorgeloos. International study on Artemia. XXIX. Nutritional evaluation of Artemia nauplii from different geographical origin for the marine crustacean Mysidopsis bahia. Marine Ecology Progress Series, 15:307–309, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/15/m015p307.pdf.

Lewis:1984:DML

[LS84b] J. B. Lewis and G. S. Storey. Differences in morphology and life history traits of the echinoid Echinometra lucunter from different habitats. Marine Ecology Progress Series, 15:207–211, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/15/
m015p207.pdf.

Levinton:1988:ESO

[LS88] J. S. Levinton and S. Stewart. Effects of sediment organics, detrital input, and temperature on demography, production, and body size of a deposit feeder. Marine Ecology Progress Series, 49:259–266, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
49/m049p259.pdf.

Liebezeit:1985:RAA

[LSB85] G. Liebezeit, M. Schumann, and F. Bohde. Residual amino acid fluxes in Kiel Bight — February to June 1980. Marine Ecology Progress Series, 27:203–207, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/27/m027p203.pdf.

Levinton:1985:FLE

[LSD85] J. S. Levinton, S. Stewart, and T. H. Dewitt. Field and laboratory experiments on interference between Hydrobia totteni and Ilyanassa obsoleta (Gastropoda) and its possible relation to seasonal shifts in vertical mudflat zonation. Marine Ecology Progress Series, 22:53–58, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/22/m022p053.pdf.

Li:1984:TRP

[LSP84] W. K. W. Li, J. C. Smith, and T. Platt. Temperature response of photosynthetic capacity and carboxylase activity in Arctic marine phytoplankton. Marine Ecology Progress Series, 17:237–243, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
17/m017p237.pdf.

Levasseur:1984:HCP

[LT84] M. Levasseur and L. Therriault, J.-C, Legendre. Hierarchical control of phytoplankton succession by physical factors. Marine Ecology Progress Series, 19:211–222, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/19/m019p211.pdf.

Levasseur:1987:PBN

[LT87a] M. E. Levasseur and J.-C. Therriault. Phytoplankton biomass and nutrient dynamics in a tidally induced upwelling. The role of the NO_{3}:SiO_{4} ratio. Marine Ecology Progress Series, 39:87– 97, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
39/m039p087.pdf.

Lohrenz:1987:PPPa

[LT87b] S. E. Lohrenz and C. D. Taylor. Primary production of protein: I. Comparison of net cellular carbon and protein synthesis with ^{14}C-derived rate estimates in steady-state cultures of marine phytoplankton. Marine Ecology Progress Series, 35:277–292, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
35/m035p277.pdf.

Lohrenz:1987:PPPb

[LTH87] S. E. Lohrenz, C. D. Taylor, and B. L. Howes. Primary production of protein: II. Algal protein metabolism and its relation to particulate organic matter composition in the surface mixed layer. Marine Ecology Progress Series, 40:175–183, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/40/m040p175.pdf.

Lavens:1986:ISA

[LTS86] P. Lavens, W. Tackaert, and P. Sorgeloos. International study on Artemia. XLI. Influence of culture conditions and specific diapause deactivation methods on the hatchability of Artemia cysts produced in a standard culture system. Marine Ecology Progress Series, 31:197–203, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/31/m031p197.pdf.

Luckenbach:1984:SEP

[Luc84] M. W. Luckenbach. Settlement and early post-settlement survival in the recruitment of Mulinia lateralis (Bivalvia). Marine Ecology Progress Series, 17:245–250, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/17/m017p245.pdf.

Luning:1988:PCS

[Lün88] K. Lüning. Photoperiodic control of sorus formation in the brown alga Laminaria saccharina. Marine Ecology Progress Series, 45: 137–144, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/45/m045p137.pdf.

Laanbroek:1986:TVB

[LV86] H. J. Laanbroek and J. C. Verplanke. Tidal variations in bacterial biomass, productivity and oxygen uptake rates in a shallow channel in the Oosterschelde basin, the Netherlands. Marine Ecology Progress Series, 29:1–5, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/29/m029p001.pdf.

Liebezeit:1987:BFB

[LvB87] G. Liebezeit and B. von Bodungen. Biogenic fluxes in the Bransfield Strait: planktonic versus macroalgal sources. Marine Ecology Progress Series, 36:23–32, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/36/m036p023.pdf.

Laanbroek:1985:DPB

[LVdVdV85] H. J. Laanbroek, J. C. Verplanke, P. R. M. de Visscher, and R. de Vuyst. Distribution of phyto- and bacterioplankton growth and biomass parameters, dissolved inorganic nutrients and free amino acids during a spring bloom in the Oosterschelde basin, the Netherlands. Marine Ecology Progress Series, 25:1–11, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/25/
m025p001.pdf.

Lough:1989:EDJ

[LVP^{+}89] R. G. Lough, P. C. Valentine, D. C. Potter, P. J. Auditore, G. P. Bolz, J. D. Neilson, and R. I. Perry. Ecology and distribution of juvenile cod and haddock in relation to sediment type and bottom currents on eastern Georges Bank. Marine Ecology Progress Series, 56:1–12, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/56/m056p001.pdf.

Lindberg:1987:PPS

[LWE87] D. R. Lindberg, K. I. Warheit, and J. A. Estes. Prey preference and seasonal predation by oystercatchers on limpets at San Nicolas Island, California, USA. Marine Ecology Progress Series, 39:105– 113, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
39/m039p105.pdf.

Lewbel:1981:SWB

[LWG^{+}81] G. S. Lewbel, A. Wolfson, T. Gerrodette, W. H. Lippincon, J. L. Wilson, and M. M. Littler. Shallow-water benthic communities on California’s outer continental shelf. Marine Ecology Progress Series, 4:159–168, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/4/m004p159.pdf.

Macpherson:1981:RPM

[Mac81] E. Macpherson. Resource partitioning in a Mediterranean demersal fish community. Marine Ecology Progress Series, 4:183–193, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
4/m004p183.pdf.

MacDiarmid:1985:SRL

[Mac85a] A. B. MacDiarmid. Sunrise release of larvae from the palinurid rock lobster Jasus edwardsii. Marine Ecology Progress Series, 21: 313–315, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/21/m021p313.pdf.

Macpherson:1985:DRF

[Mac85b] E. Macpherson. Daily ration and feeding periodicity of some fishes off the coast of Namibia. Marine Ecology Progress Series, 26:253– 260, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
26/m026p253.pdf.

MacDonald:1986:PRP

[Mac86a] B. A. MacDonald. Production and resource partitioning in the giant scallop Placopecten magellanicus grown on the bottom and in suspended culture. Marine Ecology Progress Series, 34:79–86, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
34/m034p079.pdf.

MacLulich:1986:CBR

[Mac86b] J. H. MacLulich. Colonization of bare rock surfaces by microflora in a rocky intertidal habitat. Marine Ecology Progress Series, 32: 91–96, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/32/m032p091.pdf.

MacLulich:1986:EEM

[Mac86c] J. H. MacLulich. Experimental evaluation of methods for sampling and assaying intertidal epilithic microalgae. Marine Ecology Progress Series, 34:275–280, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/34/m034p275.pdf.

MacLulich:1987:VDV

[Mac87] J. H. MacLulich. Variations in the density and variety of intertidal epilithic microflora. Marine Ecology Progress Series, 40:285–293, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
40/m040p285.pdf.

Macpherson:1988:IGD

[Mac88] E. Macpherson. Influence of geographical distribution, body size and diet on population density of benthic fishes off Namibia (South West Africa). Marine Ecology Progress Series, 50:295– 299, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
50/m050p295.pdf.

Magnhagen:1988:CFR

[Mag88] C. Magnhagen. Changes in foraging as a response to predation risk in two gobiid fish species, Pomatoschistus minutus and Gobius niger. Marine Ecology Progress Series, 49:21–26, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/49/m049p021.pdf.

Mann:1988:DBL

[Man88] R. Mann. Distribution of bivalve larvae at a frontal system in the James River, Virginia. Marine Ecology Progress Series, 50:29– 44, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
50/m050p029.pdf.

Manahan:1989:AAF

[Man89] D. T. Manahan. Amino acid fluxes to and from seawater in axenic veliger larvae of the bivalve (Crassostrea gigas). Marine Ecology Progress Series, 53:247–255, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/53/m053p247.pdf.

Mark:1983:FDC

[Mar83a] D. M. Mark. Fractal dimension of a coral reef at ecological scales: a discussion. Marine Ecology Progress Series, 14:293–294, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
14/m014p293.pdf. See [BR82].

Martens:1983:CTD

[Mar83b] Paul M. Martens. Comparison of three different extraction methods for Turbellaria. Marine Ecology Progress Series, 14:229–234, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
14/m014p229.pdf.

Marcus:1984:RCN

[Mar84] N. H. Marcus. Recruitment of copepod nauplii into the plankton: importance of diapause eggs and benthic processes. Marine Ecology Progress Series, 15:47–54, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/15/m015p047.pdf.

Mardsen:1989:LRP

[Mar89a] J. R. Mardsen. Light responses of the planktotrophic larva of the serpulid polychaete Spirobranchus polycerus. Marine Ecology Progress Series, 58:225–233, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/58/m058p225.pdf.

MZC1:1989:FMZ

[Mar89b] Marine Zooplankton Colloquium 1. Future marine zooplankton research — a perspective. Marine Ecology Progress Series, 55:197– 206, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
55/m055p197.pdf.

Mauchline:1988:EBS

[Mau88] J. Mauchline. Egg and brood sizes of oceanic pelagic crustaceans. Marine Ecology Progress Series, 43:251–258, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/43/m043p251.pdf.

McLusky:1983:DPR

[MAWM83] D. S. McLusky, F. E. Anderson, and S. Wolfe-Murphy. Distribution and population recovery of Arenicola marina and other benthic fauna after bait digging. Marine Ecology Progress Series, 11: 173–179, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/11/m011p173.pdf.

Marra:1984:LWC

[MB84a] J. Marra and D. C. Boardman. Late winter chlorophyll-a distributions in the Weddell Sea. Marine Ecology Progress Series, 19:197– 205, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
19/m019p197.pdf.

McQuaid:1984:IST

[MB84b] C. D. McQuaid and G. M. Branch. Influence of sea temperature, substratum and wave exposure on rocky intertidal communities: an analysis of faunal and floral biomass. Marine Ecology Progress Series, 19:145–151, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/19/m019p145.pdf.

McQuaid:1985:TSR

[MB85] C. D. McQuaid and G. M. Branch. Trophic structure of rocky intertidal communities response to wave action and implications for energy flow. Marine Ecology Progress Series, 22:153–161, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
22/m022p153.pdf.

Matson:1987:CET

[MB87] E. A. Matson and J. D. Buck. Community electron transport of procaryotes in euryoxic estuarine sediment. Marine Ecology Progress Series, 41:71–78, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/41/m041p071.pdf.

McArdle:1989:MDV

[MB89] B. H. McArdle and R. G. Blackwell. Measurement of density variability in the bivalve Chione stutchburyi using spatial autocorrelation. Marine Ecology Progress Series, 52:245–252, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/52/m052p245.pdf.

Montagna:1986:BMN

[MBP^{+}86] P. A. Montagna, J. E. Bauer, M. C. Prieto, D. Hardin, and R. B. Spies. Benthic metabolism in a natural coastal petroleum seep. Marine Ecology Progress Series, 34:31–40, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/34/m034p031.pdf.

McEdward:1987:VEO

[MC87] L. R. McEdward and S. F. Carson. Variation in egg organic content and its relationship with egg size in the starfish Solaster stimpsoni. Marine Ecology Progress Series, 37:159–169, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
37/m037p159.pdf.

Mayzaud:1988:ART

[MC88] P. Mayzaud and R. J. Conover. O:N atomic ratio as a tool to describe zooplankton metabolism. Marine Ecology Progress Series, 45:289–302, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/45/m045p289.pdf.

Mayzaud:1989:SCB

[MCA89] P. Mayzaud, J. P. Chanut, and R. G. Ackman. Seasonal changes of the biochemical composition of marine particulate matter with special reference to fatty acids and sterols. Marine Ecology Progress Series, 56:189–204, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/56/m056p189.pdf.

McCourt:1984:NDB

[McC84] R. M. McCourt. Niche differences between sympatric Sargassum species in the northern Gulf of California. Marine Ecology Progress Series, 18:139–148, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/18/m018p139.pdf.

McClanahan:1988:SEA

[McC88] T. R. McClanahan. Seasonality in East Africa’s coastal waters. Marine Ecology Progress Series, 44:191–199, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/44/m044p191.pdf.

McClanahan:1989:KCR

[McC89a] T. R. McClanahan. Kenyan coral reef-associated gastropod fauna: a comparison between protected and unprotected reefs. Marine Ecology Progress Series, 53:11–20, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/53/m053p011.pdf.

McCormick:1989:RET

[McC89b] M. I. McCormick. Reproductive ecology of the temperate reef fish Cheilodactylus spectabilis (Pisces: Cheilodactylidae). Marine Ecology Progress Series, 55:113–120, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/55/m055p113.pdf.

McCormick:1989:STP

[McC89c] M. I. McCormick. Spatio-temporal patterns in the abundance and population structure of a large temperate reef fish. Marine Ecology Progress Series, 53:215–225, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/53/m053p215.pdf.

McDonald:1982:DLH

[McD82] Jack McDonald. Divergent life history patterns in the co-occurring intertidal crabs Panopeus herbstii and Eurypanopeus depressus (Crustacea: Brachyura: Xanthidae). Marine Ecology Progress Series, 8:173–180, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/8/m008p173.pdf.

McGurk:1986:NMM

[McG86] M. D. McGurk. Natural mortality of marine pelagic fish eggs and larvae: role of spatial patchiness. Marine Ecology Progress Series, 34:227–242, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/34/m034p227.pdf.

McGurk:1987:NMS

[McG87] M. D. McGurk. Natural mortality and spatial patchiness: reply to gulland. Marine Ecology Progress Series, 39:201–206, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/39/
m039p201.pdf.

McGuinness:1988:EPA

[McG88] K. A. McGuinness. Explaining patterns in abundances of organisms on boulders: the failure of ’natural experiments’. Marine Ecology Progress Series, 48:199–204, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/48/m048p199.pdf.

McGuinness:1989:ESN

[McG89a] K. A. McGuinness. Effects of some natural and artificial substrata on sessile marine organisms at Galeta Reef, Panama. Marine Ecology Progress Series, 52:201–208, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/52/m052p201.pdf.

McGurk:1989:ADM

[McG89b] M. D. McGurk. Advection, diffusion and mortality of Pacific herring larvae Clupea harengus pallasi in Bamfield Inlet, British Columbia. Marine Ecology Progress Series, 51:1–18, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/51/
m051p001.pdf.

Motzkin:1982:PRS

[MCGP82] F. Motzkin, Y. Cohen, H. Gordin, and E. Padan. Productivity relations in seawater fish ponds: a comparison of stocked and unstocked ponds. Marine Ecology Progress Series, 8:203–210, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
8/m008p203.pdf.

McLachlan:1984:BFR

[MCM84] A. McLachlan, A. C. Cockcroft, and D. E. Malan. Benthic faunal response to a high energy gradient. Marine Ecology Progress Series, 16:51–63, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/16/m016p051.pdf.

Malone:1988:IRF

[MCPW88] T. C. Malone, L. H. Crocker, S. E. Pike, and B. W. Wendler. Influences of river flow on the dynamics of phytoplankton production in a partially stratified estuary. Marine Ecology Progress Series, 48:235–249, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/48/m048p235.pdf.

Massin:1986:DFE

[MD86] C. Massin and C. Doumen. Distribution and feeding of epibenthic holothuroids on the reef flat of Laing Island (Papua New Guinea). Marine Ecology Progress Series, 31:185–195, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/31/m031p185.pdf.

Menesguen:1987:STA

[MD87] A. Menesguen and L. Dreves. Sea-temperature anomalies and population dynamics variations: effects on growth and density of three bivalves. Marine Ecology Progress Series, 36:11–21, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/36/m036p011.pdf.

Mladenov:1988:DSS

[ME88] P. V. Mladenov and R. H. Emson. Density, size structure and reproductive characteristics of fissiparous brittle stars in algae and sponges: evidence for interpopulational variation in levels of sexual and asexual reproduction. Marine Ecology Progress Series, 42: 181–194, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/42/m042p181.pdf.

McCallum:1989:SDA

[ME89] H. I. McCallum and A. M. Endean, R.. Cameron. Sublethal damage to Acanthasterplanci as an index of predation pressure. Marine Ecology Progress Series, 56:29–36, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/56/m056p029.pdf.

Meehan:1985:GCM

[Mee85] B. W. Meehan. Genetic comparison of Macoma balthica (Bivalvia, Telinidae) from the eastern and western North Atlantic Ocean. Marine Ecology Progress Series, 22:69–76, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/22/m022p069.pdf.

Morris:1984:FKAb

[MERW84] D. J. Morris, I. Everson, C. Ricketts, and P. Ward. Feeding of krill around South Georgia. II. Relations between feeding activity, environment and vertical distribution. Marine Ecology Progress Series, 20:203–206, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/20/m020p203.pdf.

McManus:1988:CBS

[MF88] G. B. McManus and J. A. Fuhrman. Clearance of bacteria-sized particles by natural populations of nanoplankton in the Chesapeake Bay outflow plume. Marine Ecology Progress Series, 42:

199–206, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/42/m042p199.pdf.

McTigue:1989:FJW

[MF89a] T. A. McTigue and R. J. Feller. Feeding of juvenile white shrimp Penaeus setiferus. Periodic or continuous? Marine Ecology Progress Series, 52:227–233, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/52/m052p227.pdf.

Mitchell:1989:CSV

[MF89b] J. G. Mitchell and J. A. Fuhrman. Centimeter scale vertical heterogeneity in bacteria and chlorophyll-a. Marine Ecology Progress Series, 54:141–148, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/54/m054p141.pdf.

Meyers:1987:MIM

[MFP87] M. B. Meyers, H. Fossing, and E. N. Powell. Microdistribution of interstitial meiofauna, oxygen and sulfide gradients, and the tubes of macro-infauna. Marine Ecology Progress Series, 35:223– 241, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
35/m035p223.pdf.

Mork:1983:GVC

[MFS83] J. Mork and E. Friis-Sörensen. Genetic variation in capelin Mallotus villosus from Norwegian waters. Marine Ecology Progress Series, 12:199–205, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/12/m012p199.pdf.

Miramand:1981:AAA

[MG81] P. Miramand and J. C. Guary. Association of americium-241 with adenochromes in the branchial hearts of the cephalopod Octopus vulgaris. Marine Ecology Progress Series, 4:127–129, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
4/m004p127.pdf.

Mauchline:1985:FSD

[MG85a] J. Mauchline and J. D. M. Gordon. Foraging strategies of deep-sea fish. Marine Ecology Progress Series, 27:227–238, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/27/m027p227.pdf.

Miramand:1985:SWU

[MG85b] P. Miramand and P. Germain. Sea water uptake, sediment transfer and histo-autoradiographic study of plutonium (^{239}Pu) and americium (^{241}Am) in the edible cockle Cerastoderma edule. Marine Ecology Progress Series, 22:59–68, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/22/m022p059.pdf.

Miramand:1982:UAP

[MGC82] P. Miramand, P. Germain, and H. Camus. Uptake of americium and plutonium from contaminated sediments by three benthic species: Arenicola marina, Corophium volutator and Scrobicularia plana. Marine Ecology Progress Series, 7:59–65, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/7/m007p059.pdf.

Muschenheim:1986:FBE

[MGM86] D. K. Muschenheim, J. Grant, and E. L. Mills. Flumes for benthic ecologists: theory, construction and practice. Marine Ecology Progress Series, 28:185–196, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/28/m028p185.pdf.

Miramand:1989:HAL

[MGT89] P. Miramand, P. Germain, and J. P. Trilles. Histo-autoradiographic localisation of americium (^{241}Am) in tissues of European lobster Homarus gammarus and edible crab Cancer pagurus after uptake from labelled sea water. Marine Ecology Progress Series, 52:217– 225, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
52/m052p217.pdf.

McLachlan:1984:FRM

[MH84] A. McLachlan and P. Hesp. Faunal response to morphology and water circulation of a sandy beach with cusps. Marine Ecology Progress Series, 19:133–144, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/19/m019p133.pdf.

Murray:1989:VSS

[MH89] S. N. Murray and M. H. Horn. Variations in standing stocks of central California macrophytes from a rocky intertidal habitat before and during the 1982–1983 El Niño. Marine Ecology Progress Series, 58:113–122, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/58/m058p113.pdf.

Marra:1985:OPM

[MHL85] J. Marra, K. Heinemann, and G. Landriau, Jr. Observed and predicted measurements of photosynthesis in a phytoplankton culture exposed to natural irradiance. Marine Ecology Progress Series, 24: 43–50, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/24/m024p043.pdf.

Mahoney:1988:LHR

[MHW88] J. B. Mahoney, D. Hollomon, and R. Waldhauer. Is the lower Hudson-raritan estuary a suitable habitat for Gonyaulax tamarensis? Marine Ecology Progress Series, 49:179–186, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/49/m049p179.pdf.

Miller:1985:SUP

[Mil85] R. J. Miller. Sea urchin pathogen: a possible tool for biological control. Marine Ecology Progress Series, 21:169–174, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/21/
m021p169.pdf.

Miller:1989:AEM

[Mil89] D. C. Miller. Abrasion effects on microbes in sandy sediments. Marine Ecology Progress Series, 55:73–82, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/55/m055p073.pdf.

Miller:1990:CJA

[Mil90] R. J. Miller. Comment on Jellett et al. (1989). Marine Ecology Progress Series, 63:305–306, 1990. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/63/m063p305.pdf. See [JNCS89].

Mittelstaedt:1987:CCC

[Mit87] E. Mittelstaedt. Cyclonic cold-core eddy in the eastern North Atlantic. I physical description. Marine Ecology Progress Series, 39: 145–152, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/39/m039p145.pdf.

Moreno:1984:ESF

[MJ84] C. A. Moreno and H. F. Jara. Ecological studies on fish fauna associated with Macrocystis pyrifera belts in the south of Fueguian Islands, Chile. Marine Ecology Progress Series, 15:99–107, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/15/
m015p099.pdf.

Martin-Jezequel:1988:IIC

[MJPH^{+}88] V. Martin-Jézéquel, S. A. Poulet, R. P. Harris, J. Moal, and J. F. Samain. Interspecific and intraspecific composition and variation of free amino acids in marine phytoplankton. Marine Ecology Progress Series, 44:303–313, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/44/m044p303.pdf.

Munk:1985:FBS

[MK85] P. Munk and T. Kiøboe. Feeding behaviour and swimming activity of larval herring (Clupea harengus) in relation to density of copepod nauplii. Marine Ecology Progress Series, 24:15–21, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
24/m024p015.pdf.

Malone:1986:LVP

[MKD^{+}86] T. C. Malone, W. M. Kemp, H. W. Ducklow, W. R. Boynton, J. H. Tuttle, and R. B. Jonas. Lateral variation in the production and fate of phytoplankton in a partially stratified estuary. Marine Ecology Progress Series, 32:149–160, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/32/m032p149.pdf.

McGrath:1988:EDS

[MKG88] D. McGrath, P. A. King, and E. M. Gosling. Evidence for the direct settlement of Mytilus edulis larvae on adult mussel beds. Marine Ecology Progress Series, 47:103–106, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/47/m047p103.pdf.

Morin:1985:OPI

[MKHD85] J. G. Morin, J. E. Kastendiek, A. Harrington, and N. Davis. Organization and patterns of interactions in a subtidal sand community on an exposed coast. Marine Ecology Progress Series, 27:163–185, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
27/m027p163.pdf.

Markham:1980:ECL

[MKS80] J. W. Markham, B. P. Kremer, and K. R. Sperling. Effects of cadmium on Laminaria saccarina in culture. Marine Ecology Progress Series, 3:31–39, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/3/m003p031.pdf.

Maurer:1980:DSP

[ML80] D. Maurer and W. Leathem. Dominant species of polychaetous annelids of Georges Bank. Marine Ecology Progress Series, 3:135– 144, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
3/m003p135.pdf.

Michel:1988:PSI

[MLDT88] C. Michel, L. Legendre, S. Demers, and J.-C. Therriault. Photoadaptation of sea-ice microalgae in springtime: photosynthesis and carboxylating enzymes. Marine Ecology Progress Series, 50:177– 185, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
50/m050p177.pdf.

Makarov:1981:ELD

[MM81] R. R. Makarov and V. V. Maslennikov. Ecology of larval development of the crustacean Euphausia superba. Change in dominant larval forms as a function of environmental conditions. Marine Ecology Progress Series, 4:265–271, 1981. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p265.pdf.

MacLachlan:1986:DSB

[MM86] A. MacLachlan and L. McGwynne. Do sandy beaches accumulate nitrogen? Marine Ecology Progress Series, 34:191–195, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
34/m034p191.pdf.

Maruyama:1989:MPA

[MMS89] A. Maruyama, M. Maeda, and U. Simidu. Microbial production of auxin indole-3-acetic acid in marine sediments. Marine Ecology Progress Series, 58:69–75, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/58/m058p069.pdf.

Mahon:1987:DCS

[MN87] R. Mahon and J. D. Neilson. Diet changes in Scotian Shelf haddock during the pelagic and demersal phases of the first year of life. Marine Ecology Progress Series, 37:123–130, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/37/m037p123.pdf.

MollerJensen:1983:PRE

[Mø83] L. Møller Jensen. Phytoplankton release of extracellular organic carbon, molecular weight composition, and bacterial assimilation. Marine Ecology Progress Series, 11:39–48, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/11/m011p039.pdf.

Moerland:1985:CAN

[Moe85] T. S. Moerland. Cellulase activity in natural and temperature acclimated populations of Fundulus heteroclitus. Marine Ecology Progress Series, 26:305–308, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/26/m026p305.pdf.

Moller:1986:PFB

[Möl86] P. Möller. Physical factors and biological interactions regulating infauna in shallow boreal areas. Marine Ecology Progress Series, 30:33–47, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/30/m030p033.pdf.

Montagna:1983:LCR

[Mon83] P. A. Montagna. Live controls for radioisotope tracer food chain experiments using meiofauna. Marine Ecology Progress Series, 12: 43–46, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/12/m012p043.pdf.

Montagna:1984:LSM

[Mon84] P. A. Montagna. In situ measurement of meiobenthic grazing rates on sediment bacteria and edaphic diatoms. Marine Ecology Progress Series, 18:119–130, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/18/m018p119.pdf.

Monta:1989:MDC

[Mon89] R. Y. Monta. In memoriam: Dr Claude E. ZoBell (1904–1989). Marine Ecology Progress Series, 58:1–2, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/58/m058p001.pdf.

Mook:1981:RSP

[Moo81] D. H. Mook. Removal of suspended particles by fouling communities. Marine Ecology Progress Series, 5:279–281, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/5/m005p279.pdf.

Moore:1988:CHEb

[Moo88a] M. N. Moore. Cellular- and histopathological effects of a pollutant gradient summary. Marine Ecology Progress Series, 46:109–110, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p109.pdf.

Moore:1988:CHEa

[Moo88b] M. N. Moore. Cellular- and histopathological effects of a polutant gradient introduction MEPS 46:79. Marine Ecology Progress Series, ??:??, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/46/m046p079.pdf.

Moore:1988:CRL

[Moo88c] M. N. Moore. Cytochemical responses of the lysosomal system and NADPH-ferrihemoprotein reductase in molluscan digestive cells to environmental and experimental exposure to xenobiotics. Marine Ecology Progress Series, 46:81–89, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/46/m046p081.pdf.

Morrison:1983:SBO

[Mor83] D. Morrison. Seasonality of Batophora oersted (Chlorophyta), a tropical macroalga. Marine Ecology Progress Series, 14:235–244, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
14/m014p235.pdf.

Moran:1985:DDP

[Mor85a] M. J. Moran. Distribution and dispersion of the predatory intertidal gastropod Morula marginalba. Marine Ecology Progress Series, 22:41–52, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/22/m022p041.pdf.

Morris:1985:IMM

[Mor85b] D. J. Morris. Integrated model of moulting and feeding of Antarctic krill Euphausia superba off South Georgia. Marine Ecology Progress Series, 22:207–217, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/22/m022p207.pdf.

Morales:1987:CNC

[Mor87] C. E. Morales. Carbon and nitrogen content of copepod faecal pellets: effect of food concentration and feeding behavior. Marine Ecology Progress Series, 36:107–114, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/36/m036p107.pdf.

Moriarty:1981:DSM

[MP81] D. J. W. Moriarty and P. C. Pollard. DNA synthesis as a measure of bacterial productivity in seagrass sediments. Marine Ecology Progress Series, 5:151–156, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/5/m005p151.pdf.

Monteleone:1986:FEA

[MP86] D. M. Monteleone and W. T. Peterson. Feeding ecology of American sand lance Ammodytes americanus larvae from Long Island Sound. Marine Ecology Progress Series, 30:133–143, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/30/
m030p133.pdf.

McManus:1988:BPN

[MP88a] G. B. McManus and W. T. Peterson. Bacterioplankton production in the nearshore zone during upwelling off central Chile. Marine Ecology Progress Series, 43:11–17, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/43/m043p011.pdf.

Montgomery:1988:GAP

[MP88b] W. Linn Montgomery and Peggy E. Pollak. Gut anatomy and pH in a Red Sea surgeon-fish, Acanthurus nigrofuscus. Marine Ecology Progress Series, 44:7–13, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/44/m044p007.pdf.

Moller:1985:BFE

[MPR85] P. Möller, L. Pihl, and R. Rosenberg. Benthic faunal energy flow and biological interaction in some shallow marine soft bottom habitats. Marine Ecology Progress Series, 27:109–121, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/27/
m027p109.pdf.

Meadows:1981:MAB

[MR81] P. S. Meadows and A. A. Ruagh. Multifactorial analysis of behavioural responses of the amphipod Corophium volutator to temperature-salinity combinations. Marine Ecology Progress Series, 6:183–190, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/6/m006p183.pdf.

Myers:1983:PSN

[MR83] R. A. Myers and J. A. Runge. Predictions of seasonal natural mortality rates in a copepod population using life-history theory. Marine Ecology Progress Series, 11:189–194, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/11/m011p189.pdf.

Morris:1984:FKAa

[MR84] D. J. Morris and C. Ricketts. Feeding of krill around South Georgia. I: A model of feeding activity in relation to depth and time of day. Marine Ecology Progress Series, 16:1–7, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/16/m016p001.pdf.

Meyer-Reil:1986:MHA

[MR86] L.-A. Meyer-Reil. Measurement of hydrolytic activity and incorporation of dissolved organic substrates by microorganisms in marine sediments. Marine Ecology Progress Series, 31:143–149, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
31/m031p143.pdf.

Mimura:1988:MSR

[MR88a] T. Mimura and Y. Romano, J.-C, Souza-Lima. Microbiomass structure and respiratory activity of microneuston and microplankton in the northwestern Mediterranean Sea influenced by Rhone river water. Marine Ecology Progress Series, 49:151–162, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
49/m049p151.pdf.

Moran:1988:BAS

[MR88b] D. P. Moran and M. L. Reaka. Bioerosion and availability of shelter for benthic reef organisms. Marine Ecology Progress Series, 44: 249–263, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/44/m044p249.pdf.

Meyer-Reil:1980:UOM

[MRF80] L.-A. Meyer-Reil and A. Faubel. Uptake of organic matter by meiofauna organisms and interrelationships with bacteria. Marine Ecology Progress Series, 3:251–256, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/3/m003p251.pdf.

Morley:1984:SMW

[MS84] J. J. Morley and J. C. Stepien. Siliceous microfauna in waters beneath Antarctic sea ice. Marine Ecology Progress Series, 19: 207–210, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/19/m019p207.pdf.

Mcdonald:1988:CMI

[MS88] C. R. Mcdonald and J. B. Sprague. Cadmium in marine invertebrates and Arctic cod in the Canadian Arctic. Distribution and ecological implications. Marine Ecology Progress Series, 47:17– 30, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
47/m047p017.pdf.

Murioz:1989:DDA

[MS89] M. Murioz and B. Santelices. Determination of the distribution and abundance of the limpet Scurria scurra on the stipes of the kelp Lessonia nigrescens in Central Chile. Marine Ecology Progress Series, 54:277–285, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/54/m054p277.pdf.

Mayer:1986:MPN

[MSS86] L. M. Mayer, L. L. Schick, and F. W. Setchell. Measurement of protein in nearshore marine sediments. Marine Ecology Progress Series, 30:159–165, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/30/m030p159.pdf.

Mironov:1981:SHM

[MST81] O. G. Mironov, T. L. Shchekaturina, and I. M. Tsimbal. Saturated hydrocarbons in marine organisms. Marine Ecology Progress Series, 5:303–309, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/5/m005p303.pdf.

Maeda:1980:AAB

[MT80] M. Maeda and N. Taga. Alkalotolerant and alkalophilic bacteria in seawater. Marine Ecology Progress Series, 2:105–108, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
2/m002p105.pdf.

Morris:1984:HRR

[MT84] S. Morris and A. C. Taylor. Heart rate response of the intertidal prawn Palaemon elegans to simulated and in situ environmental changes. Marine Ecology Progress Series, 20:127–136, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
20/m020p127.pdf.

MacDonald:1985:ITFa

[MT85a] B. A. MacDonald and R. J. Thompson. Influence of temperature and food availability on the ecological energetics of the giant scallop Placopecten magellanicus. I. Growth rates of shell and somatic tissue. Marine Ecology Progress Series, 25:279–294, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/25/m025p279.pdf.

MacDonald:1985:ITFb

[MT85b] B. A. MacDonald and R. J. Thompson. Influence of temperature and food availability on the ecological energetics of the giant scallop Placopecten magellanicus. II. Reproductive output and total production. Marine Ecology Progress Series, 25:295–303, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
25/m025p295.pdf.

Martinussen:1987:UPO

[MT87] I. Martinussen and T. F. Thingstad. Utilization of N, P and organic C by heterotrophic bacteria. II. Comparison of experiments and a mathematical model. Marine Ecology Progress Series, 37: 285–293, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/37/m037p285.pdf.

Mapstone:1984:EAC

[MUC84] B. D. Mapstone, A. J. Underwood, and R. G. Creese. Experimental analyses of the commensal relation between intertidal gastropods Patelloida mufria and the trochid Austrocochlea constricta. Marine Ecology Progress Series, 17:85–100, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/17/m017p085.pdf.

Murina:1984:ES

[Mur84] G.-V. V. Murina. Ecology of sipuncula. Marine Ecology Progress Series, 17:1–7, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/17/m017p001.pdf.

Mann:1983:SBL

[MW83] R. Mann and C. C. Wolf. Swimming behaviour of larvae of the ocean quahog Arctica islandica in response to pressure and temperature. Marine Ecology Progress Series, 13:211–218, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/13/
m013p211.pdf.

MacRae:1984:PMQ

[MW84] A. F. MacRae and S. A. Woodin. A photographic method for quantifying surface defecation. Marine Ecology Progress Series, 19:233–236, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/19/m019p233.pdf.

Morris:1985:GAS

[MW85] J. T. Morris and G. J. Whiting. Gas advection in sediments of a South Carolina salt marsh. Marine Ecology Progress Series, 27: 187–194, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/27/m027p187.pdf.

Murray:1987:OPC

[MW87] L. Murray and R. L. Wetzel. Oxygen production and consumption associated with the major autotrophic components in two temperate seagrass communities. Marine Ecology Progress Series, 38:231– 239, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
38/m038p231.pdf.

Muscatine:1986:RPD

[MWM86] L. Muscatine, F. P. Wilkerson, and L. R. McCloskey. Regulation of population density of symbiotic algae in a tropical marine jellyfish (Mastigias sp.). Marine Ecology Progress Series, 32:279– 290, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
32/m032p279.pdf.

Napp:1988:VDMb

[NBMM88] J. M. Napp, E. R. Brooks, P. Matrai, and M. M. Mullin. Vertical distribution of marine particles and grazers. II. Relation of grazer distribution to food quality and quantity. Marine Ecology Progress Series, 50:59–72, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/50/m050p059.pdf.

Napp:1988:VDMa

[NBR^{+}88] J. M. Napp, E. R. Brooks, F. M. H. Reid, P. Matrai, and M. M. Mullin. Vertical distribution of marine particles and grazers. I. Vertical distribution of food quality and quantity. Marine Ecology Progress Series, 50:45–58, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/50/m050p045.pdf.

Nygaard:1988:GRB

[NBT88] K. Nygaard, K. Y. Børsheim, and T. F. Thingstad. Grazing rates on bacteria by marine heterotrophic microflagellates compared to uptake rates of bacterial-sized monodisperse fluorescent latex beads. Marine Ecology Progress Series, 44:159–165, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/44/m044p159.pdf.

Nelson:1989:SMP

[NC89] Amy L. Nelson and Bruce C. Coull. Selection of meiobenthic prey by juvenile spot (Pisces): an experiments, study. Marine Ecology Progress Series, 53:51–57, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/53/m053p051.pdf.

Nelson:1983:LTS

[NCG^{+}83] D. A. Nelson, A. Calabrese, R. A. Greig, P. P. Yevich, and S. Chang. Long-term silver effects on the marine gastropod Crepidula fornicate. Marine Ecology Progress Series, 12:155–165, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
12/m012p155.pdf.

Nelson:1981:ESD

[Nel81a] W. G. Nelson. Experimental studies of decapod and fish predation on seagrass macrobenthos. Marine Ecology Progress Series, 5:141– 149, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
5/m005p141.pdf.

Nelson:1981:IBS

[Nel81b] W. G. Nelson. Inhibition of barnacle settlement by ekofisk crude oil. Marine Ecology Progress Series, 5:41–43, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/5/m005p041.pdf.

Nelson:1989:PPM

[Nel89] J. R. Nelson. Phytoplankton pigments in macrozooplankton feces: variability in carotenoid alterations. Marine Ecology Progress Series, 52:129–144, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/52/m052p129.pdf.

Newell:1982:EBS

[NFG82] R. C. Newell, J. G. Field, and C. L. Griffiths. Energy balance and significance of microorganisms in a kelp bed community. Marine Ecology Progress Series, 8:103–113, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/8/m008p103.pdf.

Nienhuis:1986:CEZ

[NG86] P. H. Nienhuis and A. M. Groenendijk. Consumption of eelgrass (Zostera marina) by birds and invertebrates: an annual budget. Marine Ecology Progress Series, 29:29–35, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/29/m029p029.pdf.

Neilson:1986:ETP

[NGR86] M. J. Neilson, R. L. Giddins, and G. N. Richards. Effect of tannins on the payability of mangrove leaves to the tropical sesarminid crab Neosarmatium smithi. Marine Ecology Progress Series, 34: 185–186, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/34/m034p185.pdf.

Nair:1982:PDE

[NGV^{+}82] K. K. C. Nair, T. C. Gopalakrishnan, P. Venugopal, M. G. Peter, K. V. Jayalakshmi, and T. S. S. Rao. Population dynamics of estuarine amphipods in Cochin Backwaters. Marine Ecology Progress Series, 10:289–295, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/10/m010p289.pdf.

Newkirk:1982:PSG

[NH82] G. F. Newkirk and L. E. Haley. Progress in selection for growth rate in the European oyster Ostrea edulis. Marine Ecology Progress Series, 10:77–79, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/10/m010p077.pdf.

Nicol:1984:PSD

[Nic84] S. Nicol. Population structure of daytime surface swarms of the euphausiid Meganyctiphanes norvegica in the Bay of Fundy. Marine Ecology Progress Series, 18:241–251, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/18/m018p241.pdf.

Nilsson:1987:DML

[Nil87] P. Nilsson. Demography of Mesochra lilljeborgi and Amonardia normani (Copepoda. Harpacticoida) reared on two diatom diets. Marine Ecology Progress Series, 39:267–274, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/39/m039p267.pdf.

Newell:1983:PIO

[NJ83] R. I. E. Newell and S. J. Jordan. Preferential ingestion of organic material by the American oyster Crassostrea virginica. Marine Ecology Progress Series, 13:47–53, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/13/m013p047.pdf.

Novitsky:1981:PMH

[NK81] J. A. Novitsky and P. E. Kepkay. Patterns of microbial heterotrophy through changing environments in a marine sediment. Marine Ecology Progress Series, 4:1–7, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p001.pdf.

Nielsen:1982:SFE

[NK82] M. V. Nielsen and L. H. Kofoed. Selective feeding and epipsammic browsing by the deposit-feeding amphipod Corophium volutator. Marine Ecology Progress Series, 10:81–88, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/10/m010p081.pdf.

Novitsky:1986:CMA

[NK86] J. A. Novitsky and D. M. Karl. Characterization of microbial activity in the surface layers of a coastal subtropical sediment. Marine Ecology Progress Series, 28:49–55, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/28/m028p049.pdf.

Nawrocki:1989:DAT

[NK89] M. P. Nawrocki and D. M. Karl. Dissolved ATP turnover in the Bransfield Strait, Antarctica during a spring bloom. Marine Ecology Progress Series, 57:35–44, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/57/m057p035.pdf.

Noel-Lambot:1981:PIL

[NL81] F. Noel-Lambot. Presence in the intestinal lumen of marine fish of corpuscles with a high cadmium-, zinc- and copper-binding capacity: a possible mechanism of heavy metal tolerance. Marine Ecology Progress Series, 4:175–181, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p175.pdf.

Newell:1984:SMD

[NL84] R. C. Newell and E. A. S. Linley. Significance of microheterotrophs in the decomposition of phytoplankton: estimates of carbon and nitrogen flow based on the biomass of plankton communities. Marine Ecology Progress Series, 16:105–119, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/16/m016p105.pdf.

Newell:1986:DAR

[NL86] R. I. E. Newell and C. J. Langdon. Digestion and absorption of refractory carbon from the plant Spartina alterniflora by the oyster Crassosfrea virginica. Marine Ecology Progress Series, 34: 105–115, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/34/m034p105.pdf.

Noel-Lambot:1980:CZC

[NLBFD80] F. Noel-Lambot, J. M. Bouquegneau, F. Frankenne, and A. Disteche. Cadmium, zinc and copper accumulation in limpets (Patella vulgata) from the Bristol Channel with special reference to metallothioneins. Marine Ecology Progress Series, 2:81–89, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
2/m002p081.pdf.

Newell:1981:RDE

[NLL81] R. C. Newell, M. I. Lucas, and E. A. S. Linley. Rate of degradation and efficiency of conversion of phytoplankton debris by marine micro-organisms. Marine Ecology Progress Series, 6:123–136, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
6/m006p123.pdf.

Newell:1980:QSD

[NLVS80] R. C. Newell, M. I. Lucas, B. Velirnirov, and L. J. Seiderer. Quantitative significance of dissolved organic losses following fragmentation of kelp (Ecklonia maxima and Laminaria pallida). Marine Ecology Progress Series, 2:45–59, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/2/m002p045.pdf.

Neale:1985:STV

[NM85] P. J. Neale and J. Marra. Short-term variation of Pmax under natural irradiance conditions: a model and its implications. Marine Ecology Progress Series, 26:113–124, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/26/m026p113.pdf.

Novitsky:1989:MHE

[NM89] J. A. Novitsky and M. C. MacSween. Microbiology of a high energy beach sediment: evidence for an active and growing community. Marine Ecology Progress Series, 52:71–75, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/52/m052p071.pdf.

Nowicki:1985:BCM

[NN85] B. L. Nowicki and S. W. Nixon. Benthic community metabolism in a coastal lagoon ecosystem. Marine Ecology Progress Series, 22:

21–30, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/22/m022p021.pdf.

Nissen:1984:MUD

[NNA84] H. Nissen, P. Nissen, and F. Azam. Multiphasic uptake of dglucose by an oligotrophic marine bacterium. Marine Ecology Progress Series, 16:155–160, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/16/m016p155.pdf.

Nultsch:1987:FWD

[NPMW87] W. Nultsch, J. Pfau, and M. Materna-Weide. Fluence and wavelength dependence of photomhibition in the brown alga Dictyota dichotoma. Marine Ecology Progress Series, 41:93–97, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/41/
m041p093.pdf.

Neilson:1987:ICE

[NPSV87] J. D. Neilson, R. I. Perry, J. S. Scott, and P. Valerio. Interactions of caligid ectoparasites and juvenile gadids on Georges Bank. Marine Ecology Progress Series, 39:221–232, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/39/m039p221.pdf.

Neilson:1986:CAC

[NPVW86] J. D. Neilson, R. I. Perry, P. Valerio, and K. G. Waiwood. Condition of Atlantic cod Gadus morhua larvae after the transition to exogenous feeding: morphometries, buoyancy and predator avoidance. Marine Ecology Progress Series, 32:229–235, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/32/m032p229.pdf.

Nielsen:1989:FCS

[NR89a] T. G. Nielsen and K. Richardson. Food chain structure of the North Sea plankton communities: seasonal variations of the role of the microbial loop. Marine Ecology Progress Series, 56:75–87, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
56/m056p075.pdf.

Nugegoda:1989:ESC

[NR89b] D. Nugegoda and P. S. Rainbow. Effects of salinity changes on zinc uptake and regulation by the decapod crustaceans Palaemon elegans and Palaemonetes varians. Marine Ecology Progress Series, 51:57–75, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/51/m051p057.pdf.

Nugegoda:1989:SOZ

[NR89c] D. Nugegoda and P. S. Rainbow. Salinity, osmolality, and zinc uptake in Palaemon elegans (Crustacea: Decapoda). Marine Ecology Progress Series, 55:149–157, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/55/m055p149.pdf.

Nakamura:1980:MBO

[NS80] K. Nakamura and K. Sekiguchi. Mating behavior and oviposition in the pycnogonid Propallene longiceps. Marine Ecology Progress Series, 2:163–168, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/2/m002p163.pdf.

Nadarajalingam:1987:ILE

[NS87] K. Nadarajalingam and T. Subramoniam. Influence of light on endocrine system and ovarian activity in the ocypodid crabs Ocypoda platytarsis and O. macrocera. Marine Ecology Progress Series, 36: 43–53, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/36/m036p043.pdf.

Nichols:1985:PIM

[NT85] F. H. Nichols and J. K. Thompson. Persistence of an introduced mudflat community in South San Francisco Bay, California. Marine Ecology Progress Series, 24:83–97, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/24/m024p083.pdf.

Nothig:1989:OVF

[NvB89] E.-M. Nöthig and B. von Bodungen. Occurrence and vertical flux of faecal pellets of probably protozoan origin in the southeastern Weddell Sea (Antarctica). Marine Ecology Progress Series, 56:

281–289, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/56/m056p281.pdf.

Neori:1988:CCF

[NVHHH88] A. Neori, M. Vernet, O. Holm-Hansen, and F. T. Haxo. Comparison of chlorophyll far-red and red fluorescence excitation spectra with photo-synthetic oxygen action spectra for photo-system II in algae. Marine Ecology Progress Series, 44:297–302, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/44/m044p297.pdf.

Ning:1988:SSP

[NVLL88] X. Ning, D. Vaulot, Zh Liu, and Z. Liu. Standing stock and production of phytoplankton in the estuary of the Chang-jiang (Yangtse River) and the adjacent East China Sea. Marine Ecology Progress Series, 49:141–150, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/49/m049p141.pdf.

Noldt:1984:QEL

[NW84] U. Noldt and C. Wehrenberg. Quantitative extraction of living plathelminthes from marine sands. Marine Ecology Progress Series, 20:193–201, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/20/m020p193.pdf.

Overnell:1988:MML

[OA88] J. Overnell and M. I. Abdullah. Metallothionein and metal levels in flounder Platichthys flesus from four f: eid sites and n flounder dosed with water-borne copper. Marine Ecology Progress Series, 46:71–74, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/46/m046p071.pdf.

OBrien:1988:DOC

[O’B88a] D. P. O’Brien. Direct observations of clustering (schooling and swarming) behaviour in mysids (Crustacea: Mysidacea). Marine Ecology Progress Series, 42:235–246, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/42/m042p235.pdf.

OBrien:1988:SSB

[O’B88b] D. P. O’Brien. Surface schooling behaviour of the coastal krill Nyctiphanes australis (Crustacea: Euphausiacea) off Tasmania, Australia. Marine Ecology Progress Series, 42:219–233, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/42/m042p219.pdf.

Olney:1988:NIA

[OB88c] J. E. Olney and G. W. Boehlert. Nearshore ichthyoplankton associated with seagrass beds in the lower Chesapeake Bay. Marine Ecology Progress Series, 45:33–43, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/45/m045p033.pdf.

ONeil:1989:NAT

[OC89] J. M. O’Neil and D. G. Capone. Nitrogenase activity in tropical carbonate marine sediments. Marine Ecology Progress Series, 56: 145–156, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/56/m056p145.pdf.

Ojeda:1989:CSM

[OD89] F. P. Ojeda and J. H. Dearborn. Community structure of macroinvertebrates inhabiting the rocky subtidal zone in the Gulf of Maine: seasonal and bathymetric distribution. Marine Ecology Progress Series, 57:147–161, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/57/m057p147.pdf.

Ogden:1981:SCS

[OE81] J. C. Ogden and J. P. Ebersole. Scale and community structure of coral reef fishes: a long-term study of a large artificial reef. Marine Ecology Progress Series, 4:97–103, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p097.pdf.

Ohki:1982:LCP

[OF82] K. Ohki and Y. Fujita. Laboratory culture of the pelagic bluegreen alga Trichodesmium thiebautii: Conditions for unialgal culture. Marine Ecology Progress Series, 7:185–190, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/7/m007p185.pdf.

Oviatt:1982:LCA

[OFGG82] C. Oviatt, J. Frithsen, J. Gearing, and P. Gearing. Low chronic additions of no. 2 fuel oil: Chemical behavior, biological impact and recovery in a simulated estuarine environment. Marine Ecology Progress Series, 9:121–136, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/9/m009p121.pdf.

OBrien:1988:CTL

[OFGL88] P. O’Brien, H. Feldman, E. V. Grill, and A. G. Lewis. Copper tolerance of the life history stages of the splashpool copepod Tigriopus californicus (Copepoda, Harpacticoida). Marine Ecology Progress Series, 44:59–64, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/44/m044p059.pdf.

Ohman:1984:OEP

[Ohm84] M. D. Ohman. Omnivory by Euphausia pacifica: the role of copepod prey. Marine Ecology Progress Series, 19:125–131, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/19/
m019p125.pdf.

Ohman:1988:SVM

[Ohm88] M. D. Ohman. Sources of variability in measurements of copepod lipids and gut fluorescence in the California Current coastal zone. Marine Ecology Progress Series, 42:143–153, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/42/m042p143.pdf.

Ornes:1989:MST

[OK89] W. H. Ornes and D. I. Kaplan. Macronutrient status of tall and short forms of Spartina alterniflora in a South Carolina salt marsh. Marine Ecology Progress Series, 55:63–72, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/55/m055p063.pdf.

Okamura:1986:FDA

[Oka86] B. Okamura. Formation and disruption of aggregations of Mytilus edulis in the fouling community of San Francisco Bay, California. Marine Ecology Progress Series, 30:275–282, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/30/m030p275.pdf.

Oviatt:1986:PPD

[OKSB86] C. A. Oviatt, A. A. Keller, P. A. Sampou, and L. L. Beatty. Patterns of productivity during eutrophication: a mesocosm experiment. Marine Ecology Progress Series, 28:69–80, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/28/m028p069.pdf.

Olafsson:1989:CIS

[Ola89] E. B. Olafsson. Contrasting influences of suspension-feeding and deposit-feeding populations of Macoma balthica on infaunal recruitment. Marine Ecology Progress Series, 55:171–179, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/55/
m055p171.pdf.

Oliver:1984:SAR

[Oli84] J. S. Oliver. Selection for asexual reproduction in an Antarctic polychaete worm. Marine Ecology Progress Series, 19:33–38, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
19/m019p033.pdf.

Olson:1985:LSC

[Ols85] R. R. Olson. In situ culturing of larvae of the crown-of-thorns starfish Acanthaster planci. Marine Ecology Progress Series, 25: 207–210, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/25/m025p207.pdf.

Oakden:1983:BRP

[OOF83] J. M. Oakden, J. S. Oliver, and A. R. Flegal. Behavioral responses of a phoxo-cephalid amphipod to organic enrichment and trace metals in sediment. Marine Ecology Progress Series, 14:253–257, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
14/m014p253.pdf.

Oliver:1982:PAC

[OOS82] J. S. Oliver, J. M. Oakden, and P. N. Slattery. Phoxocephalid amphipod crustaceans as predators on larvae and juveniles in marine soft-bottom communities. Marine Ecology Progress Series, 7:179– 184, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
7/m007p179.pdf.

Oviatt:1984:RPE

[OPN^{+}84] C. A. Oviatt, M. E. Q. Pilson, S. W. Nixon, J. B. Frithsen, D. T. Rudnick, J. R. Kelly, J. F. Grassle, and J. P. Grassle. Recovery of a polluted estuarine system: a mesocosm experiment. Marine Ecology Progress Series, 16:203–217, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/16/m016p203.pdf.

Onuf:1983:FCC

[OQ83] C. P. Onuf and M. L. Quammen. Fishes in a California coastal lagoon: effects of major storms on distribution and abundance. Marine Ecology Progress Series, 12:1–14, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/12/m012p001.pdf.

Oviatt:1987:FES

[OQM^{+}87] C. A. Oviatt, J. G. Quinn, J. T. Maughan, J. T. Ellis, B. K. Sullivan, J. N. Gearing, P. J. Gearing, C. D. Hunt, P. A. Sampou, and J. S. Latimer. Fate and effects of sewage sludge in the coastal marine environment: a mesocosm experiment. Marine Ecology Progress Series, 41:187–203, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/41/m041p187.pdf.

Officer:1980:PIS

[OR80] C. B. Officer and J. H. Ryther. The possible importance of silicon in marine eutrophication. Marine Ecology Progress Series, 3:83– 91, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
3/m003p083.pdf.

Oresland:1986:TDS

[Øre86] V. Øresland. Temporal distribution of size and maturity stages of the chaetognath Sagitta setosa in the western English Channel. Marine Ecology Progress Series, 29:55–60, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/29/m029p055.pdf.

Oresland:1987:FCS

[Øre87] V. Øresland. Feeding of the chaetognaths Sagitta elegans and S. setosa at different seasons in Gullmarsfjorden, Sweden. Marine Ecology Progress Series, 39:69–79, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/39/m039p069.pdf.

Oviatt:1986:CSC

[ORK^{+}86] Candace A. Oviatt, David T. Rudnick, Aimee A. Keller, Peter A. Sampou, and Glen T. Almquist. A comparison of system (O_{2} and CO_{2}) and C-14 measurements of metabolism in estuarine mesocosms. Marine Ecology Progress Series, 28:57–67, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/28/m028p057.pdf.

Osterroht:1980:VTC

[OS80] C. Osterroht and V. Smetacek. Vertical transport of chlorinated hydrocarbons by sedimentation of particulate matter in Kiel Bight. Marine Ecology Progress Series, 2:27–34, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/2/m002p027.pdf.

Ojeda:1984:EDL

[OS84a] F. P. Ojeda and B. Santelices. Ecological dominance of Lessonia nigrescens (Phaeophyta) in central Chile. Marine Ecology Progress Series, 19:83–91, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/19/m019p083.pdf.

Ojeda:1984:ICH

[OS84b] F. P. Ojeda and B. Santelices. Invertebrate communities in holdfasts of the kelp Macrocystis pyrifera from southern Chile. Marine Ecology Progress Series, 16:65–73, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/16/m016p065.pdf.

Oshel:1985:APH

[OS85] P. E. Oshel and D. H. Steele. Amphipod Paramphithoe hystrix: a micro-predator on the sponge Haliclona ventilabrum. Marine Ecology Progress Series, 23:307–309, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/23/m023p307.pdf.

Oeschger:1988:AVT

[OS88] R. Oeschger and R. Schmaljohann. Association of various types of epi-bacteria with Halicryptus spinulosus (Priapulida). Marine Ecology Progress Series, 48:285–293, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/48/m048p285.pdf.

Olla:1985:BJB

[OSB85] Bori L. Olla, Anne L. Studholme, and Allen J. Bejda. Behavior of juvenile bluefish Pomatomus saltatrix in vertical thermal gradients: influence of season, temperature acclimation and food. Marine Ecology Progress Series, 23:165–177, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/23/m023p165.pdf.

Officer:1982:BFF

[OSM82] C. B. Officer, T. J. Smayda, and R. Mann. Benthic filter feeding: a natural eutrophication control. Marine Ecology Progress Series, 9: 203–210, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/9/m009p203.pdf.

Orr:1982:BGC

[OTC82] J. Orr, J. P. Thorpe, and M. A. Carter. Biochemical genetic confirmation of the asexual reproduction of brooded offspring in the sea anemone Actinia equine. Marine Ecology Progress Series, 7:227– 229, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
7/m007p227.pdf.

Ota:1980:CLY

[OTK80] T. Ota, T. Takagi, and S. Kosaka. Changes in lipids of young and adult saury Cololabis saira (Pisces). Marine Ecology Progress Series, 3:11–17, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/3/m003p011.pdf.

Orth:1987:USM

[Ov87] R. J. Orth and J. van Montîrans. Ultilization of a seagrass meadow and tidal marsh creek by blue crabs Callinectes sapidus. I. Seasonal and annual variations in abundance with emphasis on postsettlement juveniles. Marine Ecology Progress Series, 41:283–294, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
41/m041p283.pdf.

Oviatt:1981:EDM

[Ovi81] C. A. Oviatt. Effects of different mixing schedules on phytoplankton. Zooplankton and nutrients in marine microcosms. Marine Ecology Progress Series, 4:57–67, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/4/m004p057.pdf.

Osterroht:1985:CDO

[OWKG85] C. Osterroht, A. Wenck, K. Kremling, and K. Gocke. Concentration of dissolved organic copper in relation to other chemical and biological parameters in coastal Baltic waters. Marine Ecology Progress Series, 22:273–279, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/22/m022p273.pdf.

Ovlatt:1980:EAM

[OWP80] C. A. Ovlatt, H. Walker, and M. E. Q. Pilson. An exploratory analysis of microcosm and ecosystem behavior using multivariate techniques. Marine Ecology Progress Series, 2:179–191, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/2/m002p179.pdf.

Osman:1989:ERS

[OWZ89] R. W. Osman, R. B. Whitlatch, and R. N. Zajac. Effects of resident species on recruitment into a community larval settlement versus post-settlement mortality in the oyster Crassostrea virginica. Marine Ecology Progress Series, 54:61–73, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/54/m054p061.pdf.

Peterson:1984:DSF

[PA84] W. T. Peterson and S. J. Ausubel. Diets and selective feeding by larvae of Atlantic mackerel Scomber scombrus on zooplankton. Marine Ecology Progress Series, 17:65–75, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/17/m017p065.pdf.

Pawlik:1986:EDR

[PAF86] J. R. Pawlik, K. F. Albizati, and D. J. Faulkner. Evidence of a defensive role for limatulone, a novel triterpene from the intertidal limpet Collisella limatula. Marine Ecology Progress Series, 30: 251–260, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/30/m030p251.pdf.

Page:1986:DPS

[Pag86] H. M. Page. Differences in population structure and growth rate of the stalked barnacle Pollicipes polymerus between a rocky headland and an offshore oil platform. Marine Ecology Progress Series, 29:157–164, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/29/m029p157.pdf.

Pillar:1989:VMD

[PAH89] S. C. Pillar, D. A. Armstrong, and L. Hutchings. Vertical migration, dispersal and transport of Euphausia lucens in the southern Benguela Current. Marine Ecology Progress Series, 53:179–190, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
53/m053p179.pdf.

Palmer:1988:DMM

[Pal88] M. A. Palmer. Dispersal of marine meiofauna. A review and conceptual model explaining passive transport and active emergence with implications for recruitment. Marine Ecology Progress Series, 48:81–91, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/48/m048p081.pdf.

Pedros-Alio:1989:MST

[PAN89] C. Pedrós-Alió and S. Y. Newell. Microautoradiography study of thymidine uptake in brackish waters around Sapelo Island, Georgia, USA. Marine Ecology Progress Series, 55:83–94, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/55/
m055p083.pdf.

Parry:1984:CPT

[Par84] D. L. Parry. Chemical properties of the test of ascidians in relation to predation. Marine Ecology Progress Series, 17:279–282, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
17/m017p279.pdf.

Paranjape:1987:GME

[Par87a] M. A. Paranjape. Grazing by microzooplankton in the eastern Canadian Arctic in summer 1983. Marine Ecology Progress Series, 40:239–246, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/40/m040p239.pdf.

Parrish:1987:TSP

[Par87b] C. C. Parrish. Time series of particulate and dissolved lipid classes during spring phytoplankton blooms in Bedford Basin, a marine inlet. Marine Ecology Progress Series, 35:129–139, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/35/m035p129.pdf.

Parrish:1989:FCI

[Par89] J. D. Parrish. Fish communities of interacting shallow-water habitats in tropical oceanic regions. Marine Ecology Progress Series, 58:143–160, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/58/m058p143.pdf.

Patterson:1988:LHP

[Pat88] K. R. Patterson. Life history of Patagonian squid Loligo gahi and growth parameter estimates using least-squares fits to linear and von Bertalanffy models. Marine Ecology Progress Series, 47:65– 74, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
47/m047p065.pdf.

Paul:1981:NDF

[Pau81] R. K. G. Paul. Natural diet, feeding and predatory activity of the crabs Callinectes arcuatus and C. toxotes (Decapoda, Brachyura, Portunidae). Marine Ecology Progress Series, 6:91– 99, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
6/m006p091.pdf.

Paul:1983:LTN

[Pau83] A. J. Paul. Light, temperature, nauplii concentrations, and prey capture by first feeding pollock larvae Theragra chalcogramma. Marine Ecology Progress Series, 13:175–179, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/13/m013p175.pdf.

Pawlik:1989:LSH

[Paw89] J. R. Pawlik. Larvae of the sea hare Aplysia californica settle and metamorphose on an assortment of macroalgal species. Marine Ecology Progress Series, 51:195–199, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/51/m051p195.pdf.

Palmer:1981:TVS

[PB81] M. A. Palmer and R. R. Brandt. Tidal variation in sediment densities of marine benthic copepods. Marine Ecology Progress Series, 4:207–212, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/4/m004p207.pdf.

Pastorok:1985:ESP

[PB85a] R. A. Pastorok and G. R. Bilyard. Effects of sewage pollution on coral-reef communities. Marine Ecology Progress Series, 21:175– 189, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
21/m021p175.pdf.

Pisano:1985:DPI

[PB85b] E. Pisano and M. Boyer. Development pattern of an infralittoral bryozoan community in the western Mediterranean Sea. Marine Ecology Progress Series, 27:195–202, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/27/m027p195.pdf.

Pace:1987:EFM

[PB87] Michael L. Pace and Megan D. Bailiff. Evaluation of a fluorescent microsphere technique for measuring grazing rates of phagotrophic microorganisms. Marine Ecology Progress Series, 40:185–193, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
40/m040p185.pdf.

Pangallo:1988:DAB

[PB88] R. A. Pangallo and S. S. Bell. Dynamics of the aboveground and belowground structure of the seagrass Halodule wrightii. Marine Ecology Progress Series, 43:297–301, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/43/m043p297.pdf.

Perry:1989:ERB

[PB89] D. M. Perry and R. C. Brusca. Effects of the root-boring isopod Sphaeroma peruvianum on red mangrove forests. Marine Ecology Progress Series, 57:287–292, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/57/m057p287.pdf.

Pissierssens:1985:NIE

[PBP85] P. Pissierssens, M. Bergmans, and Ph. Polk. On the need for independently estimating organic tripton carbon content. Marine Ecology Progress Series, 25:107–109, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/25/m025p107.pdf.

Price:1985:TCU

[PCH85] N. M. Price, W. P. Cochlan, and P. J. Harrison. Time course of uptake of inorganic and organic nitrogen by phytoplankton in the Strait of Georgia: comparison of frontal and stratified communities. Marine Ecology Progress Series, 27:39–53, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/27/m027p039.pdf.

Poulet:1986:CAS

[PCM86] S. A. Poulet, D. Cossa, and J.-C. Marty. Combined analyses of the size spectra and biochemical composition of particles in the St. Lawrence estuary. Marine Ecology Progress Series, 30:205– 214, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
30/m030p205.pdf.

Powell:1980:ASS

[PCR80] E. N. Powell, M. A. Crenshaw, and R. M. Rieger. Adaptations to sulfide in sulfide-system meiofauna. Endproducts of sulfide detoxification in three turbellarians and a gastrotrich. Marine Ecology Progress Series, 2:169–177, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/2/m002p169.pdf.

Potter:1986:CSC

[PCW86] I. C. Potter, P. N. Claridge, and R. M. Warwick. Consistency of seasonal changes in an estuarine fish assemblage. Marine Ecology Progress Series, 32:217–228, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/32/m032p217.pdf.

Phillips:1986:DIT

[PD86] D. J. H. Phillips and M. H. Depledge. Distribution of inorganic and total arsenic in tissues of the marine gastropod Hemifusus ternatanus. Marine Ecology Progress Series, 34:261–266, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
34/m034p261.pdf.

Pasternak:1988:PDC

[PD88] A. F. Pasternak and A. V. Drits. Possible degradation of chlorophyll-derived pigments during gut passage of herbivorous copepods. Marine Ecology Progress Series, 49:187–190, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/49/
m049p187.pdf.

Potts:1985:DCC

[PDIF85] D. C. Potts, T. J. Done, P. J. Isdale, and D. A. Fisk. Dominance of a coral community by the genus Porites (Scleractinia). Marine Ecology Progress Series, 23:79–84, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/23/m023p079.pdf.

Paul:1986:ELM

[PDJ86] J. H. Paul, M. F. DeFlaun, and W. H. Jeffrey. Elevated levels of microbial activity in the coral surface microlayer. Marine Ecology Progress Series, 33:29–40, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/33/m033p029.pdf.

Pezeshki:1987:EFS

[PDP87] S. R. Pezeshki, R. D. DeLaune, and W. H. Patrick, Jr. Effects of flooding and salinity on photosynthesis of Sagittaria lancifolia. Marine Ecology Progress Series, 41:87–91, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/41/m041p087.pdf.

Parslow:1986:FZP

[PDTH86] J. S. Parslow, G. J. Doucette, F. J. R. Taylor, and P. J. Harrison. Feeding by the zooflagellate Pseudobodo sp. on the picoplanktonic prasinomonad Micromonas pusilla. Marine Ecology Progress Series, 29:237–246, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/29/m029p237.pdf.

Post:1985:PRM

[PDWF85] A. F. Post, Z. Dubinsky, K. Wyman, and P. G. Falkowski. Physiological responses of a marine planktonic diatom to transitions in growth irradiance. Marine Ecology Progress Series, 25:141– 149, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
25/m025p141.pdf.

Pineda:1989:SED

[PE89] J. Pineda and A. Escofet. Selective effects of disturbance on populations of sea anemones from northern Baja California, Mexico. Marine Ecology Progress Series, 55:55–62, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/55/m055p055.pdf.

Pearre:1980:FCR

[Pea80] S. Pearre, Jr. Feeding by Chaetognatha: The relation of prey size to predator size in several species. Marine Ecology Progress Series, 3:125–134, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/3/m003p125.pdf.

Pearre:1981:FCE

[Pea81a] S. Pearre, Jr. Feeding by Chaetognatha: Energy balance and importance of various components of the diet of Sagitta elegans. Marine Ecology Progress Series, 5:45–54, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/5/m005p045.pdf.

Pearson:1981:RRC

[Pea81b] R. G. Pearson. Recovery and recolonization of coral reefs. Marine Ecology Progress Series, 4:105–122, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p105.pdf.

Pearre:1982:FCA

[Pea82] S. Pearre, Jr. Feeding by chatognatha: Aspects of inter- and intra-specific predation. Marine Ecology Progress Series, 7:33– 45, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
7/m007p033.pdf.

Pearre:1985:RBT

[Pea85] S. Pearre, Jr. Ratio-based trophic niche breadths of fish, the Sheldon spectrum, and the size-efficiency hypothesis. Marine Ecology Progress Series, 27:299–314, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/27/m027p299.pdf.

Petraitis:1984:LEE

[Pet84] P. S. Petraitis. Laboratory experiments on the effects of a gastropod (Hydrobia totteni) on survival of an infaunal depositfeeding polychaete (Capitella capitata). Marine Ecology Progress Series, 18:263–268, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/18/m018p263.pdf.

Peterson:1986:QAG

[Pet86a] C. H. Peterson. Quantitative allometry of gamete production by Mercenaria mercenaria into old age. Marine Ecology Progress Series, 29:93–97, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/29/m029p093.pdf.

Peterson:1986:DGS

[Pet86b] W. T. Peterson. Development, growth, and survivorship of the copepod Calanus marshallae in the laboratory. Marine Ecology Progress Series, 29:61–72, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/29/m029p061.pdf.

Peterson:1988:REP

[Pet88] W. T. Peterson. Rates of egg production by the copepod Calanus marshallae in the laboratory and in the sea off Oregon, USA. Marine Ecology Progress Series, 47:229–237, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/47/m047p229.pdf.

Pett:1989:KMM

[Pet89] R. J. Pett. Kinetics of microbial mineralization of organic carbon from detrital Skeletonema costatum cells. Marine Ecology Progress Series, 52:123–128, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/52/m052p123.pdf.

Pamatmat:1983:MMN

[PF83] M. M. Pamatmat and S. Findlay. Metabolism of microbes, nematodes, polychaetes, and their interactions in sediment, as detected by heat flow measurements. Marine Ecology Progress Series, 11: 31–38, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/11/m011p031.pdf.

Paul:1986:CDT

[PF86] V. J. Paul and W. Fenical. Chemical defense in tropical green algae, order caulerpales. Marine Ecology Progress Series, 34:157– 169, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
34/m034p157.pdf.

Pawlik:1989:REI

[PF89] J. R. Pawlik and W. Fenical. A re-evaluation of the ichthyodeterrent role of prostaglandins in the Caribbean gorgonian coral Plexaura homomalla. Marine Ecology Progress Series, 52:95–98, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
52/m052p095.pdf.

Poulet:1988:SAM

[PG88] S. A. Poulet and C. W. Gill. Spectral analyses of movements made by the cephalic appendages of copepods. Marine Ecology Progress Series, 43:259–267, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/43/m043p259.pdf.

Pamatmat:1981:HPA

[PGBN81] M. M. Pamatmat, G. Graf, W. Bengtsson, and C. S. Novak. Heat production. ATP concentration and electron transport activity of marine sediments. Marine Ecology Progress Series, 4:135–143, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
4/m004p135.pdf.

Pearson:1983:OSS

[PGJ83] T. H. Pearson, J. S. Gray, and P. J. Johannessen. Objective selection of sensitive species indicative of pollution-induced change in benthic communities. 2. Data analyses. Marine Ecology Progress Series, 12:237–255, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/12/m012p237.pdf.

Prezelin:1989:BLE

[PGV^{+}89] B. B. Prézelin, H. E. Glover, B. Ver Hoven, D. Steinberg, H. A. Matlick, O. Schofield, N. Nelson, M. Wyman, and L. Campbell. Blue light effects on light-limited rates of photosynthesis: relationship to pigmentation and productivity estimates for Synechococcus populations from the Sargasso Sea. Marine Ecology Progress Series, 54:121–136, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/54/m054p121.pdf.

Pentreath:1981:PIS

[PH81] R. J. Pentreath and B. R. Harvey. The presence of ^{237}Np in the Irish Sea. Marine Ecology Progress Series, 6:243–247, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
6/m006p243.pdf.

Pocklington:1983:NRV

[PH83] P. Pocklington and M. S. Hutcheson. New record of viviparity for the dominant benthic invertebrate Exogone hebes (Polychaeta: Syllidae) from the Grand Banks of Newfoundland. Marine Ecology Progress Series, 11:239–244, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/11/m011p239.pdf.

Paul:1986:SSH

[PH86] V. J. Paul and M. E. Hay. Seaweed susceptibility to herbivory: chemical and morphological correlates. Marine Ecology Progress Series, 33:255–264, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/33/m033p255.pdf.

Pearse:1987:LTP

[PH87] J. S. Pearse and A. H. Hines. Long-term population dynamics of sea urchins in a central California kelp forest: rare recruitment and rapid decline. Marine Ecology Progress Series, 39:275–283, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
39/m039p275.pdf.

Phillips:1985:OTM

[Phi85] D. J. H. Phillips. Organochlorines and trace metals in greenlipped mussels Perna viridis from Hong Kong waters: a test of indicator ability. Marine Ecology Progress Series, 21:251–258, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
21/m021p251.pdf.

Price:1986:TEL

[PHL^{+}86] N. M. Price, P. J. Harrison, M. R. Landry, F. Azam, and K. J. F. Hall. Toxic effects of latex and tygon tubing on marine phytoplankton, zooplankton and bacteria. Marine Ecology Progress Series, 34:41–49, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/34/m034p041.pdf.

Piatt:1989:BPO

[PHL^{+}89] T. Piatt, W. G. Harrison, M. R. Lewis, W. K. W. Li, S. Sathyendranath, R. E. Smith, and A. F. Vezina. Biological production of the oceans: the case for a consensus. Marine Ecology Progress Series, 52:77–88, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/52/m052p077.pdf.

Parslow:1985:AUP

[PHT85] J. S. Parslow, P. J. Harrison, and P. A. Thompson. Ammonium uptake by phytoplankton cells on a filter: a new high-resolution technique. Marine Ecology Progress Series, 25:121–129, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/25/m025p121.pdf.

Piatkowski:1989:MCA

[Pia89] U. Piatkowski. Macroplankton communities in Antarctic surface waters: spatial changes related to hydrography. Marine Ecology Progress Series, 55:251–259, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/55/m055p251.pdf.

Pienkowski:1983:SAS

[Pie83] M. W. Pienkowski. Surface activity of some intertidal invertebrates in relation to temperature and the foraging behaviour of their shorebird predators. Marine Ecology Progress Series, 11:141– 150, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
11/m011p141.pdf.

Piersma:1987:PIB

[Pie87] T. Piersma. Production by intertidal benthic animals and limits to their predation by shorebirds: a heuristic model. Marine Ecology Progress Series, 38:187–196, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/38/m038p187.pdf.

Pihl:1985:FSC

[Pih85] L. Pihl. Food selection and consumption of mobile epibenthic fauna in shallow marine areas. Marine Ecology Progress Series, 22: 169–179, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/22/m022p169.pdf.

Palmisano:1985:SIM

[PKMS85] A. C. Palmisano, S. T. Kottmeier, R. L. Moe, and C. W. Sullivan. Sea ice microbial communities. IV. The effect of light perturbation on microalgae at the ice–seawater interface in McMurdo Sound, Antarctica. Marine Ecology Progress Series, 21:37–45, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
21/m021p037.pdf.

Piatt:1985:NMP

[PL85] H. M. Piatt and P. J. D. Lambshead. Neutral modelanalysis of patterns of marine benthic species diversity. Marine Ecology Progress Series, 24:75–81, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/24/m024p075.pdf.

Paffenhofer:1989:FBN

[PL89] G.-A. Paffenhöfer and K. D. Lewis. Feeding behavior of nauplii of the genus Eucalanus (Copepoda, Calanoida). Marine Ecology Progress Series, 57:129–136, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/57/m057p129.pdf.

Painting:1989:FHB

[PLM89] S. J. Painting, M. I. Lucas, and D. G. Muir. Fluctuations in heterotrophic bacterial community structure, activity and production in response to development and decay of phytoplankton in a microcosm. Marine Ecology Progress Series, 53:129–141, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/53/m053p129.pdf.

Painchaud:1987:BMA

[PLT87] J. Painchaud, D. Lefaivre, and J.-C. Therriault. Box model analysis of bacterial fluxes in the St. Lawrence Estuary. Marine Ecology Progress Series, 41:241–252, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/41/m041p241.pdf.

Pregnall:1988:FAS

[PM88] A. M. Pregnall and S. L. Miller. Flux of ammonium from surf-zone and nearshore sediments in Nahant Bay, Massachusetts, USA, in relation to free-living Pilayella littoralis. Marine Ecology Progress Series, 50:161–167, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/50/m050p161.pdf.

Poulet:1984:DDF

[PMJH84] S. A. Poulet, V. Martin-Jézéquel, and R. N. Head. Distribution of dissolved free amino acids in the Ushant front region. Marine Ecology Progress Series, 18:49–55, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/18/m018p049.pdf.

Perry:1988:VDT

[PN88] R. I. Perry and J. D. Neilson. Vertical distributions and trophic interactions of age-0 Atlantic cod and haddock in mixed and stratified waters of Georges Bank. Marine Ecology Progress Series, 49: 199–214, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/49/m049p199.pdf.

Pollock:1984:RBM

[Pol84] B. R. Pollock. Relations between migration, reproduction and nutrition in yellowfin bream Acanthopagrus australis. Marine Ecology Progress Series, 19:17–23, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/19/m019p017.pdf.

Posey:1986:CBC

[Pos86] M. H. Posey. Changes in a benthic community associated with dense beds of a burrowing deposit feeder, Callianassa californiensis. Marine Ecology Progress Series, 31:15–22, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/31/m031p015.pdf.

Posey:1987:IRM

[Pos87] M. H. Posey. Influence of relative mobilities on the composition of benthic communities. Marine Ecology Progress Series, 39:99– 104, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
39/m039p099.pdf.

Plante:1986:PPS

[PPCR86] R. Plante, M.-R. Plante-Cuny, and J.-P. Reys. Photosynthetic pigments of sandy sediments on the north Mediterranean coast: their spatial distribution and its effect on sampling strategies. Marine Ecology Progress Series, 34:133–141, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/34/m034p133.pdf.

Pihl:1984:FSC

[PR84] L. Pihl and R. Rosenberg. Food selection and consumption of the shrimp Crangon crangon in some shallow marine areas in western Sweden. Marine Ecology Progress Series, 15:159–168, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
15/m015p159.pdf.

Phillips:1988:BMB

[PR88] D. J. H. Phillips and P. S. Rainbow. Barnacles and mussels as biomonitors of trace elements: a comparative study. Marine Ecology Progress Series, 49:83–93, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/49/m049p083.pdf.

Probyn:1985:NUS

[Pro85] T. A. Probyn. Nitrogen uptake by size-fractionated phytoplankton populations in the southern Benguela upwelling system. Marine Ecology Progress Series, 22:249–258, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/22/m022p249.pdf.

Probyn:1987:ARM

[Pro87] T. A. Probyn. Ammonium regeneration by microplankton in an upwelling environment. Marine Ecology Progress Series, 37:53– 64, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
37/m037p053.pdf.

Platt:1982:PCP

[PRS^{+}82] T. Platt, D. V. Subba Rao, J. C. Smith, W. K. Li, B. Irwin, E. P. W. Horne, and D. D. Sameoto. Photosyntheticallycompetent phytoplankton from the aphotic zone of the deep ocean. Marine Ecology Progress Series, 10:105–110, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/10/m010p105.pdf.

Pennock:1986:PPD

[PS86] J. R. Pennock and J. H. Sharp. Phytoplankton production in the Delaware Estuary: temporal and spatial variability. Marine Ecology Progress Series, 34:143–155, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/34/m034p143.pdf.

Paffenhofer:1988:WAT

[PS88] G.-A. Paffenhöfer and D. E. Stearns. Why is Acartia tonsa (Copepoda: Calanoida) restricted to nearshore environments? Marine Ecology Progress Series, 42:33–38, 1988. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/42/m042p033.pdf.

Parsons:1981:BFS

[PSB^{+}81] T. R. Parsons, J. Stronach, G. A. Borstad, G. Louttit, and R. I. Perry. Biological fronts in the Strait of Georgia, British Columbia, and their relation to recent measurements of primary productivity. Marine Ecology Progress Series, 6:237–242, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/6/m006p237.pdf.

Palmisano:1987:SIM

[PSMS87] A. C. Palmisano, J. B. SooHoo, R. L. Moe, and C. W. Sullivan. Sea ice microbial communities. VII. Changes in under-ice spectral irradiance during the development of Antarctic sea ice microalgal communities. Marine Ecology Progress Series, 35:165– 173, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
35/m035p165.pdf.

Pengerud:1987:RIB

[PST87] B. Pengerud, E. F. Skjoldal, and T. F. Thingstad. The reciprocal interaction between degradation of glucose and ecosystem structure. Studies in mixed chemostat cultures of marine bacteria, algae, and bacterivorous nanoflagellates. Marine Ecology Progress Series, 35:111–117, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/35/m035p111.pdf.

Phillips:1984:EFP

[PT84] N. W. Phillips and K. R. Tenore. Effects of food-particle size and peptization on individual growth and larval settlement of the deposit feeding polychaete Capitella capitata type I. Marine Ecology Progress Series, 16:241–247, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/16/m016p241.pdf.

Painchaud:1989:RBB

[PT89] J. Painchaud and J.-C. Therriault. Relationships between bacteria, phytoplankton and particulate organic carbon in the Upper St. Lawrence Estuary. Marine Ecology Progress Series, 56:301– 311, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
56/m056p301.pdf.

Purcell:1984:PFL

[Pur84] J. E. Purcell. Predation on fish larvae by Physalia physalis, the Portuguese man of war. Marine Ecology Progress Series, 19:189– 191, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
19/m019p189.pdf.

Purcell:1988:QML

[Pur88] J. E. Purcell. Quantification of Mnemiopsis leidyi (Ctenophora, Lobata) from formalin-preserved plankton samples. Marine Ecology Progress Series, 45:197–200, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/45/m045p197.pdf.

Paffenhofer:1985:FRM

[PV85] G.-A. Paffenhöfer and K. B. Van Sant. The feeding response of a marine planktonic copepod to quantity and quality of particles. Marine Ecology Progress Series, 27:55–65, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/27/m027p055.pdf.

Parrish:1987:PDL

[PW87] Christopher C. Parrish and Peter J. Wangersky. Particulate and dissolved lipid classes in cultures of Phaeodactylum tricornutum grown in cage culture turbidostats with a range of nitrogen supply rates. Marine Ecology Progress Series, 35:119–128, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/35/m035p119.pdf.

Phelps:1985:FAT

[PWM85] H. L. Phelps, D. A. Wright, and J. A. Mihursky. Factors affecting trace metal accumulation by estuarine oysters Crassostrea virginica. Marine Ecology Progress Series, 22:187–197, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/22/
m022p187.pdf.

Pitcher:1989:PSR

[PWMl89] G. C. Pitcher, D. R. Walker, and B. A. Mitchell-lnnes. Phytoplankton sinking rate dynamics in the southern Benguela upwelling system. Marine Ecology Progress Series, 55:261–269, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/55/
m055p261.pdf.

Phillips:1981:CEO

[PY81] D. J. H. Phillips and W. W.-S. Yim. A comparative evaluation of oysters, mussels and sediments as indicators of trace metals in Hong Kong waters. Marine Ecology Progress Series, 6:285–293, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
6/m006p285.pdf.

Piatt:1986:PMM

[PY86] T. Piatt and N. Yamamura. Prenatal mortality in a marine cladoceran, Evadne nordmanni. Marine Ecology Progress Series, 29: 127–139, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/29/m029p127.pdf.

Quinn:1988:ECA

[Qui88] G. P. Quinn. Effects of conspecific adults, macro-algae and height on the shore on recruitment of an intertidal limpet. Marine Ecology Progress Series, 48:305–308, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/48/m048p305.pdf.

Rosso:1987:PAC

[RA87] Abbey L. Rosso and Farooq Azam. Proteolytic activity in coastal oceanic waters: depth distribution and relationship to bacterial populations. Marine Ecology Progress Series, 41:231–240, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/41/
m041p231.pdf.

Russ:1989:EIF

[RA89] G. R. Russ and A. C. Alcala. Effects of intense fishing pressure on an assemblage of coral reef fishes. Marine Ecology Progress Series, 56:13–27, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/56/m056p013.pdf.

Reichelt:1984:SPC

[RB84] R. E. Reichelt and R. H. Bradbury. Spatial patterns in coral reef benthos: multiscale analysis of sites from three oceans. Marine Ecology Progress Series, 17:251–257, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/17/m017p251.pdf.

Romeyn:1983:ECH

[RBA83] K. Romeyn, L. A. Bouwman, and W. Admiraal. Ecology and cultivation of the herbivorous brackish-water nematode Eudiplogaster pararmatus. Marine Ecology Progress Series, 12:145–153, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
12/m012p145.pdf.

Rebhun:1986:ENC

[RBA86] S. Rebhun and A. Ben-Amotz. Effect of NaCI concentration on cadmium uptake by the halophilic alga Dunaliella salina. Marine Ecology Progress Series, 30:215–219, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/30/m030p215.pdf.

Rebhun:1988:AEM

[RBA88] S. Rebhun and A. Ben-Amotz. Antagonistic effect of manganese to cadmium toxicity in the alga Dunaliella salina. Marine Ecology Progress Series, 42:97–104, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/42/m042p097.pdf.

Rice:1986:ESS

[RBR86] D. L. Rice, T. S. Bianchi, and E. H. Roper. Experimental studies of sediment reworking and growth of Scoloplos spp. (Orbiniidae: Polychaeta). Marine Ecology Progress Series, 30:9–19, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/30/
m030p009.pdf.

Roberts:1986:LGD

[RC86] D. G. Roberts and J. Caperon. Lacunar gas discharge as a measure of photosynthesis in seagrasses. Marine Ecology Progress Series, 29:23–27, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/29/m029p023.pdf.

Roman:1989:OCF

[RD89] C. T. Roman and F. C. Daiber. Organic carbon flux through a Delaware Bay salt marsh: tidal exchange, particle size distribution, and storms. Marine Ecology Progress Series, 54:149–156, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
54/m054p149.pdf.

Robertson:1988:ZDM

[RDD88] A. I. Robertson, P. Dixon, and P. A. Daniel. Zooplankton dynamics in mangrove and other nearshore habitats in tropical Australia. Marine Ecology Progress Series, 43:139–150, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/43/m043p139.pdf.

Roman:1988:PCN

[RDF^{+}88] M. R. Roman, H. W. Ducklow, J. A. Fuhrman, C. Garside, P. M. Glibert, T. C. Malone, and G. B. McManus. Production, consumption and nutrient cycling in a laboratory mesocosm. Marine Ecology Progress Series, 42:39–52, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/42/m042p039.pdf.

Redalje:1983:PCB

[Red83] D. G. Redalje. Phytoplankton carbon biomass and specific growth rates determined with the labeled chlorophyll-a technique. Marine Ecology Progress Series, 11:217–225, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/11/m011p217.pdf.

Regnault:1984:SIC

[Reg84] M. Regnault. Salinity-induced changes in ammonia excretion rate of the shrimp Crangon crangon over a winter tidal cycle. Marine Ecology Progress Series, 20:119–125, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/20/m020p119.pdf.

Reise:1981:GAA

[Rei81a] K. Reise. Gnathostomulida abundant alongside polychaete burrows. Marine Ecology Progress Series, 6:329–333, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/6/m006p329.pdf.

Reiswig:1981:POC

[Rei81b] H. M. Reiswig. Particulate organic carbon of bottom boundary and submarine cavern waters of tropical coral reefs. Marine Ecology Progress Series, 5:129–133, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/5/m005p129.pdf.

Reise:1983:BEI

[Rei83] K. Reise. Biotic enrichment of intertidal sediments by experimental aggregates of the deposit-feeding bivalve Macoma balthica. Marine Ecology Progress Series, 12:229–236, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/12/m012p229.pdf.

Reichardt:1987:DTE

[Rei87a] W. Reichardt. Differential temperature effects on the efficiency of carbon pathways in Antarctic marine benthos. Marine Ecology Progress Series, 40:127–135, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/40/m040p127.pdf.

Reise:1987:SNL

[Rei87b] K. Reise. Spatial niches and long-term performance in meiobenthic plathelminthes of an intertidal lugworm flat. Marine Ecology Progress Series, 38:1–11, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/38/m038p001.pdf.

Reichardt:1988:IBA

[Rei88] W. Reichardt. impact of bioturbation by Arenicola marina on microbiological parameters in intertidal sediments. Marine Ecology Progress Series, 44:149–158, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/44/m044p149.pdf.

Rhoads:1982:COS

[RG82] Donald C. Rhoads and Joseph D. Germano. Characterization of organism-sediment relations using sediment profile imaging: an efficient method of remote ecological monitoring of the seafloor (Remots^{T M} System). Marine Ecology Progress Series, 8:115–128, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
8/m008p115.pdf.

Riaux-Gobin:1985:LTC

[RG85] C. Riaux-Gobin. Long-term changes in microphytobenthos in a Brittany estuary after the ‘Amoco Cadiz’ oil spill. Marine Ecology Progress Series, 24:51–56, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/24/m024p051.pdf.

Rublee:1989:UFL

[RG89] P. A. Rublee and C. L. Gallegos. Use of fluorescently labelled algae (FLA) to estimate microzooplankton grazing. Marine Ecology Progress Series, 51:221–227, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/51/m051p221.pdf.

Rheinheimer:1989:VDM

[RGH89] G. Rheinheimer, K. Gocke, and H.-G. Hoppe. Vertical distribution of microbiological and hydrographic-chemical parameters in different areas of the Baltic Sea. Marine Ecology Progress Series, 52:55– 70, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
52/m052p055.pdf.

Riaux-Gobin:1987:MTS

[RGLK87] C. Riaux-Gobin, C. A. Llewellyn, and B. Klein. Microphytobenthos from two subtidal sediments from North Brittany. II. Variations of pigment compositions and concentrations determinded by HPLC and conventional techniques. Marine Ecology Progress Series, 40:275–283, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/40/m040p275.pdf.

Roy:1989:IFC

[RHP89] S. Roy, R. P. Harris, and S. A. Poulet. Inefficient feeding by Calanus helgolandicus and Temora longicornis on Coscinodiscus wailesii: quantitative estimation using chlorophyll-type pigments and effects on dissolved free amino acids. Marine Ecology Progress Series, 52:145–153, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/52/m052p145.pdf.

Ray:1989:APD

[RHS89] R. T. Ray, L. W. Haas, and M. E. Sieracki. Autotrophic picoplankton dynamics in a Chesapeake Bay sub-estuary. Marine Ecology Progress Series, 52:273–285, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/52/m052p273.pdf.

Rietema:1981:LHD

[RHvd81] H. Rietema, C. Hoek, and van den. The life history of Desmotrichum undulatum (Phaeophyceae) and its regulation by temperature and light conditions. Marine Ecology Progress Series, 4:321–335, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/4/m004p321.pdf.

Rice:1982:DNP

[Ric82] D. L. Rice. The detritus nitrogen problem: New observations and perspectives from organic geochemistry. Marine Ecology Progress Series, 9:153–162, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/9/m009p153.pdf.

Richardson:1985:PDA

[Ric85a] K. Richardson. Plankton distribution and activity in the North Sea/Skagerrak–Kattegat frontal area in April 1984. Marine Ecology Progress Series, 26:233–244, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/26/m026p233.pdf.

Richmond:1985:RMC

[Ric85b] R. H. Richmond. Reversible metamorphosis in coral planula larvae. Marine Ecology Progress Series, 22:181–185, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/22/m022p181.pdf.

Riddle:1988:PDM

[Rid88] M. J. Riddle. Patterns in the distribution of macrofaunal communities in coral reef sediments on the central Great Barrier Reef. Marine Ecology Progress Series, 47:281–292, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/47/m047p281.pdf.

Rieper:1982:FPM

[Rie82a] M. Rieper. Feeding preferences of marine harpacticoid copepods for various species of bacteria. Marine Ecology Progress Series, 7:

303–307, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/7/m007p303.pdf.

Rietema:1982:EPT

[Rie82b] H. Rietema. Effects of photoperiod and temperature on macrothallus initiation in Dumontia contorta (Rhodophyta). Marine Ecology Progress Series, 8:187–196, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/8/m008p187.pdf.

Riebesell:1989:CSS

[Rie89] U. Riebesell. Comparison of sinking and sedimentation rate measurements in a diatom winter/spring bloom. Marine Ecology Progress Series, 54:109–119, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/54/m054p109.pdf.

Riisgaard:1988:APP

[Rii88a] H. U. Riisgård. The ascidian pump: properties and energy cost. Marine Ecology Progress Series, 47:129–134, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/47/m047p129.pdf.

Riisgaard:1988:EPR

[Rii88b] H. U. Riisgård. Efficiency of particle retention and filtration rate in 6 species of Northeast American bivalves. Marine Ecology Progress Series, 45:217–223, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/45/m045p217.pdf.

Riisgaard:1989:PEC

[Rii89] H. U. Riisgård. Properties and energy cost of the muscular piston pump in the suspension feeding polychaete Chaetopterus variopedatus. Marine Ecology Progress Series, 56:157–168, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/56/m056p157.pdf.

Rivkin:1989:IIS

[Riv89] R. B. Rivkin. Influence of irradiance and spectral quality on the carbon metabolism of phytoplankton I. Photosynthesis, chemical composition and growth. Marine Ecology Progress Series, 55:291– 304, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
55/m055p291.pdf.

Rietema:1981:ECL

[RK81] H. Rietema and A. W. O. Klein. Environmental control of the life cycle of Dumontia contorta (Rhodophyta) kept in culture. Marine Ecology Progress Series, 4:23–29, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p023.pdf.

Robertson:1982:FHS

[RK82] A. I. Robertson and D. W. Klumpp. Feeding habitats of the Southern Australian garfish Hyporhamphus melanochir: a diurnal herbivore and nocturnal carnivore. Marine Ecology Progress Series, 10:197–201, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/10/m010p197.pdf.

Regnault:1983:EAN

[RL83] Nl Regnault and J.-P. Lagardere. Effects of ambient noise on the metabolic level of Crangon crangon (Decapoda, Natantia). Marine Ecology Progress Series, 11:71–78, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/11/m011p071.pdf.

Robinson:1989:CPL

[RLB89] W. E. Robinson, R. W. Langton, and C. C. Boggs. Chlorophyllous pigment and lipid stores in the digestive glands of inshore and offshore populations of the deep-sea scallop Placopecten magellanicus. Marine Ecology Progress Series, 52:181–191, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/52/m052p181.pdf.

Ryan:1988:ETB

[RM88] P. G. Ryan and C. L. Moloney. Effect of trawling on bird and seal distributions in the southern Benguela region. Marine Ecology Progress Series, 45:1–11, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/45/m045p001.pdf.

Rozas:1988:IRC

[RMO88] Lawrence P. Rozas, Carole C. Mclvor, and William E. Odum. Intertidal rivulets and creekbanks: corridors between tidal creeks and marshes. Marine Ecology Progress Series, 47:303–307, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/47/
m047p303.pdf.

Robertson:1982:MSD

[RMZ82] Mark L. Robertson, Aaron L. Mills, and Joseph C. Zieman. Microbial synthesis of detritus-like particulates from dissolved organic carbon released by tropical seagrasses. Marine Ecology Progress Series, 7:279–285, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/7/m007p279.pdf.

Reid:1989:TSC

[RN89] D. G. Reid and E. Naylor. Are there separate circatidal and circadian clocks in the shore crab Carcinus maenas. Marine Ecology Progress Series, 52:1–6, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/52/m052p001.pdf.

Riemann:1988:RPB

[RNH^{+}88] B. Riemann, T. G. Nielsen, S. J. Horsted, P. K. Bjomsen, and J. Pock-Steen. Regulation of phytoplankton biomass in estuarine enclosures. Marine Ecology Progress Series, 48:205–215, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
48/m048p205.pdf.

Riemann:1984:DCB

[RNJ^{+}84] B. Riemann, P. Nielsen, M. Jeppesen, B. Marcussen, and J. A. Fuhrman. Diel changes in bacterial biomass and growth rates in coastal environments, determined by means of thymidine incorporation into DNA, frequency of dividing cells (FDC), and microautoradiography. Marine Ecology Progress Series, 17:227–235, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
17/m017p227.pdf.

Rodrigues:1987:MPS

[RNK87] C. L. Rodrigues, S. Nojima, and T. Kikuchi. Mechanics of prey size preference in the gastropod Neverita didyma preying on the bivalve Ruditapes philippinarum. Marine Ecology Progress Series, 40:87–93, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/40/m040p087.pdf.

Roden:1984:CMA

[RO84] C. M. Roden and K. W. O’Mahony. Competition as a mechanism of adaptation to environmental stress in outdoor cultures of marine diatoms. Marine Ecology Progress Series, 16:219–227, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
16/m016p219.pdf.

Roberts:1987:HCC

[RO87a] C. M. Roberts and R. F. G. Ormond. Habitat complexity and coral reef fish diversity and abundance on Red Sea fringing reefs. Marine Ecology Progress Series, 41:1–8, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/41/m041p001.pdf.

Rozas:1987:FMU

[RO87b] L. P. Rozas and W. E. Odum. Fish and macrocrustacean use of submerged plant beds in tidal freshwater marsh creeks. Marine Ecology Progress Series, 38:101–108, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/38/m038p101.pdf.

Robertson:1982:FFF

[Rob82] D. R. Robertson. Fish feces as fish food on a Pacific coral reef. Marine Ecology Progress Series, 7:253–265, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/7/m007p253.pdf.

Roy:1989:HMC

[Roy89] S. Roy. HPLC-measured chlorophyll-type pigments during a phytoplankton spring bloom in Bedford Basin (Canada). Marine Ecology Progress Series, 55:279–290, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/55/m055p279.pdf.

Rosenberg:1982:EGS

[RR82] G. Rosenberg and J. Ramus. Ecological growth strategies in the seaweeds Gracilaria foliifera (Rhodophyceae) and Ulva sp.

(Chlorophyceae): Photosynthesis and antenna composition. Marine Ecology Progress Series, 8:233–241, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/8/m008p233.pdf.

Rodhouse:1987:CBC

[RR87] P. G. Rodhouse and C. M. Roden. Carbon budget for a coastal inlet in relation to intensive cultivation of suspension-feeding bivalve mollusks. Marine Ecology Progress Series, 36:225–236, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/36/
m036p225.pdf.

Roosenburg:1980:ECP

[RRB^{+}80] W. H. Roosenburg, J. C. Rhoderick, R. M. Block, V. S. Kennedy, S. R. Gullans, S. M. Vreenegoor, A. Rosenkranz, and C. Collette. Effects of chlorine-produced oxidants on survival of larvae of the oyster Crossostrea virginica. Marine Ecology Progress Series, 3: 93–96, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/3/m003p093.pdf.

Ryther:1980:EEZ

[RS80] J. H. Ryther and J. G. Sanders. Experimental evidence of zooplankton control of the species composition and size distribution of marine phytoplankton. Marine Ecology Progress Series, 3:279– 283, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
3/m003p279.pdf.

Rey:1987:CSA

[RS87] F. Rey and H. R. Skjoldal. Consumption of silicic acid below the euphotic zone by sedimenting diatom blooms in the Barents Sea. Marine Ecology Progress Series, 36:307–312, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/36/m036p307.pdf.

Rodhouse:1988:LCD

[RSM88] P. G. Rodhouse, R. C. Swinfen, and A. W. A. Murray. Life cycle, demography and reproductive investment in the myopsid squid Alloteuthis subulata. Marine Ecology Progress Series, 45:245–253, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
45/m045p245.pdf.

Rainbow:1980:ECA

[RSWJ80] P. S. Rainbow, A. G. Scott, E. A. Wiggins, and R. W. Jackson. Effect of chelating agents on the accumulation of cadmium by the barnacle Semihlanus balanoides, and complexation of soluble Cd, Zn and Cu. Marine Ecology Progress Series, 2:143–152, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
2/m002p143.pdf.

Ruggeri:1985:PAF

[RT85] B. Ruggeri and C. A. Thoroughgood. Prostaglandins in aquatic fauna: a comprehensive review. Marine Ecology Progress Series, 23:301–306, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/23/m023p301.pdf.

Radtke:1989:AFG

[RTK^{+}89] R. L. Radtke, T. E. Targett, A. Kellermann, J. L. Bell, and K. T. Hill. Antarctic fish growth: profile of Trematomus newnesi. Marine Ecology Progress Series, 57:103–117, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/57/m057p103.pdf.

Raimbault:1988:VSD

[RTLR88] P. Raimbault, I. Taupier-Letage, and M. Rodier. Vertical size distribution of phytoplankton in the western Mediterranean Sea during early summer. Marine Ecology Progress Series, 45:153– 158, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
45/m045p153.pdf.

Rice:1981:EDR

[RTW81] D. L. Rice, K. R. Tenore, and H. L. Windom. The effect of detritus ration on metal transfer to deposit-feeding benthos. Marine Ecology Progress Series, 5:135–140, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/5/m005p135.pdf.

Rubin:1985:MAC

[Rub85] J. A. Rubin. Mortality and avoidance of competitive overgrowth in encrusting Bryozoa. Marine Ecology Progress Series, 23:291– 299, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
23/m023p291.pdf.

Rudnick:1988:TLB

[Rud88] D. T. Rudnick. Time lags between the deposition and meiobenthic assimilation of phytodetritus. Marine Ecology Progress Series, 50: 231–240, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/50/m050p231.pdf.

Ruger:1989:BSN

[Rüg89] H.-J. Rüger. Benthic studies of the northwest African upwelling region: psychrophilic and psychrotrophic bacterial communities from areas with different upwelling intensities. Marine Ecology Progress Series, 57:45–52, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/57/m057p045.pdf.

Rumrill:1989:PSS

[Rum89] S. S. Rumrill. Population size-structure, juvenile growth, and breeding periodicity of the sea star Asterina miniata in Barkley Sound, British Columbia. Marine Ecology Progress Series, 56:37– 47, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
56/m056p037.pdf.

Russell:1983:FEE

[Rus83a] G. Russell. Formation of an ectocarpoid epiflora on blades of Laminaria digitata. Marine Ecology Progress Series, 11:181–187, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
11/m011p181.pdf.

Russell:1983:PGP

[Rus83b] G. Russell. Parallel growth patterns in algal epiphytes and Laminaria blades. Marine Ecology Progress Series, 13:303–304, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/13/
m013p303.pdf.

Russ:1984:DAHa

[Rus84a] G. Russ. Distribution and abundance of herbivorous grazing fishes in the central Great Barrier Reef. I. Levels of variability across the entire continental shelf. Marine Ecology Progress Series, 20:23– 34, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
20/m020p023.pdf.

Russ:1984:DAHb

[Rus84b] G. Russ. Distribution and abundance of herbivorous grazing fishes in the central Great Barrier Reef. II. Patterns of zonation of midshelf and outershelf reefs. Marine Ecology Progress Series, 20:35– 44, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
20/m020p035.pdf.

Russo:1987:RHC

[Rus87] A. R. Russo. Role of habitat complexity in mediating predation by the gray damselfish Abudefduf sordidus on epiphytal amphipods. Marine Ecology Progress Series, 36:101–105, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/36/m036p101.pdf.

Russell:1984:ESS

[RV84] G. Russell and C. J. Veltkamp. Epiphyte survival on skin-shedding macrophytes. Marine Ecology Progress Series, 18:149–153, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/18/
m018p149.pdf.

Reed:1986:RMP

[RW86] R. H. Reed and P. J. Wright. Release of mannitol from Pilayella littoralis (Phaeophyta. Ectocarpales) in response to hypoosmotic stress. Marine Ecology Progress Series, 29:205–208, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/29/m029p205.pdf.

Rabsch:1983:ILC

[RWK83] U. Rabsch, K. Wolter, and P. Krischker. Influence of low cadmium and zinc concentrations on batch culture Prorocentrum micans (Dinophyta) containing low levels of dissolved organic carbon. Marine Ecology Progress Series, 14:275–285, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/14/m014p275.pdf.

Roesijadi:1984:BTM

[RYD84] G. Roesijadi, J. St Young, and J. M. Drum, A. S, Gurtisen. Behavior of trace metals in Mytiius edulis during a reciprocal transplant field experiment. Marine Ecology Progress Series, 18:155– 170, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
18/m018p155.pdf.

Ryer:1988:PFE

[Rye88] C. H. Ryer. Pipefish foraging: effects of fish size, prey size and altered habitat complexity. Marine Ecology Progress Series, 48:37– 45, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
48/m048p037.pdf.

Rygg:1985:ESC

[Ryg85] B. Rygg. Effect of sediment copper on benthic fauna. Marine Ecology Progress Series, 25:83–89, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/25/m025p083.pdf.

Rylaarsdam:1983:LHA

[Ryl83] K. W. Rylaarsdam. Life histories and abundance patterns of colonial corals on Jamaican reefs. Marine Ecology Progress Series, 13: 249–260, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/13/m013p249.pdf.

Ruzzante:1985:SCT

[RZ85] D. E. Ruzzante and H. E. Zaixso. Settlement of Chlamys tehuelchus (D’Orb.) on artificial collectors. Seasonal changes in spat settlement. Marine Ecology Progress Series, 26:195–197, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
26/m026p195.pdf.

Spotte:1981:PRC

[SA81] S. Spotte and G. Adams. Pathogen reduction in closed aquaculture systems by UV radiation: Fact or artifact? Marine Ecology Progress Series, 6:295–298, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/6/m006p295.pdf.

Spotte:1982:EAU

[SA82] S. Spotte and G. Adams. Estimation of the allowable upper limit of ammonia in saline waters. Marine Ecology Progress Series, 10: 207–210, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/10/m010p207.pdf.

Simon:1989:PCP

[SA89] M. Simon and F. Azam. Protein content and protein synthesis rates of planktonic marine bacteria. Marine Ecology Progress Series, 51:201–213, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/51/m051p201.pdf.

Smith:1988:API

[SACC88] R. E. H. Smith, J. Anning, P. Clement, and G. Cota. Abundance and production of ice algae in Resolute Passage, Canadian Arctic. Marine Ecology Progress Series, 48:251–263, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/48/m048p251.pdf.

Sammarco:1982:PBE

[Sam82] P. W. Sammarco. Polyp bail-out: an escape response to environmental stress and a new means of reproduction in corals. Marine Ecology Progress Series, 10:57–65, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/10/m010p057.pdf.

Sammarco:1983:EFG

[Sam83] P. W. Sammarco. Effects of fish grazing and damselfish territoriality on coral reef algae. I. Algal community structure. Marine Ecology Progress Series, 13:1–14, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/13/m013p001.pdf.

Sameoto:1988:FLF

[Sam88] D. D. Sameoto. Feeding of lantern fish Benthosema glaciale off the Nova Scotia Shelf. Marine Ecology Progress Series, 44:113– 129, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
44/m044p113.pdf.

Sibbald:1987:CRH

[SAS87] M. J. Sibbald, L. J. Albright, and P. R. Sibbald. Chemosensory responses of a heterotrophic microflagellate to bacteria and several nitrogen compounds. Marine Ecology Progress Series, 36:201–204, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
36/m036p201.pdf.

Schaffner:1982:STR

[SB82] L. C. Schaffner and D. F. Boesch. Spatial and temporal resource use by dominant benthic Amphipoda (Ampeliscidae and Corophiidae) on the Middle Atlantic Bight Outer Continental Shelf. Marine Ecology Progress Series, 9:231–243, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/9/m009p231.pdf.

Stepien:1985:NAN

[SB85a] C. A. Stepien and R. C. Brusca. Nocturnal attacks on nearshore fishes in southern California by crustacean zooplankton. Marine Ecology Progress Series, 25:91–105, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/25/m025p091.pdf.

Stoddart:1985:CGP

[SB85b] J. A. Stoddart and R. Black. Cycles of gametogenesis and plantation in the coral Pocillopora damicornis. Marine Ecology Progress Series, 23:153–164, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/23/m023p153.pdf.

Stocker:1987:IAT

[SB87] L. J. Stocker and P. R. Bergquist. Importance of algal turf, grazers, and spatial variability in the recruitment of a subtidal colonial invertebrate. Marine Ecology Progress Series, 39:285–291, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
39/m039p285.pdf.

Spies:1988:NRM

[SBK88] A. Spies, U. H. Brockmann, and G. Kattner. Nutrient regimes in the marginal ice zone of the Greenland Sea in summer. Marine Ecology Progress Series, 47:195–204, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/47/m047p195.pdf.

Sloan:1984:CDT

[SBR84] N. A. Sloan, S. M. Bower, and S. M. C. Robinson. Cocoon deposition on three crab species and fish parasitism by the leech Notostomum cyclostoma from deep fjords in northern British Columbia. Marine Ecology Progress Series, 20:51–58, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/20/
m020p051.pdf.

Shaw:1987:STA

[SBR87] P. W. Shaw, J. A. Beardmore, and J. S. Ryland. Sagartia troglodytes (Anthozoa: Actiniaria) consists of two species. Marine Ecology Progress Series, 41:21–28, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/41/m041p021.pdf.

Smith:1982:BWA

[SC82] John Clegg Smith and Chi Keung Chong. Body weight, activities of cytochrome oxidase and electron transport system in the liver of the American plaice Hippoglossoides platessoides. Can these activities serve as indicators of metabolism? Marine Ecology Progress Series, 9:171–179, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/9/m009p171.pdf.

Shafee:1984:EPP

[SC84a] M. S. Shafee and G. Conan. Energetic parameters of a population of Chlamys varia (Bivalvia: Pectinidae). Marine Ecology Progress Series, 18:253–262, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/18/m018p253.pdf.

Slater:1984:EMN

[SC84b] J. Slater and D. G. Capone. Effects of metals on nitrogen fixation and denitrification in slurries of anoxic salt-marsh sediment. Marine Ecology Progress Series, 18:89–95, 1984. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/18/m018p089.pdf.

Sanders:1985:ABE

[SC85] J. G. Sanders and S. J. Cibik. Adaptive behavior of euryhaline phytoplankton communities to arsenic stress. Marine Ecology Progress Series, 22:199–205, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/22/m022p199.pdf.

Southward:1986:DPW

[SC86] Eva C. Southward and James K. Culter. Discovery of Pogonophora in warm shallow waters of the Florida Shelf. Marine Ecology Progress Series, 28:287–289, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/28/m028p287.pdf.

ScottodiCarlo:1982:FES

[SCF^{+}82] B. Scotto di Carlo, G. Costanzo, E. Fresi, L. Guglielmo, and A. lanora. Feeding ecology and stranding mechanisms in two lanternfishes, Hygophum benoiti and Myctophum punctatum. Marine Ecology Progress Series, 9:13–24, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/9/m009p013.pdf.

Scheibling:1980:DFA

[Sch80a] R. E. Scheibling. Dynamics and feeding activity of high-density aggregations of Oreaster reticulatus (Echinodermata: Asteroidea) in a sand patch habitat. Marine Ecology Progress Series, 2:321– 327, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
2/m002p321.pdf.

Schleyer:1980:PEH

[Sch80b] M. H. Schleyer. A preliminary evaluation of heterotrophic utilisation of a labelled algal extract in a subtidal reef environment. Marine Ecology Progress Series, 3:223–229, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/3/m003p223.pdf.

Schleyer:1981:MDW

[Sch81a] M. H. Schleyer. Microorganisms and detritus in the water column of a subtidal reef of Natal. Marine Ecology Progress Series, 4:307– 320, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
4/m004p307.pdf.

Schneider:1981:SSP

[Sch81b] D. C. Schneider. Size-selective predation on mysids by birds (plovers). Marine Ecology Progress Series, 5:223–224, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/5/
m005p223.pdf.

Schmidt:1982:RAS

[Sch82a] G. H. Schmidt. Random and aggregative settlement in some sessile marine invertebrates. Marine Ecology Progress Series, 9:97–100, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
9/m009p097.pdf.

Schneider:1982:FSA

[Sch82b] D. Schneider. Fronts and seabird aggregations in the southeastern Bering Sea. Marine Ecology Progress Series, 10:101–103, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
10/m010p101.pdf.

Schmidt:1983:HTL

[Sch83a] G. H. Schmidt. The hydroid Tubularia larynx causing bloom of the ascidians Ciona intestinalis and Ascidiella aspersa. Marine Ecology Progress Series, 12:103–105, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/12/m012p103.pdf.

Schnack:1983:FCT

[Sch83b] S. B. Schnack. On the feeding of copepods on Thalassiosira partheneia from the Northwest African upwelling area. Marine Ecology Progress Series, 11:49–53, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/11/m011p049.pdf.

Schwinghamer:1983:GEH

[Sch83c] P. Schwinghamer. Generating ecological hypotheses from biomass spectra using causal analysis: a benthic example. Marine Ecology Progress Series, 13:151–166, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/13/m013p151.pdf.

Schleyer:1984:HUL

[Sch84a] M. H. Schleyer. Heterotrophic utilisation of a labelled algal extract in water samples from a subtidal reef. Marine Ecology Progress Series, 16:39–44, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/16/m016p039.pdf.

Schuhmacher:1984:RBP

[Sch84b] H. Schuhmacher. Reef-building properties of Tubastraea micranthus (Scleractinia, Dendrophylliidae), a coral without zooxanthellae. Marine Ecology Progress Series, 20:93–99, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/20/m020p093.pdf.

Scheltema:1987:RRG

[Sch87a] A. H. Scheltema. Reproduction and rapid growth in a deep-sea aplacophoran mollusc, Prochaetoderma yongei. Marine Ecology Progress Series, 37:171–180, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/37/m037p171.pdf.

Schneider:1987:RAD

[Sch87b] G. Schneider. Role of advection in the distribution and abundance of Pleurobrachia pileus in Kiel Bight. Marine Ecology Progress Series, 41:99–102, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/41/m041p099.pdf.

Schalk:1988:RET

[Sch88a] P. H. Schalk. Respiratory electron transport system (ETS) activities in zooplankton and micronekton of the Indo–Pacific region. Marine Ecology Progress Series, 44:25–35, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/44/m044p025.pdf.

Schwinghamer:1988:IPAb

[Sch88b] P. Schwinghamer. Influence of pollution along a natural gradient and in a mesocosm experiment on biomass-size spectra of benthic communities. Marine Ecology Progress Series, 46:199–206, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p199.pdf.

Schwinghamer:1988:IPAa

[Sch88c] P. Schwinghamer. Influence of pollution along a natural gradient and in a mesocosm experiment on sediment microbial numbers and biomass. Marine Ecology Progress Series, 46:193–197, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p193.pdf.

Schleyer:1989:MAW

[Sch89a] M. H. Schleyer. Microbial activity in water overlying the nearshore sand substratum in Natal, South Africa. Marine Ecology Progress Series, 53:101–103, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/53/m053p101.pdf.

Shapiro:1989:IRP

[SCH89b] L. P. Shapiro, L. Campbell, and E. M. Haugen. Immunochemical recognition of phytoplankton species. Marine Ecology Progress Series, 57:219–224, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/57/m057p219.pdf.

Scranton:1983:RCO

[Scr83] M. I. Scranton. The role of the cyanobacterium Oscillatoria (Trichodesmium) thiebautii in the marine hydrogen cycle. Marine Ecology Progress Series, 11:79–87, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/11/m011p079.pdf.

Swartz:1986:ECS

[SCSD86] R. C. Swartz, F. A. Cole, D. W. Schults, and W. A. DeBen. Ecological changes in the Southern California Bight near a large sewage outfall: benthic conditions in 1980 and 1983. Marine Ecology Progress Series, 31:1–13, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/31/m031p001.pdf.

Sole-Cava:1987:FGE

[SCT87] A. M. Sole-Cava and J. P. Thorpe. Further genetic evidence for the reproductive isolation of green sea anemone Actinia prasina Gosse from common intertidal beadlet anemone Actinia equina (L.). Marine Ecology Progress Series, 38:225–229, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/38/m038p225.pdf.

Schneider:1985:SDV

[SD85] David C. Schneider and David Cameron Duffy. Scale-dependent variability in seabird abundance. Marine Ecology Progress Series, 25:211–218, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/25/m025p211.pdf.

Suteau:1988:PME

[SDMN88] P. Suteau, M. Daubeze, M. L. Migaud, and J. F. Narbonne. PAHmetabolizing enzymes in whole mussels as biochemical tests for chemical pollution monitoring. Marine Ecology Progress Series, 46:45–49, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/46/m046p045.pdf.

Seiderer:1984:UBN

[SDRN84] L. J. Seiderer, C. L. Davis, F. T. Robb, and R. C. Newell. Utilisation of bacteria as nitrogen resource by kelp-bed mussel Choromytilus meridionalis. Marine Ecology Progress Series, 15: 109–116, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/15/m015p109.pdf.

Seiderer:1987:DEA

[SDRN87] L. J. Seiderer, C. L. Davis, F. T. Robb, and R. C. Newell. Digestive enzymes of the anchovy Engraulis capensis in relation to diet. Marine Ecology Progress Series, 35:15–23, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/35/m035p015.pdf.

Simard:1986:AEA

[SdT86] Y. Simard, R. de Ladurantaye, and J. C. Therriault. Aggregation of euphausiids along a coastal shelf in an upwelling environment. Marine Ecology Progress Series, 32:203–215, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/32/m032p203.pdf.

Sakshaug:1987:TOP

[SDY87] E. Sakshaug, S. Demers, and C. M. Yentsch. Thalassiosira oceanica and T. pseudonana: two different photoadaptational responses. Marine Ecology Progress Series, 41:275–282, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/41/
m041p275.pdf.

Stoecker:1985:EPH

[SE85] D. K. Stoecker and G. T. Evans. Effects of protozoan herbivory and carnivory in a microplankton food web. Marine Ecology Progress Series, 25:159–167, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/25/m025p159.pdf.

Shick:1981:PED

[SED81] J. M. Shick, K. C. Edwards, and J. H. Dearborn. Physiological ecology of the deposit-feeding sea star Ctenodiscus crispatus: Ciliated surfaces and animal-sediment lnteractions. Marine Ecology Progress Series, 5:165–184, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/5/m005p165.pdf.

Seitzinger:1987:NBU

[Sei87] S. P. Seitzinger. Nitrogen biogeochemistry in an unpolluted estuary: the importance of benthic denitrification. Marine Ecology Progress Series, 41:177–186, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/41/m041p177.pdf.

Selmer:1988:ARE

[Sel88] J.-S. Selmer. Ammonium regeneration in eutrophicated coastal waters of Sweden. Marine Ecology Progress Series, 44:265–273, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
44/m044p265.pdf.

Shafir:1980:ISC

[SF80] A. Shafir and J. G. Field. Importance of a small carnivorous isopod in energy transfer. Marine Ecology Progress Series, 3:203– 215, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
3/m003p203.pdf.

Shannon:1985:FSF

[SF85] L. V. Shannon and J. G. Field. Are fish stocks food-limited in the southern Benguela pelagic ecosystem? Marine Ecology Progress Series, 22:7–19, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/22/m022p007.pdf.

Suzuki:1986:CDS

[SF86] R. Suzuki and Y. Fujita. Chlorophyll decomposition in Skeletonema costatum: a problem in chlorophyll determination of water samples. Marine Ecology Progress Series, 28:81–85, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/28/m028p081.pdf.

Shpigel:1989:HPB

[SF89] M. Shpigel and L. Fishelson. Habitat partitioning between species of the genus Cephalopholis (Pisces, Serranidae) across the fringing reef of the Gulf of Aqaba (Red Sea). Marine Ecology Progress Series, 58:17–22, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/58/m058p017.pdf.

Sharp:1982:MNR

[SFHU82] J. H. Sharp, A. C. Frake, G. B. Hillier, and P. A. Underhill. Modeling nutrient regeneration in the ocean with an aquarium system. Marine Ecology Progress Series, 8:15–23, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/8/m008p015.pdf.

Stuart:1982:EAK

[SFN82] V. Stuart, J. G. Field, and R. C. Newell. Evidence for absorption of kelp detritus by the ribbed mussel Aulacomya ater using a new ^{51}Cr-labelled microsphere technique. Marine Ecology Progress Series, 9:263–271, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/9/m009p263.pdf.

Stoner:1984:GVM

[SG84] A. W. Stoner and H. S. Greening. Geographic variation in the macrofaunal associates of pelagic Sargassum and some biogeographic implications. Marine Ecology Progress Series, 20:185–192, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
20/m020p185.pdf.

Schmidt:1985:SAD

[SG85] G. H. Schmidt and D. J. Garbutt. Species abundance data from fouling communities conform to the gamma distribution. Marine Ecology Progress Series, 23:287–290, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/23/m023p287.pdf.

Seitzinger:1987:NFC

[SG87a] S. P. Seitzinger and J. H. Garber. Nitrogen fixation and ^{15}N2 calibration of the acetylene reduction assay in coastal marine sediments. Marine Ecology Progress Series, 37:65–73, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/37/m037p065.pdf.

Sorensen:1987:IBF

[SG87b] J. Sørensen and E. Glob. Influence of benthic fauna on trimethylamine concentrations in coastal marine sediments. Marine Ecology Progress Series, 39:15–21, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/39/m039p015.pdf.

Shanks:1988:CDS

[SG88a] A. L. Shanks and W. M. Graham. Chemical defense in a scyphomedusa. Marine Ecology Progress Series, 45:81–86, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/45/
m045p081.pdf.

Sundback:1988:IMN

[SG88b] K. Sundbäck and W. Granéli. Influence of microphytobenthos on the nutrient flux between sediment and water- a laboratory study. Marine Ecology Progress Series, 43:63–69, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/43/m043p063.pdf.

Siron:1989:CFA

[SGB89] R. Siron, G. Giusti, and B. Berland. Changes in the fatty acid composition of Phaeodactylum tricornutum and Dunaliella tertiolecta during growth and under phosphorus deficiency. Marine Ecology Progress Series, 55:95–100, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/55/m055p095.pdf.

Schneider:1987:PEM

[SGG87] David C. Schneider, Jean-Marc Gagnon, and Kent D. Gilkinson. Patchiness of epibenthic megafauna on the outer Grand Banks of Newfoundland. Marine Ecology Progress Series, 39:1– 13, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
39/m039p001.pdf.

Spiro:1986:SRM

[SGSD86] B. Spiro, P. B. Greenwood, A. J. Southward, and P. R. Dando. 13C/^{12}C ratios in marine invertebrates from reducing sediments: confirmation of nutricional importance of chemoautotrophic endosymbiotic bacteria. Marine Ecology Progress Series, 28:233– 240, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
28/m028p233.pdf.

Schmitt:1984:OPS

[SH84] R. J. Schmitt and S. J. Holbrook. Ontogeny of prey selection by black surfperch Embiotoca jacksoni (Pisces: Embiotocidae): the roles of fish morphology, foraging behavior, and patch selection. Marine Ecology Progress Series, 18:225–239, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/18/m018p225.pdf.

Snelgrove:1985:SDD

[SH85] P. V. R. Snelgrove and R. L. Haedrich. Structure of the deep demersal fish fauna off Newfoundland. Marine Ecology Progress Series, 27:99–107, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/27/m027p099.pdf.

Smith:1989:CCN

[SH89] S. V. Smith and J. T. Hollibaugh. Carbon-controlled nitrogen cycling in a marine ’macrocosm’- an ecosystem-scale model for managing cultural eutrophication. Marine Ecology Progress Series, 52: 103–109, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/52/m052p103.pdf.

Shanks:1983:SSA

[Sha83] A. L. Shanks. Surface slicks associated with tidally forced internal waves may transport pelagic larvae of benthic invertebrates and fishes shoreward. Marine Ecology Progress Series, 13:311– 315, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p311.pdf.

Shaffer:1988:CBM

[Sha88a] G. P. Shaffer. A comparison of benthic microfloral production on the West and Gulf coasts of the United States: an introduction to the dynamic K-systems model. Marine Ecology Progress Series, 43:55–62, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/43/m043p055.pdf.

Shaffer:1988:KSA

[Sha88b] G. P. Shaffer. K-systems analysis for determining the factors influencing benthic microfloral productivity in a Louisiana estuary, USA. Marine Ecology Progress Series, 43:43–54, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/43/m043p043.pdf.

Shafee:1989:RPP

[Sha89] M. S. Shafee. Reproduction of Perna picta (Mollusca: Bivalvia) from the Atlantic coast of Morocco. Marine Ecology Progress Series, 53:235–245, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/53/m053p235.pdf.

Smigielski:1983:SED

[SHB^{+}83] A. S. Smigielski, T. A. Halavik, L. J. Buckley, S. M. Drew, and G. C. Laurence. Spawning, embryo development and growth of the American sand lance Ammodytes americanus in the laboratory. Marine Ecology Progress Series, 14:287–292, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/14/m014p287.pdf.

Sieracki:1987:EFP

[SHCL87] M. E. Sieracki, L. W. Haas, D. A. Caron, and E. J. Lessard. Effect of fixation on particle retention by microflagellates: underestimation of grazing rates. Marine Ecology Progress Series, 38:251–258, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
38/m038p251.pdf.

Sheppard:1980:CCZ

[She80] C. R. C. Sheppard. Coral cover, zonation and diversity on reef slopes of Chagos Atolls, and population structures of the major species. Marine Ecology Progress Series, 2:193–205, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/2/m002p193.pdf.

Sheppard:1982:CPR

[She82] C. R. C. Sheppard. Coral populations on reef slopes and their major controls. Marine Ecology Progress Series, 7:83–115, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
7/m007p083.pdf.

Sheppard:1985:USC

[She85] C. R. C. Sheppard. Unoccupied substrate in the central Great Barrier Reef: role of coral interactions. Marine Ecology Progress Series, 25:259–268, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/25/m025p259.pdf.

Stock:1989:CLO

[SHG89] M. Stock, R. F. M. Herber, and H. M. A. Geron. Cadmium levels in oystercatcher Haematopus ostralegus from the German Wadden Sea. Marine Ecology Progress Series, 53:227–234, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/53/m053p227.pdf.

Siebers:1981:MIU

[SHH81] D. Siebers, M. Hoppenheit, and U. Harms. Mercury influences uptake of amino acids by marine bacteria. Marine Ecology Progress Series, 6:101–113, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/6/m006p101.pdf.

Stretch:1988:FBA

[SHH^{+}88] J. J. Stretch, P. P. Hamner, W. M. Hamner, W. C. Michel, J. Cook, and C. W. Sullivan. Foraging behavior of Antarctic krill Euphausia superba on sea ice microalgae. Marine Ecology Progress Series, 44: 131–139, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/44/m044p131.pdf.

Shimura:1980:EAP

[Shi80] S. Shimura. Exological aspects of Proctoeces ichiharai (Trematoda: Digenea) parasitic in Batillus cornutus (Gastropoda). Marine Ecology Progress Series, 3:145–149, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/3/m003p145.pdf.

Smith:1986:MCE

[SHIP86] R. E. H. Smith, W. G. Harrison, B. Irwin, and T. Piatt. Metabolism and carbon exchange in microplankton of the Grand Banks (Newfoundland). Marine Ecology Progress Series, 34:171– 183, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
34/m034p171.pdf.

Schwinghamer:1986:PPR

[SHPH86] P. Schwinghamer, B. Hargrave, D. Peer, and C. M. Hawkins. Partitioning of production and respiration among size groups of organisms in an intertidal benthic community. Marine Ecology Progress Series, 31:131–142, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/31/m031p131.pdf.

Shulenberger:1980:FAH

[Shu80] E. Shulenberger. Factor analyses of a hyperiid amphipod assemblage from the North Pacific central gyre. Marine Ecology Progress Series, 2:109–120, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/2/m002p109.pdf.

Siegismund:1982:LCP

[Sie82] H. R. Siegismund. Life cycle and production of Hydrobia ventrosa and H. neglecta (Mollusca: Prosobranchia). Marine Ecology Progress Series, 7:75–82, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/7/m007p075.pdf.

Siebeck:1988:EIU

[Sie88] O. Siebeck. Experimental investigation of UV tolerance in hermatypic corals (Scleractinia). Marine Ecology Progress Series, 43: 95–103, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/43/m043p095.pdf.

Sinclair:1986:NTS

[Sin86] D. F. Sinclair. A new test for species sequencing. Marine Ecology Progress Series, 30:283–286, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/30/m030p283.pdf.

Sand-Jensen:1988:MLR

[SJ88a] K. Sand-Jensen. Minimum light requirement for growth in Ulva lactuca. Marine Ecology Progress Series, 50:187–193, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/50/
m050p187.pdf.

Sand-Jensen:1988:PRU

[SJ88b] K. Sand-Jensen. Photosynthetic responses of Ulva lactuca at very low light. Marine Ecology Progress Series, 50:195–201, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/50/
m050p195.pdf.

Sundback:1989:IAD

[SJNL89] K. Sundbäck, B. Jönsson, P. Nilsson, and I. Lindström. Impact of accumulating drifting macroalgae on a shallow-water sediment system: an experimental study. Marine Ecology Progress Series, 58: 261–274, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/58/m058p261.pdf.

Seidel:1982:ISA

[SJSO82] C. R. Seidel, D. M. Johns, P. S. Schauer, and C. E. Olney. International study on Artemia. XXVI. Food value of nauplii from reference Artemia cysts and four geographical collections of Artemia for mud crab larvae. Marine Ecology Progress Series, 8:309–312, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
8/m008p309.pdf.

Shumway:1982:OCA

[SK82] S. E. Shumway and R. K. Koehn. Oxygen consumption in the American oyster Crassostrea virginica. Marine Ecology Progress Series, 9:59–68, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/9/m009p059.pdf.

Stuart:1984:EFR

[SK84] V. Stuart and D. W. Klumpp. Evidence for food-resource partitioning by kelp-bed filter feeders. Marine Ecology Progress Series, 16:27–37, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/16/m016p027.pdf.

Stillwell:1985:FFE

[SK85] C. E. Stillwell and N. E. Kohler. Food and feeding ecology of the sword-fish Xiphias gladius in the western North Atlantic Ocean with estimates of daily ration. Marine Ecology Progress Series, 22: 239–247, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/22/m022p239.pdf.

Spies:1989:SIR

[SKIR89] R. B. Spies, H. Kruger, R. Ireland, and D. W. Rice, Jr. Stable isotope ratios and contaminant concentrations in a sewagedistorted food web. Marine Ecology Progress Series, 54:157–170, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
54/m054p157.pdf.

Sorokin:1985:ADM

[SKM85] Yu. I. Sorokin, A. I. Kopylov, and N. V. Mamaeva. Abundance and dynamics of microplankton in the central tropical Indian Ocean. Marine Ecology Progress Series, 24:27–41, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/24/m024p027.pdf.

Solbakken:1984:IFL

[SKS^{+}84] J. E. Solbakken, A. H. Knap, T. D. Sleeter, C. E. Searle, and K. H. Palmork. Investigation into the fate of ^{14}C-labelled xenobiotics (naphthalene, phenan-threne, 2,4,5,2’,4’,5’-hexachlorobiphenyl, octachlorostyrene) in Bermudian corals. Marine Ecology Progress Series, 16:149–154, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/16/m016p149.pdf.

Small:1989:RPC

[SLE^{+}89] L. F. Small, M. R. Landry, R. W. Eppley, F. Azam, and A. F. Carlucci. Role of plankton in the carbon and nitrogen budgets of Santa Monica Basin, California. Marine Ecology Progress Series, 56:57–74, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/56/m056p057.pdf.

Stoner:1988:SEM

[SLMB88] A. W. Stoner, R. N. Lipcius, L. S. Marshall, Jr., and A. T. Bardales. Synchronous emergence and mass migration in juvenile queen conch. Marine Ecology Progress Series, 49:51–55, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
49/m049p051.pdf.

Stuart:1981:HUP

[SLN81] V. Stuart, M. I. Lucas, and R. C. Newell. Heterotrophic utilisation of particulate matter from the kelp Laminaria pallida. Marine Ecology Progress Series, 4:337–348, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p337.pdf.

Steiner:1982:ASU

[SLO82] W. W. Steiner, J. J. Luczkovich, and B. L. Olla. Activity, shelter usage, growth and recruitment of juvenile red hake Urophycis chuss. Marine Ecology Progress Series, 7:125–135, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/7/m007p125.pdf.

Sloan:1985:LHC

[Slo85] N. A. Sloan. Life history characteristics of fjord-dwelling golden king crabs Lithodes aequispina. Marine Ecology Progress Series, 22:219–228, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/22/m022p219.pdf.

Shaw:1983:DPI

[SLP83] K. M. Shaw, P. J. D. Lambshead, and H. M. Piatt. Detection of pollution-induced disturbance in marine benthic assemblages with special reference to nematodes. Marine Ecology Progress Series, 11:195–202, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/11/m011p195.pdf.

Segner:1984:EMI

[SM84] H. Segner and H. Moller. Electron microscopical investigations on starvation-induced liver pathology in flounders Platichthys flesus. Marine Ecology Progress Series, 19:193–196, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/19/m019p193.pdf.

Smith:1985:RLH

[SM85] W. G. Smith and W. W. Morse. Retention of larval haddock Melanogrammus aeglefinus in the Georges Bank region, a gyreinfluenced spawning area. Marine Ecology Progress Series, 24:1– 13, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
24/m024p001.pdf.

Sullivan:1986:FCS

[SM86] B. K. Sullivan and L. T. McManus. Factors controlling seasonal succession of the copepods Acartia hudsonica and A. tonsa in Narragansett Bay, Rhode Island: temperature and resting egg production. Marine Ecology Progress Series, 28:121–128, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/28/m028p121.pdf.

Sainte-Marie:1987:MSF

[SM87a] B. Sainte-Marie. Meal size and feeding rate of the shallow-water lysianassid Anonyx sarsi (Crustacea: Amphipoda). Marine Ecology Progress Series, 40:209–219, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/40/m040p209.pdf.

Stacey:1987:CEN

[SM87b] B. M. Stacey and B. M. Marcotte. Chronic effect of no. 2 fuel oil on population dynamics of harpacticoid copepods in experimental marine mesocosms. Marine Ecology Progress Series, 40:61–68, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
40/m040p061.pdf.

Smalley:1983:PEI

[Sma83] T. L. Smalley. Possible effects of intraspecific competition on the population structure of a solitary vermetid mollusk. Marine Ecology Progress Series, 14:139–144, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/14/m014p139.pdf.

Sainte-Marie:1985:SGS

[SMB85] B. Sainte-Marie and P. Brunei. Suprabenthic gradients of swimming activity by cold-water gammaridean amphipod crustacea over a muddy shelf in the Gulf of Saint Lawrence. Marine Ecology Progress Series, 23:57–69, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/23/m023p057.pdf.

Smith:1981:RSD

[Smi81] R. C. Smith. Remote senslng and depth distribution of ocean chlorophyll. Marine Ecology Progress Series, 5:359–361, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/5/
m005p359.pdf.

Smith:1986:IPD

[Smi86] B. D. Smith. Implications of population dynamics and interspecific competition for harvest management of the seaweed Laminaria. Marine Ecology Progress Series, 33:7–18, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/33/m033p007.pdf.

Short:1986:REW

[SMN86] F. T. Short, A. C. Mathieson, and J. I. Nelson. Recurrence of the eelgrass wasting disease at the border of New Hampshire and Maine, USA. Marine Ecology Progress Series, 29:89–92, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
29/m029p089.pdf.

Santelices:1981:CAC

[SMO81] B. Santelices, S. Montalva, and P. Oliger. Competitive algal community organization in exposed intertidal habitats from Central Chile. Marine Ecology Progress Series, 6:267–276, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/6/m006p267.pdf.

Sasaki:1981:VFP

[SN81] H. Sasaki and S. Nishizawa. Vertical flux profiles of particulate material in the Sea of Sanriku. Marine Ecology Progress Series, 6: 191–201, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/6/m006p191.pdf.

Seiderer:1985:RSP

[SN85] L. J. Seiderer and R. C. Newell. Relative significance of phytoplankton, bacteria and plant detritus as carbon and nitrogen resources for the kelp bed filter-feeder Choromytilus mendionalis. Marine Ecology Progress Series, 22:127–139, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/22/m022p127.pdf.

Seiderer:1988:EPF

[SN88] L. J. Seiderer and R. C. Newell. Exploitation of phytoplankton as a food resource by the kelp bed ascidian Pyura stolonifera. Marine Ecology Progress Series, 50:107–115, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/50/m050p107.pdf.

Seki:1989:GCT

[SN89] H. Seki and T. Naganuma. Growth characteristics of Thiobacterium sp from the plume of hydrothermal vents of the North Fiji Basin. Marine Ecology Progress Series, 54:199–202, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/54/m054p199.pdf.

Stuart:1982:CKD

[SNL82] V. Stuart, R. C. Newell, and M. I. Lucas. Conversion of kelp debris and faecal material from the mussel Aulacomya ater by marine micro-organisms. Marine Ecology Progress Series, 7:47– 57, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
7/m007p047.pdf.

Santelices:1983:ECR

[SO83a] B. Santelices and F. P. Ojeda. Effects of canopy removal on the understory algal community structure of coastal forests of Macrocystis pyrifera from southern South America. Marine Ecology Progress Series, 14:165–173, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/14/m014p165.pdf.

Santelices:1983:PDC

[SO83b] B. Santelices and F. P. Ojeda. Population dynamics of coastal forests of Macrocystis pyrifera in Puerto Toro, Isla Navarino, Southern Chile. Marine Ecology Progress Series, 14:175–183, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
14/m014p175.pdf.

Santelices:1984:RGS

[SO84] B. Santelices and F. P. Ojeda. Recruitment, growth and survival of Lessonia nigrescens (Phaeophyta) at various tidal levels in exposed habitats of central Chile. Marine Ecology Progress Series, 19:73–82, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/19/m019p073.pdf.

Shaffer:1985:REE

[SO85] G. P. Shaffer and C. P. Onuf. Reducing the error in estimating annual production of benthic microflora: hourly to monthly rates, patchiness in space and time. Marine Ecology Progress Series, 26: 221–231, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/26/m026p221.pdf.

Shulman:1987:WCT

[SO87] M. J. Shulman and J. C. Ogden. What controls tropical reef fish populations: recruitment or benthic mortality? An example in the Caribbean reef fish Haemulon flavolineatum. Marine Ecology Progress Series, 39:233–242, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/39/m039p233.pdf.

Sogard:1984:UMF

[Sog84] S. M. Sogard. Utilization of meiofauna as a food source by a grassbed fish, the spotted dragonet Callionymus pauciradiatus. Marine Ecology Progress Series, 17:183–191, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/17/m017p183.pdf.

Somiya:1980:FES

[Som80] H. Somiya. Fishes with eye shine: Functional morphology of guanine type tapetum lucidum. Marine Ecology Progress Series, 2:9– 26, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
2/m002p009.pdf.

Sorgeloos:1980:ARA

[Sor80] P. Sorgeloos. Availability of reference Artemia cysts. Marine Ecology Progress Series, 3:363–364, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/3/m003p363.pdf.

Sorokin:1985:PMP

[Sor85] Yu. I. Sorokin. Phosphorus metabolism in planktonic communities of the eastern tropical Pacific Ocean. Marine Ecology Progress Series, 27:87–97, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/27/m027p087.pdf.

Smith:1984:CET

[SP84a] R. E. H. Smith and T. Piatt. Carbon exchange and ^{14}C tracer methods in a nitrogen-limited diatom, Thalassiosira pseudonana. Marine Ecology Progress Series, 16:75–87, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/16/m016p075.pdf.

Summerson:1984:RPO

[SP84b] H. C. Summerson and C. H. Peterson. Role of predation in organizing benthic communities of a temperate-zone seagrass bed. Marine Ecology Progress Series, 15:63–77, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/15/m015p063.pdf.

Smith:1985:TRR

[SP85] John Clegg Smith and Trevor Platt. Temperature responses of ribulose bisphosphate carboxylase and photosynthetic capacity in Arctic and tropical phytoplankton. Marine Ecology Progress Series, 25:31–37, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/25/m025p031.pdf.

Schneider:1986:SDC

[SP86a] David C. Schneider and John F. Piatt. Scale-dependent correlation of seabirds with schooling fish in a coastal ecosystem. Marine Ecology Progress Series, 32:237–246, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/32/m032p237.pdf.

Shailaja:1986:RBP

[SP86b] M. S. Shailaja and A. Pant. Relationship between particulate and extracellular carbon compounds of phytoplankton photosynthesis in a tropical estuary. Marine Ecology Progress Series, 32:161– 167, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
32/m032p161.pdf.

Stephenson:1988:SVM

[SP88] R. L. Stephenson and M. J. Power. Semidiel vertical movements in Atlantic herring Clupea harengus larvae: a mechanism for larval retention? Marine Ecology Progress Series, 50:3–11, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
50/m050p003.pdf.

Steinberg:1989:FFC

[SP89] P. D. Steinberg and V. J. Paul. Fish feeding and chemical defenses of tropical brown algae in Western Australia. Marine Ecology Progress Series, 58:253–259, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/58/m058p253.pdf.

Spaargaren:1980:HPB

[Spa80] D. H. Spaargaren. Hydrodynamic properties of benthic marine crustacea. II. Energy requirements for maintaining vertical position. Marine Ecology Progress Series, 2:153–156, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/2/m002p153.pdf.

Sogard:1987:EFC

[SPH87] S. M. Sogard, G. V. N. Powell, and J. G. Holmquist. Epibenthic fish communities on Florida Bay banks: relations with physical parameters and seagrass cover. Marine Ecology Progress Series, 40:25–39, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/40/m040p025.pdf.

Spies:1987:GRA

[Spi87] A. Spies. Growth rates of Antarctic marine phytoplankton in the Weddell Sea. Marine Ecology Progress Series, 41:267–274, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
41/m041p267.pdf.

SooHoo:1987:SLA

[SPK^{+}87] J. B. SooHoo, A. C. Palmisano, S. T. Kottmeier, M. P. Lizotte, S. L. SooHoo, and C. W. Sullivan. Spectral light absorption and quantum yield of photosynthesis in sea ice microalgae and a bloom of Phaeocystis pouchetii from McMurdo Sound, Antarctica. Marine Ecology Progress Series, 39:175–189, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/39/m039p175.pdf.

Smith:1984:AMP

[SPL^{+}84] John Clegg Smith, Trevot Piatt, W. K. W. Li, E. P. W. Home, W. G. Harrison, D. V. Subba Rao, and B. D. Irwin. Arctic marine photoautotrophic picoplankton. Marine Ecology Progress Series, 20:207–220, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/20/m020p207.pdf.

Sprung:1984:PEMc

[Spr84a] M. Sprung. Physiological energetics of mussel larvae (Mytiius edulis). III. Respiration. Marine Ecology Progress Series, 18:171– 178, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
18/m018p171.pdf.

Sprung:1984:PEMd

[Spr84b] M. Sprung. Physiological energetics of mussel larvae (Mytiius edulis). IV. Efficiencies. Marine Ecology Progress Series, 18:179– 186, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
18/m018p179.pdf.

Sprung:1984:PEMa

[Spr84c] M. Sprung. Physiological energetics of mussel larvae (Mytilus edulis). I. Shell growth and biomass. Marine Ecology Progress Series, 17:283–293, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/17/m017p283.pdf.

Sprung:1984:PEMb

[Spr84d] M. Sprung. Physiological energetics of mussel larvae (Mytilus edulis). II. Food uptake. Marine Ecology Progress Series, 17:295– 305, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
17/m017p295.pdf.

Sorensson:1989:FNM

[SPSS89] F. Sörensson, K. Pettersson, J.-S. Selmer, and E. Sahlsten. Flows of nitrogen in a mesocosm experiment in the Baltic Sea. Marine Ecology Progress Series, 58:77–88, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/58/m058p077.pdf.

Springer:1985:CBF

[SR85] A. M. Springer and D. G. Roseneau. Copepod-based food webs: auklets and oceanography in the Bering Sea. Marine Ecology Progress Series, 21:229–237, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/21/m021p229.pdf.

Schubert:1986:PEN

[SR86] A. Schubert and K. Reise. Predatory effects of Nephtys hombergii on other polychaetes in tidal flat sediments. Marine Ecology Progress Series, 34:117–124, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/34/m034p117.pdf.

Scott:1987:EGS

[SR87] F. J. Scott and G. R. Russ. Effects of grazing on species composition of the epilithic algal community on coral reefs of the central Great Barrier Reef. Marine Ecology Progress Series, 39:293–304, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
39/m039p293.pdf.

Springer:1986:SRF

[SRL^{+}86] A. M. Springer, D. G. Roseneau, D. S. Lloyd, C. P. McRoy, and E. C. Murphy. Seabird responses to fluctuating prey availability in the eastern Bering Sea. Marine Ecology Progress Series, 32:1– 12, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
32/m032p001.pdf.

Sherr:1989:BPC

[SRS89] E. B. Sherr, F. Rassoulzadegan, and B. F. Sherr. Bacterivory by pelagic choreotrichous ciliates in coastal waters of the NW Mediterranean Sea. Marine Ecology Progress Series, 55:235–240, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
55/m055p235.pdf.

Sherman:1983:RND

[SRTM83] K. M. Sherman, J. A. Reidenauer, D. Thistle, and D. Meeter. Role of a natural disturbance in an assemblage of marine freeliving nematodes. Marine Ecology Progress Series, 11:23–30, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
11/m011p023.pdf.

Santos:1980:MSB

[SS80a] S. L. Santos and J. L. Simon. Marine soft-bottom community establishment following annual defamation: Larval or adult recruitment? Marine Ecology Progress Series, 2:235–241, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/2/m002p235.pdf.

Santos:1980:RSB

[SS80b] S. L. Santos and J. L. Simon. Response of soft-bottom benthos to annual catastrophic disturbance in a South Florida estuary. Marine Ecology Progress Series, 3:347–355, 1980. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/3/m003p347.pdf.

Spotte:1982:HRT

[SS82] S. Spotte and P. E. Stake. Hand-rearing of twin Gray Seals (Halichoerus grypus) from birth to weaning. Marine Ecology Progress Series, 9:181–189, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/9/m009p181.pdf.

Sale:1986:RPD

[SS86a] P. F. Sale and W. J. Steel. Random placement and the distribution of fishes among coral patch reefs. Marine Ecology Progress Series, 28:165–174, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/28/m028p165.pdf.

Sieracki:1986:SIG

[SS86b] M. E. Sieracki and J. McN. Sieburth. Sunlight-induced growth delay of planktonic marine bacteria in filtered seawater. Marine Ecology Progress Series, 33:19–27, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/33/m033p019.pdf.

Sorensson:1987:NDC

[SS87] F. Sörensson and E. Sahlsten. Nitrogen dynamics of a cyanobacteria bloom in the Baltic Sea: new versus regenerated production. Marine Ecology Progress Series, 37:277–284, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/37/m037p277.pdf.

Sale:1989:TVP

[SS89a] P. F. Sale and W. J. Steel. Temporal variability in patterns of association among fish species on coral patch reefs. Marine Ecology Progress Series, 51:35–47, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/51/m051p035.pdf.

Shirley:1989:IVD

[SS89b] S. M. Shirley and T. C. Shirley. Interannual variability in density, timing and survival of Alaskan red king crab Paralithodes camtschatica larvae. Marine Ecology Progress Series, 54:51–59, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
54/m054p051.pdf.

Sherr:1986:TIB

[SSA^{+}86] B. F. Sherr, E. B. Sherr, T. L. Andrew, R. D. Fallon, and S. Y. Newell. Trophic interactions between heterotrophic protozoa and bacterioplankton in estuarine water analyzed with selective metabolic inhibitors. Marine Ecology Progress Series, 32:169–180, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
32/m032p169.pdf.

Subramanian:1980:ESS

[SSK80] P. Subramanian, S. Samasivam, and K. Krishnamurthy. Experimental study on the salinity tolerance of Macrobrachium idae larvae. Marine Ecology Progress Series, 3:71–73, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/3/m003p071.pdf.

Sherman:1984:SSF

[SSM^{+}84] K. Sherman, W. Smith, W. Morse, M. Berman, J. Green, and L. Ejsymont. Spawning strategies of fishes in relation to circulation, phytoplankton production, and pulses in zooplankton off the north-eastern United States. Marine Ecology Progress Series, 18: 1–19, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/18/m018p001.pdf.

Sano:1987:LTE

[SSN87] M. Sano, M. Shimizu, and Y. Nose. Long-term effects of destruction of hermatypic corals by Acanthaster plana infestation on reef fish communities at Iriomote Island, Japan. Marine Ecology Progress Series, 37:191–199, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/37/m037p191.pdf.

Sherr:1989:SMB

[SSPA89] B. F. Sherr, E. B. Sherr, and C. Pedros-Alio. Simultaneous measurement of bacterio-plankton production and protozoan bacterivory in estuarine water. Marine Ecology Progress Series, 54: 209–219, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/54/m054p209.pdf.

Stoecker:1989:OMF

[SST89] D. K. Stoecker, N. Swanberg, and S. Tyler. Oceanic mixotrophic flatworms. Marine Ecology Progress Series, 58:41–51, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/58/
m058p041.pdf.

Shin:1982:SDI

[ST82] P. K. S. Shin and G. B. Thompson. Spatial distribution of the infaunal benthos of Hong Kong. Marine Ecology Progress Series, 10:37–47, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/10/m010p037.pdf.

Stearns:1986:CGB

[Ste86] D. E. Stearns. Copepod grazing behavior in simulated natural light and its relation to nocturnal feeding. Marine Ecology Progress Series, 30:65–76, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/30/m030p065.pdf.

Stergiou:1989:CMV

[Ste89] Konstantinos I. Stergiou. Capelin Mallotus villosus (Pisces: Osmeridae), glaciations, and speciation: a nomothetic approach to fisheries ecology and reproductive biology. Marine Ecology Progress Series, 56:211–224, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/56/m056p211.pdf.

Stephenson:1986:USC

[STM86] R. L. Stephenson, F. C. Tan, and K. H. Mann. Use of stable carbon isotope ratios to compare plant material and potential consumers in a seagrass bed and a kelp bed in Nova Scotia, Canada. Marine Ecology Progress Series, 30:1–7, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/30/m030p001.pdf.

Stoecker:1988:AAM

[STM88] D. K. Stoecker, A. Taniguchi, and A. E. Michaels. Abundance of autotrophic, mixotrophic and heterotrophic planktonic ciliates in shelf and slope waters. Marine Ecology Progress Series, 50:241– 254, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
50/m050p241.pdf.

Stoecker:1980:RBC

[Sto80a] D. Stoecker. Relationships between chemical defense and ecology in benthic ascidians. Marine Ecology Progress Series, 3:257–265, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
3/m003p257.pdf.

Stoner:1980:PCS

[Sto80b] A. W. Stoner. Perception and choice of substratum by epifaunal amphipods associated with seagrasses. Marine Ecology Progress Series, 3:105–111, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/3/m003p105.pdf.

Stoner:1985:PCA

[Sto85] A. W. Stoner. Penicillus capitatus: an algal island for macrocrustaceans. Marine Ecology Progress Series, 26:279–287, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/26/
m026p279.pdf.

Stocker:1986:AEC

[Sto86] L. J. Stocker. Artifactual effects of caging on the recruitment and survivorship of a subtidal colonial invertebrate. Marine Ecology Progress Series, 34:305–307, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/34/m034p305.pdf.

Stoner:1988:NGF

[Sto88] A. W. Stoner. A nursery ground for four tropical Penaeus species: Laguna Joyuda, Puerto Rico. Marine Ecology Progress Series, 42: 133–141, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/42/m042p133.pdf.

Stoner:1989:WMM

[Sto89] A. W. Stoner. Winter mass migration of juvenile queen conch Strombus gigas and their influence on the benthic environment. Marine Ecology Progress Series, 56:99–104, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/56/m056p099.pdf.

Solbakken:1984:UEA

[STP84] J. E. Solbakken, S. Tilseth, and K. H. Palmork. Uptake and elimination of aromatic hydrocarbons and a chlorinated biphenyl in eggs and larvae of cod Gadus morhua. Marine Ecology Progress Series, 16:297–301, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/16/m016p297.pdf.

Spanier:1988:SSS

[STPA88] E. Spanier, M. Tom, S. Pisanty, and G. Almog. Seasonality and shelter selection by the slipper lobster Scyllarides latus in the southeastern Mediterranean. Marine Ecology Progress Series, 42: 247–255, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/42/m042p247.pdf.

Stuart:1986:FME

[Stu86] V. Stuart. Feeding and metabolism of Euphausia lucens (Euphausiacea) in the southern Benguela current. Marine Ecology Progress Series, 30:117–125, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/30/m030p117.pdf.

Stearns:1989:DCE

[STW89] D. E. Stearns, P. A. Tester, and R. L. Walker. Diel changes in the egg production rate of Acartia tonsa (Copepoda, Calanoida) and related environmental factors in two estuaries. Marine Ecology Progress Series, 52:7–16, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/52/m052p007.pdf.

Santelices:1987:ALH

[SU87] B. Santelices and R. Ugarte. Algal life-history strategies and resistance to digestion. Marine Ecology Progress Series, 35:267–275, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
35/m035p267.pdf.

Sulkin:1984:BBD

[Sul84] S. D. Sulkin. Behavioral basis of depth regulation in the larvae of brachyuran crabs. Marine Ecology Progress Series, 15:181–205, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
15/m015p181.pdf.

Sloan:1980:DFS

[Sv80a] N. A. Sloan and B. von Bodungen. Distribution and feeding of the sea cucumber Isostichopus badionotus in relation to shelter and sediment criteria of the Bermuda platform. Marine Ecology Progress Series, 2:257–264, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/2/m002p257.pdf.

Sulkin:1980:EES

[Sv80b] S. D. Sulkin and W. F. van Heukelem. Ecological and evolutionary significance of nutritional flexibility in planktotrophic larvae of the deep sea red crab Geryon quinquedens and the stone crab (Menippe mercenaria). Marine Ecology Progress Series, 2:91–95, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
2/m002p091.pdf.

Staples:1985:EJB

[SV85a] D. J. Staples and D. J. Vance. Emigration of juvenile banana prawns Penaeus merguiensis from a mangrove estuary and recruitment to offshore areas in the wet-dry tropics of the Gulf of Carpentaria, Australia. Marine Ecology Progress Series, 27:239– 252, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
27/m027p239.pdf.

Staples:1985:STL

[SV85b] D. J. Staples and D. J. Vance. Short-term and long-term influences of the immigration of postlarval banana prawns Penaeus merguiensis, into a mangrove estuary of the Gulf of Carpentaria, Australia. Marine Ecology Progress Series, 23:15–29, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/23/m023p015.pdf.

Sybesma:1981:ETC

[SvDB81] J. Sybesma, F. C. van Duyl, and R. P. M. Bak. The ecology of the tropical compound ascidian Trididemnum solidum. III. Symbiotic association with unicellular algae. Marine Ecology Progress Series, 6:53–59, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/6/m006p053.pdf.

Sulkin:1983:BBD

[SVK83] S. D. Sulkin, W. F. Van Heukelem, and P. Kelly. Behavioral basis of depth regulation in hatching and post-larval stages of the mud crab Eurypanopeus depressus. Marine Ecology Progress Series, 11: 157–164, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/11/m011p157.pdf.

Sammarco:1982:DTI

[SW82a] P. W. Sammarco and A. H. Williams. Damselfish territoriality: Influence on Diadema distribution and implications for coral community structure. Marine Ecology Progress Series, 8:53–59, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
8/m008p053.pdf.

Scheltema:1982:STL

[SW82b] R. S. Scheltema and I. P. Williams. Significance of temperature to larval survival and length of development in Balanus ebumeus (Crustacea: Cirripedia). Marine Ecology Progress Series, 9:43– 49, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
9/m009p043.pdf.

Schmidt:1984:ECD

[SW84] G. H. Schmidt and G. F. Warner. Effects of caging on the development of a sessile epifaunal community. Marine Ecology Progress Series, 15:251–263, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/15/m015p251.pdf.

Simenstad:1985:EOF

[SW85] C. A. Simenstad and R. C. Wissmar. δ ^{13}C evidence of the origins and fates of organic carbon in estuarine and near-shore food webs. Marine Ecology Progress Series, 22:141–152, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/22/m022p141.pdf.

Schmidt:1986:SCB

[SW86a] G. H. Schmidt and G. F. Warner. Spatial competition between colonial ascidians: the importance of stand-off. Marine Ecology Progress Series, 31:101–104, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/31/m031p101.pdf.

Skjoidal:1986:SPO

[SW86b] H. R. Skjoidal and P. Wassmann. Sedimentation of particulate organic matter and silicium during spring and summer in Lindåspollene, Western Norway. Marine Ecology Progress Series, 30:49–63, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/30/m030p049.pdf.

Steffe:1989:HSD

[SWB89] A. S. Steffe, M. Westoby, and J. D. Bell. Habitat selection and diet in two species of pipefish from seagrass: sex differences. Marine Ecology Progress Series, 55:23–30, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/55/m055p023.pdf.

Stegeman:1988:ACP

[SWG88] J. J. Stegeman, B. R. Woodin, and A. Goksoyr. Apparent cytochrome P-450 induction as an indication of exposure to environmental chemicals the flounder Platichthys flesus. Marine Ecology Progress Series, 46:55–60, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/46/m046p055.pdf.

Smith:1984:HPE

[SWH84] D. F. Smith, W. J. Wiebe, and H. W. Higgins. Heterotrophic potential estimates: an inherent paradox in assuming Michaelis– Menten kinetics. Marine Ecology Progress Series, 17:49–56, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/17/
m017p049.pdf.

Scheltema:1981:GSL

[SWSL81] R. S. Scheltema, I. P. Williams, M. A. Shaw, and C. Loudon. Gregarious settlement by the larvae of Hydroides dianthus (Polychaeta: Serpulidae). Marine Ecology Progress Series, 5:69–74, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
5/m005p069.pdf.

Tedengren:1988:ESR

[TAK88] M. Tedengren, M. Arner, and N. Kautsky. Ecophysiology and stress response of marine and brackish water Gammarus species (Crustacea, Amphipoda) to changes in salinity and exposure to cadmium and diesel-oil. Marine Ecology Progress Series, 47:107– 116, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
47/m047p107.pdf.

Tallmark:1980:PDN

[Tal80] B. Tallmark. Population dynamics of Nassarius reticulatus (Gastropoda, Prosobranchia) in Gullmar Fjord, Sweden. Marine Ecology Progress Series, 3:51–62, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/3/m003p051.pdf.

Tamaki:1985:ZSA

[Tam85] A. Tamaki. Zonation by size in the Armandia sp. (Polychaeta: Opheliidae) population on an intertidal sand flat. Marine Ecology Progress Series, 27:123–133, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/27/m027p123.pdf.

Tamaki:1987:CRT

[Tam87] A. Tamaki. Comparison of resistivity to transport by wave action in several polychaete species on an intertidal sand flat. Marine Ecology Progress Series, 37:181–189, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/37/m037p181.pdf.

Tamminen:1989:DOP

[Tam89] T. Tamminen. Dissolved organic phosphorus regeneration by bacterioplankton: 5’-nucleotid-ase activity and subsequent phosphate uptake in a mesocosm enrichment experiment. Marine Ecology Progress Series, 58:89–100, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/58/m058p089.pdf.

Tande:1988:ADM

[Tan88] K. S. Tande. Aspects of developmental and mortality rates in Calanus finmarchicus related to equiproportional development. Marine Ecology Progress Series, 44:51–58, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/44/m044p051.pdf.

Targett:1981:TES

[Tar81] T. E. Targett. Trophic ecology and structure of coastal Antarctic fish communities. Marine Ecology Progress Series, 4:243–263, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
4/m004p243.pdf.

Tegner:1985:DTS

[TB85] M. J. Tegner and R. A. Butler. Drift-tube study of the dispersal potential of green abalone (Haliotis fulgens) larvae in the southern California Bight: implications for recovery of depleted populations. Marine Ecology Progress Series, 26:73–84, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/26/m026p073.pdf.

Talbot:1986:DPC

[TB86] M. M. B. Talbot and G. C. Bate. Diel periodicities in cell characteristics of the surfzone diatom Anaulus birostratus: their role in the dynamics of cell patches. Marine Ecology Progress Series, 32: 81–89, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/32/m032p081.pdf.

Tipton:1988:FPT

[TB88] K. Tipton and S. S. Bell. Foraging patterns of two syngnathid fishes: importance of harpacticoid copepods. Marine Ecology Progress Series, 47:31–41, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/47/m047p031.pdf.

Thresher:1989:EII

[TB89] R. E. Thresher and E. B. Brothers. Evidence of intra- and inter-oceanic regional differences in the early life history of reefassociated fishes. Marine Ecology Progress Series, 57:187–205, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
57/m057p187.pdf.

tenHoopen:1983:PRF

[tBB83] A. ten Hoopen, S. Bos, and A. M. Breeman. Photoperiodic response in the formation of gametangia of the long-day plant Sphacelaria rigidula (Phaeophyceae). Marine Ecology Progress Series, 13:285–289, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/13/m013p285.pdf.

Tornabene:1985:LLC

[TBRBA85] T. G. Tornabene, T. F. Bourne, S. Raziuddin, and A. Ben-Amotz. Lipid and lipopolysaccharide constituents of cyanobacterium Spirulina platensis (Cyanophyceae, Nostocales). Marine Ecology Progress Series, 22:121–125, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/22/m022p121.pdf.

Tester:1981:EIG

[TC81] P. A. Tester and J. D. Costlow, Jr. Effect of insect growth regulator Dimilin(R) (TH 6040) on fecundity and egg viability of the marine copepod Acartia tonsa. Marine Ecology Progress Series, 5:297– 302, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
5/m005p297.pdf.

Trevelyan:1987:LIS

[TC87] G. A. Trevelyan and E. S. Chang. Light-induced shell pigmentation in post-larval Mytilus edulis and its use as a biological tag. Marine Ecology Progress Series, 39:137–144, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/39/m039p137.pdf.

Tutschulte:1988:FBA

[TC88] Theodore C. Tutschulte and Joseph H. Connell. Feeding behavior and algal food of three species of abalones (Haliotis) in southern California. Marine Ecology Progress Series, 49:57–64, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
49/m049p057.pdf.

Tyler:1982:EDE

[TCA82] M. A. Tyler, D. W. Coats, and D. M. Anderson. Encystment in a dynamic environment: Deposition of dinoflagellate cysts by a frontal convergence. Marine Ecology Progress Series, 7:163–178, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
7/m007p163.pdf.

Thayer:1987:URM

[TCH87] G. W. Thayer, D. R. Colby, and W. F. Hettler, Jr. Utilization of the red mangrove prop root habitat by fishes in south Florida. Marine Ecology Progress Series, 35:25–38, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/35/m035p025.pdf.

Tegner:1981:PSR

[TD81a] M. J. Tegner and P. K. Dayton. Population structure, recruitment and mortality of two sea urchins (Strongylocentrotus franciscanus and S. purpuratus) in a kelp forest. Marine Ecology Progress Series, 5:255–268, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/5/m005p255.pdf.

Todd:1981:RSM

[TD81b] C. D. Todd and R. W. Doyle. Reproductive strategies of marine benthic invertebrates: a settlement-timing hypothesis. Marine Ecology Progress Series, 4:75–83, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/4/m004p075.pdf.

Tomasko:1989:EPI

[TD89] D. A. Tomasko and C. J. Dawes. Evidence for physiological integration between shaded and unshaded short shoots of Thalassia testudinum. Marine Ecology Progress Series, 54:299–305, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
54/m054p299.pdf.

Tallmark:1986:SRD

[TE86a] B. Tallmark and S. Evans. Substrate-related differences in antipredator behaviour of two gobiid fish species and the brown shrimp, and their adaptive value. Marine Ecology Progress Series, 29:217–222, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/29/m029p217.pdf.

Tournie:1986:MCA

[TE86b] T. Tournie and H. El Mednaoui. Metal-complexing agents released into the marine environment by the deposit feeder Hydrobia ulvae (Gastropoda: Prosobranchia): characterization and regulation processes. Marine Ecology Progress Series, 34:251–259, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
34/m034p251.pdf.

Tenore:1982:WCA

[Ten82] K. R. Tenore. What controls the availability to animals of detritus derived from vascular plants: Organic nitrogen enrichment or caloric availability? Marine Ecology Progress Series, 10:307– 309, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
10/m010p307.pdf.

Torreton:1989:BBP

[TGA89] J.-P. Torréton, D. Guiral, and R. Arfi. Bacterioplankton biomass and production during destratification in a monomictic eutrophic bay of a tropical lagoon. Marine Ecology Progress Series, 57:53– 67, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
57/m057p053.pdf.

Taylor:1984:FMI

[TH84] P. R. Taylor and M. E. Hay. Functional morphology of intertidal seaweeds; adaptive significance of aggregate vs. solitary forms. Marine Ecology Progress Series, 18:295–302, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/18/m018p295.pdf.

Thistle:1981:NPD

[Thi81] D. Thistle. Natural physical disturbances and communities of marine soft bottoms. Marine Ecology Progress Series, 6:223–228, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
6/m006p223.pdf.

Thingstad:1987:UPO

[Thi87] T. F. Thingstad. Utilization of N, P, and organic C by heterotrophic bacteria. I. Outline of a chemostat theory with a consistent concept of ’maintenance’ metabolism. Marine Ecology Progress Series, 35:99–109, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/35/m035p099.pdf.

Thompson:1984:PRE

[Tho84] R. J. Thompson. Production, reproductive effort, reproductive value and reproductive cost in a population of the blue mussel Mytilus edulis from a subarctic environment. Marine Ecology Progress Series, 16:249–257, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/16/m016p249.pdf.

Thresher:1982:ECD

[Thr82] R. E. Thresher. Environmental correlates of the distribution of planktivorous fishes in the One Tree Reef Lagoon. Marine Ecology Progress Series, 10:137–145, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/10/m010p137.pdf.

Thrush:1986:SHS

[Thr86] S. F. Thrush. Spatial heterogeneity in subtidal gravel generated by the pit-digging activities of Cancer pagurus. Marine Ecology Progress Series, 30:221–227, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/30/m030p221.pdf.

Thomas:1980:TMT

[THSW80] W. H. Thomas, J. I. Hollibaugh, D. L. R. Seibert, and G. I. Wallace, Jr. Toxicity of a mixture of ten metals to phytoplankton. Marine Ecology Progress Series, 2:213–220, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/2/m002p213.pdf.

Thursby:1984:NRS

[Thu84a] G. B. Thursby. Nutritional requirements of the submerged angiosperm Ruppia maritima in algae-free culture. Marine Ecology Progress Series, 16:45–50, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/16/m016p045.pdf.

Thursby:1984:REO

[Thu84b] G. B. Thursby. Root-exuded oxygen in the aquatic angiosperm Ruppia maritima. Marine Ecology Progress Series, 16:303–305, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
16/m016p303.pdf.

Taylor:1985:IBG

[TIS85] G. T. Taylor, R. Iturriaga, and C. W. Sullivan. Interactions of bactivorous grazers and heterotrophic bacteria with dissolved organic matter. Marine Ecology Progress Series, 23:129–141, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/23/
m023p129.pdf.

Tiselius:1989:CAC

[Tis89] P. Tiselius. Contribution of aloricate ciliates to the diet of Acartia clausi and Centropages hamatus in coastal waters. Marine Ecology Progress Series, 56:49–56, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/56/m056p049.pdf.

Taylor:1986:IBZ

[TKP86] G. T. Taylor, D. M. Karl, and M. L. Pace. Impact of bacteria and zooflagellates on the composition of sinking particles: an in situ experiment. Marine Ecology Progress Series, 29:141–155, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
29/m029p141.pdf.

Twilley:1985:NEE

[TKS^{+}85] Robert R. Twilley, W. Michael Kemp, Kenneth W. Staver, J. Court Stevenson, and Walter R. Boynton. Nutrient enrichment of estuarine submersed vascular plant communities. 1. Algal growth and effects on production of plants and associated communities. Marine Ecology Progress Series, 23:179–191, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/23/m023p179.pdf.

Theilacker:1986:UEI

[TKT86] G. H. Theilacker, A. S. Kimball, and J. S. Trimmer. Use of an ELISPOT immunoassay to detect euphausiid predation on larval anchovy. Marine Ecology Progress Series, 30:127–131, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/30/
m030p127.pdf.

Todd:1984:ELA

[TL84] C. D. Todd and J. R. Lewis. Effects of low air temperature on Laminaria digitata in south-western Scotland. Marine Ecology Progress Series, 16:199–201, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/16/m016p199.pdf.

Taguchi:1985:ASC

[TL85a] S. Taguchi and E. A. Laws. Application of a single-cell isolation technique to studies of carbon assimilation by the subtropical silicoflagellate Dictyocha perlae vis. Marine Ecology Progress Series, 23:251–255, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/23/m023p251.pdf.

Turley:1985:DMB

[TL85b] C. M. Turley and K. Lochte. Direct measurement of bacterial productivity in stratified waters close to a front in the Irish Sea. Marine Ecology Progress Series, 23:209–219, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/23/m023p209.pdf.

Taggart:1987:STMb

[TL87a] C. T. Taggart and W. C. Leggett. Short-term mortality in postemergent arval capelin Mallotus villosus. II. Importance of food and predator density, and density-dependence. Marine Ecology Progress Series, 41:219–229, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/41/m041p219.pdf.

Taggart:1987:STMa

[TL87b] C. T. Taggart and W. C. Leggett. Short-term mortality in postemergent larval capelin Mallotus villosus. I. Analysis of multiple in situ estimates. Marine Ecology Progress Series, 41:205–217, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
41/m041p205.pdf.

Turner:1980:STB

[TM80] R. L. Turner and C. E. Meyer. Salinity tolerance of the brackishwater echinoderm Ophiophragmus filograneus (Ophiuroidea). Marine Ecology Progress Series, 2:249–256, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/2/m002p249.pdf.

Tyler:1985:RBD

[TMBG85] P. A. Tyler, A. Muirhead, D. S. M. Billett, and J. D. Gage. Reproductive biology of the deep-sea holothurians Laetmogone violacea and Benthogone rosea (Elasipoda: Holothurioidea). Marine Ecology Progress Series, 23:269–277, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/23/m023p269.pdf.

Tsuchiya:1985:IMH

[TN85] M. Tsuchiya and M. Nishihira. Islands of Mytilus as a habitat for small intertidal animals: effect of island size on community structure. Marine Ecology Progress Series, 25:71–81, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/25/m025p071.pdf.

Tsuchiya:1986:IME

[TN86] M. Tsuchiya and M. Nishihira. Islands of Mytilus edulis as a habitat for small intertidal animals: effect of Mytilus age structure on the species composition of the associated fauna and community organization. Marine Ecology Progress Series, 31:171–178, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
31/m031p171.pdf.

Turley:1986:SST

[TNR86] C. M. Turley, R. C. Newell, and D. B. Robins. Survival strategies of two small marine ciliates and their role in regulating bacterial community structure under experimental conditions. Marine Ecology Progress Series, 33:59–70, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/33/m033p059.pdf.

Tanabe:1988:LVS

[TO88] K. Tanabe and T. Oba. Latitudinal variation in shell growth patterns of Phacosoma japonicum (Bivalvia: Veneridae) from the Japanese coast. Marine Ecology Progress Series, 47:75–82, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
47/m047p075.pdf.

Tully:1989:IGB

[TÓ89] O. Tully and P. Ó Ceidigh. The ichthyoneuston of Galway Bay (west of Ireland). II. Food of post-larval and juvenile neustonic and pseudo-neustonic fish. Marine Ecology Progress Series, 51:301– 310, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
51/m051p301.pdf.

Todd:1985:STH

[Tod85] C. D. Todd. Settlement-timing hypothesis: reply to Grant and Williamson. Marine Ecology Progress Series, 23:197–202, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/23/
m023p197.pdf.

Townsend:1984:CIZ

[Tow84] D. W. Townsend. Comparison of inshore zooplankton and ichthyoplankton populations of the Gulf of Maine. Marine Ecology Progress Series, 15:79–90, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/15/m015p079.pdf.

Taraschewski:1981:DSP

[TP81] H. Taraschewski and I. Paperna. Distribution of the snail Pirenella conica in Sinai and Israel and its infection by Heterophyidae and other trematodes. Marine Ecology Progress Series, 5:193–205, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
5/m005p193.pdf.

Thingstad:1985:FEA

[TP85] T. F. Thingstad and B. Pengerud. Fate and effect of allochthonous organic material in aquatic microbial ecosystems an analysis based on chemostat theory. Marine Ecology Progress Series, 21:47–62, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
21/m021p047.pdf.

Trendall:1989:SPS

[TP89] J. T. Trendall and J. Prescott. Severe physiological stress associated with the annual breeding emigration of Panulirus ornatus in the Torres Strait. Marine Ecology Progress Series, 58:29–39, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
58/m058p029.pdf.

Toudal:1987:ASE

[TR87] Kaisa Toudal and H. U. Riisgård. Acute and sublethal effects of cadmium on ingestion, egg production and life-cycle development in the copepod Acartia tonsa. Marine Ecology Progress Series, 37: 141–146, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/37/m037p141.pdf.

Tan:1989:BSN

[TR89] Tjhing Lok Tan and Hans-Jürgen Rüger. Benthic studies of the Northwest African upwelling region: bacteria standing stock and ETS-activity, ATP-biomass and Adenylate Energy Charge. Marine Ecology Progress Series, 51:167–176, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/51/m051p167.pdf.

Tracey:1988:FRR

[Tra88] G. A. Tracey. Feeding reduction, reproductive failure, and mortality in Mytilus edulis during the 1985 ‘brown tide’ in Narragansett Bay, Rhode Island. Marine Ecology Progress Series, 50:73–81, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
50/m050p073.pdf.

Tremblay:1989:GSI

[TRL89] C. Tremblay, J. A. Runge, and L. Legendre. Grazing and sedimentation of ice algae during and immediately after a bloom at the ice–water interface. Marine Ecology Progress Series, 56:291– 300, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
56/m056p291.pdf.

Townsend:1989:RIL

[TRMF89] D. W. Townsend, R. L. Radtke, M. A. Morrison, and S. D. Folsom. Recruitment implications of larval herring overwintering distributions in the Gulf of Maine, inferred using a new otolith technique. Marine Ecology Progress Series, 55:1–13, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/55/m055p001.pdf.

Taylor:1988:FSP

[TS88a] A. C. Taylor and J. I. Spicer. Functional significance of a partialemersion response in the intertidal prawn Palaemon elegans (Crustacea: Palaemonidae) during environmental hypoxia. Marine Ecology Progress Series, 44:141–147, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/44/m044p141.pdf.

Trutschler:1988:SGA

[TS88b] K. Trutschler and C. Samtleben. Shell growth of Astarte elliptica (Bivalvia) from Kiel Bay (Western Baltic Sea). Marine Ecology Progress Series, 42:155–162, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/42/m042p155.pdf.

Tseitlin:1980:DGD

[Tse80] V. B. Tseitlin. Duration of gastric digestion in fishes. Marine Ecology Progress Series, 2:277–280, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/2/m002p277.pdf.

Tsutsumi:1987:PDC

[Tsu87] H. Tsutsumi. Population dynamics of Capitella capitata (Polychaeta; Capitellidae) in an organically polluted cove. Marine Ecology Progress Series, 36:139–149, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/36/m036p139.pdf.

Thomas:1983:GTA

[TTH83] T. E. Thomas, D. H. Turpin, and P. J. Harrison. γ-glutamyl transferase activity in marine phytoplankton. Marine Ecology Progress Series, 14:219–222, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/14/m014p219.pdf.

Turner:1984:ZFE

[Tur84] J. T. Turner. Zooplankton feeding ecology: contents of fecal pellets of the copepods Eucalanus pileatus and Paracalanus quasimodo from continental shelf waters of the Gulf of Mexico. Marine Ecology Progress Series, 15:27–46, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/15/m015p027.pdf.

Tempel:1980:UID

[TW80] D. Tempel and W. Westheide. Uptake and incorporation of dissolved amino acids by interstitial Turbellaria and Polychaeta and their dependence on temperature and salinity. Marine Ecology Progress Series, 3:41–50, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/3/m003p041.pdf.

Tararam:1981:MFE

[TW81] A. S. Tararam and Y. Wakabara. The mobile fauna — especially Gammaridea of Sargassum cymosum. Marine Ecology Progress Series, 5:157–163, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/5/m005p157.pdf.

Thorman:1982:SOF

[TW82] S. Thorman and A.-M. Wiederholm. Seasonal occurrence and food resource use of an assemblage of nearshore fish species in the Bothnian Sea, Sweden. Marine Ecology Progress Series, 10:223– 229, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
10/m010p223.pdf.

Trotter:1983:FPS

[TW83] D. B. Trotter and J. M. Webster. Feeding preferences and seasonality of free-living marine nematodes inhabiting the kelp Macrocystis integrifolia. Marine Ecology Progress Series, 14:151–157, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
14/m014p151.pdf.

Tunnicliffe:1988:BPD

[TW88] V. Tunnicliffe and K. Wilson. Brachiopod populations: distribution in fjords of British Columbia (Canada) and tolerance of low oxygen concentrations. Marine Ecology Progress Series, 47:117– 128, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
47/m047p117.pdf.

Uchima:1988:GCA

[Uch88] M. Uchima. Gut content analysis of neritic copepods Acartia omorii and Oithona davisae by a new method. Marine Ecology Progress Series, 48:93–97, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/48/m048p093.pdf.

Ueda:1987:SSO

[Ued87] H. Ueda. Small-scale ontogenetic and diel vertical distributions of neritic copepods in Maizuru Bay, Japan. Marine Ecology Progress Series, 35:65–73, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/35/m035p065.pdf.

Ueda:1983:UOC

[UKTA83] H. Ueda, A. Kuwahara, Nl Tanaka, and M. Azeta. Underwater observations on copepod swarms in temperate and subtropical waters. Marine Ecology Progress Series, 11:165–171, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/11/m011p165.pdf.

Underwood:1988:TEF

[UP88] A. J. Underwood and C. H. Peterson. Towards an ecological framework for investigating pollution. Marine Ecology Progress Series, 46:227–234, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/46/m046p227.pdf.

Veldhuis:1985:TPP

[VA85a] M. J. W. Veldhuis and W. Admiraal. Transfer of photosynthetic products in gelatinous colonies of Phaeocystis pouchetii (Haptophyceae) and its effect on the measurement of excretion rate. Marine Ecology Progress Series, 26:301–304, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/26/m026p301.pdf.

Vink:1985:HDC

[VA85b] S. Vink and M. J. Atkinson. High dissolved C: P excretion ratios for large benthic marine invertebrates. Marine Ecology Progress Series, 21:191–195, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/21/m021p191.pdf.

Vangenechten:1983:UAT

[VAF83] J. H. D. Vangenechten, S. R. Aston, and S. W. Fowler. Uptake of americium-241 from two experimentally labelled deep-sea sediments by three benthic species: a bivalve mollusc, a polychaete and an isopod. Marine Ecology Progress Series, 13:219– 228, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p219.pdf.

VanMoorsel:1983:RST

[Van83] G. W. N. M. Van Moorsel. Reproductive strategies in two closely related stony corals (Agaricia, Scleractinia). Marine Ecology Progress Series, 13:273–283, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/13/m013p273.pdf.

VanAlstyne:1986:EPT

[Van86a] K. L. Van Alstyne. Effects of phytoplankton taste and smell on feeding behavior of the copepod Centropages hamatus. Marine Ecology Progress Series, 34:187–190, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/34/m034p187.pdf.

vanderVeer:1986:ISD

[van86b] H. W. van der Veer. Immigration, settlement, and densitydependent mortality of a larval and early postlarval 0-group plaice (Pleuronectes platessa) population in the western Wadden Sea. Marine Ecology Progress Series, 29:223–236, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/29/m029p223.pdf.

VanMoorsel:1988:EMG

[Van88a] G. W. N. M. Van Moorsel. Early maximum growth of stony corals (Scleractinia) after settlement on artificial substrata on a Caribbean reef. Marine Ecology Progress Series, 50:127–135, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
50/m050p127.pdf.

Vance:1988:ESC

[Van88b] R. R. Vance. Ecological succession and the climax community on a marine subtidal rock wall. Marine Ecology Progress Series, 48:125– 136, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
48/m048p125.pdf.

VanAlstyne:1989:ABH

[Van89] K. L. Van Alstyne. Adventitious branching as a herbivore-induced defense in the intertidal brown alga Fucus distichus. Marine Ecology Progress Series, 56:169–176, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/56/m056p169.pdf.

VandenBerghe:1981:DFP

[VB81] W. Vanden Berghe and M. Bergmans. Differential food preferences in three co-occurring species of Tisbe (Copepoda. Harpacticoida). Marine Ecology Progress Series, 4:213–219, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/4/m004p213.pdf.

VanWambeke:1985:BBPa

[VB85a] F. Van Wambeke and M. A. Bianchi. Bacterial biomass production and ammonium regeneration in Mediterranean sea water supplemented with amino acids. 1. Correlations between bacterial biomass, bacterial activities and environmental parameters. Marine Ecology Progress Series, 23:107–115, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/23/m023p107.pdf.

VanWambeke:1985:BBPb

[VB85b] F. Van Wambeke and M. A. Bianchi. Bacterial biomass production and ammonium regeneration in Mediterranean sea water supplemented with amino acids. 2. Nitrogen flux through heterotrophic microplankton food chain. Marine Ecology Progress Series, 23: 117–128, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/23/m023p117.pdf.

vanderVeer:1987:PCN

[vB87] H. W. van der Veer and M. J. N. Bergman. Predation by crustaceans on a newly settled 0-group plaice Pleuronectes platessa population in the western Wadden Sea. Marine Ecology Progress Series, 35:203–215, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/35/m035p203.pdf.

VivesRego:1985:FAP

[VBFS85] J. Vives Rego, G. Billen, A. Fontigny, and M. Somville. Free and attached proteolytic activity in water environments. Marine Ecology Progress Series, 21:245–249, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/21/m021p245.pdf.

Virnstein:1986:CAS

[VC86] R. W. Virnstein and M. C. Curran. Colonization of artificial seagrass versus time and distance from source. Marine Ecology Progress Series, 29:279–288, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/29/m029p279.pdf.

Vanhaecke:1981:ISA

[VCS81] P. Vanhaecke, A. Cooreman, and P. Sorgeloos. International study on Artemia. XV. Effect of light intensity on hatching rate of Artemia cysts from different geographical origin. Marine Ecology Progress Series, 5:111–114, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/5/m005p111.pdf.

Vasquez:1984:DPD

[VCS84] J. A. Vásquez, J. C. Castilla, and B. Santelices. Distributional patterns and diets of four species of sea urchins in giant kelp forest (Macrocystis pyrifera) of Puerto Toro, Navarino Island, Chile. Marine Ecology Progress Series, 19:55–63, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/19/m019p055.pdf.

vanDuyl:1981:ETC

[vDBS81] F. C. van Duyl, R. P. M. Bak, and J. Sybesma. The ecology of the tropical compound ascidian Trididemnum solidum. I. Reproductive strategy and larval behaviour. Marine Ecology Progress Series, 6:35–42, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/6/m006p035.pdf.

Verheye-Dua:1988:SBF

[VDL88] F. Verheye-Dua and M. I. Lucas. Southern Benguela frontal region. I. Hydrology, phytoplankton and bacterioplankton. Marine Ecology Progress Series, 47:271–280, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/47/m047p271.pdf.

Vague:1989:PCB

[VDM89] D. Vagué, C. M. Duarte, and C. Marrasé. Phytoplankton colonization by bacteria: encounter probability as a limiting factor. Marine Ecology Progress Series, 54:137–140, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/54/m054p137.pdf.

vanderVeer:1985:ICP

[vdV85] H. W. van der Veer. Impact of coelenterate predation on larval plaice Pleuronectes platessa and flounder Platichthys flesus stock in the western Wadden Sea. Marine Ecology Progress Series, 25: 229–238, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/25/m025p229.pdf.

vanderVeer:1986:DTR

[vdVB86] H. W. van der Veer and M. J. N. Bergman. Development of tidally related behaviour of a newly settled 0-group plaice (Pleuronectes platessa) population in the western Wadden Sea. Marine Ecology Progress Series, 31:121–129, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/31/m031p121.pdf.

Velimirov:1986:DDM

[Vel86] B. Velimirov. DOC dynamics in a Mediterranean seagrass system. Marine Ecology Progress Series, 28:21–41, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/28/m028p021.pdf.

Verity:1986:GPN

[Ver86a] P. G. Verity. Grazing of phototrophic nanoplankton by microzooplankton in Narragansett Bay. Marine Ecology Progress Series, 29: 105–115, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/29/m029p105.pdf.

Verity:1986:GRN

[Ver86b] P. G. Verity. Growth rates of natural tintinnid populations in Narragansett Bay. Marine Ecology Progress Series, 29:117–126, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
29/m029p117.pdf.

Vermeij:1989:IDA

[Ver89] G. J. Vermeij. Interoceanic differences in adaptation: effects of history and productivity. Marine Ecology Progress Series, 57:293– 305, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
57/m057p293.pdf.

Vanzella:1989:ECL

[VGJ89] A. Vanzella, M. A. Guerrero, and R. D. Jones. Effect of CO and light on ammonium and nitrite oxidation by chemolithotrophic bacteria. Marine Ecology Progress Series, 57:69–76, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/57/m057p069.pdf.

Videau:1987:PPP

[Vid87] C. Videau. Primary production and physiological state of phytoplankton at the Ushant tidal front (west coast of Brittany, France). Marine Ecology Progress Series, 35:141–151, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/35/m035p141.pdf.

Vogt:1985:SPR

[VK85] N. B. Vogt and H. Knutsen. SIMCA pattern recognition classification of five infauna taxonomic groups using non-polar compounds analysed by high resolution gas chromatography. Marine Ecology Progress Series, 26:145–156, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/26/m026p145.pdf.

VanVeld:1988:IGT

[VL88] P. A. Van Veld and R. F. Lee. Intestinal glutathione s-transferase activity in flounder Platichthys flesus collected from contaminated and reference sites. Marine Ecology Progress Series, 46:61–63, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p061.pdf.

vanderMerwe:1987:SFF

[vM87] D. van der Merwe and A. McLachlan. Significance of free-floating macrophytes in the ecology of a sandy beach surf zone. Marine Ecology Progress Series, 38:53–63, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/38/m038p053.pdf.

Viarengo:1988:ICS

[VMM^{+}88] A. Viarengo, G. Mancinelli, G. Martino, M. Pertica, L. Canesi, and A. Mazzucotelli. Integrated cellular stress indices in trace metal contamination: critical evaluation in a field study. Marine Ecology Progress Series, 46:65–70, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/46/m046p065.pdf.

Verhagen:1982:SMP

[VN82] J. H. G. Verhagen and P. H. Nienhuis. A simulation model of production, seasonal changes in biomass and distribution of eelgrass (Zostera marina) in Lake Grevelingen. Marine Ecology Progress Series, 10:187–195, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/10/m010p187.pdf.

Varela:1985:PPB

[VP85] M. Varela and E. Penas. Primary production of benthic microalgae in an intertidal sand flat of the Ria de Arosa, NW Spain. Marine Ecology Progress Series, 25:111–119, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/25/m025p111.pdf.

Vezina:1988:FWD

[VP88] A. F. Vézina and T. Piatt. Food web dynamics in the ocean. I. Best-estimates of flow networks using inverse methods. Marine Ecology Progress Series, 42:269–287, 1988. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/42/m042p269.pdf.

Vanhaecke:1980:ISA

[VS80] P. Vanhaecke and P. Sorgeloos. International study on artemia. XIV. Growth and survival of Artemia larvae of different geographical origin in a standard culture test. Marine Ecology Progress Series, 3:303–307, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/3/m003p303.pdf.

Varadarajan:1982:RCB

[VS82] S. Varadarajan and T. Subramoniam. Reproduction of the continuously breeding tropical hermit crab Clibanarius clibanarius. Marine Ecology Progress Series, 8:197–201, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/8/m008p197.pdf.

Vinogradov:1984:SME

[VS84] M. E. Vinogradov and E. A. Shushkina. Succession of marine epipelagic communities. Marine Ecology Progress Series, 16:229– 239, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
16/m016p229.pdf.

Vranken:1981:ALC

[VTHVd81] G. Vranken, L. K. Thielemans, Carlo Heip, and M. Van-dycke. Aspects of the life-cycle of Monhystera parelegantul a (Nematoda; Monhysteridae). Marine Ecology Progress Series, 6:67–72, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
6/m006p067.pdf.

Wacasey:1987:EVM

[WA87] J. W. Wacasey and E. G. Atkinson. Energy values of marine benthic invertebrates from the Canadian Arctic. Marine Ecology Progress Series, 39:243–250, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/39/m039p243.pdf.

Wahl:1985:MSD

[Wah85] M. Wahl. Metridium senile: dispersion and small scale colonization by the combined strategy of locomotion and asexual reproduction (laceration). Marine Ecology Progress Series, 26:271–277, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
26/m026p271.pdf.

Wahl:1989:MEF

[Wah89a] M. Wahl. Marine epibiosis. I. Fouling and antifouling: some basic aspects. Marine Ecology Progress Series, 58:175–189, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/58/
m058p175.pdf.

Wilson:1989:PRJ

[WAH89b] K. A. Wilson, K. W. Able, and K. L. Heck, Jr. Predation rates on juvenile blue crabs in estuarine nursery habitats: evidence for the importance of macroalgae (Ulva lactuca). Marine Ecology Progress Series, 58:243–251, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/58/m058p243.pdf.

Wallace:1985:SPA

[Wal85a] C. C. Wallace. Seasonal peaks and annual fluctuations in recruitment of juvenile scleractinian corals. Marine Ecology Progress Series, 21:289–298, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/21/m021p289.pdf.

Walline:1985:GLW

[Wal85b] P. D. Walline. Growth of larval walleye pollock related to domains within the SE Bering Sea. Marine Ecology Progress Series, 21:197– 203, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
21/m021p197.pdf.

Winter:1984:QEE

[WAN84] J. E. Winter, M. A. Acevedo, and J. M. Navarro. Quempillen estuary, an experimental oyster cultivation station in southern Chile. Energy balance in Ostrea chilensis 1 . Marine Ecology Progress Series, 20:151–164, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/20/m020p151.pdf.

Ward:1984:RAM

[War84] T. J. Ward. Role of acute metal toxicity in structuring seagrass fauna near a lead smelter. Marine Ecology Progress Series, 17:117– 124, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
17/m017p117.pdf.

Warwick:1988:ACA

[War88a] R. M. Warwick. Analysis of community attributes of the macrobenthos of Frierfjord/Langesundfjord at taxonomic levels higher than species. Marine Ecology Progress Series, 46:167–170, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/46/
m046p167.pdf.

Warwick:1988:ECSa

[War88b] R. M. Warwick. Effects on community structure of a pollutant gradient introduction MEPS 46:149. Marine Ecology Progress Series, ??:??, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/46/m046p149.pdf.

Warwick:1988:ECSb

[War88c] R. M. Warwick. Effects on community structure of a pollutant gradient summary. Marine Ecology Progress Series, 46:207–211, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p207.pdf.

Wassmann:1983:SOI

[Was83] Paul Wassmann. Sedimentation of organic and inorganic particulate material in Lindåspollene, a stratified, land-locked fjord in western Norway. Marine Ecology Progress Series, 13:237–248, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
13/m013p237.pdf.

Wassmann:1985:SPM

[Was85] P. Wassmann. Sedimentation of particulate material in nordasvannet, a hypertrophic, landlocked fjord in western Norway. Marine Ecology Progress Series, 22:259–271, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/22/m022p259.pdf.

Watkins:1986:VSA

[Wat86] J. L. Watkins. Variations in the size of Antarctic krill, Euphausia superba Dana, in small swarms. Marine Ecology Progress Series, 31:67–73, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/31/m031p067.pdf.

Watling:1988:SSF

[Wat88] L. Watling. Small-scale features of marine sediments and their importance to the study of deposit-feeding. Marine Ecology Progress Series, 47:135–144, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/47/m047p135.pdf.

Wallis:1980:GEN

[WB80] G. P. Wallis and J. A. Beardmore. Genetic evidence for naturally occurring fertile hybrids between two goby species, Pomatoschistus minutus and P. Ioszanoi (Pisces, Gobiidae). Marine Ecology Progress Series, 3:309–315, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/3/m003p309.pdf.

Walker:1983:NDM

[WB83a] T. A. Walker and G. D. Bull. A newly discovered method of reproduction in gorgonian coral. Marine Ecology Progress Series, 12: 137–143, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/12/m012p137.pdf.

Weinstein:1983:CEN

[WB83b] M. P. Weinstein and H. A. Brooks. Comparative ecology of nekton residing in a tidal creek and adjacent seagrass meadow, community composition and structure. Marine Ecology Progress Series, 12:15– 27, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
12/m012p015.pdf.

Walters:1986:DPA

[WB86] K. Walters and S. S. Bell. Diel patterns of active vertical migration in seagrass meiofauna. Marine Ecology Progress Series, 34:95– 103, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
34/m034p095.pdf.

Williams:1985:GSD

[WBAN85] S. L. Williams, V. A. Breda, T. W. Anderson, and B. B. Nyden. Growth and sediment disturbances of Caulerpa spp. (Chlorophyta) in a submarine canyon. Marine Ecology Progress Series, 21: 275–281, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/21/m021p275.pdf.

Williams:1981:CCP

[WBJA81] W. T. Williams, J. S. Bunt, R. D. John, and D. J. Abel. The community concept and the phytoplankton. Marine Ecology Progress Series, 6:115–121, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/6/m006p115.pdf.

Wilson:1986:DSM

[WBVS86] J. O. Wilson, R. Buchsbaum, I. Valiela, and T. Swain. Decomposition in salt marsh ecosystems: phenolic dynamics during decay of litter of Spartina alterniflora. Marine Ecology Progress Series, 29: 177–187, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/29/m029p177.pdf.

Wright:1983:PBE

[WC83] R. T. Wright and R. B. Coffin. Planktonic bacteria in estuaries and coastal waters of northern Massachusetts: spatial and temporal distribution. Marine Ecology Progress Series, 11:205–216, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
11/m011p205.pdf.

Williams:1988:NLP

[WC88] S. L. Williams and R. C. Carpenter. Nitrogen-limited primary productivity of coral reef algal turfs: potential contribution of ammonium excreted by Diadema antillarum. Marine Ecology Progress Series, 47:145–152, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/47/m047p145.pdf.

Warwick:1988:MEE

[WCC^{+}88] R. M. Warwick, M. R. Carr, K. R. Clarke, J. M. Gee, and R. H. Green. A mesocosm experiment on the effects of hydrocarbon and copper pollution on a sublittoral soft-sediment meiobenthic community. Marine Ecology Progress Series, 46:181–191, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/46/
m046p181.pdf.

Warwick:1986:SSD

[WCGG86] R. M. Warwick, N. R. Collins, J. M. Gee, and C. L. George. Species size distributions of benthic and pelagic metazoa: evidence for interaction? Marine Ecology Progress Series, 34:63–68, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
34/m034p063.pdf.

Widdows:1984:RIE

[WDS^{+}84] J. Widdows, P. Donkin, P. N. Salkeld, J. J. Cleary, D. M. Lowe, S. V. Evans, and P. E. Thomson. Relative importance of environmental factors in determining physiological differences between two populations of mussels (Mytilus edulis). Marine Ecology Progress Series, 17:33–47, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/17/m017p033.pdf.

Widbom:1988:RBM

[WE88] B. Widbom and R. Elmgren. Response of benthic meiofauna to nutrient enrichment of experimental marine ecosystems. Marine Ecology Progress Series, 42:257–268, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/42/m042p257.pdf.

Wassef:1985:SFP

[WEHF85] E. Wassef, A. Ezzat, T. Hashem, and S. Faltas. Sardine fishery by purse-seine on the Egyptian Mediterranean coast. Marine Ecology Progress Series, 26:11–18, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/26/m026p011.pdf.

Weikert:1982:VDZ

[Wei82] H. Weikert. The vertical distribution of zooplankton in relation to habitat zones in the area of the Atlantis II deep, central Red Sea. Marine Ecology Progress Series, 8:129–143, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/8/m008p129.pdf.

Weinberg:1989:PPA

[Wei89a] J. R. Weinberg. Predicting population abundance and age structure: testing theory with field data. Marine Ecology Progress Series, 53:59–64, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/53/m053p059.pdf.

Weisse:1989:MLR

[Wei89b] T. Weisse. The microbial loop in the Red Sea. dynamics of pelagic bacteria and heterotrophic nanoflagellates. Marine Ecology Progress Series, 55:241–250, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/55/m055p241.pdf.

Woo:1980:SBR

[WF80] N. Y. S. Woo and A. C. Y. Fung. Studies on the biology of the Red Sea bream Chrysophrys major. I. Temperature tolerance. Marine Ecology Progress Series, 3:121–124, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/3/m003p121.pdf.

Wulff:1983:IDT

[WF83] F. V. Wulff and J. G. Field. Importance of different trophic pathways in a nearshore benthic community under upwelling and downwelling conditions. Marine Ecology Progress Series, 12:217–228, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
12/m012p217.pdf.

Warwick:1984:CSE

[WG84] R. M. Warwick and J. M. Gee. Community structure of estuarine meiobenthos. Marine Ecology Progress Series, 18:97–111, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
18/m018p097.pdf.

Williams:1982:SFC

[WH82] D. McB. Williams and A. I. Hatcher. Structure of fish communities on outer slopes of inshore, mid-shelf and outer shelf reefs of the Great Barrier Reef. Marine Ecology Progress Series, 10:239–250, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
10/m010p239.pdf.

Wright:1985:URD

[WH85] D. A. Wright and E. W. Hetzel. Use of RNA: DNA ratios as an indicator of nutritional stress in the American oyster Crassostrea virginica. Marine Ecology Progress Series, 25:199–206, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
25/m025p199.pdf.

Wear:1987:NDC

[WH87] R. G. Wear and M. Haddon. Natural diet of the crab Ovalipes catharus (Crustacea, Portunidae) around central and northern New Zealand. Marine Ecology Progress Series, 35:39–49, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps/35/
m035p039.pdf.

Wikner:1988:ETC

[WH88] J. Wikner and Å. Hagstrom. Evidence for a tightly coupled nanoplanktonic predator–prey link regulating the bacterivores in the marine environment. Marine Ecology Progress Series, 50:137– 145, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
50/m050p137.pdf.

Whitfield:1989:IIM

[Whi89] A. K. Whitfield. Ichthyoplankton interchange in the mouth region of a southern African estuary. Marine Ecology Progress Series, 54: 25–33, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/54/m054p025.pdf.

Williams:1982:DMY

[Wil82] D. McB. Williams. Daily, monthly and yearly variability in recruitment of a guild of coral reef fishes. Marine Ecology Progress Series, 10:231–237, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/10/m010p231.pdf.

Wilson:1983:REP

[Wil83] J. H. Wilson. Retention efficiency and pumping rate of Ostrea edulis in suspensions of Isochrysis galbana. Marine Ecology Progress Series, 12:51–58, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/12/m012p051.pdf.

Wilson:1985:JPS

[Wil85] R. P. Wilson. The jackass penguin (Spheniscus demersus) as a pelagic predator. Marine Ecology Progress Series, 25:219–227, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
25/m025p219.pdf.

Williams:1986:TVS

[Wil86a] D. McB. Williams. Temporal variation in the structure of reef slope fish communities (central Great Barrier Reef): short-term effects of Acanthaster planci infestation. Marine Ecology Progress Series, 28:157–164, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/28/m028p157.pdf.

Wilson:1986:IPI

[Wil86b] W. H. Wilson, Jr. Importance of predatory infauna in marine softsediment communities. Marine Ecology Progress Series, 32:35–40, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
32/m032p035.pdf.

Williams:1987:CBS

[Wil87] S. L. Williams. Competition between the seagrasses Thalassia testudinum and Syringodium filiforme in a Caribbean lagoon. Marine Ecology Progress Series, 35:91–98, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/35/m035p091.pdf.

Williams:1988:DRD

[Wil88] S. L. Williams. Disturbance and recovery of a deep-water Caribbean seagrass bed. Marine Ecology Progress Series, 42:63– 71, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
42/m042p063.pdf.

Wilber:1989:RBD

[Wil89] D. H. Wilber. Reproductive biology and distribution of stone crabs (Xanthidae, Menippe) in the hybrid zone on the northeastern Gulf of Mexico. Marine Ecology Progress Series, 52:235–244, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
52/m052p235.pdf.

Wright:1987:FPM

[WJ87] S. W. Wright and S. W. Jeffrey. Fucoxanthin pigment markers of marine phytoplankton analysed by HPLC and HPTLC. Marine Ecology Progress Series, 38:259–266, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/38/m038p259.pdf.

Widdows:1988:PEM

[WJ88] J. Widdows and D. Johnson. Physiological energetics of Mytilus edulis: Scope for growth. Marine Ecology Progress Series, 46:113– 121, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
46/m046p113.pdf.

Wildish:1988:GRG

[WK88] D. J. Wildish and D. D. Kristmanson. Growth response of giant scallops to periodicity of flow. Marine Ecology Progress Series, 42: 163–169, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/42/m042p163.pdf.

Warheit:1984:PDL

[WLB84] K. I. Warheit, D. R. Lindberg, and R. J. Boekelheide. Pinniped disturbance lowers reproductive success of black oystercatcher Haematopus bachmani (Aves). Marine Ecology Progress Series, 17:101–104, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/17/m017p101.pdf.

Wainwright:1982:EAS

[WM82] P. F. Wainwright and K. H. Mann. Effect of antimicrobial substances on the ability of the mysid shrimp Mysis stenolepis to digest cellulose. Marine Ecology Progress Series, 7:309–313, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
7/m007p309.pdf.

Wolter:1982:BIO

[Wol82] K. Wolter. Bacterial incorporation of organic substances released by natural phytoplankton populations. Marine Ecology Progress Series, 7:287–295, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/7/m007p287.pdf.

Wolff:1988:SRP

[Wol88] M. Wolff. Spawning and recruitment in the Peruvian scallop Argopecten purpuratus. Marine Ecology Progress Series, 42:213–217, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
42/m042p213.pdf.

Weis:1987:BBF

[WP87] J. S. Weis and J. Perlmutter. Burrowing behavior by the fiddler crab Uca pugilator. Inhibition by the insecticide diflubenzuron. Marine Ecology Progress Series, 38:109–113, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/38/m038p109.pdf.

Wylie:1988:FPS

[WP88] C. Ft Wylie and V. J. Paul. Feeding preferences of the surgeonfish Zebrasoma flavescens in relation to chemical defenses of tropical algae. Marine Ecology Progress Series, 45:23–32, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/45/m045p023.pdf.

Whitehead:1989:FSS

[WPS89] H. Whitehead, V. Papastavrou, and S. C. Smith. Feeding success of sperm whales and sea-surface temperature off the Galapagos Islands. Marine Ecology Progress Series, 53:201–203, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/53/m053p201.pdf.

Webb:1987:FMS

[WPW87] P. Webb, R. Perissinotto, and T. H. Wooldridge. Feeding of Mesopodopsis slabberi (Crustacea, Mysidacea) on naturally occurring phytoplankton. Marine Ecology Progress Series, 38:115–123, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
38/m038p115.pdf.

Webb:1988:DFG

[WPW88] P. Webb, R. Perissinotto, and T. H. Wooldridge. Diet and feeding of Gastrosaccus psammodytes (Crustacea, Mysidacea) with special reference to the surf diatom Anaulus birostratus. Marine Ecology Progress Series, 45:255–261, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/45/m045p255.pdf.

Williams:1981:SVA

[WR81] R. Williams and D. Robins. Seasonal variability in abundance and vertical distribution of Parathemisto gaudichaudi (Amphipoda: Hyperiidea) in the North East Atlantic Ocean. Marine Ecology Progress Series, 4:289–298, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/4/m004p289.pdf.

White:1982:RAC

[WR82] S. L. White and P. S. Rainbow. Regulation and accumulation of copper, zinc and cadmium by the shrimp Palaemon elegans. Marine Ecology Progress Series, 8:95–101, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/8/m008p095.pdf.

White:1984:RZC

[WR84a] S. L. White and P. S. Rainbow. Regulation of zinc concentration by Palaemon elegans (Crustacea: Decapoda): zinc flux and effects of temperature, zinc concentration and moulting. Marine Ecology Progress Series, 16:135–147, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/16/m016p135.pdf.

White:1984:ZFP

[WR84b] S. L. White and P. S. Rainbow. Zinc flux in Palaemon elegans (Crustacea: Decapoda): moulting, individual variation and tissue distribution. Marine Ecology Progress Series, 19:153–166, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
19/m019p153.pdf.

White:1986:ACP

[WR86] S. L. White and P. S. Rainbow. Accumulation of cadmium by Palaemon elegans (Crustacea: Decapoda). Marine Ecology Progress Series, 32:17–25, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/32/m032p017.pdf.

White:1987:HMC

[WR87] S. L. White and P. S. Rainbow. Heavy metal concentrations and size effects in the mesopelagic decapod crustacean Systeilaspis debilis. Marine Ecology Progress Series, 37:147–151, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/37/m037p147.pdf.

Wiadnyana:1989:SFA

[WR89] N. W. Wiadnyana and F. Rassoulzadegan. Selective feeding of Acartia clausi and Centropages typicus on microzooplankton. Marine Ecology Progress Series, 53:37–45, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/53/m053p037.pdf.

Wright:1983:TTE

[WRC83] D. A. Wright, W. H. Roosenburg, and M. Castagna. Thermal tolerance in embryos and larvae of the bay scallop Argopecten irradians under simulated power plant entrapment conditions. Marine Ecology Progress Series, 14:269–273, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/14/m014p269.pdf.

Wolter:1984:ILC

[WRKD84] K. Wolter, U. Rabsch, P. Krischker, and A. G. Davies. Influence of low concentrations of cadmium, copper and zinc on phytoplanktor of natural water samples. Marine Ecology Progress Series, 19:167– 173, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
19/m019p167.pdf.

Wrobleswki:1984:FGM

[Wro84] J. S. Wrobleswki. Formulation of growth and mortality of larval northern anchovy in a turbulent feeding environment. Marine Ecology Progress Series, 20:13–22, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/20/m020p013.pdf.

Wilkinson:1983:EFG

[WS83] C. R. Wilkinson and P. W. Sammarco. Effects of fish grazing and damselfish territoriality on coral reef algae. II. Nitrogen fixation. Marine Ecology Progress Series, 13:15–19, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/13/m013p015.pdf.

Witman:1984:MFD

[WS84] J. D. Witman and T. H. Suchanek. Mussels in flow: drag and dislodgement by epizoans. Marine Ecology Progress Series, 16: 259–268, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/16/m016p259.pdf.

Wolaver:1988:CTB

[WS88] T. G. Wolaver and J. D. Spurrier. Carbon transport between a euhaline vegetated marsh in South Carolina and the adjacent tidal creek: contributions via tidal inundation, runoff and seepage. Marine Ecology Progress Series, 42:53–62, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/42/m042p053.pdf.

Wiencke:1989:TRG

[WtD89] C. Wiencke and I. torn Dieck. Temperature requirements of growth and temperature tolerance of macroalgae endemic to the Antarctic region. Marine Ecology Progress Series, 54:189–197, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
54/m054p189.pdf.

Warwick:1980:DBM

[WU80] R. M. Warwick and R. J. Uncles. Distribution of benthic macrofauna associations in the Bristol Channel in relation to tidal stress. Marine Ecology Progress Series, 3:97–103, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/3/m003p097.pdf.

Wikfors:1982:GAE

[WU82] G. H. Wikfors and R. Ukeles. Growth and adaptation of estuarine unicellular algae in media with excess copper, cadmium or zinc, and effects of metal-contaminated algal food on Crassostrea virginica larvae. Marine Ecology Progress Series, 7:191– 206, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
7/m007p191.pdf.

Worrall:1983:PCF

[WW83] C. M. Worrall and J. Widdows. Physiological changes following transplantation of the bivalve Scrobicularia plana between three populations. Marine Ecology Progress Series, 12:281–287, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
12/m012p281.pdf.

Wethey:1986:QSP

[WW86] D. S. Wethey and L. J. Walters. Quantifying spatial patterns of overgrowth in epibenthic communities. Marine Ecology Progress Series, 29:271–278, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/29/m029p271.pdf.

Walker:1988:QSS

[WW88a] D. I. Walker and Wm. J. Woelkerling. Quantitative study of sediment contribution by epiphytic coralline red algae in seagrass meadows in Shark Bay, Western Australia. Marine Ecology Progress Series, 43:71–77, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/43/m043p071.pdf.

Wooldridge:1988:PPI

[WW88b] T. H. Wooldridge and P. Webb. Predator–prey interactions between two species of estuarine mysid shrimps. Marine Ecology Progress Series, 50:21–28, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/50/m050p021.pdf.

Wallace:1986:RJC

[WWB86] C. C. Wallace, A. Watt, and G. D. Bull. Recruitment of juvenile corals onto coral tables preyed upon by Acanthaster plana. Marine Ecology Progress Series, 32:299–306, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/32/m032p299.pdf.

Wolaver:1988:BCE

[WWD^{+}88] T. G. Wolaver, G. J. Whiting, R. F. Dame, T. M. Williams, and J. D. Spurrier. Bly Creek ecosystem study — nitrogen exchange within a euhaline salt marsh basin of North Inlet, South Carolina. Marine Ecology Progress Series, 49:107–116, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/49/m049p107.pdf.

Worrall:1983:PET

[WWL83] C. M. Worrall, J. Widdows, and D. M. Lowe. Physiological ecology of three populations of the bivalve Scrobicularia plana. Marine Ecology Progress Series, 12:267–279, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/12/m012p267.pdf.

Ward:1982:EST

[WY82] T. J. Ward and P. C. Young. Effects of sediment trace metals and particle size on the community structure of epibenthic seagrass fauna near a lead smelter, South Australia. Marine Ecology Progress Series, 9:137–146, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/9/m009p137.pdf.

Williams:1985:ADP

[WYG85] S. L. Williams, S. M. Yarish, and I. P. Gill. Ammonium distributions, production, and efflux from backreef sediments, St. Croix, US Virgin Islands. Marine Ecology Progress Series, 24:57– 64, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
24/m024p057.pdf.

Witte:1983:MTF

[WZ83a] J. IJ. Witte and J. J. Zijlstra. The meiofauna of a tidal flat in the western part of the Wadden Sea and its role in the benthic ecosystem. Marine Ecology Progress Series, 14:129–138, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/14/m014p129.pdf.

Wolaver:1983:EWC

[WZ83b] T. G. Wolaver and J. Zieman. Effect of water column, sediment and time over the tidal cycle on the chemical composition of tidal water in a mesohaline marsh. Marine Ecology Progress Series, 12: 123–130, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/12/m012p123.pdf.

Whitlatch:1985:BIA

[WZ85] R. B. Whitlatch and R. N. Zajac. Biotic interactions among estuarine infaunal opportunistic species. Marine Ecology Progress Series, 21:299–311, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps/21/m021p299.pdf.

Yanez-Arancibia:1988:ETS

[YALD88] A. Yànez-Arancibia and A. L. Lara-Dominguez. Ecology of three sea catfishes (Ariidae) in a tropical coastal ecosystem — Southern Gulf of Mexico. Marine Ecology Progress Series, 49:215–230, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
49/m049p215.pdf.

Yang:1982:TAT

[Yan82] J. Yang. A tentative analysis of the trophic levels of North Sea fish. Marine Ecology Progress Series, 7:247–252, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/7/m007p247.pdf.

Yamada:1984:SAR

[YD84] S. S. Yamada and C. F. D’Elia. Silicic acid regeneration from estuarine sediment cores. Marine Ecology Progress Series, 18:113– 118, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
18/m018p113.pdf.

Young:1986:SSD

[YLH86] P. C. Young, J. M. Leis, and H. F. Hausfeld. Seasonal and spatial distribution of fish larvae in waters over the North West continental shelf of Western Australia. Marine Ecology Progress Series, 31: 209–222, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-

1599 (electronic). URL https://www.int-res.com/articles/
meps/31/m031p209.pdf.

YoneshigueBraga:1980:GAC

[Yon80] Y. Yoneshigue Braga. Growth of Acetabularia calyculus in three different media. Marine Ecology Progress Series, 3:157–161, 1980. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
3/m003p157.pdf.

Yoshioka:1986:CCE

[Yos86a] P. M. Yoshioka. Competitive co-existence of the dorid nudibranchs Doridella steinbergae and Corambe pacifica. Marine Ecology Progress Series, 33:81–88, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps/33/m033p081.pdf.

Yoshioka:1986:LHP

[Yos86b] P. M. Yoshioka. Life history patterns of the dorid nudibranchs Doridella steinbergiae and Corambe pacifica. Marine Ecology Progress Series, 31:179–184, 1986. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/31/m031p179.pdf.

Young:1981:TCV

[You81] P. C. Young. Temporal changes in the vagile epibenthic fauna of two seagrass meadows (Zostera capricorni and Posidonia australis). Marine Ecology Progress Series, 5:91–102, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/5/m005p091.pdf.

Yoshioka:1989:MMA

[YY89] P. M. Yoshioka and B. B. Yoshioka. A multispecies, multiscale analysis of spatial pattern and its application to a shallow-water gorgonian community. Marine Ecology Progress Series, 54:257– 264, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
54/m054p257.pdf.

Zehr:1989:RIN

[ZCF89] J. P. Zehr, D. G. Capone, and P. G. Falkowski. Rapid incorporation of ^{13}NO_{3} by NH_{4}^{−} limited phytoplankton. Marine Ecology Progress Series, 51:237–241, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/51/m051p237.pdf.

Zwarts:1989:VMC

[ZE89] L. Zwarts and P. Esselink. Versatility of male curlews Numenius arquata preying upon Nereis diversicolor deploying contrasting capture modes dependent on prey availability. Marine Ecology Progress Series, 56:255–269, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/56/m056p255.pdf.

Zeldis:1985:EMG

[Zel85] J. R. Zeldis. Ecology of Munida gregaria (Decapoda, Anomura) distribution and abundance, population dynamics and fisheries. Marine Ecology Progress Series, 22:77–99, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/22/m022p077.pdf.

Zeller:1988:STE

[Zel88] D. C. Zeller. Short-term effects of territoriality of a tropical damselfish and experimental exclusion of large fishes on invertebrates in algal turfs. Marine Ecology Progress Series, 44:85–93, 1988. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
44/m044p085.pdf.

Zmarzly:1981:RFT

[ZH81] D. L. Zmarzly and N. D. Holland. Rates of food transport down the ambulacra grooves and through the gut of Comanthus bennetti (Echinodermata: Crinoidea) observed in Situ. Marine Ecology Progress Series, 6:229–230, 1981. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/6/m006p229.pdf.

Zieman:1984:HET

[ZIO84] J. C. Zieman, R. L. Iverson, and J. C. Ogden. Herbivory effects on Thalassia testudinum leaf growth and nitrogen content. Marine Ecology Progress Series, 15:151–158, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/15/m015p151.pdf.

Zaroogian:1983:CUL

[ZJ83] G. E. Zaroogian and M. Johnson. Chromium uptake and loss in the bivalves Crassostrea virginica and Mytilus edulis. Marine Ecology Progress Series, 12:167–173, 1983. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://

www.int-res.com/articles/meps/12/m012p167.pdf.

Zimmerman:1985:LSG

[ZK85] Richard C. Zimmerman and James N. Kremer. In situ growth and chemical composition of the giant kelp, Macrocystis pyrifera: response to temporal changes in ambient nutrient availability. Marine Ecology Progress Series, 27:277–285, 1985. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/27/m027p277.pdf.

Zimmermann:1984:EDD

[ZM84] C. F. Zimmermann and J. R. Montgomery. Effects of a decomposing drift algal mat on sediment pore water nutrient concentrations in a Florida seagrass bed. Marine Ecology Progress Series, 19:299– 302, 1984. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
19/m019p299.pdf.

Zimmerman:1987:GEN

[ZSA87] R. C. Zimmerman, R. D. Smith, and R. S. Alberte. Is growth of eelgrass nitrogen limited? A numerical simulation of the effects of light and nitrogen on the growth dynamics of Zostera marina. Marine Ecology Progress Series, 41:167–176, 1987. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/41/m041p167.pdf.

Zajac:1982:REIa

[ZW82a] R. N. Zajac and R. B. Whitlatch. Responses of estuarine infauna to disturbance. I. Spatial and temporal variation of initial recolonization. Marine Ecology Progress Series, 10:1–14, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/10/m010p001.pdf.

Zajac:1982:REIb

[ZW82b] R. N. Zajac and R. B. Whitlatch. Responses of estuarine infauna to disturbance. II. Spatial and temporal variation of succession. Marine Ecology Progress Series, 10:15–27, 1982. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/10/m010p015.pdf.

Zajac:1989:NDI

[ZW89a] R. N. Zajac and R. B. Whitlatch. Natural and disturbanceinduced demographic variation in an infaunal polychaete Nephtys incise. Marine Ecology Progress Series, 57:89–102, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL

https://www.int-res.com/articles/meps/57/m057p089.pdf.

Zhou:1989:CCC

[ZW89b] X. Zhou and P. J. Wangersky. Changes in copper-complexing organic ligands during spring blooms in the coastal waters of Nova Scotia, Canada. Marine Ecology Progress Series, 53:277– 284, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
53/m053p277.pdf.

Zhou:1989:DCC

[ZW89c] X. Zhou and P. J. Wangersky. Distribution of copper-complexing organic ligands in western North Atlantic waters. Marine Ecology Progress Series, 53:285–291, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps/53/m053p285.pdf.

Zajac:1989:EIS

[ZWO89] R. N. Zajac, R. B. Whitlatch, and R. W. Osman. Effects of interspecific density and food supply on survivorship and growth of newly settled benthos. Marine Ecology Progress Series, 56:127– 132, 1989. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/meps/
56/m056p127.pdf.